A Complete Bibliography of Publications in Technology and Culture: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 September 2023 Version 1.00

Title word cross-reference

e [Kra85, de 82b, Smi86c].

1 [Lan83a]. 100 [Mil88a]. 1000 [Gui84]. 100th [Jac84]. 1100 [Wes86b]. 12 [Slo81]. 1500 [Elb87]. 15th [Coh89]. 1600 [Whi82b]. 1607 [Kel88]. 1640 [Col84]. 1650 [And82]. 1660 [Thi80]. 1695 [Rob80]. 16th [Als81, Oer89]. 1700 [Mye89]. 1736 [Sta87]. 1759/1760 [Sta88a]. 1780 [Bed89c]. 1783 [Sta83a]. 1789 [Bus87, Lan87c]. 1792 [Fol82]. 17th [San84, Mul85]. 17th-Century [Mul85]. 1800 [McK80, Ran82b]. 1800-talet [San81a]. 1804 [Ken86]. 1810 [Bed89b]. 1812 [Pre87]. 1815 [Ung88, Wri83]. 1820 [Har89a, Ste85]. 1825 [Mus83]. 1830s [Mir82]. 1831 [Smi83b]. 1835 [Dan80]. 1840 [Kid85]. 1843 [Ell87]. 1844 [Lav84]. 1848 [Joh83, Nic81]. 1850 [Ben88, Jan86, V88, Sch85b]. 1852 [Rob85]. 1854 [Cha82b]. 1857 [Kam81]. 1858 [Mor87e]. 1860 [Blu82, Car80, Eat88, Fol80b, Gor85a, Gro87a, Haf80, Hin85, Jer88, Lan83b, Lau87, Pre83b, The83]. 1866 [Cas88b, Rei80a]. 1868 [Mor87e]. 1870 [Gei80, Mal86b, Mul84a, Scr87b]. 1875 [Ata82]. 1876 [Por89a]. 1879 [Goe82]. 1880 [Cas88b, Bog85, Fol80a, Tin87]. 1880s [Mor87c]. 1881 [Joh87a]. 1884 [Ola85]. 1885 [Ber89, Lub86a]. 1887 [Dun89]. 1888 [Sha84]. 1889 [Sta89]. 1890 [Cru87, Dan87, Lub87a]. 1891 [Moh87c]. 1896 [Tan81]. 18th [Boy85, KR80, Pri87]. 18th-Century [KR80, Pri87]. 1900 [Lub85, Wri86, Bed83a, Ben86, Cum86, Hok82, Kie89, Lub85, Mac88b, McD86]. 1901 [McG85b]. 1902 [Jac83]. 1905 [Har82b, Hic85a, Rol82]. 1907 [Dan81b]. 1909 [Cro86, Isr87]. 1910 [Jac80, Mar89a]. 1912 [Hil81]. 1913 [Arg88]. 1914 [Gei80, Jer85, Bur86, Cai88, Coo84c, Cro82, Ger84b, Has86, Kli86b, McI84, Ran82c, Ran86a, Smi82b, Sti85b, Tod84a, Tod87, Web81c]. 1915 [Mel85, Nas89, Sti85a]. 1916 [Rot84]. 1917 [Gra88, Wor88]. 1918 [Bil87c, The85b]. 1919 [Moo86]. 1920 [Hal82, Kra86b, Per85a]. 1921 [Sch82a]. 1922 [Fri89a, O’C88]. 1923 [Mey85b]. 1926 [Bro87a]. 1927 [Lev87a]. 1928 [Cha84]. 1929 [Seg89a, Smi85d]. 1930 [Car82c, Che84, Cos85, Gor89a, Law89, Per81b, Rey84, Sin87, Tay89, Tul85]. 1930s [Bus83, Fie84, Hod89, Huf82]. 1931 [Jac86]. 1932 [Bos85, Hun87b]. 1934 [Wik82]. 1935 [Cri80, Mul82]. 1937 [Och88, Sto89]. 1938 [Bil80a]. 1939 [Bri89, Gar81a, Hal89b, Mei82, Ryd89, Sla88, Tuc82a]. 1939/40 [Mei82]. 1940 [Bil87b, Blu86, Cam89, McD85a, Pel83, Smi85a, Vey89, Wal87b, War84, Whi85, Wil87c]. 1941 [Lie81a, Sau80, Sch85a, Smu87]. 1945 [Mor86, Coc81, Gre83, Hac86, Rae80a, Smi84b, Whi86b, Wil85b]. 1946 [All84, All87, Feu88, Tit89]. 1947 [Hal82]. 1948 [Tuc80a]. 1949 [Wal82]. 1950 [Sho88, Dea84, Lew85, Mor89e, Rey87, Rol83]. 1951 [Sin86a]. 1952 [Rab81]. 1953 [Dav89b]. 1954 [Pit86]. 1956 [Fos80]. 1958 [Com86, Kop89]. 1959 [Per80d]. 1960 [Smi88b, Web80a]. 1961 [Bil80a]. 1962 [Hac87]. 1964 [Rea81]. 1967 [Dob81a]. 1968 [Con81a]. 1970 [Bay80, McK86]. 1971 [Rey89a]. 1972 [Fee85, Wri80]. 1974 [Bro83]. 1975 [Sma80, Pyn84, Win82]. 1976 [Com87b, Con82]. 1977 [Gie85, Jac81, Sch81b]. 1978 [KT80, Wer81, Bri81a, Con81b, Dan83b, Gru86, Koh80, Run81, San80]. 1979 [CL80, GK81, HBR^{+}80, Bra83b, Bro80, Dav89c, Sch82d, Sha81c, Tod84b]. 1980 [Bil81a, FBW^{+}81, GS81, Har81b, Jon81, McD85a, Tuc82b, Wri86, Cal85, Dan83a, Dav85, Fro89, Har83a, Hul88, Lan89b, Lev83, McS83, Row82b, Sal84]. 1980s [Pap82, Fie84, Tei82]. 1981 [DR83, HRP^{+}82, Lau82, MD82, SB82, Wos82, Bec86, Mel83a, Mit85, Pin85, Tod85, De 84a, Smi83e]. 1982 [Buc83c, Pfe83, TLK^{+}83, Tom83b, Kli87b]. 1983 [Cas88a, Fra84, Haf85b, Hin84, Kel84c, LS84, LHH^{+}84, Seg87, Tar85a, Wei84, Kie86, Lau86, Mar86d, McM86, Wil89a, Wor86b, Zim84]. 1984 [Buc85, Car85, Hei86, Hob85, Hol87a, LR85, TOE^{+}85, Wor86a, Boy89, Bro86c, Gol87, Gre88b, Mul87]. 1985 [DKP^{+}86, GR86, Mel87, Mor87a, Scr86b, Smi87b, Web85b, Wes87a, Win87, Fla89, Fle88, Jac87, Jac88, Mol88, Sto88]. 1986 [But87, DHW^{+}87, FW87, McC88, War87b, Smi88a]. 1987 [Ano88i, Lew89, Mac88a, Mor89c, Rec88]. 1988 [Ano89n, Mor89d, Sch89]. 19th [Con83b, Ful80, Haf85b, Mar88, San81a, Wal88, Whi83a, Esp82, Eva81, Fli80, Lan87d, Lub87b, Pet83, Ste89]. 19th-Century [Ful80, Wal88, Esp82, Eva81, Fli80, Lan87d, Lub87b, Pet83, Ste89].

2 [Mil82b, Nel87]. 2000 [Fer80b, Fri82a]. 2001 [Seg80]. 20th [Fri86b, Haf85b, Rey86]. 21st [Web85a]. 24 [vB86].

3-D [Hin83].

4 [Pyc82]. 40 [Mei82]. 48 [Smi80b].

57 [Tob85]. 58 [Hab85].

60 [Coo86a].

7 [Buc82a]. 70 [Wis84]. 75 [Hol87b, McM86]. ’78 [KT80]. ’79 [GK81, Fri81c].

’80 [GS81]. ’80’s [De 81b]. 80th [She84]. ’81 [SB82]. ’82 [Pfe83]. ’83 [Kel84c]. ’84 [Hei86]. ’85 [Wes87a]. ’86 [McC88]. ’87 [Rec88].

A.D [Col84, Gui84, Mic84a]. ångans [San81b]. Aarhus [Wer81]. Aaron [Bel82]. Abatement [Fli80]. Abbott [Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89a, McG86]. ABC [Fow82]. Abella [Pal80]. Abir [Dav89c]. Abir-Am [Dav89c]. Aborted [Bro87b]. Abraham [For84, Pow83]. Abrams [Sin82]. Absorption [De 84a]. Abundance [Sti87]. Academic [Car88, War87b]. Academy [Ove85, Tom85]. Access [Lev87c, Ain89]. Accident [Mac87b, Maz84b]. Accidents [Sim86]. Account [San84]. Accounting [McG85a]. Accounts [Sta88a]. Achievement [Win87]. Achse [Bul89]. Acquisition [Mit85]. across [Gar87b, Roe82]. Act [Bil80a]. Acting [De 81b]. Action [Bij88, Lay88]. Activists [Hod89]. Activities [Mei89]. Activity [Rob85, Spi82a]. Adam [Wes83, How80a]. Adams [Per80c, Gar88, Lic85, Row83, TL88]. Adapting [Kie87]. Addington [Buc86d]. Adeline [Pia84]. Adirondack [Wil86a]. Adler [Wei81a]. administration [Smi86d]. Admiral [Mor80a]. Adoption [Bar89b]. Adrian [Mei88b]. Adrift [Cut87]. Advance [LM87, Vey89]. Advances [Wis81]. Adventure [Jac89a]. Advertising [SL89]. Advice [Hol89]. Advisory [Cal85, Com86]. AEG [Kli86b]. Aerial [Lev87a]. Aerodynamics [Whe83]. Aeronautical [Kop89, Luk89, Smi85d]. Aeronautics [Com86, Hol86b, Luk86]. aéronautique [Mor89e]. Aeroplane [Smi81a]. Aerospace [Bil87a, Smi85e, Wri80]. Aesthetes [Kem86c]. Aesthetic [Wil80]. Aesthetics [Kem86a]. AF [Sta86]. Affair [Emm80b, Hol87a]. Affairs [Har81b]. Affleck [Qui89]. Africa [Goo86, Lec88, Hea89, Tod86].

African [Kil88]. After [Fin89, Vin86b, Wei81b, War88, Klu82, Gra88]. Again [Fri85, Kid87]. against [Bee83, How85a, Scl88, How85a]. Age [Bil81b, Bil87b, Cam89, Col84, Cri80, Cru87, Duf80, Eva88, Fer80a, Fri82a, Gol80a, Gol81, Gre88a, Hac82, Hen86, Hin80a, Huy82, Kie87, Kul88, Lan87c, Lev87c, Man81, Mar86c, McD85b, McG88c, Mei87, Par82, Pet88, Pre87, Row82a, Sin88c, Smi81a, Wal82, Yor88, Cur84b, Wes89, Car83b, Cer85, Cor89, Cut81a, Gra87a, Gut85, Hal87, Hea88, Hin81, Hof85, Moh87a, Rum88, Sch82b, Shu84b, Wal87a]. Agenda [Sin89]. Ages [Fit81, Rui80, Bac85a, Bac86, Mul89c, Ung83]. Agonistes [Gol80b]. Agrarian [Est85]. Agribusiness [Jac86]. Agricultural [Ber85, Dan87, De 84d, Pur83, Sch81b, Dan87, Spi82a, Wes86b]. Agriculturalists [Sto89]. Agriculture [Bog85, Hab85, Hea89, Muh81, Wei81b, Wer81, Win82, Mét86]. Ahmad [Gli88]. Ahsan [Mic84a]. AI [Lub86b]. Air [Bil80a, Cla80, Fli89a, Gre83, Hal81, Hal84b, Hal84a, Han88, Hig80c, Hol86b, Kel89, Lev87a, MD82, McM81b, Mor86, Bri81a, Gra87a]. Air-Pump [Han88]. Aircraft [Bil87c, Hal82, Hol87b, Bil82, Han87b]. Airfoil [Vin86a]. Airline [Bri80b]. Airliner [Smi83d]. Airlines [Moo86]. Airmail [TL88]. Airplane [Rol82, Smi83d]. Airplanes [Vin84]. Airway [Dav89b]. Aitken [Hun87b]. Akin [Bla89]. Akira [Ino81]. al [Beb80, Bed80b, Coa83, Cut81b, Dow85, Gli88, Hal80a, Kel80, Smi83b]. al-Hassan [Gli88]. al. [Bed89b, Bed89c, Bel84b, Bra80b, Bri81b, Cha87, Coz89, De 80b, De 84d, Dow85, Ell87, Fee85, Fer88, Fow80a, Fri83, Ger84a, Gie81b, Hac80b, Har89a, Har86a, Har89b, Her88, Hey88, Hil86, Hor88, Hyd86, Hyd87, Kap82, Kil88, Kon84, LK80, Lan83b, Lat86, Ler88, Lyt88, Mar86a, Maz84b, Mel83a, Mil84a, Mil88a, Mol88, New83, Ovi88, Pet88, Por80, Pyc82, Ras80a, Rei80a, Rei87, Ros81, Rot85b, Sch81c, Sch82d, Sin88c, Smi82a, The84a, Tod87, War89a, War82, Wil84d, Wil81, Wol83, vB86]. Alabama [Lew89]. Alain [Hey88, dSP81]. Alamos [Tre85]. Alan [Coa83, Dan87, Eis80a, Gei80, Hil86, LK80, Pur82b]. Alaska [Run84]. Alastair [Buc81b]. Albemarle [Tro84b]. Albert [Dob81a, Eis80b, Eve86, For84, Wil86b, Woo86, Wil84c]. Alberta [Tur85, Tur85]. Alberto [Mar84a]. Albuquerque [Arn89, Jac87]. Aldcroft [Bau86, Gar81a]. Ale [Ful81]. Aleksandr [Kam81]. Alex [Com86, Joh87b]. Alexander [Eva88, Tro84b, How86, Yat81]. Alexandra [Gre81]. Alfred [All87, Bat82, Bri81a, Cro86, Mar89a]. Alice [Kar81, Sla88, War83, McG86]. alla [Bul81a]. Allan [Blu82, Web85a]. Allen [Smi84a, Hew83]. Allied [Mau84]. Allis [Mey88]. Allis-Chalmers [Mey88]. Aloft [Gil85a]. along [Upt88, Wes89]. Alonso [San89]. Alternative [De 82a]. Alternatives [Wor81]. Altshuler [Hil86]. Alva [Rhe82]. Alvares [Rah82]. Am [Dav89c]. Amann [Cam88, Coo83]. Amazing [McL85]. Amazon [Per85a, Wei87b]. Ambassador [Kea89]. Ambroise [Kra88a]. America [Abb86, Bus87, Cor85, Dav87, Dun89, Eat88, Fol82, Gal87, Gor89a, Gre88c, Hac80a, Hin80b, Hof85, Hor88, Hou83, Hyd88b, Hyd88a, Ino82, Kam81, Kap82, Lan88, Lan89b, Lew85, Mei87, Mir82, Moo86, Mul84a, Pas85, Qui89, Rei87, Sch81b, Scr87b, Sla89b, Tuc82a, War84, Whi81, Wil87c, Wri85a, Ano83f, DR83, Hai84, Lub85, Pur81, Sha89b, Sin86b, Ste89, Twe86, Bla88, Bur83, Coc82, Cor87, Dub88, Gil85a, Hil85, Hod89, How84, Jaf84, Joh88, Keh84, Mar87b, Maz84a, Mei88a, Mei88c, Mil86, Pur83, Sie81, Soc89, Sti86b, Sti87, Upt83a, Wel87b, Whe83, Win89, Yat82, Yor88, dlB88, Fle88, Kem84b, Lub85]. American [Ada89, Ata82, Bar88, Ber83, Bez85, Bla81, Bli82, Bow87, Bri81c, Bro87a, Cas88b, Con80a, Con85b, Coo85, Cor88a, Cow83, Dav89b, Dav85, Don86a, Doo87, Dou84, Eng87, Fah89, Fol80b, Fri80a, Fri89a, Fri81b, Fri86c, Ger84a, Goo86, Goo89, HH87a, Hac88, Han85, Har82a, Hev89, Hil84, Hin86b, Hob85, Hof80, Hug88b, Hun87a, Hun87b, Hyd89, Ino82, Isr87, Jac83, Jac89a, Jer88, Kie89, Kli86c, Lau87, Les86b, Mar85a, Mar83b, Mar83a, Mar86c, McD84, McL85, McM86, Mei81, Mer81b, Mey86, Mic84b, Mil89, Mor80a, Mul84a, Mul83b, Ola85, Ovi86a, Pap82, Per81b, Per85b, Por89a, Por89b, Pre87, Pur82a, Pur84b, Qui89, Reu89, Rhe82, Rhe88, Ros85, Rot82, Row82b, Row89, San80, Sch81b, Sch85b, See87, Seg89b, Sei80, Sha81c, Sla89a, Sla89b, Smi83b, Sta83a, Sti81, Sto89]. American [Tho84, Tra85, Tur89, Upt88, Vey89, Wal87a, War88, Wes87b, Wil88, Woo86, Wri85a, Wri86, Car85, Doo87, Fri85, Hok80, Hou80b, McD85b, Nic80, Sch81a, Smi86b, Sta82a, Vin84, Bat80, Bla87, Con82, Fit83, Gro87a, Hud86, Kar87a, Rei80b, Rot83, Sha83a, Whi85]. Americana [Smi83b]. Americans [Hev86, Rol82, Rae80a]. Americas [Hin81]. Ames [Com87b]. among [Fri86c]. Amos [Coo85]. Amoskeag [Hof80]. Amritananda [Wei81a]. Amsterdam [Sch82d, Mul85]. Analysis [Dur83a, Far80, Lue86, Tei88, Vin82, Coa83, Ste80]. Analytical [Gri81a]. Analytische [Bra82b]. Analyzer [Owe86]. Anatomy [Isr87, Bra80b]. Anchorage [Wos84]. Ancient [Har89b, LK80, Phi83, Ste88b, Wil84b, dM85, Had80a, Pet89, Tom85]. Ancients [Kin81]. Andalusian [Mye89]. Andean [Lec84]. Anderson [Bil80b, De 84d, Mor87e, Sal84]. Andors [tZS82]. André [Fro87]. Andrew [Hil85, Wes86b, Reu85]. AnElissa [Fie84]. Anesthesia [Mar87b]. Angela [Tan82a]. Angele [Ble83]. Angeles [Bar89a, Pow83]. Anglo [Bul82]. Anglo-Norman [Bul82]. Animal [Rit88]. Animated [Cha84]. Anita [Bed83a, Had80a]. Ann [Bha87, Con88b, Gar87b, Upt88]. Anna [Har89b]. Annali [Bed80a]. Anne [Arg88, Pla86b, Tei88]. Anniversary [Jac84]. Annotated [Bee85, Bul86b, Edg84, Gre88a, Kra88b, Och89, Pin85]. Announcement [Ano81b, Ano82b, Ano83b]. Announcements [Ano83c]. Annual [Ber80, Car82a, Fra84, Fre80, Har86a, Lau86, Smi88a, De 80a]. Anomalous [Sch81a]. Anonymous [Kel84b]. Ansätze [vW83]. Anteriores [Cum86]. Anthony [Bed87, Blu86, Hev89, McC85, Met83, Top80, Tro84a, Wri86, dM85, Hea89]. Anti [Per86, Pit86]. Anti-Biography [Per86]. Anti-Submarine [Pit86]. Antique [Goo85a, Hok83]. Antiquities [Bed85a]. Antiquity [Hab85, Och89, Sim85, Ste88b, Zee86]. Anton [Asp84]. Antonio [Joh87a]. Apollo [All86, Hal80a, Sha81b, Zig85, Zig88, Gol80b]. Apollo-Soyuz [Sha81b]. Apostle [TL88]. Appalachian [Tul85]. Appel [Ger81]. Apple [Och89]. Applebaum [Fre82]. Application [Goo80a]. Applications [Web85a]. Applied [Str85b]. Appraisal [Man83, Cou88, Gre81]. Appraising [Lyt88]. Apprentice [Yor88]. Approach [Bro85a, Fow80b, Mor89b]. Approaches [Mei89]. Appropriate [Gre81, Lew80, Wei81a]. April [But87, Hob85, Jon81, Lau82, Lew89, Mac88a, Smi83e, War87b, Wei84, Wos82, Ano80i, Ano81h, Ano82g, Ano83j, Ano84h, Ano85i, Ano86k, Ano87i, Ano88l, Ano89k]. Arab [aH80]. Arabic [SS84]. Arbeitsteilung [Web84]. arbetsmetodik [San81b]. Arch [Dan80]. Archaeological [Hab85]. Archaeology [Buc82c, Buc83a, Buc86b, Jan86, Lin81b, Har80a]. Archeology [Atk89, Fra84, Lan83a, LR85, Pen82]. Archibald [Elb87]. Archimedes [Sim87, Sim89]. Architect [Con88a, Hal88b, Kle87, Hal88a]. Architector [She81b]. Architects [Con86a, She81b]. Architectural [The87]. Architecture [Ada84, Con82, Con86b, Cou88, Gar88, Kap80, Lev88, Mor87a, O’B81, V88, She81a, Zuk80, Big88, Con80a, Con81d, Mar87a, Eat88, Mar84a]. Archives [Fri80b]. Arctic [Mar88]. Arditti [Rot85b]. Area [Kru85, Wil80]. Areas [Dan85, Far80]. Argaman [Tra89]. Arguments [Sha89a]. Aridity [Hun87a]. Arithmetic [Coh89]. Arizona [Cro83]. Arkwright [Hew87]. Armageddon [Dea85]. Armed [Gui84]. Arms [Woo86]. Armstrong [Coa83, Tay89, Rot84]. Army [Cai82a, Cal85, Cas85, Cas86, Dob81a, Dob81b, Goe82, Hal89b, Lev83, Rot84, Sma80, Sta83a, Sti84, Swe81, Reu85, Cas85]. Arne [Lun88]. Arnold [Cow86, Han87c, Huy82, Kam81, Mac86b, Rei87, Rei89, Zig85, Ber87, Mye87]. Arrive [Lan83c]. Arrow [Hig82]. Ars [Fle84]. Art [Bur82, Con80b, Eis80b, Fah89, Koe84, Lan85, Lev88, Lew80, Mei84, Sal84, Tra85, Zee82, Bed80b, Jac89b, Atk89, BM85, Fri81a, Had80b, Mei88a]. Art/Blei [Zee82]. Arther [Eva88]. Arthur [Arl84, Bow87, Chr89, Fli89b, Gie85, Har82b, Har87b, Hin86b, McG88b, Mer88, SPV87, Whi85, Mol84a]. Artifact [Wor86a]. Artifacts [Seg87, Tho84]. Artificial [Jas87, Lub86b, Wil87b]. Artillery [Sim87]. Artisans [Hin80a]. Artistic [Wri86]. Arts [Bah87, Con88a, Dav87, Fle84, Har81b, Mor87c, Sta88a, Wil81, Hal88b, Spr81]. Asa [Bry81, Car83b, Hin86a]. Ashby [She80b]. Asher [Rum87]. Ashes [Mil80b, Sla88]. Ashworth [Str88]. Asia [McL84, Muh81, De 84d, Far80]. Asian [Rut89]. Aspects [Mor81a, Swe88a, Uss88]. Assembly [Sta83c]. Assessment [Coa83, KR80, Rei80a, TMMY84, Fri86a, Por80]. Assessments [Coa83]. Assistant [LGWK80]. Associates [Baa89]. Association [Car85, Hob85, Mor87a, San84, Str85b]. Astrolabe [Bed87]. Astrolabes [Bed87, Ovi86a]. Astrolabes/Astrolabe [Bed87]. Astronautical [Sha81c]. Astronautics [Hol86b, Han89c]. Astronauts [Cor85]. Astronomical [Bed89c]. Astronomy [Edg84, War89b, DeV85, Hir85, Kin81]. Astrophysics [Edg84]. Atchison [McD86]. Atlanta [Cri80, McC88, Rec88, Wes87a, Hei86]. Atlantic [Bul86a, Joh89, Mar88]. Atlantis [Big88]. Atmosphere [Kop82].

Atom [Hac87, SF82]. Atomic [Arn89, Boy89, Dea84, Hac88, Maz82, Wal87a, Wan89, Wal89]. Attitudes [Ovi86b, Wil87c]. Aubrey [Row80a]. Audience [Bir87]. Audio [Per80a]. Audio-Visual [Per80a]. Audrey [Ger81, Law83]. Auffay [Ben88]. Augarten [Hen87]. August [Coc81, FW87, Haf85b, Pur80, Sch82d]. Augustine [DeG87]. Ausonius [Sim83]. Austerman [Joh87a]. Austin [Sau80]. Australia [Gar87b, Bha87, Roh84]. Austrian [Hym84]. Auswirkung [Neu88]. Authentication [Wil81]. Authenticity [Sim83]. Authority [Seg86, Smi80b, Han87a]. Auto [Rae80a]. Auto-Camp [Rae80a]. Autobiography [Bri85, Luk89]. Automata [And82]. Automatic [Han87a]. Automation [Nel81, Rum88, Les86a]. Automobile [Bar89a, Bel84b, Che84, Cor89, Har82a, Hil84, Hil86, Hyd89, Les86b, Rae87, Cri80, Hyd87]. av [San81a]. avait [Day89]. Avebury [Row80a]. Aveni [dM85]. Averch [Tei88]. Avery [McG84]. Aviation [Bil80a, Bil88, Cro83, Dav89b, Han89b, Lan85, Lew85, Bil81b, Hig83a, Lan87b, Smi81a, Hal89b, Tri83]. Aviv [Kel86]. Award [Ano88d, Ano89c]. Awareness [Cal85, Ste89]. Axel [PS82, Spi82a]. Axes [Fri89b]. Azudes [Cum86]. Azzaroli [Wel87a].

B [Aga82, Bat80, Bed85a, Bir81, Bro88b, Con87, Gar81b, Gei80, Hal86b, Hew83, Hin80a, Hin81, Hol87a, Jac83, Jer85, Kas84, Law83, Les86b, Lic85, Lub86a, Mau84, McG86, McI81, McL84, Mor80c, New83, Ola85, Ros87b, Sin88c, Wal82, Wal87b, Wal81, Whi86b, dlB88]. B.C [Col84, Mye89, Rol89]. B.C. [Gui89]. Baark [Qur86]. Babbage [Bül88, Tro84a]. Back [Bry80, Bry82, De 82a, Tro87, Fos82]. Back-to-Nature [Bry80, Bry82, De 82a]. Backyard [Hac88]. Bacon [Con88a]. Badash [Hac80a]. Bailes [Lie81a, She87]. Bailey [Och89]. Bajío [Sel88]. Baker [McG84, Con81c, Mil88b, Smi85a, Wes87b]. Ball [Pal81]. Balloon [Gil85a, Zie89a]. Ballooning [Lan85]. Baltimore [Fre89, Pre87, War87b, Fri81b]. Ban [Web80a, Wal83]. Banū [Kel80]. Bang [Whi82a]. Banham [Big88]. Banjeli [Kil88]. Barbara [Bin89, Coo89, Das85, Dor83, Fri88b, Lau87, Mar85a, McG86, Nei88, Per85a, Sta85a]. Barcelona [Scr86b, Gli86]. Bardou [Bel84b]. Barker [Fos89]. Barlett [dlB88]. Barnard [Lan86]. Barnes [Pye84]. Baron [Bla89]. Barraclough [War86b]. Barrett [Che84, O’C88, Way89]. Barrie [Bry81]. Barrington [Fow82]. Barry [Pye84]. Barth [Neu88]. Barton [Tit89]. Barty [Gre88b]. Barty-King [Gre88b]. Basic [Hod81]. Basin [Hab85, Wea83]. Basis [Dur83a]. Bass [Mar86c]. Batch [Nuw88]. Bates [Fle88]. Bath [Sta88a]. Baton [Kra81c, Pos82]. Battery [Sch81a]. Battle [Cas88a, Hal81, Num88]. Battles [Bac85a, Est85]. Baudoin [Sch82c]. Bauer [Vol88]. Baum [Pri82]. Beacons [Bil80a]. Bear [SB80]. Beatrice [Hun84]. Beaune [Kra85, Kra85]. Beauty [How85b, Fla89]. Beaux [Con88a]. Beaux-Arts [Con88a]. Bechtel [Pur89a]. Becker [Rol83]. Beckett [Buc86a, Emm82]. Bedini [Bru88, How87, McC85]. Beeckman [Mul84b]. Beetle [Whi84b]. before [Bed83a, Boy88, Mac88b, McI84, Smi83b, The83, Tuc80a, War86b, vS80, Cha84]. Beginnings [All87, Col84, Hac87, Kou81, May83]. Behavior [Fow80a]. Behavioral [Tar81]. Behram [Dow85]. Behrens [Kli86b]. Being [Dan81b, Cut81a]. Beiträge [Bra83b, Fis82, Kra81a, vW83]. Bela [Wol83]. Belasco [Rae80a]. Bell [Bro87a, Fli89c, Gre88c, Isr87, Kli86a, Maz84a, Hod81, McG83]. Bells [RBD84, Wes87c]. Bellwood [Bly88]. Belt [Sau86, Dib81]. Beniger [Cos88]. Benjamin [Bed89b, Hal84b, Har89a, Ken86, McM81a, Smu85, Per84b, Ket85]. Bennett [Car82c, War89b]. Bensman [Pre87]. Benson [Lif85, Vig83, Mac86a, Ran82a]. Bentley [Mer86]. Bereich [Web85b]. Berg [Kul88, Nic81]. Bergbau [Mul89a]. Bergbaus [Wer82]. Berger [Wik81]. Berghahn [Tod84b]. Bergshistoria [San81b]. Berkebile [Ryd80]. Berkel [Mul84b]. Berkshire [Ber89, McG85a]. Berlin [Bra83b, Lev88]. Bern [Bed89a]. Bernard [Bur87, Jak87, DeV88]. Bernardo [Bru88]. Bernd [Rap81]. Bernice [Lew82a]. Bert [Kel84b]. Bertolotti [Bro85a]. Bertrand [Hug88a, Sle81, Smi81b, Dau81]. Beschloss [Hol87a]. Besieged [SF82]. besoin [Day89]. Bessemer [War86b]. Best [Cas88b]. Beth [Aga82]. Bethlehem [Cas88a, Hyd86]. Betsy [Atk89]. Better [Tra85]. Betts [Ras80b]. between [Das85, Dub83, Gar81a, Gre88b, Ino82, Joh89, LeM82, Mir82, Smu87, Vin82, Rei83b]. Beverly [Spr81]. Beville [Bir87]. Bewertung [Web87b]. Beyer [Leh83]. Beyond [Les82, Tre85, Tre87, Bol88, Hod89, Koh80, Kop82, Kro86, Sta83b]. Bezilla [Con81a, Seg83]. Biblical [Tra89]. Bibliographic [Och89]. Bibliographical [Wri86]. Bibliography [CRA^{+}89, CRM84, CRM85, CRA86, CRA87, CRA^{+}88, Goo82a, Goo80b, Goo81, Goo82b, Goo83, Maa81, Man89, Pin85, Run81, Smu85, Och89, Bee85, Bul86b, Con85d, Edg84, Gre88a, Kra88b, Rei80b]. Bibliotheca [Fer82]. Bibliothek [Mis87]. Bicentennial [Rei80b]. Biennial [Scr86b]. Big [Cas88a, Fli89a, Run84]. Bijdragen [vB86]. Bijker [Coz89]. Bill [dB80]. Billingham [Str85a]. Billington [Con80b, Mei84]. Bilstein [Cor85, Smi85d]. bin [Kel80]. Binary [Asp84]. Bing [Blo81, Lev88]. Binnie [Car84b]. Biographical [Och89, Pre81, Fri80a]. Biography [Fer88, Mah82, McG86, Per86, DeV88, Har87a, Ket85, The84b]. Biological [Ham88, Hoc88, Mét86]. Biology [Mét86]. Birdzell [Kul88]. Birmingham [Lew82a, Lew89, Vol87, Str85a]. Birth [Sla88, Tut81, War86b, Wil84d]. Birthday [She84]. bis [Mul89b]. Bisenzio [Wes89]. Bit [Hen87, Hen87]. Black [Ale85, Bin89, Har84, Mor87e, Rei80b, Roe85a, All86, Roh84]. Blacker [Smi85b]. Blacksmith [Mul82]. Blacksmithing [Lig87]. Blake [Hof84]. Blaming [McD82]. Blank [Rot85b]. Blätter [Web87a]. Blei [Zee82]. Blériot [Mor89e]. Blister [War86b]. Bloodletting [Ger81]. Blooms [Kil88]. Blowing [Web80a]. Blue [Fre82, Tra89]. Blunt [Top80]. Bo [San81a]. Board [Cal85]. Bob [Nel87]. Bobrick [Vig83]. Bochum [Web81c]. Bockstoce [Mar88]. Bodnar [McD85a]. Bolivian [Bur80b]. Bologna [Bul81a]. Bolter [Cer85]. Boltzmann [Bru85]. Bomb [Dea84, Wal87a, Wan89]. Bombard [GDNCH88]. Bombs [Hac88]. Bonavia [Dav89a]. Bond [Ste89]. Bonds [Hof85]. Bonelli [Bed80a, Bed82]. Bonfires [Bil80a]. Bonneville [Rey89b]. Book [Ano84b, Ano85b, Ano86b, Ano88b, Bed83c, Hou80a, Kel80, LGWK80, Ras80b, Sta88c, Bra81b]. Books [Hor84, Ovi85, She81b]. Boom [Per85a]. Boomers [Dub88]. Booms [Pre83a]. Boorstin [Coh86]. Booth [Hic85b]. Booty [Kra81b]. Borch [Hal86a]. Borchet [Cut81a]. Borgmann [Wil86b]. Boris [Roe89, Mei88a]. Borisovich [Kam81]. Born [She81a]. Borrichii [Hal86a]. Borromini [Top80]. Boryczka [McD85a]. Bos [Sch82d]. Boserup [McN83]. Boss [Rae85]. Bosses [Hul88]. Boston [Baa89, Bla84, Hil81, LR85, Lub86c]. Botanic [De 81a]. Bottled [Bel84c]. Bottles [Bar89a]. Boulton [Mus83]. Bound [Con88b, For88a]. Boundaries [Rey86, Scr88b]. Bourgeois [Ols89]. Bowater [Smi83a]. Bowers [Hal82]. Box [Har84]. Boy [Hor84]. Boyer [Wal87a]. Boyle [Han88, Lat86]. Boyne [Bil81b, Smi81a, Hig82]. Bradford [Cas88b, Cla86]. Bradley [Mor80a]. Branch [All84]. Brandenburg [Hey88]. Brannigan [DeG87]. Brass [Ger84a]. Brattleboro [Coo87]. Braun [Car84a, Hev84, Por80, Hoc85]. Bray [Mét86, Rut89]. Breakers [Kah85]. Breaking [Tin87]. Brecher [Ger84a, Kin81]. Breeder [Maz82]. Breeding [Rit88]. Brenner [Maz83]. Brentjes [Sch81c]. breview [Bur82]. Brewing [Mei81]. Brewster [Kas83, Mor87e]. Brian [All88, Has86, Sem84]. Brians [Boy89]. Bridbury [Jer85]. Bridge [BM84a, BM84b, Bro88a, Bry81, Con81e, Con84, Dor80, Hin84, Jac84, Kea89, Kou82, McI81, Mei84, Pet83, Wat81, Wos84, Bry81, Hin84, Mar86a, Wat80, Mar86a]. Bridges [Con80b, Dan80, Law89, Mar86b, Mil89, Pet89, Bil83, Kem86a, Mil83]. Bridgewater [Tan89b]. Bridson [Wri86]. Brief [Bul81c, Hil85, Tur89]. Briggs [Bry81, Car83b, Hin86a]. Bright [Bha87, Wri80]. Brimblecombe [Fli89a]. Bringing [Whe83]. Bristol [Sta88a]. Britain [Bri89, Buc89, Cro86, Gre88b, Mir82, Tuc80a, Wri86, Eva81, Fli80, KR80, Kul88, LeM82, Nic80, Str85a, Buc86a, Emm82, LeM86, Mil84b, Rhe86, Tan89a]. Britannia [Dor80]. British [Bed85a, Bos85, Bow80, Buc88b, Bus87, Car80, De 81a, Gor88b, Har85, Has86, Jer85, Lie83, Lin81b, Mor87d, Sal84, Sau80, Smi85e, Ste81b, Str88, Wil83, Buc86c, Pec86, Sha89b, Gar81a, Ran82a]. Briton [Gor86]. Broadcasters [Wik82]. Broadcasting [Fri89a, Kie86]. Brock [Hil84]. Brockway [De 81a]. Broda [Bru85]. Bromberg [Whe84]. Bronowski [Cha80]. Bronwen [Smi80a]. Bronx [Wos82]. Bronze [Col84, Gre88a]. Brooches [You87a]. Brooke [Jen84, Jer88, Kas84, Sch82b, Woo82]. Brooklyn [Hin84, LS84, Mar86a, Mei87, Mel85, Jac84, Wos84]. Brooks [Hal80a]. Brothers [Cro86, Gar81c, Lan85, Rol80, Rig81, Ste88a]. Brow [Tan82a]. Brown [Tro84b, Fer83, Fol82, McK80, Pen84, Swe81, Wel89, Whi81]. Bruce [Ame89, Bal85, Cor88a, Joh88, Koe84, Kro86, Per80a, Reu89, Row82b, Por89a]. Bruchey [Bla81, Por89b]. Bruna [Ada84]. Brunel [Emm80b, Emm82]. Brunilde [Gui89]. Bruno [Bij88, Hal80a]. Brunot [Smi86d]. Bruton [And81]. Bubble [Zuk80]. Buchanan [Emm83b, Mil80a]. Bucher [She81b].

Buchwald [Swe88a]. Buckland [Vac84]. Buckminster [Seg87]. Bud [Rhe86]. Budapest [Bro88a]. Buddensieg [Kli86b]. Buedeler [Ree81]. Buffalo [Sti84]. Buford [Gia81]. Bugliarello [Mor80c]. Build [Klu82]. Builders [Dan80, Gra88, Leh83, Mar83a, Met83]. Building [Big88, Cro82, Dob81a, Fri81b, Iri89b, Lan85, Mor89b, See87, Ste81a, Sti85a, Wor88, Eve86, She80a, Boy88, Mar86b, Mil88b, Reu89, Upt83a]. Buildings [Con81d, Hud86, See87, Wri85a, Kle87]. Bull [Sta82b]. Bullets [Rot84]. Bundesrepublik [Bra82a]. Bunker [Wei87b]. Burchard [Sch81c]. Bureau [Don81a, Jac81, See84]. Bureaucracy [Seg86]. Bureaucratic [Joh88]. Bureaucrats [Rot84]. Burgess [Row82a]. Burke [Cut81b, Gra87b, Buc81a]. Burl [Row80a]. Burnett [Bay80]. Burns [Gie81b]. Burton [Met83]. Bush [Owe86]. Bushworkers [Hul88]. Business [Bla81, Bra81b, Bro87a, Bro88b, Cas88a, Con89, Coo85, Eng87, Gut84, Isr87, Lub86b, Maz86b, McG85a, Seg89a, Jas87, Coc81, Col81]. Butt [Lin81b]. Buttenwieser [Con88b]. Butti [Kap80]. Buxton [Gar81a]. Byatt [Bow80]. Bynum [Lan84]. Byron [Bac82a].

C [Bed89b, Bow80, Bri81a, Bro86c, Bur83, Car88, Cas88a, Cas88b, Cha82b, Con81f, Cou86, Cut81b, Dan83a, Emm80b, Eng87, Fle88, Gie86, Gli89, Gor86, Gre84, Gru86, Hal80a, Han89a, Har89a, Hin81, Hor88, How86, Jac81, Kan85, Ken86, Koh80, Lav84, Lew80, Maz84a, McD82, Mel83b, Mil89, Mol88, Mor89a, Mor81b, Mor87e, Mye89, Pur84b, Ran82b, Rea81, Rot82, Saw88, Seg86, Sha81b, Sha83a, Sha83b, Sin88a, Smi84b, Soc89, Ste81b, Tit89, Tre87, Tre85, Ung88, War86b, Wil81, Wis81, Buc82a]. ca [Ada84, Col84, Ran86b, Vin84]. Cable [Rei80a, Bir87]. Cadillac [Row89]. Cady [Bil81a]. Cahalin [Ano80f, Ano81e, Ano82e, Ano83h, Ano84e, Ano85f, Ano86f, Ano87e, Ano88g, Ano89g]. Calculating [Jaf84]. Calculus [Mar87b]. Caldwell [Pap82]. Calgary [Tur85]. California [Hob85, Jac86, Jac88, Kah85, Nas89, Pau83, Sho88, Hob85, Neu89, Whi83c]. called [Kel84b]. Cambridge [LR85, Mul89c, TOE^{+}85, The83]. Came [Keh84]. Camp [Rae80a]. Campaign [McG85b]. Can [Jor89, Law85]. Canada [All84, Rab81, Sin86a, Luk86, Mil80b, Mar83b]. Canadian [Mac86b, Pal80, Pal81, SPV87, Tay89]. canadienne [Pal81]. canadiennes [Pal81]. Canal [Klu82, Met83, Pia84, Zim84, Bau86, Tro84b, Gro81, Tro84b, Mor87d]. Canals [Eva81]. Canned [O’B87]. Canned-Salmon [O’B87]. Cannery [Sho88, Sho88]. Cantelon [Gre84]. Cantrell [Tan89b]. Cape [Gro81]. Capital [Lin85]. Capitalism [Lub86a, New83, Coh89]. Capitalist [Lub85, Nel81]. Caplan [McG88b]. Capper [Fee85]. Captains [Cas88b]. Car [Eis80a, Fos82, Sau80]. Carding [Gro87b]. Cardot [Fro89]. Care [Mel85]. Career [Rot85a]. Careers [Dav89c]. Cargill [Han89c]. Cargo [Ung81]. Carl [Fol80a, Hok87, Kou81, Lue86, May83, Iri89a, Kra89b, Man89]. Carla [Elb88]. Carlo [Bra83a, Est85, Kel82]. Carlos [Har87d]. Carlson [Fos82, Wes83]. Carlyn [Kil88]. Carmela [Tei82]. Carnet [Hey88]. Carnot [Gil82, Kid88]. Carol [Lin85]. Carolina [Ano88i, Mac88a, GF87]. Caroline [Och89]. Carolingian [She81a]. Carolyn [Jer82, Nye89]. Carpenters [Gro88a]. Carpet [Moh87a]. Carriage [Ryd80]. Carroll [Hou83, Ska88]. Carrow [Bro86b]. Cars [Wil88]. Carstensen [Bla87]. Carter [Cla88, Hor88, Hyd86]. Carter-Edwards [Cla88]. Cartographic [Har87b]. Cartography [Kar87a, Arl84, Kar87b, Hat89]. Carus [Jer85]. Carus-Wilson [Jer85]. Carver [Cre83]. Cary [McD85a]. Casado [Har87d]. Case [Buc88b, Coc81, Cum88, Got83, Hob80, How85a, Ino81, Law87, Lev87b, Mac85b, Maz82, McG88b, Nic80, Por80, Ran86b, Roe82, Rol83, Smi85e, Str84, Twe86, Wei81a, Zie89b]. Caspian [Bul86a]. Cassedy [Gro87a]. Cast [RBD84]. Cast-Iron [RBD84]. Castle [Per85b]. Catalog [Tur89]. Catálogo [Cum86]. Catalogue [Bed85a, Bed87, Har80b]. Catalonia [Gli86]. Cathedral [BM84a, BM84b, Con84, She80a]. Catherine [Smu87]. Cathode [Car84a]. Cathode-Ray [Car84a]. Catlett [Goo89]. Cattell [Sha84]. Cause [Con89]. Cecelia [Mei88c]. Celebrating [Car86]. Celebration [Mar86a]. Celebrations [Jac84]. Celestial [Hil88]. Celluloid [Bud84]. Cent [Coo88]. Centenary [Gol87, Han80]. Centennial [Car85, Car86, Hin84, Jac84, Jac86, Mar86a, Mor81b, Row82b, Wil84c]. Center [Ano80d, Com87b, Har89b, Jon81, Kel86, Smi86b, Lau82]. Central [Bez83, May83, Mul88b, Str84]. Centurie [Gar81c]. Centuries [Gil85a, Haf85b, Har82b, Qui89, Est85, Fri80a]. Century [Als81, Arm86, Ber85, Ber83, Bil80a, Boy85, Buc82a, Car86, Cla86, Con81b, Con88b, Con81f, Con83b, Das85, Dub88, Fri86b, Ful80, Gar88, Gol87, Goo86, Gor89a, Har83b, Jer85, Kel82, Kel89, Kle87, Kli87b, Leh86, Mar87b, Mil86, Mis84, Mye89, Och89, Ran82a, Rey86, San84, Seg83, Tan89a, Tuc82a, Wal88, War86a, Wil84d, Wri86, San81a, Esp82, Eva81, Fli80, Isr87, KR80, Lan87d, Lub87b, Mar88, Mul85, Oer89, Pet83, Pre87, Pri87, Ste89, Bra83a, Buc86d, Coh89, Elb88, Gra84, Joh89, Kar87a, Kem84a, Koe84, Lan84, Lin85, Mil82c, Mul83d, Nye89, Pal80, Ros87b, Ska88, Sti82, Swe88a, Whi83a, Bed86, Hin80b, War87a]. Ceramic [Mye87]. Ceramics [Mil82c, Mye89]. Certainty [Das85]. Ceruzzi [Wil85b]. Cetina [Wil85a, Pin84]. Chadeau [Mor89e]. Chain [Bro88a]. Challenge [Fer80b, Gie85, Tuc80b, Whi82c]. Challenges [Eng87]. Chalmers [Mey88]. Chamberlain [How80a, Hac82, Keh84]. Chambers [Bin83]. Chance [Bra81b, Hal80a]. Chancery [Hew87]. Chandler [Bat82, Koe84]. Chandra [Gut86]. Change [Ber89, Bij86, Bur80a, Bur83, Con81b, Con89, Coo80, Cor87, Cum88, Cut81b, Dan85, Gie85, Gor83, HH87a, HH87b, Har87c, Hob80, Jer88, Lau82, Law87, Lif85, Mac85b, McG85a, McN83, Mer86, Moh87c, Mor87d, O’B87, Pap84, Pec86, Pre83a, Qur86, Seg89a, SF87, Tan89a, Twe86, Cut81b, DeG87, Fah89, Mer88, Rae82, Rum87, Ste88a, Tay81]. Changed [Gre88b, Sch81b]. Changes [Far80, Ros87a]. Changing [De 84c, Nev80, Sta88b]. Channel [Dav89a]. Chanter [Ger84b]. Chants [Sch85b]. Chapman [Jer85, Tan82b]. Character [Bed85b, Dun89, Mei89, Wil86b, Woo86]. Characteristics [Kob80]. Charcoal [Bro85b]. Charge [Kop89]. Chariots [Hal80a, Zig88].

Charitable [Mel85]. Charles [Bal80, Bül88, Car84a, Cum88, Fle88, Fri81a, Hil81, HS88, Kar81, Lan85, Lan87b, Lan89a, Lat86, Mac86a, Mol84b, Pre87, Ran82c, Scr87a, Sim86, Wal82, Wri80, Zig88, Abr80, Jor89, Kas84, Smi85b, Tro84a]. Charlton [Kem84a]. Chassagne [Jer85]. Chauncey [Mor81b]. Chaussées [Smi86d, Tar85a]. Cheape [Hil81, Scr87a]. Checklist [Smi83b]. Cheese [Kel82]. Chemical [Bee85, Bel84c, Gué81, Hei88, Hoc86, Hol88, McM86, Mis84, Mul83a, Mul88a, Pan87, Rey86, Smi86b]. Chemicals [Mul83b]. Chemistry [Ful80, Mor87e, Mul88a, Pan87, Rhe86, Smi86b, De 84b, Rei87]. Cheney [Car83a]. Cherniack [Lan88]. Cherry [Lev87c]. Chesapeake [Tro84b, Fee85]. Chew [Ste83]. Chicago [Cai82a, Car85, Che84, Hoy88, O’C88, Sau86, Sin88b, Ska88]. Chicano [Sho88]. Chicken [Qur86]. China [Mét86, Mit85, Mul88a, Och88, Ovi88, Pan87, Qur86, Rah82, Tom85, Qur86, RBD84, Tom83a, Tom84, tZS82, Vol88]. Chinese [Pet89, Ros87b, Sto89, Tom85, Bra81a, Fie84]. Chip [Cos85, Hev86, Ros87a]. Chips [Ler88, Ros87a, Fri89b]. Choice [Tei88, San81a]. Choices [Eva81, Ros84]. Cholera [Dav85]. Chopra [Qur86]. Chris [Kar86]. Christiaan [Sch82d]. Christian [Lue86, Rui80, Ovi86b]. Christians [Lov82]. Christie [Mor87e, Sti86a, Tod86]. Christienne [Lan87b]. Christine [Bla88]. Christoph [De 83]. Christopher [Har87c, Kie86, Lan87c, Tay89, Thi80]. Chronicles [Bot80]. Chronological [Man89]. Chubin [Pin85]. Church [Sau80]. Ciencia [Mau82]. Cigar [Coo88]. Cincinnati [Lub87a]. Cinema [Wes84]. Circle [War89b]. Citation [Goo80a]. Cities [McK86, McS83, Per81b, War86a, Bar88, Cow83, Row83, Kir80, Sti81, The87]. Citizens [Rei82]. City [Bla88, Blu82, Con86b, Dob81a, Hin80a, Hin84, Kir80, Kra86b, Les82, Moh87c, Mor83, Per85b, Ros84, Ros89, Sch85b, Tar81, Tar85a, Wal87b, Wei84, Maz86b, Bah87, Bar89a, Bez85, Fri81b, Hof80, Seg89b, Bil88, Lun88, Sti85a]. City-Systems [Blu82]. Civil [Bil80a, Bil82, Cas88b, Kra88b, McD89, Smi85e]. Civilian [Bro83]. Civilisation [Mét86, Mul88a, Pan87, Smi85f, Mil80a]. civilisations [Smi81b]. Civilization [Smi83b, Tom83a, aH80]. Civilizations [Hug88a]. Civilized [Mor89a]. Clair [Bar88]. Clara [Sho88]. Clark [Fer82, Fri80a, Hir86, Rot82, Ket85, Smi89a, Spi81]. Clarke [Gol87, Maj81, Seg80, Sha89a]. Class [Fol80b, Fri86c, Mar84b, Nel81, Ran82c, Sch85b, Smi80b, Sti86b, Tod86, Mey85a]. Classical [Lan87a]. Claude [Her88, Kra85, Rah82]. Claudia [Bul86b, Wal81]. Clay [Phi83, Har87c]. Clayton [Han85, Swe81, Wel89, Whi81]. Clear [Les85, Gar87b]. Clercq [War87a]. Clerk [The84b]. Cleveland [Mil83]. Clifford [Lov82]. Clifton [Nut81]. Cline [Jaf84]. Clipped [McD84]. Clips [Ler88]. Clive [Coo84c, Cro87b]. Clock [Har80b]. Clockmaker [HS88]. Clocks [And81, And82, Bed89c, May85, Bed80b, Bed87, SS84, Dav83]. Clockwork [And82, Fri84a, Bru88]. Closure [McG88b]. Cloth [Jer85].

Clothiers [LGWK80]. Clothing [Jer85, Ran86a]. Clothmaking [Jer85]. Cloud [Hac88, Hol88]. Coal [Bin89, Don81b, Don86a, Gar81b, Har85, Has86, Sal84, Str84, Str85a, Str88, Tan82a, San81b, Sal84, Och88, Sal84, Str84]. Coalition [Don81b]. Coalminers [Ran82a]. Coast [Bul86a, Gol87, Gro88b, O’B87]. Coat [For88b]. Coates [Rei80a]. Coatsworth [Bee83]. Cobb [Bro86c, Dan83a]. Cochran [Bur83, Eng87]. Cockburn [Ste88a]. Cod [Gro81]. Coeur [Lan86]. Cohen [Jaf84, Mul81]. Cold [Dea84]. Cole [Cut85]. Coleman [Fri81a]. Coletta [Mor80a]. Colglazier [Dow84b]. Colin [Bed83c, Cos87, Lev87c, Row82a, Tom83a]. Collaboration [Bro86b, McG83]. Collected [Str80]. Collection [Bed87, Fer82, Mor87e, Qui89, Rae80b, Seg87, Dav80]. Collections [Fri80b, Goo89, Tur89, Bar85, Maa81]. Collector [Hok83]. Collectors [Wis84]. College [Gor87, Mac85c, Wos82]. Colleges [Dan87, Pin85]. Collegiate [Mor87a]. Collins [Sta88b]. Colloquium [Bil81a, But87]. Colonial [Bed83b, Bha87, Coc82, De 81a, Fol82, Hin80b, Hob80, Lub85, Sel88, Way89, Yat82]. Colonialism [Buc88a]. Colorado [Smi88b]. Colossus [Upt88]. Colour [Mul87]. Colourists [Mul87]. Colton [Por89b]. Combining [Soc83]. Combs [Rig81]. Comedy [Per81a]. Comes [Hac82, Hen86, Wal82]. Coming [Hin85, Par82, Joh89]. Command [Mor86]. Commemorating [Hin84, Jac84]. Commemoration [You87a]. Commemorative [Mor87e]. Comment [Bar91, Buc83b, Flo86]. Commentary [Ras80b, Bed80b]. Commerce [Bil80a, Don81a]. Commercial [Gor89a, Gre88b, Kar87a, Lub86b, Ste82b]. Commercialization [Wil84d]. Commission [Pau83, Wal89]. Committee [Com86, Fri81c, Ove85, Bra81a]. Committees [Coo86b]. Communication [Beb80, Joh89, Nye89, Pyn84]. Communications [Ber85, Nev80, Per84b, Smu85, Sta83b, Cur84a, Smu85]. Community [Haf80, McD85a, O’C88, Sie81, Joh89, Dub83]. comp [Smu85]. Companies [Gor86]. Companion [Hil85]. Company [Bos85, Bot80, Gie85, Har82b, Hor84, McK80, Rea81, Sch82a, Scr87a, Tod80, Jac86, Lub87b, Maz82, Eis80a, Fit82, Hei88, Ino81, Kor86, Mol84b, Sau86]. Comparative [Fie84, Rui80, She87, Uss88]. Comparato [Bot80]. Comparison [Got83, Ino82, Buc88a, Qur86]. Compass [McC85]. Competition [Bus83, Vie83b]. Competitiveness [Bro85b]. Compilation [Fre80]. Complexity [Per84a]. Comprehensive [Nee83, Seg87]. Compton [Mac86a]. Compulsory [Ros85]. Computer [Cer85, Hen86, Kli87b, Man81, Owe86, Ros87a, Rum88, Tom83b, Wil85b, Lub86c, Ste81b, Tro84a, Ler88]. Computerization [dSP81]. Computerized [Gol80b]. Computers [Blo81, Kap86, Ste82b, Tro87, Ste81b, Hen87, Man83, Tay81]. Computing [Bra83a]. Comstock [Fow80a]. Comune [Bul81a]. Concept [Bir83, Car81, Tom83b, Wil85b]. Conceptions [Jam89]. Conceptual [Ste82b]. Concerning [Bir83, Dan83c, Goe83]. Concert [Roe85a]. Concise [Con80a]. Concluding [Siv80]. Concrete [Big88, Bil81a]. Condit [Iri89a, Kou81, Kra89b, Man89, May83]. Condition [Sok81]. Conditions [Ras80a]. Condom [Bul81c]. Conducting [Coa83]. Confederate [Rol87]. Conference [Bah87, CL80, De 84c, Fra84, Haf85b, Hob85, LR85, Lew89, Mac88a, Mel83a, Mor87a, Pal81, Siv80, Smi87b, Tar85a, Tom83b, Vic84, Wei84, Pur80, Siv80, Ber80, Har86a]. Conflict [Ble83, Dub88, Mar85a, Str84, McD84]. Conflicting [Mil87]. Congrès [Pal81]. Congress [Gor82, Hol89, Wri85b, Ano80l]. Conjectures [Tra85]. Conkin [Seg86]. Connecticut [Coo84a, Gor83, Vic87, Coo84a, Rum87]. Connection [Kra82]. Connections [Buc81a]. Conquering [Abb86]. Conqueror [Bac85b]. Conscience [Fer80a]. Consequence [Con89]. Consequences [Bra80b, Ras80a]. Conservation [Fri82b, Gor82, LR85, Ste88c, Vic84, Cen80]. Conservative [Thr82]. Constance [Hea89]. Constant [Lew82b]. Constantinople [Bac82a]. Construction [BBM84, Boy88, Coz89, Fre82, Hil88, Lig87, Mor87b, V88, Mar86b, See87, Mor87b, Mor89b]. Constructivist [Pin84]. Constructor [Pri87]. Consumer [Mei89, Wil84d, Wil87c]. Consumption [Har83b, Mei89]. Contained [Gre84]. Contamination [Kro86]. Contamine [Bac86]. Contemporary [San84, Sla84, Wil86b]. Context [Bri81c, Ovi86b, Owe86, Pye84, Had80a]. Contextual [Pin84]. Continental [Coo84c, Gei80]. Continuing [Ros85]. Continuities [Fie84]. Continuity [SF87]. Continuous [Hun87b]. Contraception [Bul81c]. Contributions [Wis81]. Control [Bin83, Car82c, Cos88, Cro82, Fli89b, Gie81b, Sch82a, Vin82, Fin83]. Control-Volume [Vin82]. Controlling [Kro86, Hac87]. Controversies [McG88b]. Controversy [Bro83, Buc83b, Emm83b, Hew83, Ste81a, Tre87, Pow83]. Convergent [Hug81]. Cooke [Sti86a]. Cooker [Kap81]. Cooking [Bus83]. Cooper [Lev88, Cam88, Coo83, Fri85, Law89, Sla84]. Cooper-Hewitt [Lev88, Fri85]. Cooperative [Car88]. Copenhagen [Nye88]. Coping [Sti87]. Copper [Jac86, Lan83c, LM87, Lec88, Way89]. Copying [How87]. Coquand [Smi86d]. Corliss [Bel84a]. Corn [Sau86, Lew85, Sla89a]. Cornelius [Baa87]. Cornell [Hir85]. Cornford [Rob88]. Cornwall [Ste88b, Cla88]. Corporate [Bat82, Rei83a, Sin87, Lin85]. Corporation [Gie86, Pur89a, Roe85b]. Corps [Cai82a, Cal85, Cas85, Cas86, Dob81b, Lev83, Reu85, Sma80, Smi86d, Sti84]. Correspondence [Bed89b, Har89a, Ken86]. Corridor [Con85b]. Cort [Pay85]. Cosa [Har89b]. Cosman [Koe84]. Cosmic [Swe88b, Zie89b]. Cosmology [Keh84]. Cosmos [Bru88, Dav83, Kel82]. Cossons [Bry81]. Costandina [Hac88]. Costello [Wes86a]. Costin [McG86]. Costs [War82]. Cotton [McG85b, Tan81, Tin87, Bro80]. Cottons [Qui89]. Coulston [Lan85]. Council [All84, Rab81, Sin86a, Har81b]. countries [Bed80b, Dow85, Sta81, Wil84a]. Country [Cai88, Coo88, Gle84, Mar84b, V88, Wik81]. County [Wil80, McG85a]. Course [Car88, Ros81, Wil89b]. Courtly [Mor81a]. Courtney [Hal80a]. Covers [Lew88a]. Covert [Smu87]. Cowan [Hac85]. Cowdrey [Dob81a]. Cox [Bow87]. CPR [Leh87, Tur85]. Cradle [Sma80]. Craft [And88, Gar81c, See87, Yor88]. Crafton [Cha84]. Crafts [Fol80b, Mor87c]. Craftsman [Mei88a, Wel87b]. Craig [Mar86d]. Craven [Pyn84]. Creation [Hal84b]. Creative [Eis80b]. Creativity [Bro87b]. Creator [Pay85]. Creighton [Lan84]. Creps [Shu84a]. Cressy [Joh89]. Creveld [Rol89]. Crew [Web81c]. Crewe [Sem84]. Crises [Fow80b]. Crisis [Gil85b, Joh88, Mar84a, Bez83, Cas88a, Gre84]. Cristofori [Roe85a]. Critical [Con85d, Dan83b, Dur83b, Dur83a, Gor84, Gre81, Kid88, McM81b, Pur80, Sha81a, She87, Mac86b]. Crom [Hok82]. Cromford [Hew87]. Crops [Wes86b]. Crosby [Tro84b]. Cross [Buc88a, Mey85a]. Cross-Cultural [Buc88a]. Crossroads [Web85a]. Crouch [Gil85a, Rol82]. Crucible [War86b]. Cruise [Sah87]. Crusade [Dan83a, Jac86]. Crystal [Bry81, Mac87b]. Cultivated [LeM82]. Cultivation [Kid85]. Cults [Est85]. Cultural [Buc88a, Dun81, Mye87, Mye89, Ovi86b, Rui80, Smu87, Wil84c]. Culture [And86, Ben86, Bul81b, Cer85, Col84, Cow86, Fre82, Ger84b, Har83a, Hob85, How84, Lub87b, Mar85a, Meh87, Mei88c, Mer87, Mic84a, Qur86, Rah82, Sch82b, Sin86b, Sla89c, Smi83e, Spi82b, The85b, Ung83, Wal87a, Ame89, Bol88, Cor88a, Dal88, Har82a, Hyd87, Koh87, Lec88, Nev82, Rhe88, Upt83b, Ano80e, Ano81d, Ano82d, Ano83g, Ano84d, Ano85e, Ano86e, Ano87d, Ano88f, Ano89f, Mei89]. Cultures [Och87, Tin87, Tuc80b, Bed80b]. Cumulative [Goo82a]. Current [CRA^{+}89, CRM84, CRM85, CRA86, CRA87, CRA^{+}88, Goo80b, Goo81, Goo82b, Goo83, Sha81a]. Curricula [Pri82]. Curriculum [Har81b]. Curtis [Don86a]. Curtiss [Hal82]. Custom [Soc83]. Cusumano [Rae87]. Cut [Coo85, Wil89a, Str84]. Cutcliffe [Smi83b]. Cutler [Mar87c]. Cutting [Wal81]. Cynthia [Aga82, Gui89, Har87a, Ste88a]. Cyril [Bur82]. Czechoslovakia [Buc83c]. Czitrom [Dou84].

D [Ame89, Arm86, Bat82, Bea81, Beb80, Bed80b, Bil80b, Bro86b, Cla88, Gar81a, Gil85a, Hal86a, Har86b, Hil85, Hin80b, Hin83, Hoc86, Hym84, Jer85, Kas83, LK80, Leh83, Mac86a, McI83, Mol84b, Mor87e, Och89, Pin84, Rol82, Sal84, Smi85a, Smu87, Soc83, Swe81, Tan81, Tan82b, Tul85, Wei80b, Wei81a, Wel89, Whi81, Wil85a, Wri80, Yao88, Beb80, Wit83]. D-Day [Wit83]. D.C [Dob81a, LHH^{+}84, Mor87a, Sch89, Wor86a]. D.C. [Doo87, Ste87, War88, Wri85b]. Dabieshan [Ros87b]. Dabney [Hig80a]. Daedalus [Vac81]. Daems [Bat82]. Dahmus [Bac85a]. Dakota [Gra88]. dal [Bul81a]. d’Alenes [Lan86]. Dalrymple [BM85]. Dalzell [Baa89]. Dam [Jac89a, Joh88, Kru85, Ada84, Rey89b]. Dams [Mil89]. Damsites [Cas86]. Dan [Hok83]. Dane [Bez85]. Daniel [Bra81b, Cam89, Coh86, Dou84, Hab82, Rob88, Sti82, Wil87c, Tin87]. Daniels [Hir84, And83]. Danish [Hal86a]. Danly [Fah89]. dans [Cen80, Fro87, Ger84b]. Dark [Gen81]. d’Arts [Lan89a]. Darwin [Kra88b, McM81a, The83, Win89]. Daryl [Pin85]. Dascher [Tod87]. Dassault [Mor89e]. Daughter [Gor86]. Daumas [Her85, Smi85f]. Dave [Hyd86]. Davenport [Hal80b]. David [Bar88, Bau86, Bin83, Bly88, Bri80b, Cam85, Cer85, Con80b, Cor88b, Cut81a, Dav89b, Fer82, Fol85, Gli89, Har82a, Hyd86, Isr87, Jas87, Joh89, Les86a, Mac86a, May85, Maz84b, Mei84, Mel85, Mey86, Mir82, Moh87c, Mor87e, Pal80, Pen82, Per86, Pre87, Pye84, Rot84, Sha89a, Sin87, Ste88b, Tre85, Web81c, Wik80, Wre89, Wri83]. Davids [Pre83a, Ung88]. Davidson [Kan85, Les85, McL85]. Davies [Cos85, Zie89a, Moo86]. Davis [Ger81, Dur83a, Gra87b, Law83, Vin86a]. Davison [Jac83]. Dawn [Baa87, Wal87a, Mei82]. Day [Hac86, Mei82, New85, Sch82a, Vin86b, Lan89a, Kle87, Rah82, Wit83]. Days [Lan84, Twe88]. DDR [Rap81]. DDT [Rei82]. Dea [Sta85b]. Deal [Ale86, Mar87c]. Dean [Mor80c, Mye87]. Dearborn [Hyd88b, Hyd89, DKP^{+}86]. Death [Ale85, Jer82, Ste82a]. Debate [Tod84b, Web80a, Con80c, Ros85]. Debates [Kas83]. Debeir [Her88]. Deborah [Cut87]. Debris [Pau83]. Decade [Blo81]. Decades [Hac88, Sch81b]. Decay [Hac80a]. Decazeville [Bin87]. December [Bri81b, Car85, Ell87, Pyc82, Tar85a]. Decision [Hal80a, Wit83, Nel89]. Decision-Makers [Nel89]. Decisions [Por80, Tei88]. Decisive [Bac85a]. Decline [Coo85, Har83a, Hev89, Hil84, McK86]. Decorate [Hin80b]. Decorated [Sin88b]. Defeat [Mor86]. Defense [Bac82a, Elb88, Wei81b]. Defied [Col84]. Defining [Rey86]. définition [Kra85]. Deified [Col84]. Deindustrialization [Bin87]. Del [Bro83, Hac82, Keh84, San81b, Kel84a]. Delaware [Ano89n, Smi87b]. della [Bed80a]. dell’Istituto [Bed80a]. Delta [Bri80b]. Democracy [Bot80]. Democratic [Mer86, Sch85b]. Dempsey [Smi88b, Leh87]. Denis [Gut87, Mor85]. Denkmale [Mul89a]. Denkmäler [Bra82a]. Denmark [Wer81]. Dennis [War86a]. Denver [Ros84]. Department [Bed85a, Gre84]. Dependency [Bur80b]. Deployment [Seg87]. Depth [Wes84]. Derek [Bau86, Hic85b, Hit85, Bed84]. Derrick [Buc86a, Emm82]. Dertouzos [Man81]. Desagüe [Hob80]. Desert [Jac87, Mul88b, Row89]. Design [Cha89, Fri88a, Kem86c, Lev88, Lub87b, Mac87b, Mei88b, Mor89b, San89, Tuc82a, Vin86a, Whi85, Har87b, Kem86a, Kem86b, Pla86b, Smi83e, Whi86b]. Designers [Duf80]. Designing [Fli89b, Kou82]. Designs [Ada84, Cow83, Seg87]. Desire [Mei88b]. Desmond [Cal81]. Despair [Sta83b]. Detectors [Ait81]. Detroit [Hol87b]. Deutsche [Bed80b]. deutschen [Bed80b]. Deutscher [Tod85, Web85b]. Deutschland [Bra82a]. Developed [Dow85, Smi81a]. Developing [Bos85, Con88b, Dow85, Wil84a]. Development [All84, Bee83, Bez85, Cas85, Cro87a, Dan83a, De 80b, Feu88, Fri83, Gar81a, Gei80, Gor85b, Gué81, Has86, Hof85, Kli87a, Kob80, Lec84, Maz82, McG83, Mey86, Mil82a, Mor81a, Pel83, Reu85, Ros87a, Rut89, Sch84, Smi85d, Tuc80a, Vie83b, Wei80b, Wei81a, Cam88, De 84c, Law85, McI83, Pap84, SF87, War82, Wor81]. Devendra [Fri84b, Mil82a]. Devices [Kel80, Ovi85]. Devil [Rig81, Wik81]. DeVore [Dan81a]. DeVorkin [Edg84]. Dexter [Ano80b, Ano81c, Ano82c, Ano83d, Ano84c, Ano85c, Ano86c, Ano87b, Ano88c, Ano89b]. Diamond [Tuc82b]. Diana [Pre83a]. Diane [Cos85, Qui89]. Dianne [You87b]. Diaphragm [Bul81c]. Diaspora [Buc86c].

Dibdin [Ham88]. Dibner [Ano88d, Ano89c, Bed89a]. Dickinson [Klu82, Rob80]. Dickson [Ame89]. Dictionary [Fri80a, Och89, Ryd80, Mei81]. Did [Rae84b]. Diebold [Pap82]. Diego [HS88, Ovi85]. Diehl [Con87]. Diesel [Kra89a]. Dieselize [Wit83]. Dieter [Mul89b]. Difference [Bül88, LL82, Vin82]. Different [War88]. Differential [Owe86]. Differentials [Aga82]. Diffusion [Bir83, De 84a, Dor80, McI84, Mir82, Pen84, Wes86b, Mac85a]. DiFilippo [Hev89]. Digging [Bin89]. Digital [Wil85b]. Dignity [Fer80a]. Dilemma [McD85b]. Dimension [BM85, Per80c, Buc89]. Dimensions [Dow84a, Maz84b]. Diplomat [Rob80]. Directing [Coa80]. Directions [Coz89]. Directory [Ano83e, Ano85d, Ano87c, Ano89d, Per80a]. Dirk [Mul83d]. Disappearing [Row89]. Disaster [Ale85, Scl88, Lan88]. disciplinary [Wil81]. Disciplines [Och89]. Discourse [Bed85b]. Discoverer [Bea81]. Discoverers [Coh86]. Discovering [Rig81]. Discovers [Fis88]. Discovery [Bro87b, Zie89b, Bur82]. Discussion [Wei81a]. Disease [Bul81d, Bin89]. Disorder [Tra85]. Display [Kid89]. Disposer [Hoy85]. Dispossessed [Kru85]. Disputes [McG88b]. Dissent [Rob88, SF82]. Distance [Isr87]. Distant [Law89]. Distilling [Mei81]. District [Cai82a, Cas85, Cas86, Dob81b, Jac87, Sma80, Smi88b, Sti84]. Divided [War89b]. Divine [Web80a]. Division [Ben88, Wel89]. Divisions [Tan89a]. Diyala [Hab85]. Dizer [Hor84]. Djen [Kid87]. Dobney [Sma80]. Dobson [Ran82b]. Docks [Bil83]. Documentary [Sta83a]. Documentation [Hir86, Maa81]. Documents [Lev87b, Per86]. Dogwatch [Lan84]. Dollar [Sch82a]. Dolmetsch [Hin80b]. Dolores [Cow83]. Dome [Zuk80]. Domestic [Bus83, Cow83, Gre88b, Sch86b]. Dominance [Ste88a]. Don [Dur85, Res81, Ryd80, Sto88]. Donald [Arm86, Bil81b, Bin87, Cha84, Con86b, Cut81a, Gli88, Hoc87, Hol83, Hun87a, Kra89a, Lan87a, Mau84, Mil89, Mol89, V88, Ros87b, SS84, Smi81a, Sti86b, Wei81a, dlB88]. Doner [Mor80c]. Donnachie [Lin81b]. D’Onofrio [Pap84]. D’Onofrio-Flores [Pap84]. Donovan [Hin86b]. Doodlebug [Mar84b]. Dordick [Ber80]. Dorf [Lew80]. Dorinda [Dav89c]. Dorothy [Lew88a, Mil87, SF82]. Dott.ssa [Bed80a]. Douglas [Fit83, Hal82, Fri89a]. Douglass [Rot85a]. Douhet [Mor86]. Dover [SL80]. Down [Keh84]. Downard [Mei81]. Dr [Dan82]. Dr. [Bry82, Emm83b]. Draft [LGWK80]. Drafting [Wal81]. Dragon [Hal84b, Tit89]. Dragons [Roe85a]. Draw [Mul82]. Drawings [McM81a, Seg87]. Dream [Hol87b, Hor86, Rol82, Upt83a, Har83b]. Dredge [Kah85]. Drepper [Pet87]. Drescher [Sti84]. Dresden [FW87]. Dressmaker [Wal81]. Drew [Zuk80]. drie [vB83]. Drillers [Cla89]. Drills [Lan83c]. Dritten [Fis82, Kra81a]. driven [Sim83, Sim85]. Driver [Boy85]. Drugs [Num88]. Dryden [Bec86]. Duane [Lan89b]. Dublin [Haf80]. Duby [Bac82b]. Ducker [McD86]. Dudley [Bat80, DeV88]. Duffy [Lan87c, Thi80]. Dugald [Car88]. Duhem [Cha82a]. Duke [Mac88a]. Dummy [Whi86a]. Dumping [Kro86]. Dunbaugh [Bra84]. Duncan [Nel81]. Dunlap [Rei82]. Dunstan [Moh87c]. Durbin [Smi86a, Bul81b, Fer81, The85a]. Durham [Mac88a]. during [Mul83d, Rob85, Rup88, San81a]. Düsseldorf [Web85b]. Dutch [Car89, Mul83d, vS80]. Dutra [Kah85]. Dutton [Rob85]. Dyers [Mul87]. Dymaxion [Seg87]. Dynamic [Moy80]. Dynamics [Rup88, Spi82b]. Dynamite [Han84, Sha85]. Dynamo [Jor89]. Dynamometers [Con83b]. Dynamos [Coo80]. Dynastie [Mul89b]. Dyson [Wri86].

Eads [Wat80, Kou82]. Eagle [Gil85a]. Eagles [vdL83]. Eakin [Cut81b]. Earl [Hig80a]. Earliest [Bul86a, Mar88]. Early [Bri85, Bur83, Cla88, Col81, Dav87, Elb88, Gre88c, Gué81, Gut85, Han87a, Hyd86, Jaf84, Joh87b, Kas83, Kir80, Kop82, New83, Owe86, Pre83b, Qui89, Rey86, Rit88, Row80b, Rui80, Sal83, Sch81a, Sch86a, Tan89b, Thi80, Wal87a, Wel87a, Wel87b, Wer81, Wes86b, Win89, Ful80, Oer89, Ovi86b, Sta82a, Ste81b, Ait81, Arl84, Car84b]. Early-16th-Century [Oer89]. earning [War83]. Earth [Gor84, Hal87, Keh84, Sto89, Del83]. East [Gol87, Hea89, Muh81, Pap82, Wer81, Fle84]. East-West [Pap82]. Eastern [Emm80b, O’B81, Roh84, Buc83b, Emm83b, Phi83, Buc82b]. Eastwood [Per80a]. Eat [Bur80b]. Eberhard [Mul89a, Wer82]. Echoes [Kid87]. Eckert [Man83, Ste82b]. Eckert-Mauchly [Man83]. École [Tar85a, Kra88a]. Écoles [Lan89a]. Ecological [Cro87a, Sok81]. Ecology [Cou86, Jer82, Per80b]. Economic [Bee83, Bow80, Bra81a, Cai82b, Cam88, Cos88, De 80a, Fos89, Fre80, Gei80, Gri81a, Jer85, Kul88, Lin81a, Lin88, Lub85, McI83, Mit85, Mul84a, Mul89c, Rot83, War82, Bra80b, Har87c, Has86]. Economics [Bro84, Mor89b, She80a, Shu84a, Wor81, Wei81a, Har84]. Economies [Aga82, Gei80, Rut89]. Economists [HN89]. Economy [Ata82, Bra81a, Bra80b, Bus87, Coc82, Lub85, McD89, McL84, McL85, Mor80b, Nic81, Och88, Por89b, Sch85a, She81a, Whi82b, Mor87d]. ed [Ain89, Ali89, Ame89, And86, And82, Ata82, Atk89, Bah87, Bat82, Bau86, Bed80a, Bed83c, Bed89b, Ber87, Ber80, Ber84, Bij86, Bil81b, Bin83, Bla81, Bli82, Blo81, Blu86, Bra81b, Bra83a, Bra83b, Bri81a, Bri81b, Bri81c, Bro86b, Buc81b, Buc82a, Buc88a, Bul81a, Bul89, Bul81b, Cai82b, Cam88, Car82a, Cer84, Cha87, Cla80, Coa80, Col81, Coo80, Coo89, Coo83, Cor88a, Coz89, Cur84a, Cut81a, Cut81b, Dan85, Dan82, Dav89c, De 80a, De 81b, De 83, De 84b, De 84d, De 84c, DeG87, Dow84a, Dow84b, Dow85, Dun81, Ell87, EE81, Fah89, Fel87, Fer81, Fis82, Fos89, Fre80, Fri83, Fri88b, Fro87, Fro89, Ger84a, Gia81, Gie81b, Gol80b, Gol81, Gol89, Goo82a, Goo89, Gor82, Gre88c, Gre81, Hac82, Hac84, HH87a, Had81, Had82, Haf85a, Hal86a, Hal88a]. ed [Hal81, Han80, Han89c, Har82a, Har89a, Har86a, Hea88, Hig80c, Hig83a, Hin80b, Hin81, Hin86b, Hir84, Hir85, Hir86, Hoc87, Hoc88, Hok83, Hok87, Hol83, Hol89, Hor88, Hou83, Huf82, Hug88a, Hyd88a, Jer85, Joh87b, Kam81, Kan85, Kar81, Kas83, Ken86, Kin81, Koe84, Koh87, Kon84, Kra81a, Kro86, LK80, Lan84, Lan83b, Lau86, Leh87, Ler88, Lev87c, Lew80, Lic85, Lin81a, Lin88, Lov82, Lub86b, Lue86, Mac86b, Man81, Mar86a, Mar80, Maz84b, McG88b, McL84, McL85, McM81a, Mei82, Mel83a, Mel83b, Mer87, Mer81b, Mil82a, Mil84a, Mil87, Mil88a, Mis84, Mol89, Mor87b, Mor80c, Mor81b, Mor87e, Mul81, Mul83a, Mul89c, Mul89a, Mus83, Nas80, Nee83, Nel87, Nev82, Och89, Ovi87, Ovi88, Pal80, Pal81, Pap84, Par82, Per84b, Per81b, Por89b, Pur82b, Pur84b]. ed [Pyc82, Pye84, Rae82, Ras80b, Rav81, Rei89, Rol80, Ros81, Ros89, Ros85, Rot85b, Rot83, Row80b, Sal83, San84, San80, Sar81, Sch82b, Sch82c, Sch80, Sch81c, Sch82d, See87, Seg86, Seg87, Sei80, Sha88, Sha84, Sha81a, Sha81c, She84, She81b, She87, Shu84a, Sla89a, Slo81, Smi81a, Smi82b, Smi82a, Smi83b, Smi85f, Smi86a, Smi88a, Smu87, Spi82b, Sta83a, Sta88a, Sta81, Sta85b, Tan82b, Tan89a, Tei82, The84a, The85a, Tod85, Tod87, Tra89, Tro87, Tuc80b, Uss88, Vac81, Vic84, Wal87b, War87a, Web85a, Web81b, Web84, Web87a, Wei82, Wei80b, Wei81a, Wel87b, Wer82, Whi86b, Wil84c, Wil85a, Wil89b, Wil81, Win82, Wis81, Woo86, Wor81, Yao88, Zim84, aH80, vB83, vB86, vW83]. Edgar [Kas84]. Edge [Lub87a, Pye84]. Edificii [Kel88]. Edinburgh [Mor87e]. Edison [Cha87, Jac86, Mor81b, Rhe82, Per86, Fer88]. Edith [Abb86, McG86, Tre85]. Edition [Bed80b, Cha87, Kid88, Lan80a]. Editor [Pos82]. editors [Fri82a]. Edmondson [Lev87c]. Edmund [Mar84b]. Educate [Hin80b]. Education [Car88, Gol89, Gor87, Scr86a, Seg83, Sha84, The83, Wei86, Emm83a, Lan89a]. Educational [Bez85]. Educator [Woo82]. Edvard [Bül88]. Edward [Bla89, Bul82, Cut81b, Gru86, Han89a, Hic85a, Hor88, Kah85, Kel89, Kid89, Lew82b, Mul89c, Sha81b, Smi84b, Tei88, The86, Tre87, Vin86b, Wei80a, Wes87c]. Edwards [Cla88]. Edwin [Bra84, Coc82, Kie89, Kli86c, Ras80b, Roe85a]. Effects [Fos89]. Efficiency [Jor89]. Egbert [Whi82c]. Egg [Qur86]. Eggert [Mey85b]. Ehrhart [Bri81a]. Ehud [Tra89]. Eight [Kra86b]. Eighteenth [Buc86d, Jer85, Wil84d, Hin80b]. Eighteenth-Century [Wil84d, Hin80b]. Eighth [Bri81a, Lau86]. Eileen [Mei88a]. Einar [Wes83]. Einführung [Lan80a]. Einstein [Wil84c, For84, Han80]. Eiseman [Gui89]. Eisen [Mis87, Tod84a]. Eisen- [Tod84a]. Eisen-Bibliothek [Mis87]. Eisenbud [Per80b]. Eisenhower [Bil80a, Hol87a]. Eleazar [Ano86i, Ano87g, Ano88j, Ano89i]. Electric [Cla88, Coo86a, Fer88, Isr87, Nye89, Sin87, Smi87a, Wis80, Con81a]. Electrical [Ano80d, Ano88e, Ano89e, Arm86, Bel84c, Bow80, Coo86a, Fro88, Kli87b, Mil86, Rey89a, Sch81a, Sch86b, Tod80]. Electricals [Car86]. électriciens [Fro89]. Electricity [But87, Cla88, HN89, Mil84b, Tuc80a, Tod86, Tuc80a, Whi81]. Electrification [May83, Rey84, San81a]. Electro [Mul83b]. Electro-Chemicals [Mul83b]. Electrochemistry [Ser83]. Electromagnetic [Swe88a]. Electron [Wis81]. Electronic [All87, Pyn84, Shu84b, Wis84, Cur84b, Kie86]. Electronics [Fri82a, Qur86, Wis81, Fri82a, Hev84]. Electrons [Arm86]. Elegance [Hai84]. Elektrifieringen [San81a]. Eleventh [Smi88a]. Eli [Vic87]. Elinore [Way89]. Elite [Baa89]. Elizabeth [Cas88a, Com87b, Kid85, Maz81, Tro84b, War82]. Elkana [Wil84c]. Ellen [Bin89]. Elliott [Fri80a, Hir86]. Ellul [Lov82, Del82]. Elmer [Dib81]. Else [Och89]. Elster [Mac85b]. Elton [Bil83]. Emergence [Bol88, Bur86, DeV85, Fow80b, Hod81, Mul84a, Wil83]. Emerging [Ata82, Eis80b]. Emilie [Hil88]. Emmanuel [Mor89e, de 82b]. Emme [Fri86d, Hac84, Sha81c]. Emmerson [Buc82b]. Emory [Fri88b]. Empire [Cou88, Hun87a, McG85b, Smi85a, Sti82, Rob88]. Employment [Cos85, Ler88]. Emulation [Jen84]. Enchanted [How86]. Encounter [Luk86, Sha89b]. Encounters [Web80b]. Encyclopedia [Hig83a, Nee83, Goo85a, Mar83b]. Endangered [Ste88c]. Endless [Fri81c]. Energie [Hol86a]. Energy [Bel84c, Bra81b, Buc85, Gre84, Hin86b, Hol86a, Maz86a, Ros84, Sti87, Str85b, Wal89, Whe84, Dow85, Car81, Bra81b, Dow85, Gri81a, Hir84, Ros81]. Energy/Solar [Str85b]. Engelbert [Bru85]. Engend’ring [Rit88]. Engewald [Wer82]. Engine [LL82, Mil82b, Mus83, Con81f]. Engineer [Cas86, Flo86, Hal88b, Hal88a, Kem86b, Sti85b, Bal80, Buc81b, Luk89, Mor89a, Sha83b, Smi85b, Sti86a, Kop89, Leh86]. Engineer/Damsites [Cas86]. Engineered [Pur89a]. Engineering [Ano80d, Bel84c, Bil80b, Bro86a, Buc86c, Buc89, Car82c, Car88, Cha82c, Cha88b, Coo84a, Don86b, Emm83a, Fri86b, Gol87, Hat89, Kli86a, Kli86c, Kli87b, Kli87a, Kon89, Kra88b, Lan87a, Lay88, Leh86, McM81a, Mer81b, Mil86, Mul81, Nic80, Ove85, Pri82, Reu85, Rey86, Sch81a, See84, Sin86b, Sin86a, Smi86b, Tan89b, Vin82, Vin86a, Wei86, Wor81, Wri85a, For88b, Con80b, Lan89a, Mei84, Mis84, Mul83a, Tod80, Kan85, Mil84a, Mol88, Seg83, The84a, Wei86]. Engineers [Bij88, Cai82a, Cal85, Cas85, Dob81a, Goe82, Gol87, HN89, Kem86c, Kli86c, Lev83, McM86, Mei88d, Reu89, Rey89a, Row82b, Sch81b, Sma80, Sta83a, Swe81, Wis84, Reu85, Buc88b, Cai82a, Car84b, Cas85, Dob81b, Fri86c, Sma80, Sti84, Arm86, Buc89, Cas86, Lev83, Mil84b, Sta83a, Sti84]. Engines [Bül88, KR80, Jer88]. England [Buc82c, Das85, Dub83, Gut85, Har87c, Joh89, Lan83b, Scr87a, Sha84, Ste85, Ful80, Ran86b, Ste89, Tob85, Vol87, Hin86a, Kel89, Rit88, Wil84d]. Engle [Had80a]. Englehardt [McG88b]. English [Bed80b, Jer85, Ran82c, Tan81, Joh89, Har83a, War86a]. Engraving [Wri86]. ENIAC [Man83]. Enid [Gle84]. Enough [Arn89]. Enterprise [Bla87, Col81, Doo87, Fri85, Gor86, Gut84, HH87a, Hof85, Isr87, Maj81, Mel85, Pur84b, Bat82, Bow87, Nel81, Scr87a, Tro87]. Enterprising [Baa89]. Enthusiasts [Bil87a]. Entrepreneur [Mol84a, Ino81]. Entrepreneurship [Car88, Tan89b]. Entrepreneuse [Ben84]. Entropy [Gri81b]. Entwicklung [Web85b, Neu88]. Envelopes [Phi83, SB80]. Environment [Cru87, Fri88a, Lan89b, Lew80, McS83, Nas89, Nev80, Rei83a, Sti87, Wal89, Fow82, Mul83b, She80b, Per80b]. Environmental [Cal85, Fla89, Mac88a, Per81a, Rey87, Tar81, Vie83a, Don81b, She87]. Environments [Ros84]. Environs [Ska88]. Ephemeral [Ryd89]. Epic [Jac80, Kle89, Mul88a]. Équatoires [de 82b]. Era [Bra84, Col84, Mar83a, Pas85, Sta83b, Zig85]. Eric [And81, Mar89b, She80b]. Erik [Ber87, McG84, Qur86, San81b]. Erlande [Hey88]. Erlande-Brandenburg [Hey88]. Erlich [Gro88a]. Ernest [All86, Hev84, Nel89, Por80, Rot85a, She81a, Wei82, Con88a]. Ernst [Hoc85, Lan80a]. Eroberung [Pet87]. Erratum [Ano80c]. Erwin [Seg86]. Esmond [Hut88]. Esnault [EE81]. Esnault-Pelterie [EE81]. Esno [EE81]. Esno-Pel’tri [EE81]. Español [Kel84a]. espãnola [Mau82]. Españoles [Cum86]. Especially [Bed85b]. Espionage [Sta88a]. Essay [Pin84]. Essays [Bur82, Gre88c, Iri89a, Jer85, Koh87, Lue86, She84, Sla89c, Wei80b, Bli82, Gri81a, Lov82, Str80, Bah87, Kon84, Mel83a, Mor82]. Essence [Smi83d]. Essential [Hab85]. Establish [Ano80d]. Establishment [Lun88]. Ester [McN83]. Ethic [Whi85, Cro87a, Per81a]. Ethics [Mil84a, Gol81, Pri82]. Ethnology [Sha83b]. Ethnotechnologie [Smi83e]. Ethyl [Gie86]. Euclidean [BM85]. Eugene [Bol88, Dan85, Dan81b, Goo80a, Hac84, Mor82, Pur82a, Sha81c, Fri86d, GD81]. Eugenia [Lec88]. Euro [McD85b]. Euro-Gaullism [McD85b]. Europa [Bra82c]. Europe [Gei80, Jer85, Mul89c, Kra81b, McD85b, Ung81, Ale85, Han87a, Kir80, PS82, Row80b, Row86, Wal88]. European [Bed85a, Big88, Mic84a, San84, Sti82, Elb87, Jer85]. Éva [Ain89]. Evald [Smi83b]. Evaluating [Wol83]. Evans [Pur82a, Wei81a]. Evelyn [Sta85a]. Evenson [Con81b]. Everett [All86, Mac85a]. Everyday [Lic85]. Everywhere [Dav89b]. Evidence [Wor86a]. Evolution [Bel84c, Cha82a, Gor85a, Han87b, Isr87, Mul81, Sah87, Zie89a, Wis81, Bud82]. ex [Sta85b]. Examination [Har87b]. Exchange [Wei87b]. Exegesis [Lue86]. Exhaustion [Sin88c]. Exhibit [Ano86d, Car86, Fin89, Hyd88b, Hyd89, Pen80, Per85b, Whi83b, Hyd88a]. Exhibition [Gli86, Hal88b, Mei87, Moh87b, Ste87, Wos84, Wri85b, Han80]. Exhibitions [Ryd89, War88]. Exhibits [Doo87, Jac84]. Existential [Dur85]. Expansion [Dan83c, De 81a, Goe82, Goe83, Har87c, Law87, Rot83]. Expect [Rae84b]. Expectation [Seg80]. Expeditions [Goo86]. Experience [Dav85, Spu89, HH87a, Hug88b, Och87, Por89b, Tra85, Vie83a]. experiences [Ino82]. Experiment [Bro80, Dan87, Seg87, Wor88]. Experimental [Han88]. Experimentation [Moh88]. Experiments [Con83a]. Expert [Jas87]. Expertise [Fer80a]. Experts [Gil85b, Seg86]. Explaining [Mac85b]. Explanation [Law87]. Exploitation [Bur80b, Gal87]. Exploration [Gru86, Joh89, Swe82, Nee83]. Explorers [Goo86]. Exploring [Com87a]. Explosion [Hac86]. Explosives [Ola85]. Export [Nic80]. Exposition [Tur85]. Expositions [Ryd89]. Express [Fos80]. Extraction [Wei87b]. Extracts [Ras80b]. Extraparliamentary [SF82]. Exzerpte [Web84, Web84]. Eye [Cha80]. Ezell [Gru86, Sha81b, Gru86, Sha81b, Smi84b, Tre87].

F [Baa87, Baa89, Bed85a, Bol88, Bri81a, Cas85, Cum88, Dan85, Dav89b, DeV85, Dow84a, Fle88, Fre89, Gol87, Gor88b, Hic85a, Hol88, How80a, Lan84, Les86a, Maj81, Mis84, Mul83a, Pur84a, Qui89, Rey89b, Rhe86, Rol87, Rui80, Seg80, Smi83c, Smu85, dM85]. Faber [Rah82]. Fabienne [Fro89]. Fabric [Tha86]. Facini [Bru88]. Factories [Twe88]. Factors [Mul84a]. Factory [Gli86, Hin85, Hof80, Hor86]. Factory-made [Hor86]. FAES [Sta86]. Fahlman [Atk89]. Fahrstuhl [Pet87]. Failed [Hyd86]. Failure [Bül88, Gar87a, Kem86b, Wat81, Wei87b]. Failures [Con81e, Mil81, Pet83]. Fair [Mei82, Sta89, Num88]. faire [Pet88]. Fairmount [Jac89b]. Fairs [Ryd89]. Faler [Pre83b]. Falk [How85a]. Fall [Hil84, Ols89]. Families [Sho88]. Family [Dub83, Jac86, Kor80, Scr86a, Dub83, Scr87a]. Fannia [Hyd88a]. Far [Com87a]. Farm [Fit83, Jac86, Ras80b, Soc89]. Farmall [Soc89]. Farmers [Dan87, Sha83a]. Farming [Buc82c, Far80, Wes86b]. Farmland [Koh80]. Farnese [Bru88]. Farnie [Tan81]. Farrall [Bin83]. Farson [Gro81]. Farwell [Hun84]. Fashionable [Wal81]. Fast [Maz82]. Fatal [Cru87]. Fatalities [LM87]. Fate [Del83]. Faulkner [Ber87]. Favor [Sha89a]. FDR [Chr89]. Fe [McD86]. Feat [Had80b]. Feather [De 81b]. February [CL80, LS84, Smi83e, Web85b]. Fedden [dB80]. Federal [Bil80a, Buc85, Cas86, Wik82]. Feeling [Sta85a]. Feibleman [Fel84]. Feirtag [Kin81]. Feld [Tod87]. Feldenkirchen [Tod84a]. Fell [HS81, Smi88b]. Fellowship [Gar81c]. Feminist [Cow83, Pur89b, Sta85b]. Fence [Koh80]. Ferdinand [Car84a]. Ferenc [Hac86]. Ferguson [Pur82a]. Ferkiss [Gia81]. Fern [Bla84]. Fernández [Cum86, GD81, Har87d]. Ferracute [Bow87]. Ferrarotti [Cha88a]. Ferrier [Bos85]. Ferrill [Eva88]. Ferrner [Sta88a]. Ferry [Gor86, Hyd86]. Fertilizer [Gal87]. Festschrift [Dan85]. Feudal [Bac82b]. Few [Mor85, Wre89]. Fiasco [Hol87b]. Fiction [Boy89, Hac84, LB82]. Fictions [Smi82a]. Fidelity [McG83]. Field [Bro86a, Pen82, Spi82a, Wes87b]. Fields [Cla89, Eis80b]. Fiery [Gor88a]. Fifteen [Mil84b, Per80c]. Fifth [Car82a]. Fifty [Seg86, Str84, Wri85b]. Fight [Whi81]. Fighter [Bil87c, Hol87b, Mor85, vdL83]. Film [Cha84]. Fine [Hal88b]. Finer [Pay85]. Finishers [Pre87]. Fink [Arm86]. Fire [Cas88a, Gen81, Kid88, Don86a]. Fire-Clocks [Bed87]. Firearms [Fol82, Gar88]. Fireline [Pyn84]. Firenze [Bed80a]. Fires [Bla88]. Firm [Scr86a, Scr87a, Wor86b, Kor80]. First [But87, Fri81a, Hol88, Lew82a, LL82, Lun88, Mil84b, Mil82c, Mor85, Num88, Nut81, Pal81, Per80c, SB80, Str84, Web85a, Wri85b]. Fischer [vB86]. Fish [Ada84]. Fishery [Har89b]. Fishman [Ols89]. Fiske [Mor80a]. Fit [Wal81]. Fitchen [Boy88]. Fite [Koh80, Sha83a]. Fitz [Goo85a]. Five [Coo88, Pre83a, Sch82a, Spi82a, Sta84, Tro87, Fer83, Sha81c]. Five-Cent [Coo88]. Fix [Flo84]. Flagg [Con88a]. Flagler [Bla89]. Flame [Gre88b]. Fleet [Wil88]. Fleets [Wil84b]. Fleming [How86]. Flight [Bec86, Bil80b, Bil88, Hig83b, Per80c, Cro87b, Swe82, Tri89, Cor85, Smi85d]. Flink [Cor89]. Flinn [Har85]. Floods [Wes89]. Floor [Scr86a]. Florence [Mye89, Scr87b]. Flores [Pap84, Gli82]. Florida [Bla89, Hof84, Dor83]. Florman [McD82, Mor89a]. Flow [Nuw88]. Flush [Vin84]. Flying [Bil85b, Lew88b]. Folger [Ste87]. Follow [Bij88]. Folly [Bot80]. Folsom [Kas83]. fonctions [Cen80]. Food [De 84b, McG88a, Num88, Sho88]. Fools [McI81]. Forbes [Hof85, Wes83]. Force [Gui84, Nuw88, Smi80b]. Forces [Les86a]. Ford [Fli89b, Hai84, Hyd88b, Hyd88a, Hyd89, Lan80b, Sch82a, Wik80, Sta88c]. Fordson [Soc89]. Forecasting [Fow80b, Fri86a]. Foreign [Bla87, Mit85, Wil84a]. Forest [Pyn84]. Forests [Mac88a]. Forevermore [dlB88]. Forge [Lig87, Smi86c]. Forges [Kem84b]. Forgotten [Hym84, Sch86b, Whi80]. Form [Lay88, MB89]. Formation [Ada89, Che84, Rui80]. Formative [Tob85]. Forms [San84]. Formulation [Mal86a]. Formwalt [Ken86]. forskning [San81b]. Forsyth [Kle87]. Fort [Wos82]. Fortifications [Bac82a, Gar88]. Fortress [Lan87c, Thi80]. Fortschritts [Web87b]. Forty [Bil81b, Smi81a, Mei88b]. Fought [Fos82]. Foundation [Tob85]. Foundations [Wei81a, Wes84]. Founding [Bir81]. Foundry [Tan89b]. Four [Rol83]. Fourcy [Kra88a]. Fourteenth [Koe84]. Fourth [BM85]. Fowlers [Min81]. Fox [Smi82b, Sta85a, Vac84]. Fram [Fri82b]. Framework [Beb80]. francaise [Gil82, Fro87, Smi86d]. France [Ber85, Bur86, Fro89, Gut87, Hat89, Joh83, Mey85a, Mor89e, SF82, Smi80b, Smi82b, Ber85, Gué81, Har83b]. Frances [Mar87a]. Francesca [Mét86, Rut89]. Francillon [Hal82]. Francisco [Lev83]. Franco [Cha88a]. Francois [Cen80, She81b, Smi86c]. Frank [Bil87b, Bot80, Coh89, Dav80, De 81b, Goe82, Hok87, Kan85, Les85, Seg81, Smi83c, Tri83]. Frankel [Roe89]. Franklin [Gre81, Ket85, LK80, Yao88, Yor88, Dan82, Hut88]. Fraser [Ovi88]. Fraunces [Dav87]. Fred [Bla87, Dea85, Fos82, Luk89]. Frederick [Con82, Emm81, Kap82, Lan87c, Sha83b, Sma80, Hab82]. Free [Shu84b]. Freedom [Shu84b]. Freeland [Vac82]. Freeman [Bau86, Lan84, Mar83b]. Freiberger [Mul89a]. French [Ben86, Ben88, Don81a, Emm83a, Lan87b, Lan89a, Fro88, Kra82, Pec86, Pri87, Hun84, Mar87a]. Frese [Ata82, Col81, Rot83]. Fresh [Hab85]. Fresh-Water [Hab85]. Freyssinet [GD81]. Fried [Bur80a]. Friedel [Fer88, Bud84]. Friedland [Ras80a]. Friedrich [Bra82b, Car84a, Smi86a, May84]. Friends [McG84, Pur89a]. Fritz [Kem86a]. Fritzie [Seg81]. Frontier [Bec86, Cru87, Cut87, Fri81c, Mil80b, New85, Rol83, Sti81, You87b]. Frontiers [De 84b, Bur83, Ove85]. Frontiersmen [Hig83b]. Frühindustrialisierung [Bra80a]. Fry [Whi84b]. Fuel [Per80d]. Fuller [Seg87, Fle88, Seg87]. Fulton [Har87a]. Fundamenta [Sch82c]. Fünftausend [Bul89]. Furnace [Gor88a, Hyd86]. Furnaces [Kem84b]. Furniture [Sei80]. Furter [Mis84, Mul83a]. Fusfeld [Tei82]. Fusion [Whe87, Whe84]. Future [Bea81, Bra81b, Bro86b, Hil86, Hin84, Kan85, Mar86a, Rot85b, Web80b, Wei82, Whi82c, Ber84, Hig82, Les85, Sla89a]. Future/The [Bea81]. Futures [Hac84].

G [Arm86, Bea81, Bri89, Car80, Cut81b, Dav89c, Dor80, Eat88, Hal80a, Har87c, Hea89, Hud86, Hun87b, Jan86, Jer85, Lov82, Mar85b, McI84, Mey85b, Moo86, Mor87e, Pre83b, Pre87, Sch85a, Sei80, Sla88, Slo81, Smi89a, Tan82b, Wei87b, Wel87b]. Gabler [Kie89]. Gabriele [vW83]. Gail [Vac84]. Galaxy [Bol88]. Gale [Rea81]. Galerie [Sta89]. Galileo [Wes89]. Gall [She81a]. Gallagher [Wei80a]. Galleons [Elb88]. Gallie [Nel81]. Galluzzi [Hal88a]. Game [Run84, Bil85a]. Gamst [Sha83b]. Gandhian [Wei81a]. Gandy [Har86a]. Gansvindt [EE81]. Ganswindt [EE81]. Garbage [Hoy85, McS83]. Garbrecht [Sha88]. García [HS88, Ovi85]. Garden [Lan83c, Pla86b]. Gardens [PS82, De 81a]. Gardet [Bed80b]. Gardner [Lic85]. Garfield [Goo80a, Mor82]. Garland [The83]. Garnet [Isr87]. Garrett [Smi80a]. Gary [Lan83b, LB82, Mey85a, Pyn84]. Gas [Bud82, Gre88b, Lie83, Gol80a]. Gasworks [Lun88]. Gate [Lev83]. Gateshead [Fit82]. Gathering [Kie87]. Gatland [Nee83]. Gauldie [Gle84]. Gaullism [McD85b]. Gavin [McD89, Wri86]. Gay [Ben88]. GE [Bro87a, Car88]. Gears [Mei88c]. Gebhard [Wis81]. Gedenkbibliothek [Web81a]. Gee [Sta81]. Gehlen [Huy82]. Geiger [Vey89]. Gender [Kil88, Rup88]. Gene [Dan85]. Genealogy [Bin87]. General [Coo86a, Fli89b, Rae85, Sin87, Smi87a, Wis80]. générale [Smi85f]. Generation [Dor80, Dub88]. Generations [Kor80, Wei82]. Genius [Bot80, Pur82a, aH80, Cam85, Hew83]. Gentleman [Ino81]. Geodesic [Seg87]. Geoffrey [Car84b, Gor89a, Hym84, McL84, Wil83, Wri86]. Geography [Hit85, Scr86b, War86a]. Geological [Con85a]. Geologists [Run81]. Geology [Run81]. Geomagnetic [Bed83a]. Geomagnetism [Tur89]. Geometry [BM85]. Georg [Bül88]. George [And83, Buc82b, Bur86, Cum88, Dal88, Fow80a, Hac87, Hir84, Hok87, Isr87, Koh87, Lic85, McI83, Mil82b, Mor80c, Rei80b, Rot85a, Smi82b, Smi87a, Sti85b, Upt83b, Whi86b, Woo82, Mac87b, Mul84a, Cre83]. Georges [Bac82b]. Georgescu [Gri81a]. Georgescu-Roegen [Gri81a]. Georgia [GK81, GS81, Hei86, Kel84c, KT80, McC88, Mol88, Pfe83, Rec88, SB82, Wes87a]. Geraedts [Mit85]. Gerald [Cla89, Hyd87, Mey85b, Per84b, Rav81, Wil84c]. Gerard [Bed86, Coo84a]. Gerhard [Bro84]. Germ [Per81c]. German [And82, Bed80b, Hyd86, McK86, Mer86, Bro85b, Mor81a, Hal84a]. German-speaking [Bed80b]. Germanic [Row86]. Germany [Buc85, Sha84, Kra89a, SF82]. Gerrylynn [Rhe86]. Gertrude [Gut85]. Geschichte [Bra82c, Bra83b, Hol86a, Pet87, Tod85, Wer82, Zee82, Mul89a, Ree81, Sch81c, vW83]. Geschichte/Energy [Hol86a]. Gesellschaft [Bra83b, Mul89b, Tod85, Web81b]. Get [Kra84]. Ghost [Coo87]. Gian [Bra83a]. Gian-Carlo [Bra83a]. Giant [Hym84, Sta88c]. Gibbs [Ovi86a]. Gibson [Bir81]. Giebelhaus [Coc81]. Giffard [Kra82]. Gift [KP80]. Gijitsu [Ino82]. Gilbert [Hor86, Hyd86, Koh80, Sha83a, Whi84a]. Gilded [Moh87a]. Gill [Cos87]. Gille [Dau81, Hug88a, Sle81, Smi81b]. Gillette [Roe85b]. Gillispie [Lan85]. Ginzburg [Est85, Kel82]. Giraffes [Roe85a]. Girouard [The87]. Gisela [Wer82]. Gisela-Ruth [Wer82]. Giulio [Mor86]. Glaser [Asp84]. Glass [Eng80, Had80a, Had81, Had82, Lay87, Mac87b]. Glasses/Fire [Bed87]. Gleanings [Lev87b]. Glen [Gor84, Sch86a]. Glenbow [Tur85]. Glenn [McK86, Wal83]. Glick [Rui80]. Glimpsing [DeV85]. Global [Dow85]. Globally [De 81b]. Globes [Hil88]. Gloeilampenfabrieken [Car89]. Glory [Bül88]. Glover [Buc89]. GM [Fli89b]. Go [Hal80a]. goals [San81b]. God [Wik81]. Goddard [EE81, Web85a]. Goddess [Eis80b]. Godet [Fow80b]. Goetzeler [Tod80]. Goff [Ung83]. Gohen [Bur87]. Gold [Jac88, Lec88, Ras80a, Smi81a, Wol83, She80a]. Golden [Kap80, Lev83, Mul83d, Hol89]. Goldenberg [DeG87]. Goldian [Lew80]. Goldstein [Lew88b]. Golley [Daw89]. Gollin [Cro86]. Goman [EE81]. Gómez [Mar84a]. González [Kel84a]. Good [Coo88, Hoy85, Sti84, Wre89, Roe85a, Tur89]. Goodman [Fri83, Gli89]. Goods [Mei89]. Goonatilake [Bro87b]. Göran [San81a]. Gordon [Bro80, Buc86b, Pia84, Sin88c, Bez85, Spr81]. Gorenstein [Hir85]. Gorman [Whi84a]. Gospel [Lew85]. Gothic [Con83a, San89]. Gottfried [Ale85]. Governing [Moh87c]. Government [Coc81, Gli89, Han87c, Wik82, Har87c, Smi85e]. Governmental [Mer88]. Governments [Pur84a]. Grace [McG86]. Graduate [Ano83f]. Graham [Dre89, Gué81, Hoc86, Bro88b, Rei83b]. Grahame [Spi81]. grains [Cen80]. Gramophone [LeM82]. Grand [Cow83, May83, Roe85a, Seg89a]. Granite [Pla86b]. Grant [Bed83c, Mei89, Seg83, Sau86, SPV87]. Grapevine [Nev80]. Grass [Seg86]. Grass-Roots [Seg86]. Grassland [Cou86]. Graubard [Vac81]. Graven [Gut86]. Graves [Bin89]. Gray [Pyn84]. Great [Bla88, Boy86, Buc82b, Cla86, Con87, Kle89, Lan87c, Mar85c, Pay85, Ryd89, Sau86, Soc83, Tre87, Tur85, War89a, Wil88, Buc83b, Emm80b, Emm83b, Str85a, Eva88, Bla84, Mil89, Rey87]. Greatest [Buc82b, Ste87]. grecs [Sle81]. Greek [Gre88a, Har86b]. Green [Ras80a, Wei81b]. Greenhalgh [Ryd89]. Greenhill [Leh86]. Greenwald [Nel82]. Gregg [Dea84]. Gregori [Bry82, SF87]. Gregory [Dur83a, Mor89b, Tur89, Hit85]. Grenander [Gol80b]. Gresley [Bly88]. Grew [Kul88]. Grice [Whi86a]. Griliches [Fel87]. Grith [Spi82a]. Grocer [McK80]. Gropius [Hor86]. Grote [Lue86]. Ground [Luk89]. Group [San84, Scr86b, Mor87b]. Growth [Bla88, Blu82, Cla86, Con80c, Gor85a, Hac80a, Hun87a, Mul84a, Och87, Sti85a, Vey89, Vin86a, War86b, Wor81, Lin88, Bee83, Wor86b]. Grundlinien [Lan80a]. Guerlac [Swo89]. Guide [Bar85, Bul81b, Dun81, Goo89, Hok83, Hol83, Kie86, Mel83b, Mil88b, Och89, Pen82, Smi83b, Sta83b, Wes86a, Wes87b, Wis84, Wri86, Fri80b, Jac83, Jak87, Lyt88, Mer87, Pur89b, Fri80b]. Guidebook [Coa83, Con85a]. Guilds [vS80]. Guinea [PS82, Smi81a]. Gullickson [Ben88]. Gummett [Coa80]. Gun [Gar87a, Han84, Rot84, Sha85]. Gunnar [Wor86b]. Gunpowder [Mul88a, Sim87]. Guns [Gui83]. Gunston [dB80]. Günther [Sha88]. Gutenberg [Bol88]. Gwei [Kid87]. Gwei-Djen [Kid87]. Gwendolyn [Upt83a]. Gwyn [How86]. Gyro [Hal80b].

H [Arg88, Arl84, Bau86, Bee83, Bel84c, Bil87b, Bli82, Buc86d, Bud82, Coa83, Coo86b, Cum88, Dav80, De 81a, Don81b, Dow84a, Dur83a, Edg84, Fin83, Fos80, Fri81b, Gar81a, Gil85b, Got83, Gre80, Gro87a, Gro81, Gui84, Had80b, Had81, Had82, Hal86a, Hal84a, Har83b, Har87b, Har86a, Hig83a, Hir84, How80a, Huf82, Isr87, Joh83, Kie86, Kra88b, Lub86b, Mar87a, Mar83a, Mar86d, McD86, McG88b, McL85, McM81a, Mil84a, Mil82c, Mye89, Och87, Ras80a, Rob80, Ros89, Rot85b, Ryd80, Sau86, Sch82d, See87, Smi83b, Smi85b, Smi89b, Sok81, Sta82b, Sta88b, Ste81b, Sti85b, Tay89, The84a, Web85a, Whi86b, Wil88, Wil81, Win89, de 81c]. H. [Bea81]. Ha’aretz [Kel86]. Haas [Hei88, Lyt88]. Haber [Hol88]. Habitat [Les82]. Habitats [Mac88a]. Habs [Wri85b]. Hacker [Tit89]. Hackmann [Pit86]. Hadfield [Bal80]. Hagan [San80]. Hagley [Smi87b]. Hagwood [Lev83, Pau83]. Hainsworth [McL84]. Haklisch [Tei82]. Half [Mil82c, Tra85]. Hall [Bed89c, Car82a, Lan80b, Sch80, Gor89b, Han89c, Kel84b, Mac85c]. Hallion [Bec86, Bil87c, Hig83b, Hol86b, Rol80]. Hamden [Vic87]. Hamilton [McG86]. Hamm [Dow84a]. Hand [Fit83, Mar85a]. Handbook [Swe82, Swe82]. Handcraft [Mil82c]. Handguns [Smi84b]. Handloom [Moh88]. Hands [Gro88a]. Hanle [Hac82, Whe83]. Hannah [Mil84b, Tuc80a]. Hannover [Bra80a]. Hans [Han87c, Lan80a, Mar86b, Neu88, Web84]. Hans-Martin [Lan80a]. Hans-Peter [Web84]. Hansen [Kop89, Luk89]. Hanson [Bai83]. Harbor [Jac88]. Harbours [Bil83]. Hard [Jac80]. Hardeman [Jac88]. Hardesty [Gre83]. Hardware [Hen86]. Hareven [Hof80, Dub83]. Hargrove [Seg86, LGWK80]. Haritash [Mal86a]. Harley [Rum88]. Harman [Coa83]. Harmon [Jas87]. Harmony [Cha82c]. Harnessing [Lin88]. Harold [Don81a, Moy80, Rob85, Smi83c, Sti85a]. Harpers [Hyd86]. Harpoons [Mar85b]. Harrington [Tri83]. Harris [War83, Zie89a]. Harrison [Tan89a, Mei82]. Harry [Rig81]. Hart [Lan86, Cro87b]. Harte [Jer85]. Hartmann [Ler88]. Hartmut [Wal88]. Harvard [Bra81b]. Harvest [Ras80a]. Harvester [Bla87]. Harvesting [Kel89]. Harvey [Tei88, Wor81, Buc82c]. Haskell [Seg86]. Hassan [Gli88]. Hassinger [Lew82a]. Hath [Sta84]. Haugeland [Wil87b]. Hawk [Lan88]. Hawkins [Tre85]. Hawthorn [Maj81]. Hayami [Bra80b]. Hayden [Cow83]. Hayes [aH80]. Haynes [Hic85a, Bea81]. Hays [Fla89]. Hayward [Emm80a, Smi85e]. Heacock [Web85a]. Headrick [Cam89, Sti82]. Headwaters [Dob81b]. Healing [Dav87]. Health [Fee85, Fla89, Hoy85, Mel85, Mor83, Och89, Per80b, Ros85, Mil87, Tar81]. Healthiest [Mor83]. Hearden [McG85b]. Hearth [Hac85, Upt83b]. Heartland [Row83]. Heartlands [Jac88]. Heat [Kra82, Lan85]. Heather [Spi82b]. Heavenly [Bed89c]. Heavens [Hal87]. Heavy [Gar87a, Gué81]. Hechler [Fri81c]. Heerding [Car89]. Heidenstam [Jan87]. Heidi [Ler88]. Heims [Ste82a]. Heinz [Tod85]. Heirs [Rol80]. Heitmann [Mar89a]. Held [De 84a, Win82]. Helen [Mei82]. Helena [Bar85]. Henan [Ros87b]. Henderson [BM85, Lew82a]. Hendrik [Web87b]. Henry [Abb86, Bed89b, Bri81b, Car81, Ell87, Fer83, Har89a, Hyd88b, Hyd88a, Hyd89, Kem86b, Ken86, Mac87a, McM81a, Mit85, Pyc82, She80a, Swo89, Hai84, Lan80b, Nut81, Sta88c, Pay85, Wik80]. Heppenheimer [Whe87]. Herbert [Ber80, Dur83b, Fer82, Fis82, Fre82, Kra81a, Row86, Sti85b, Tei82, Tod80, Abr81b, Hor86, Lec88, Rae80b, Sta86, Sau80]. Heredity [Rit88]. Herf [Meh87]. Heritage [Fri82b, LR85, Lub85, Vic84, Wil80, Hud86]. HERK [Hig80a]. Herken [Dea84]. Herman [Bat82, Hym84]. Hermann [Neu88]. Hero [Hig80a]. Herzenberg [Och89]. Hewitt [Fri85, Lev88]. Heyningen [Dav85]. Hidráulica [Har87d]. Hidy [War89a]. Hierarchies [Bat82]. Hiesinger [Whi86b]. High [Luk86, McG83, McG88c, Pur89a, Roe82, Rol83, Sim86, Str85a, Tei88]. High-Fidelity [McG83]. High-Pressure [Str85a]. High-Risk [Sim86]. High-Speed [Luk86, Rol83]. High-Tech [Tei88]. Higham [Hol83]. Highway [Fos80, Kem86c, Reu89, See84, Upt88]. Hilgartner [Maz84a]. Hill [Jon81, Rig81, Gli88, Lan87a, SS84]. Hills [Leh83]. Himmelfarb [Gut85]. Himself [Coh86]. Hince [Str84]. Hindle [Jer88, Kas84, Jen84, Sch82b]. Hirsch [Fol80b]. Hirsh [DeV85]. Hirsimaki [Mar89b]. Hiskes [Tei88]. Histoire [Mor89e, Smi86d, Her88, Pal81, Kra88a, Smi81b, Smi85f]. historia [San81b, Gli82]. Historian [Dun80, Tho84]. Historians [Bil87a, Wis84, Sta88a, Wes86a]. Historic [Coo84a, Mer81b, Mil88b, Moh87b, Pen80, See87, Wri85a, Wri85a]. Historical [And86, Bro85a, Car85, Col84, Goo85b, HN89, Hir84, Maa81, Mar85c, Maz86a, Mor89b, Mor89d, Per80b, Rei87, Ryd80, Sal83, San84, Sha88, Sha81c, Sok81, Tod80, Wil84c, Hoc88]. Histories [Con86b, Rol83, Wei81a]. Historiography [Jac83]. Historische [vW83]. Historischen [Web85b, Web84]. History [All87, And81, Ano80d, Ano80l, Ano83e, Ano83f, Ano85d, Ano87c, Ano88e, Ano89d, Ano89e, Arl84, Bah87, Bay80, Bee85, Bil81a, Bil88, Bla84, Bos85, Bow87, Boy85, Bra83a, Bri80b, Bri81a, Bro80, Buc82a, Buc85, Buc86b, Bul82, But87, Cai82a, Car82c, Car85, Car89, Car83b, Cas85, Cas86, Cha82b, Cla88, Coh86, Com87b, Con80a, Con81f, Con85d, Con85c, Coo86a, Cow83, Coz89, CRA^{+}89, CRM84, CRM85, CRA86, CRA87, CRA^{+}88, Dan83b, Dan81b, Das85, Dav89b, De 80a, Dob81b, Doo87, Dun81, Edg84, Eis80a, Elb87, Eve86, Feu88, Fit82, Fli89a, Fol82, FW87, Fre80, Fri80b, Fri89b, Fri81c, Fri88b, Ful81, Gar87b, Ger81, Gie86, Gle84, Gol87, Goo80b, Goo81, Goo82b, Goo83, Gor88a, Hac82, Haf85b, Hal80a, Hal87, Han85, Han80, Han89b, Har85, Hei88]. History [Hen87, Hev84, Hil88, Hil85, Hin86a, Hir86, Hof84, Hok87, Hol86a, Hol83, Hou83, Hug81, Hug88a, Ino81, Jac86, Jac88, Kem84a, Kie89, Kli86a, Kni81, Koh80, Kop89, Kou81, Kra88b, Kra86a, Lan87a, Lan87b, LR85, Lec88, Les86a, Lev87b, Lev83, Lie83, Lin81a, Mac88b, Mac85c, Mac86b, MD82, Mac86a, Mac88a, Mac89, Maj81, Mar86b, Mar83a, Mar88, Mau84, May83, McD85a, McM86, Mei81, Mel83a, Mel87, Mel83b, Mol84b, Mor87b, Mor89d, Mor80c, Mul87, Mul88b, Mul89c, Nee83, Nel89, Och89, Pal81, Pas85, Per80a, Pon80, Pur80, Pur83, Pyn84, Rea81, Rey87, Rey89a, Rey89b, Roe85a, Roh84, Ros89, Row82b, Run81, Run84, Sal84, San80, Sau86, Saw88, Sch81b, Scr88b, Seg86, Ser83, Sha81b, She87, Sie81, Sin86b, Sla89a, Sma80, Smi86b, Smi89a, Soc89, Sta83a]. History [Sta84, Sta88c, SL89, Ste81b, Sti81, Sti84, Str88, Tod84b, Tom85, Tro84b, Tur89, Upt83a, Vig83, War88, War83, War89b, Wea83, Web81c, Wei84, Wel87a, Wes87c, Whi83b, Whi85, Wil88, Wil89b, Wor86b, Yat82, Zee82, Zim84, Ino82, San81b, vB83, Bur82, Lan84, Bil80b, Cai82b, Coo80, Cro83, Gli88, Har81a, Hig83b, Hin86b, Kon84, Lic85, Mar89b, May89, Mor80d, Nas80, Ovi86a, Pau83, Qui89, Ran82a, Smi83a, Soc83, Tan82b, The86, The87, Tur89, War89a, Wil87a, Asp84, Bog85, Car82a, Cer84, Cou86, Fre89, Gie81a, Han89c, Hig80b, Lau86, Mil80a, Mul83a, Sch80, Smi88a]. Hiyal [Kel80]. HMS [Emm80a]. Hobbes [Han88]. Hochheiser [Hei88]. Hoe [Bot80]. Hoffecker [Lin85]. Hoffman [Woo86, Pow83]. Hofsommer [Sto88]. Hogg [Wei81a]. Hoghead [Sha83b]. Hohenemser [De 83]. Hohmann [EE81]. Holdings [Sta82a]. Hole [Don86a]. Hollerith [Hym84]. Holley [Hal86b, Yat81]. Holliday [McI83]. Hollister [Bil81a]. Holocausts [Boy89]. Holton [Rav81, Wil84c]. Home [Upt83b, Tro84b]. Homer [Kop82]. Homes [Cow83, Bel82]. Homesteaders [Gra88]. Homo [Rah82]. Honeyman [Gut84]. Honnecourt [Hey88]. Honor [Gre88c, Iri89a]. Honour [She84, Wei80b]. Hook [Lov82]. Hoover [Fer82, Rae80b, Sti85b, Sta86, Jac89a]. Hopes [Sla88]. Hopewell [Hyd86]. Hopkins [War87b]. Horace [Vac82]. Horgan [Bil88]. Horikoshi [vdL83]. Horizon [Com87b]. Horizontal [Isr87]. horlogère [Smi86c]. horlogerie [de 82b]. Horn [She81a]. Horne [Bed89b, Har89a, Ken86]. Horological [Hok82]. Horowitz [Wil87c]. Horse [Bac85b]. Horsecars [Eis80a]. Horsemanship [Wel87a]. Horwitch [McD84]. hos [Jan87]. Hosmer [Wal82]. Hospital [Feu88]. Hospitals [Mel85]. Hounshell [Mey86]. Hours [Kra86b]. House [Hor86, Coo86a]. Household [BBM84, Coo87, Hac85, Thr82]. Housework [BBM84]. Housing [Bay80, Upt83a]. Houston [Sti85a, Tom83b]. Howard [Cri80, Hin80a, Rhe88, Wat80]. Howe [Fri88b, Lub87b]. Howell [Wes83]. Howlett [Bra83a]. Hoy [Kon84, Mel83b]. Høyrup [Och89]. Hoyt [Kle89]. Hsuin [Pan87]. Huancheng [Pet89]. Hubbard [Mar83b]. Hubbel [Emm83a]. Hudson [Har80a]. Hugh [Bir87, Gre88b, Hun87b, Kni81, Leh87]. Hughes [Bed83b, Rey84]. Hugill [Ame89]. Hugins [Hyd86]. Human [Ale85, Cut81a, Fow80a, Har81b, Jon81, Kap86, Kem86b, Maz84b, Mor81b, Och87, Per80b, Pla86b, Rot85b, SF87, Sok81, Tha86, Wei80a]. Humane [Law85]. Humanism [Dur83a]. Humanist [Sla82]. Humanities [CL80, GK81, GS81, Gol80b, Hei86, Kel84c, KT80, McC88, Pfe83, Rec88, SB82, Wes87a, Wos82, Goo80a, The84a]. Humphreys [Reu85]. Hundred [Con85c, Fer83, Kar86]. Hungary [Haf85b, Win82]. Hunt [Mau84, Con81c]. Hunter [Lew80, Con81f, Saw88]. Hurley [Bri81a]. Hurt [Fit83]. Husbandry [PS82]. Hussite [Kel84b]. Hutslar [V88]. Hüttenwesens [Wer82]. Huxford [Smu85]. Huygens [Mul83d, Sch82d]. Hyde [Mol84b]. Hydraulic [Reu85, Sha88]. Hydraulics [Sha88]. Hydro [Fri82b, Fri82b]. Hydroelectric [Fri83]. Hydrogen [Per80d]. Hydrogenation [Str85a].

Hydrological [Gor85b]. Hyman [Tro84a].

I.T [Law85]. I.W.W [Arg88]. IA [Atk89, Atk89, Lan83a]. Ian [Buc89, Hul88, Joh89, Lif85, Lin81b, Mor87d, Sei80, Slo81, Vie83a, Wel87b]. IBM [Mei88b]. Icarus [Tri89]. Ice [Mar88]. ICOHTEC [Buc83c, Buc85, FW87]. Iconography [LB82]. Idea [Gut85, Ham88, Wil87b]. Ideal [Gor88c, Mei88a, The83, Jan87]. Idealist [Wik80]. Ideals [Cha89, Jan87]. Ideas [Ame89, Hou83, Hal86b]. idées [Gil82]. Identities [Sin87]. Ideological [Ste80]. Ideologie [Fis82, Kra81a]. Ideology [Kem86c]. idyll [Jan87]. IEEE [Ano80d, Ano88e, Ano89e]. IEEE’s [Car86]. Ignorant [Wik80]. Ihde [Dur85, Res81]. ihrer [Mul89b]. II [Buc82a, Gli89, Hor88, Min82, Tre87, Lew82b, Muh81, Siv80, Tut81, Hal84b, Rot82, Rup88, Swo89]. III [Emm81, Fos82, Sch82a, Vin86b]. Illustrated [Car83b, Eis80a, Gli88, Har81a, Hen87, Hig83b, McD85a, Mil88b, Mol84b, Nee83]. Illustrating [Bur82]. Illustrations [Kel84b]. Im [Web85b, Bed80b, Bra80a, Mul89b, Web81b]. Image [Gun82, Lay87, Sin87]. Imagery [Hun84]. Images [Bry81, Gut86, Ber83]. Imaginary [Dur83b]. Imagination [Smi82a]. Imagined [Bac82b]. Imagining [Sla89a]. Immigrant [Mey85a]. Impact [Bee83, Bel84b, Bro86b, Bry81, Coa83, Cos85, Cut81b, Fol82, Gor88a, Hab85, Hev84, Lif85, Moh88, Nel82, Ros87a, Soc89, Smu85]. Imperative [MB89]. Imperatives [Tei88]. Imperial [Bla87, Elb88, Mac85c]. Imperialism [Cam89, Sti82]. Importance [Lan85]. Important [Con86b]. Imported [De 84a]. Improve [Jor89, Kon89]. Improved [Jor89]. Incandescent [Car89]. Incident [Lan88]. Increase [Bir83]. Incubator [Coo84a]. Indefensible [How85a]. Independence [Sta83a, Woo86, McG85b]. Independent [Ata82]. Index [Ano80e, Ano81d, Ano82d, Ano83g, Ano84d, Ano85e, Ano86e, Ano87d, Ano88f, Ano89f, Och89]. Indexing [Goo80a]. India [Mor87d, Qur86, Rah82, Qur86]. Indian [And86, Mic84a, Wei81b]. Indiana [Law89, Sal83]. Indians [Keh84]. Indicators [Rei87]. Individual [Ano83e, Ano85d, Ano87c, Ano89d, Lew80]. Individuals [Hou83]. Induction [DAK89, Kli87a, Sta88b]. industria [Bul81a]. Industrial [And88, Ano83f, Atk89, Bah87, Bat82, Bau86, Bez85, Big88, Bro87a, Bro84, Bry81, Buc82c, Buc83a, Con81f, Coo84a, Coo86b, Dan83a, Dub83, Esp82, Feu88, Fol80a, Fra84, Gar81a, Ger84a, Gor82, Gre88b, Gut84, Gut85, Har80a, Har82b, Har83a, Har85, Hev89, Hin85, Hit85, Jac80, Jer88, Joh88, Kra89a, Kra86b, Kul88, Lan89a, Lan83a, LR85, Lan88, Les86a, Lif85, Maj81, McD85a, Mel83a, Mel87, Mer81b, Mey85a, Mil80a, Mir82, Mul84a, New83, Pas85, Pen82, Pre83b, Pur82a, Ran82b, Rob85, Roe89, Saw88, Sch86a, Sha83b, Smi80b, Smi87a, Sta88a, Sti87, Str84, tZS82, Tuc82a, Uss88, Vic84, War86a, Wei84, Whi80, Whi85, Win89, Wis80, Wri86, Pel83, vB83, Coo83, Gar81b, Gor85a, Lin81b]. Industrialisation [Car80, McI84, Pel83, And88, Wal88]. Industrialism [Bah87]. Industrialism/The [Bah87]. Industrialization [Bur83, Coo84c, Cru87, Fol80b, Kas83, LR85, Lew89, Moh87a, Rot83, Tul85, Uss88, Bro86c]. Industrialized [Sta81]. Industrializing [Lub87a]. Industrie [Mor89e, Web84]. Industriekultur [Kli86b]. industriële [vB83]. industrielle [Smi85f]. Industries [Coo86b, Hou80b, Koh80, Pec86, Ran86a, Mei81, Hil83]. Industry [Ada89, Ben88, Bow80, Bro85b, Coc81, Coo85, Fis88, Gar81a, Gor83, Gor85a, Gué81, Har85, Har87c, Has86, Hyd86, Jer85, Kul88, Lan89b, Lau87, Lub87b, Mac85c, Mar89a, McG85a, Mil82b, Mil82c, Mul83b, Mul89c, Neu89, Nic80, Nuw88, O’B87, Ola85, Rae87, Rey89c, Roh84, Sau80, Sch85a, Scl88, Sho88, Smi85a, Smi86b, Str84, Tan81, Tod86, Tuc80a, Tut81, Twe88, Wal87b, War86b, Wis80, Wri80, Str88, Hoy88, Isr87, Sal84, Sin88b, Ada89, Bel84b, Bud82, Car89, Don81a, Don86a, Dre89, Gar81b, Hac80b, Hev89, Hil84, Hit85, Hoc86, HS81, Les86b, Lie83, Lif85, Smi89b, Str88, Tan89b, Vie83b, Way89, Wil83]. Inequality [Wal88]. Inevitable [Cha88a]. Infantry [Tre87]. Infection [Bul81d, Per81c]. Infernal [Rol87]. Influence [Bea81, Lan84]. Influences [Tar81]. Informal [Beb80]. Information [Cos88, Dun81, Hym84, Lev87c, Mor82, Nev82, Qur86, Smi83b, Kid89, Ber84, Law85, Shu84a]. Infrastructures [Qur86]. Ingeniería [Har87d]. Ingenieure [Tod85, Web85b, Bra80a]. Ingenious [Kel80]. Ingmar [San81b]. Injector [Kra82]. Inland [Abb86, Jac88]. Inmos [Tro87]. Innis [SPV87]. Innis/McLuhan/Grant [SPV87]. Innovates [Got83]. Innovation [All87, Bez85, Boy85, Bro84, Coo83, Fri82a, Fro88, Haf85a, Isr87, Jam89, Kul88, McG85a, Ros84, Sha81a, Vin84, Wei87a, Wes86b, Fri84b, Mil82a, Dre89, Pen84]. Innovations [Mac85a, Wol83]. Innovator [Ste87]. Innovators [Gra87b]. Inquirers [Swe88b]. Inquiry [Mar80, Rav81, Wil86b]. Insecticide [Gil85b]. Insects [Gil85b]. Instability [Gar81a]. Institute [Bez85, Bra80b, GK81, GS81, Har81b, Kel84c, KT80, McM86, Mor89d, Ove85, Pfe83, SB82, Tom85, Tur85, Win87, Far80]. Institution [Gol87, Hok80, Rey89a]. Institutional [Gri81a, Lev87b]. Institutions [HH87b, Han87c, Goo82a]. Instrument [McK80, SL80]. Instrument-Makers [McK80]. Instruments [Bed83a, Bed85a, Bed87, Bed89c, Fin83, Ger81, de 82b, War87a, War89b, Bed86, Goo85a]. Insurance [Ros85]. Insuring [Scl88]. Integrated [Sar81]. Integration [Nel81]. Intelligence [Jas87, Wil87b, Lub86b]. Intelligentsia [Lie81b, Lie81a]. Interactions [Wei86]. Interchangeable [How80b]. Intercontinental [Smi83d]. Interdependence [Gor89a]. Interface [Hei86, GK81, GS81, Kel84c, KT80, McC88, Pfe83, Rec88, SB82, Wes87a]. Interfacing [Haf85a]. International [Ano80l, Bla87, Bra80b, Bro86b, But87, De 84c, Far80, Gor82, Haf85b, Haf85a, Hil86, Hoc86, LR85, Mor89c, Mor80d, Run81, Scr86b, Spi82a, Tar85a, Tod84b, Uss88, Vic84]. Internationales [Wer82]. Interpretation [Cur84b]. Interpretations [San80]. Interpretive [Lov82, Far80]. Interrelationships [Coo85]. Interstate [Fos80]. Intimate [Dav89c]. Introduction [Lan80a, Sta87, Dan81a, Fri88b, Bil80b]. Introductions [Bed80b]. Introductory [Wer81]. introuvable [Kra85]. Invaded [McG88a]. Invasion [Nic80]. invención [Gli82]. Invented [Hev86, Per86]. Inventing [Fri89a]. Invention [Als81, Ber87, Doo87, Fri89b, Isr87, Kli87a, Lan85, Mac87b, New85, Sim87, Smi85c, Vac84, Whe84, Fer88, Jen84, Hou80a]. Inventions [Moh87b, Seg87]. Inventive [Pur82a, Rob85]. Inventor [Abr80, Car84a, Mol84a, Rob80]. Inventors [Pur81]. Inventory [Coo84a, Mer81b]. Inversion [Bry82, De 82a]. Investment [Wil84a]. Invisible [DeV85, Pin85]. Invited [Sch82d]. Iowa [Gia81]. Ira [Smi89a]. Ireland [Buc83a]. Iron [Bro85b, Bry81, Buc82b, Con81e, Gor83, Gor85a, GR86, Leh87, Mar87a, Pet83, Rey89c, RBD84, Ros87b, Sin88c, Ste81a, Vig83, Wes83, Bry81, Jac86, Par82, Pay85, Law89, San84]. Iron-Production [Ros87b]. Ironies [Hac85]. Ironmaking [Kil88]. Ironmaster [San84]. Ironworks [Gor86]. Irrational [McD82, Wei81b]. Irrigation [Hab85, Jac86, Mor87d, Sel88]. Irving [Mor80d, Pal80, Rei83a]. Isaac [Mah82, Mul84b, Mul84b]. Isacson [And88]. Isern [Soc83]. ISIS [Goo82a]. Islam [Fle84]. Islamic [Wes86b, Gli88, Rui80]. Islamicate [Hil88]. Island [Bra84, Cro86, LS84, Maz84b, Per80c, Moh87b, Moh88, Gre84]. Isles [Gor88b, Lin81b]. Issadore [Moy80]. issue [Smi83e]. Issues [Bro86b, Mac86b, McL85, Pur80, She87, Hir84]. iteu [Ino82]. Ithiel [Fri86a, Shu84b]. Itinerarium [Hal86a]. Ivan [The84b]. Ivar [Lan86]. Izbranniye [EE81].

J [Ait81, Ali89, Ame89, Bar88, Bau86, Bed80b, Bed84, Bed87, Bil81b, Bil88, Bli82, Bow87, Bra83a, Bro86b, Bul82, Bus87, Cai82b, Cer85, Cha82c, Cla88, Coc82, Coh86, Coh89, Coo85, Cor89, Das85, Dou84, Duf80, Elb87, Fli89b, Fri89a, Fri83, Fri88b, Gar81b, Gol87, Goo89, Gre88c, Gui89, Hac87, Hal82, Har83a, Hic85b, Hig82, Hig83a, Hod89, Hol83, How80a, How84, Hug88b, Hun87b, Jac86, Jan86, Joh88, Kel89, Kem84b, Kon89, Kor86, Kru85, Lan84, Lav84, Lev83, Lew85, Lub87a, Maj81, Maz83, McG85b, McI81, Mer87, Mil88a, Mir82, Mor87e, Mul83d, Ovi88, Pau83, Pri82, Pur82b, Rey89a, Sch82d, Sie81, Sla89a, Sma80, Smi81a, Smi83a, Ste82a, Tan89b, Tho84, Tob85, War89b, Wei80a, Whi85, Woo86, Yor88, aH80, vB86]. Jack [Cai88, Cla86]. Jackson [Buc86b, Car88, Mil89]. Jacob [Ata82, Bur80a, Cha80, Col81, Rot83, Web87b]. Jacobs [Wil80]. Jacobsen [Hab85]. Jacques [Del82, Sar81, Ung83, Lov82]. Jaffe [Wil81]. Jahre [Bul89]. Jahrhundert [Web81b]. Jakle [Ber83]. James [Bil88, Bro86c, Buc81a, Cas88b, Cor89, Cos88, Cou86, Dan83a, Dun89, Fel84, Gor87, Gro87a, Hir85, HS81, Kop89, Law89, Law85, Lic85, Luk89, McD86, Mil84a, O’C88, Par82, Rae80a, Ran86a, Seg87, Sha84, Smi81a, Smi88b, The84a, Tri83, Wri86, Yat82, dlB88, Bir83, Pre81, Car81, Tan89b, The84b]. Jan [Haf85a, Mic84a]. Jandl [See87]. Jane [Hig83a]. Janis [Day89]. January [Bri81b, De 84a, Ell87, Pyc82, Ano80h, Ano81g, Ano82f, Ano83i, Ano84g, Ano85h, Ano86j, Ano87h, Ano88k, Ano89j]. Japan [Got83, McL85, Min82, Kob80]. Japanese [Ino81, Ino82, McL85, Rae87]. Jaramillo [Gor84]. Järnvägar [San81a]. Jarrell [Mac86b, Pal81]. Jarrett [Tri89]. Javanese [Smi80a]. Jay [Hic85a, How85a, Lov82, Mol84a]. Jean [Bel84b, Her88, Kra85, Mar88, Sti86a, Tuc80b, Wil84b]. Jean-Claude [Her88, Kra85]. Jean-Pierre [Bel84b, Mar88]. Jeanette [Mul82]. Jeanne [De 83, Yat81]. Jeannot [Pet87]. Jed [Swe88a]. Jefferson [Bed85b, How87, Ras80b, Hin80a]. Jeffrey [Cha87, May89, Meh87, Tuc82a]. Jenkins [Jer85]. Jenkinson [Bly88]. Jennifer [Mus83]. Jequier [Smi86c]. Jeremy [Ger84a, Gri81b, LeM86, Mir82]. Jerome [Zee86]. Jerry [Whi84b]. Jersey [Per85b, Win87]. Jerusalem [Wil84c]. Jessop [Bal80]. Jet [Bil81b, Han85, Smi81a, Wri80]. Jim [Lue86]. Jimmy [Ful81, Wil89a]. Jiro [vdL83]. Joan [Ano80f, Ano81e, Ano82e, Ano83h, Ano84e, Ano85f, Ano86f, Ano87e, Ano88g, Ano89g, Haf85a, Hin80b, Lyt88, Pur89b, Sta85b, Whe84]. Job [Rup88]. Jobs [Cos85]. Jocelyn [Kor80]. Joe [Coa83]. Joel [How85a, Kli86a, Man81, Nas80, Por89b, Ros89]. Johan [Web87b]. Johann [Bül88]. John [Arm86, Bay80, Bed83c, Bed89b, Bee83, Ber83, Bil80b, Bil85a, Boy88, Bus87, Cai82a, Car81, Cas88a, Con85b, Cut81b, Dav85, Daw89, Emm83a, Eve86, Fee85, Feu88, Fle88, Fre89, Gil85b, Goo85a, Goo82a, Gre88a, Gué81, Had81, Had82, Hal84a, Har89a, Hof85, Hor84, How86, Kah85, Kap80, Ken86, Kle89, Klu82, Kru85, LGWK80, Les86b, Lif85, Lin81b, Mal86b, Mar89a, Mar83a, Mar88, May89, McD85a, Mil83, Per80d, Ran82a, Rol83, Scl88, Shu84a, Smu87, Sta82b, Sti81, Tha86, Tra85, Web85a, Wil88, Wil87b, Wis84, Fli89c, Gor89b, Kar83, Lub87b, Sin82, SL80, Tan82a, Ste82a, Woo82]. Johnny [Soc89]. Johns [War87b]. Johnson [Mac87a, Dob81b, Gie85, Scl88, Swe82]. Johnston [Coa80]. Joint [Bra81a]. Jon [Blo81, Mac85b]. Jonathan [Ada89, Del83, Hin85, Tan89a]. Jones [Bog85, Wil83]. Jordan [Hud86]. José [Cum86, HS88, Mau82, Ovi85]. Josef [Hat89, Kir80]. Joseph [Ata82, Bac85a, Bed89c, Bil82, Bri81b, Col81, Cro87a, Ell87, Gie86, Hig80a, Jac89a, Kid87, Lev83, Lew85, Mul88a, Pau83, Per81a, Per80c, Pyc82, Rot83, Win82, Mor87e]. Joshua [Zig88]. Joule [Car81, Pre81]. Journal [Atk89, Gia81, Hal86a, Lan83a, Mor87b, Sta88a, Eng80, Had81, Had82, Hig80b, Nas80, Ros89]. Journals [Sha84]. Journey [Kam81, Whi83a]. Journeymen [Ran82b]. Joy [Bra84]. Joyce [Kam81, McK80]. JPL [Han85]. Jr [Cas88b, Cum88, Har86a, Kul88, McG88b, Mol88, Wes86a, Str80]. Jr. [Baa89, Bil80b, Bol88, Coo85, Dal88, Dow84b, For88a, Gor87, Gra87a, Hal89a, Hal84a, Hal86b, Han87b, HS81, Hou83, Joh83, Kli86a, Kli86c, Koh87, Kra89a, Lev83, Mar86c, McG84, Pap82, Pau83, Per84a, Rei80b, Shu84a, Smi88b, Sta82b, Tei88, Wal82, Wea83, Wil84a, Woo82, Gre88c]. Juanelo [HS88, Ovi85]. Jubilee [Tuc82b]. Judd [Ata82, Col81, Rot83]. Judiciary [Hol89]. Judith [Ber89, Mor83, Seg89a]. Judson [Vac82]. Judy [Hoc87]. Julia [Bil83]. Julian [Cam88, Coo83]. July [Hac86, Ano80j, Ano81i, Ano82h, Ano83k, Ano84i, Ano85j, Ano86l, Ano87j, Ano88m, Ano89l]. June [Bil81a, Buc83c, Bur80b, Har81b, LR85, Scr86b, Tom83b]. Jung [San81b]. Jungle [O’C88]. Juniper [Tro84b]. Jupiter [dB80]. Juptner [Bil82]. Jurdant [Sch82c]. Just [All86, Qui89]. Justice [McG86].

Kaelble [Wal88]. Kahrl [Ble83]. Kaijser [Lun88]. Kali [Bra82a]. Kali-und [Bra82a]. Kansas [Ros84]. Kaplan [Dea85]. Kapp [Lan80a]. kara [Ino82]. Karal [Upt88]. Karen [Cla88, Lub86b, Swe88b, Wor88]. Kargon [Gre80, Mic84b]. Karin [Pin84, Wil85a]. Karl [For88a, Kli87b, She84, Tod85, Web84, Web84]. Karl-Heinz [Tod85]. Kasperson [De 83]. Kasschau [Ber84]. Kathleen [Nev82]. Kathryn [Whi86b]. Katologe [Tod87]. Katrina [Gut84]. Katz [New83]. Katzen [Cur84a]. Kax [Pon80]. Kealey [Bul82, Kel89]. Kebabian [Bat80, Hin81]. Keilty [Mar84b]. Keith [Mar80, Smi85e]. Keller [Sta85a]. Kelly [Sha81a, Buc89]. Kelp [Neu89]. Kelsey [Hal84b]. Kelvin [Moy80]. Kemp [Fri88b]. Ken [Fri81c, Kap80]. Kendall [Lie81a, She87]. Kennedy [Eat88, Wal83]. Kenneth [All86, Ber84, Har80a, Nee83, Sah87, San80, Wes83]. Kent [Ste81b]. Kenwood [McI84]. Keris [Smi80a]. Kern [The85b]. Kessler [War83]. Kessler-Harris [War83]. Kettering [Rae85]. Kevin [Str84]. Kevles [Sie81]. Khrushchev [Hol87a, Wal83]. Kidwell [Wal81]. Kill [Rig81]. Killian [Gor87]. Kimball [Kar81]. King [Elb88, Gre88b, Jas87]. Kingdom [Vie83a]. Kingsport [Seg89b]. Kingston [Smi83c, Tri83]. Kinship [McD85a]. Kitāb [Kel80]. Kitchen [Coo87]. Klaas [Mul84b]. Klaus [And82, Bed80b]. Klawonin [Sma80]. Klein [Gre88c]. Klemm [May84]. Klosterman [Mil88a]. Knorr [Wil85a, Pin84]. Knorr-Cetina [Wil85a, Pin84]. Know [Coh86]. Knowledge [Bij86, Bin83, Bro86a, Fin83, Pin84, Vey89, Vin84, Vin86a, Dor80, Sha81a]. Knowles [All84, Sin86a]. Known [LB82]. Knut [Blo81]. Kobayashi [Ino82]. Koeper [Con82]. Kolefs [San81b]. Komons [Bil80a]. König [Tod85]. Königreich [Bra80a]. Konrad [Hor86]. Konvitz [Hat89, Kir80]. Koppes [Han85]. Korean [Bed89c]. kosatsu [Ino82]. Kouwenhoven [Tra85]. Kovel [How85a]. Kozo [McL85]. Kraft [Ali89]. Kranzberg [Ros81, Gol81, Gol89, Sha81a]. Kranzberg’s [Kra86a]. Kraus [She80a]. Kren [Bul86b]. Kroker [Gor82, SPV87]. Krysko [Zee82]. Kuehn [Pur82b]. Kuhn [Fli89b, Mul89b]. Kuhner [Fer82]. Kula [Zup88]. Kulik [Lan83b]. Kultur [Bra82c, Bul89, Mul89b]. Kultur- [Bul89]. Kunst [Bed80b, Zee82]. Kursunoglu [Dow85]. Kurylo [Car84a]. Kusum [Wei81b]. Kuznick [Hod89]. Kylhammar [Jan87]. Kytle [Tro84b].

L [Bar89a, Bed80b, Ben88, Bla84, Ble83, Boy86, Bra84, Bro83, Cas88b, Cer84, Con88b, Coo84a, Cri80, Dav83, De 84b, Fol82, Fri88b, Gie81a, Gre84, Hac88, Har82a, Hin86b, Hol88, Kli87b, Kra88b, Kul88, Law89, Leh83, Lew80, Man81, Maz84b, McG88b, McI84, Mei81, Mol89, Nas89, Och89, Per80d, Pur82b, Qui89, Ros85, Seg86, Sho88, Sla84, Smu87, Spu89, Sti85a, Sto88, Str84, Swe82, Tei88, Tuc82a, Vey89, Wes86a, Wik81, Wil81, Wis81, Wor88, dlB88, Sta86]. L. [Emm80b]. Labor [Ben88, Coo86a, Dal88, Dub88, Fro88, HH87b, Joh83, Mei88a, Mey85b, Moh87a, Ovi86b, Per85b, Ran86a, Rum88, Sch82a, Scr88a, Scr88b, Smi80b, Moh87a]. Laboratories [McG83, Han87c]. Laboratory [Hod89, Kop89, The86, Wis81, Han85]. Laboring [Mey85a]. Labour [Bli82, Cro82, Lif85, Nel87, Tan89a]. labyrint [vB83]. labyrinth [vB83]. Labyrinths [Vig83]. L’Accueil [Gil82]. Lachman [Ber84]. Lackawanna [Cas88a]. Ladrière [Tuc80b]. Ladurie [Dun80]. Lag [Sch81a]. Lake [Lay87, Lan83c]. Lakes [Mar85c]. Lakier [Kam81]. Lamb [McC85]. Lambert [Per85b]. Lambright [Mac87a]. Lammot [Ola85]. Lamp [Car89, Mul85]. Lamp-Lanterns [Mul85]. Lampert [Lie81b]. Lance [Zim84]. Land [Bed83b, Bha87, Har87c, Hof84, Iri89b, Maj81, Mul83d, Seg83, Tin87, Hof84]. Landau [Lin88]. Landes [May85]. Landscape [Han89a, Mar87c]. Lane [Hew87, Min81]. Langdon [McG88c]. Langenbach [Hof80]. Langins [Day89]. Langley [Kop89]. Langmuir [Rei83a]. Language [Joh83, Mil87]. Languages [Cer84]. Lankton [Mol84b]. Lanterns [Mul85]. Larder [McG88a]. Larry [Buc86d, Dow84a, Mol84b, Web85a]. Lars [And88, Bra80a]. Larson [Cai82a, Hof85]. Lasansky [Mul82]. Laser [Bro86a]. Lasers [Bro85a]. Lásló [Web81a]. Last [Bra81b, Seg84, Swe88a]. Late [Con87, Har83b, Nye89, Qui89, Sim85]. Later [Bed85a]. Latham [Bed85b]. Latimer [Mar86a]. Latin [Moo86]. Laton [Pur89a]. Latour [Bij88]. Latrobe [Bed89b, Har89a, Ken86, McM81a, Hor88]. Laudan [Bij86]. Laughery [Ber84]. Launched [Hev86]. Launching [Por89a]. Laurence [Han87b, Lew88b]. Laurent [Ger84b]. Lauretis [Smi82a]. Laurie [Wei81a]. Lauritz [Hof85]. L’Automate [Kra85]. Laver [Tay81]. Lavington [Ste81b]. Law [Bro86b, Das85, Ful80, Bro86b, Blo81, Rot83]. Lawrence [Hac80a, Hal80b, Jac83, Kan85, Kem84b, Pap82, Pen84]. Laws [Kra86a]. Layton [Kli86c]. Lead [Har82b, Law85, Mac87b, Zee86, Zee82, Zee86]. Lead-Crystal [Mac87b]. Leadership [Gut84, Jac83]. Learning [Scr86a]. Leary [Cas88a, Lev87a]. Leavitt [Mor83]. Leben [Neu88]. Lechtman [Spi82b]. Lee [Jac83, Ken86, McD85a]. Leeds [Cla88, Min81]. Leeson [Gar81c]. Legacy [Gor86, Smi83c]. Legend [Hok87, HS88]. legende [Gil82]. Legislators [Sch85a]. Legitimacy [Dan87]. Leif [Mor80d]. Leistritz [Dow84a]. Leisure [Kra86b, Lub87a]. Lela [McG86]. Leland [Con80a, Con86a, Dob81b]. Lemonnier [Mul83c]. l’Empire [Gut87]. l’énergie [Her88]. Lenin [Lie81a]. Lenny [Wes84]. Leo [Fah89, Mil88a, Sla89c]. Leonard [Bro87a, Cut85, For88a, Got83, Hal80a, Ste85]. Leonardo [Ano80g, Ano81f, Ano84f, Ano85g, Ano86g, Ano87f, Ano88h, Ano89h, Car82b, FRCL83, Hal88b, Hal88a]. Leonhardt [Kem86a]. Lerbach [Buc85]. Lerche [Spi82a]. LeRoy [Dun80]. Leslie [Bog85, Mau84, Mil84b, Tuc80a, Rae85, Maj81]. Less [Lew80]. Let [Kra84]. Letter [Abr81a]. Letters [Kar81, Sha84, McG86]. Lettuce [Ras80a]. Level [McL84]. Leviathan [Han88]. Levidow [Nel87]. Levine [Han87c, Moh87a, Pia84, Zig85]. Levinson [Ano86i, Ano87g, Ano88j, Ano89i, Ano86h, She84]. Lewis [Boy86, Bri80b, Bri85, Coo86b, Dav89b, Elb87, Hyd86, Mol89, Win87, Har82a, Mor89c, Pel83]. Lex [Mer81a]. l’histoire [Cen80, Gil82]. Liberal [Har81b, The83]. Liberation [Gie81b, Bin83]. Liberty [Han87a, Lan84]. Library [Goo89, Mor89d, Smi87b, Ste87, Wri85b, Sta82a]. Licht [Gra84]. Liebenau [Ada89]. Life [All86, Ano88e, Ano89e, Boy86, Bri85, Car84a, Dav83, For84, Gre88b, Hal80b, Hin85, Hof80, Hoy85, Hyd89, Lan84, Lic85, McC85, McG86, Mic84b, Rot85b, Sch82d, She81a, Sta85a, Ste82a, Sti85b, Wil86b, dB80, Per85b, Bla81, Han88]. Lifton [How85a]. Light [Cla88, Fer88, Hew87, Mul85, Wal87a]. Lighted [Dan81b]. Lighting [Gol80a, Mul85]. Lights [Lun88]. Like [Rit88, Rit88]. Lilienfeld [Ste80]. Lilleshall [Rea81]. Lillian [Kas84]. Lima [Mar89b]. Limited [Gor83, Rea81, Tuc82a]. Limits [Bla88, Con80c, McG88c, Con80c, Rav81]. Lind [Eis80a]. Linda [BM85, Cre83, Gru86, Sha81b]. Lindenberg [Rap81]. Lindgren [Bül88, Kli87b]. Lindqvist [Dan83b, Sta87]. Lindsay [Mil88b, Wes87b]. L’Industrie [Gut87]. Line [Bra84, How85b, Sta83c]. L’Ingénieur [Fro87]. Linhart [Sta83c]. Lipke [Hin81]. Lipton [Wes84]. Liquid [Per80d]. Lisa [Sla89b, Whe84]. Lissarrague [Lan87b]. List [Pur83]. Listener [Kle87]. Listening [Kid87]. Lister [Mye89, Mye89]. Lit [Kar86]. Litchfield [Hyd86]. Literature [Hol86b, LS84, Mei88c, Seg84, Sla89c, Das85, Hun84, Lew88b, Mel83b, Mor80b]. Lithographic [Hun84]. Lithographs [Hun84]. Live [GF87, Ros81, Gol84]. Lives [Dav89c, Ger84a, Mar86c, Sho88, Ber87]. Livestock [Wil89a]. Living [Sim86, Mac86a]. Llewellyn [Sti86a]. Lloyd [Aga82, Lif85]. Loaders [Smi84b]. Lobstering [Gro88b]. Local [New85, Sin86b, Wes86a]. Locally [De 81b]. Lock [FRCL83, Rey89b]. Locomotive [Leh83, Mar83a, Sem84, Sha83b, Cor88b, Sta82b, Duf80]. Locomotives [Spu89, Whi80, Whi83c]. Lodge [She81b]. Loftin [Han87b]. Log [Hud86, V88]. Logging [Hul88]. Logique [Mul83c]. Lokken [Dan82]. Lombard [Cum88]. Lon [Gra87a]. London [Fli89a, Har82b, Ran86a, Tuc82b]. Lone [Mil88b]. Long [Bra84, Dow84a, Gre81, Isr87, McN83, Whi86a, Cen80, LS84, Ovi87]. Long-Distance [Isr87]. Long-Term [Dow84a, McN83]. Longer [Cro86]. Longley [Cor87]. Look [Han87c, Har87b]. Looking [Lay87, Tro87]. Looks [Kam81]. Lopez [Smi81a, Mau82, Bil81b]. Lord [For84, Wes86a, Moy80]. Loren [Rei83b]. Lorin [McD85a]. l’Orme [San89]. Lorraine [Con87]. Lost [Cut87]. Lougheed [McI84]. Louis [Con81f, Eis80a, Fri83, Saw88, Sma80, Wil84a, Wis81, Bil88]. Louise [Ska88]. Louisiana [Mar89a]. Louvre [Whi83c]. Love [Pia84]. Loveday [Coo85]. Lovejoy [Mul88b]. Lovell [DeV88]. Low [Mor87e]. Lowell [Bah87, Haf80, LR85, Mel83a, Wei84, LR85, Mel87, Wei84]. Lowitt [Ale86]. Lu [Kid87]. Lubar [Jer88, Kas83]. Lucas [Fie84]. Lucile [De 81a]. Luddism [Ran86b]. Ludwig [Tod85, Bru85]. Luftwaffe [Mor86, Mor88]. Luisa [Bed80a, Bed82]. Lumber [Nei88, Sch85a, Smi85a]. Lumberjacks [Sch85a]. Lunar [Hal80a, Swe82, Zig88]. Lundstrom [Wre89]. Lung [Bin89].

Lyle [Wor81]. Lynch [Cla89]. Lyndsay [Bin83]. Lynn [Pre83b, Str80, Got83, Hal89a]. Lytle [Mar85b, TL88].

M [Ben86, Bla88, Bli82, Bro86b, Cha82c, Con80a, Con86a, Cro83, Cut81a, EE81, Fol82, Fri86d, Gie85, Gol80b, Gra88, Hac84, Hac86, Hal82, Jer85, Kem84a, Kin81, Kon84, Kor86, Lau87, Lev87a, Lov82, Mac88b, Mac85a, Mar80, May89, Mel83b, Mer88, Mul89c, Pap84, Pre83a, Rap81, Roe85a, Sch82d, Scr87b, Sei80, Sha84, Sha81c, Shu84a, Sla89b, Slo81, Sta88b, Ste81b, Tha86, Way89, Web80b, Wei80b, Wel87b, Wes86b, Wil89a, Woo82]. Mūsà [Kel80]. Maarten [vB83]. maatschappijgeschiedenis [vB83]. Macaulay [Fol85]. Macchine [Bul81a]. MacDonald [Cou88, Hev84]. Macey [Dav83]. Macfarlane [How86]. MacGregor [Wri83]. Machaut [Koe84]. Machiavelli [Hea88]. Machina [Sta85b]. Machine [Ber89, Cum88, Hev89, Jan87, Kel88, Lan83c, Lew88b, Mac84, Mei87, Moh87b, Pen80, Rot84, Sin88b, Wil89b, Wos82, Yor88, Iri89b]. Machinery [Han87a, Nic81]. Machines [Rol87, Sta89, Com87a, How87, Ovi85, Ste83, Har81a]. MacKenzie [Hoc87, May89]. MacMackin [The83]. Macro [Kan85, Les85]. Macro-Engineering [Kan85]. Made [Gie86, Whe87, Hor86, Baa89, Sla89b]. Madeleine [Koe84]. Magazine [Twe88]. Magic [For88b]. Magnificent [Wor88]. Magnusson [And88]. Mail [Joh87a, Kie87, Lev87a]. Maillart [Con80b]. Maine [Gro88b]. Maize [Hea89]. Major [Har81b]. Makepeace [Kar86]. Maker [SL80]. Makers [Cas88b, McK80, Wri80, Yat81, Har80b, Nel89, Reu89, Sei80, Swe88b, War87a, Hea88]. Makes [Iri89b]. Making [Bar89a, Ber89, Bro87a, Bud84, Cor88b, Cri80, Emm83a, Hai84, Leh87, May85, Mey85a, Mil86, Nei88, Nic81, Wan89, Zig88]. Mal [San81b]. Maladies [Bul87]. Malcolm [Bir87, Bud82, Tar81]. Male [Ste88a]. Malim [Bow87]. Malin [Cou86]. Man [Bro86b, Bru85, Car83a, Cer85, Coh86, Emm83a, HS88, How86, Kel86, She80b, Spu89, Whe87, Huy82]. Man-Made [Whe87]. Management [Coo85, Gil85b, Han87c, Hof84, Kan85, Mac87a, Mer88, Rae87, Sch82a, Shu84a, Smi89a, tZS82, Fri83, Hab82]. Managerial [Bat82]. Managers [Don86a, Mil84b]. Managing [Hoy88, Hab85, Zig85]. Manchester [Bau86, Feu88, Kar86, Gre80]. Manfred [Fer80a, Hof84]. Manhattan [Con88b, Tit89, Con88b]. Mankind [Gor88a]. Manned [Hal80a, Swe82]. Manners [Vie83a]. Manning [All86]. Manual [Mue89, Ovi86b]. Manuel [Seg81, Seg81]. Manufacture [Coo84b, Lub86a, Phi83, Pin84, Scr86a, Smi86c]. Manufacturers [Pre83b]. Manufactures [Hok80, Kas83, Kul88, Pur84b]. Manufacturing [Gor83, Gor85b, Gor89b, Lub87b, Mey86, Ras80a]. Manuscript [Bar85, Fri80b]. Manuscripts [Kid88]. Mao [Bra81a]. Mapmakers [Kar87a]. Mapping [Arl84]. Maps [Har87b, Kar87a]. Marc [Buc88a, Con85d, Mor80b, Row89]. March [Zim84, Hok80]. Marcus [Con82, Dan87, Whi86b]. Marcuse [Dur83b]. Mardges [Con88a]. Mares [Hok87]. Margaret [Bro88b, Car83a, Lan84, Mar86a, Pur83, Seg89b, War84].

Margery [Cos85]. Margiotta [Yao88]. Marguerite [Har89b]. Maria [Bed80a, Mau82, Bed82]. Marie [Cha82a, Klu82, Pla86a, Vic84]. Marietta [GK81, GS81, Kel84c, KT80, Pfe83, SB82]. Marilyn [Och89, Smu85]. marine [San81b]. Marinturbinens [San81b]. Mario [Bro85a, Con81d]. Marion [Sma80]. Maritime [Elb87, Gro88b, Lan84, Law87, Wos82]. Mark [Fri82b, Gro88a, Han87c, Hir84, Jac80, Kar87b, Mor87b, Ros89, Str88, The87, Con83a]. Marked [Dan81b]. Market [Ben86, Ber85, Bus83, San84, Tan81]. Marketing [San84]. Marketplace [Num88]. Markets [Bla87]. Marling [Upt88]. Marmin [Kar83]. Marquis [Sla88]. Mars [Gru86]. Marshall [Abr81b, Cut81b]. Marshes [Cro87a]. Martha [Coo80, Mul83b, The83, Wes86a]. Martin [Cal85, Con86b, Har83a, Jan87, Lan80a, Lan88, Mar87b, McS83, Per81b, Rol89, Sti87, Had81, Had82]. Marton [Wis81]. Marty [Ger84b]. Martyr [Mor87e]. Marvin [Nye89]. Marx [For88a, Web84, Fah89, Mac84, Sla89c]. Marxist [Nel87]. Mary [Vie83b, Kid85]. Maryland [War87b]. Maschinerie [Web84]. Masers [Bro85a]. Masi [Hok87]. Maskin [Jan87]. Mason [Shu84a, Kea89]. Mass [Hai84, Har83b, Hol87b, Mey86, Smu87]. Mass-Produced [Hai84, Hol87b]. Massachusetts [Fol80a, Haf80, Har81b, Hin85, LR85, Mel87, Pre83b, Lub86c, Gro88a]. Masses [Hil81]. Masters [Cro82, Ran82b]. Masturbation [Bul87]. Material [Ame89, Koh87, How84, Mer87, Sch82b, Spi82b]. Materialism [Gut86]. Materials [Goo85b, Gor83, See87, Hin86b, Mul81]. Math [Coh89]. Mathematical [McK80]. Mathematics [Ste82a]. Maths [And88]. Matthew [Coo84a]. Mature [Lub85]. Mauchly [Man83, Ste82b]. Maurer [Cla80, Hal81, Hal89b, Hig80c, Cla80, Hal81, Hal89b, Hig80c]. Maurice [And82, Cha82a, Goo85a, Mul87, Per84a, Smi85f, Bed80b, Her85]. Maurine [Nel82]. Maxine [Kul88, Nic81]. Maxwell [Smi85a, Swe88a, The84b]. May [Cur84a, Fra84, GR86, Hin84, MD82, Mac88a, Sch89, Nel89]. Mayday [Hol87a]. Mayr [Pur84b, And82, Han87a]. Mazuzan [Hac87]. McC [Row83]. McCall [Abb86]. McCann [Har89b, Sti84]. McCartney [Pur89a]. McClintock [Sta85a]. McConnell [Bed83a]. McCord [Gar82]. McCormick [Gar81b]. McCracken [Mei89]. McCusker [Bus87]. McCutcheon [Buc83a, Con80c]. McDaniel [Upt83b]. McDonald [Kru85, Mau84, Joh88]. McDonnell [Hal82]. McDougall [Hal87]. McElroy [Web85a]. McGaw [Ber89]. McHugh [Yat81]. McKeen [Sha84]. McKendrick [Wil84d]. McKim [Con86a]. McLean [Tro87]. McLuhan [SPV87, Abr81b, Dou84]. McMahon [Mil86]. McMains [Wor81]. McMath [Mol88]. McMurry [Cre83]. McNeill [Gui84, Sok81]. Me [Hig82]. Me-262 [Hig82]. Mead [Con86a]. Meador [Kan85]. Meal [Gle84]. means [San81b]. Measurement [Mac88b]. Measures [Zup88]. Measuring [Bed83b, Bed87, Fin83]. Meat [Vie83b]. Meatpacking [Wil89a]. Mechanical [Fer83, Fol80a, Gor88c, Min82, Row82b, San81a, Bed80b, Cor87, Sin86a, Sta88a]. Mechaniciens [Sle81]. Mechanics [Pre87, Cha82a, Fri86c, Pre83b]. mechanischen [Bed80b]. mechanisering [Mul84b]. Mechanization [Ber89, O’B87, See87, Mul84b, Boy88, Ras80a, Hea89]. Mechanized [Gor89b]. Medal [Ano80g, Ano81f, Ano84f, Ano85g, Ano86g, Ano87f, Ano88h, Ano89h, Car82b]. medel [San81b]. Media [Kie86, LeM86, Smu87, All88, Cur84a, Dou84]. Medical [Ada89, Fie84, Ada89]. Medicine [Gro87a, Haf85b, Hug81, Och89, Per80a, Ove85, Bul81b, Bul82, Mac86b, Feu88, Law83]. Medicus [Bul82]. Medieval [Ale85, Bed85a, Dal88, Fli89a, GR86, Jer85, Lan87a, Ovi87, She81b, Whi82b, Jer85, Ung81, Bul82, Bul86b, Str80]. Mediterranean [Gui89, Wil84b]. Meeker [Per81a]. Meet [Mar85c, Dan82]. Meeting [Ano88i, Ano89n, Car85, DKP^{+}86, DHW^{+}87, FBW^{+}81, HBR^{+}80, HRP^{+}82, LHH^{+}84, Scr86b, TLK^{+}83, TOE^{+}85]. Megaera [Emm80a]. Mehrtens [Fis82, Kra81a]. Meikle [Tuc82a]. Mein [Tod87]. Mekaniska [San81a]. Mel [McD84]. Melbourne [Moh87c]. Mellars [Row80b]. Mell’kumov [EE81]. Melosh [Mar85a]. Melosi [McS83, Per81b, Sti87]. Melvin [Gol81, Gol89, Ros81, Sha81a]. Members [Ano83e, Ano85d, Ano87c, Ano88e, Ano89d, Ano89e]. Memoir [Gor87]. Memorial [Web85a]. Memory [Jer85]. Men [Con85c, Cro82, Eat88, Jac86, Mar88, Wre89, Sem84, Zup88, McD86, Wes83]. Menard [Bus87]. Merchandise [Sta88a]. Merchant [Gui89, Jer82]. Merehant [Wri83]. Mergen [Jak87]. Merrill [Per80b, Spi82b]. Merrimack [Bar85]. Merritt [HH87a]. Merton [Tob85, Zie89a]. Mesolithic [Spi81]. Mess [Tut81]. Message [Pos82]. Messerschmitt [Hig82]. mestiere [Bul81a]. Metal [Law89]. Metallica [Fer82]. Metallurgical [Esp82]. Metallurgy [Fer82, Gor88b, Lec84, San84, Twe86, San81b, Rae80b]. Metals [Gor88a, HS81, Yat82, Mau84]. Metamorphosis [Cor87]. Methods [San89, San81b]. Métiers [Lan89a]. Metropolis [Bra83a, Cri80, Cud88, Moh87c, Blu86]. Metropolitan [Mil83, Con85b]. Metz [Zim84]. Mexican [Sho88, Way89]. Mexico [Smi89a, Wea83, Arn89, Hob80, Bee83, Sel88, dM85, Gli82]. Meyer [Lub85, Sch82a]. Miami [Seg89a]. Michael [Ali89, Bau86, Bel82, Buc89, Bül88, Cal81, Con81a, Dav89a, Har85, Hig83a, Hol87a, Jac81, Jac87, Joh88, Kas83, Kle87, Kru85, Man81, Maz83, Mel83b, Min81, Nas89, New83, Rae87, Seg83, Sel88, Sha84, SF82, Wik81, Wil85a]. Michal [Mil86]. Michel [Fow80b, Sch82c]. Michigan [DKP^{+}86, Mer81b, DR83, Hyd88b, Hyd89, LM87, Rey89c]. Mick [Tro87]. Mickey [Cha84]. Microchip [Hev86]. Microcomputers [Bol88]. Microelectronics [Cos85]. Microfilm [Cha87]. Microphysics [Swe88a]. Microscope [Nut81]. Microscopes [Dav80]. Microwave [Hac85]. Mid [Bil80a, Mar88]. Mid-19th [Mar88]. Mid-Century [Bil80a]. Middle [Bac85a, Bac86, Fri86c, Mul89c, Rui80, Sma80, Smi80b, Ung83, Phi83]. Middleton [All84, Rab81, Sin86a]. Migration [Joh89, V88]. Mihiel [Hal81]. Mile [Gre84, Maz84b, Smi88b]. Militarism [Tod84b]. Militarization [Hol87a]. Military [Bri81a, Gar88, HH87b, Hol83, Lan87b, Mul88a, Smi84b, War87b, Woo86, HH87a, Hev89]. Milkman [Rup88]. Mill [DD88, Hyd86, Lan83b, McG85b, Moh87b, Pen80, Fol85]. Millar [Pal80]. Miller [Abb86, Gle84, Hac88, Kas84, Mol89, Mul89c, Wat80, Kel82]. Millhands [Tul85]. Millikan [Mic84b]. Milling [FRCL83, Gle84]. Mills [Hyd86, Qui89, Goo85a, Wes86a]. Milton [Bro86b, Rol87]. Milward [Gei80]. Milwaukee [Mor83, HRP^{+}82]. Minc [dSP81]. Mind [Dou84, LeM82, SPV87]. Minds [Och87]. Mine [Rol87]. Miner [Mol84a, Mar86d]. Mineral [Roh84]. Miners [Bin87, Bin89, Don86a, Jac80, Tul85]. Mines [Bur80b, Lan83c, LM87, Bur80b, Tan82a]. Mingay [Bri89]. Miniature [Hev84]. Mining [Fer82, Mol84b, New85, Och88, Rae80b, Roh84, Sal84, Ste88b, Str84, Wil86a, You87b, You87a, San81b, Lan86, Lan89b, Smi88b]. Minneapolis [Fra84]. Minnesota [Jon81, Mar85c, Mar85c]. Minority [Sha83a]. Mintz [Wil87a]. Mirror [Lay87]. Miscellaneous [Bed89b, Har89a, Ken86]. Miskolc [Web81a]. Missile [Cas86, Gra87a, Sah87]. Mission [Jac87, Kra81b]. Mississippi [Goe82, Rey87, Sma80]. Missouri [Mar86d]. Misunderstanding [All88]. MIT-GE [Car88]. Mitcham [Lue86]. Mitchell [Emm81, Num88, Has86]. Mitchill [Bed85b]. Mitsubishi [vdL83]. Mitterling [Dun81]. mobiles [Kra85]. Mobilization [Coo86b]. Mobro [Hoy88]. Model [Bur80a, Com86]. modèles [Kra85]. Modeling [Sin88c]. Models [Bij86, Doo87, Fri85, Lau82]. Modem [Hea88]. Modern [Bar89a, Bat82, BM85, Big88, Bry82, Bur86, Col84, Con80c, DR83, Dur83a, Edg84, Gut86, Han89a, Han87a, Han87b, Hig80b, How85b, Kap82, Kir80, Lew88b, Lyt88, Mar84a, May85, Por89a, Rit88, Roe85a, Scr87a, Sie81, Sla88, Swe88b, Thi80, Thr82, Wei87b, Wik82, Wil87a]. Modernism [Meh87]. Modernist [Mei88c, Sla89b]. Modernization [Ber85, Fie84, Mar89a, McL84, Win82, Vol88]. Modesto [Kel84a]. Module [Zig88]. Moen [War82]. Molnar [Bur80a]. Moment [Tay89]. Monastery [She81a]. Money [Eat88]. Monmonier [Kar87b]. Monopoly [Tay89]. Monson [Rav81]. Montgolfıer [Lan85]. Montgomery [Scr87b]. Montreal [Hal88b]. Monument [Kea89]. Monuments [Gor82, Law89, Vic84]. Moon [Hal80a]. Moore [Coo80, Mul83b]. Morality [Wil87c]. Moreau [Hen86]. Morehouse [Qur86, Tuc80b]. Morgan [Chr89, De 80b, Run84, O’B81]. Morland [Rob80]. Morley [Roh84]. Morris [Con81c, Mei88a, Mor87c, Mul81, Ras80b, Sti86a]. Morrison [Cas88b, Mor87e]. Morrison-Low [Mor87e]. Morrow [Hal84a]. Morse [Car84b, Dou84]. Mortar [She80a]. Morton [Dur83b, Kor80]. Mosella [Sim83]. Moses [Man81]. Most [Bar91, Pur89a, Ste89, Ber89]. Motel [Rae80a]. Mother [Hac85]. Motherhood [Rot85b]. Motiv [San81a]. Motive [Kid88, Spu89]. motives [San81a]. Motor [DAK89, Fos89, Kli87a, Sau80, Sch82a]. Motorcar [Mar84b]. Motorcycle [Vol87]. Motorization [Bar88]. Motors [Fli89b, Rae85]. Mott [Pay85]. Mountain [HS81]. Mountaineers [Tul85]. Mountains [Wil86a].

Movement [Bli82, Bry80, Bry82, De 82a, Fli80, Jac83, Ste88c]. Movements [Fer83]. Moving [Hil81]. Moyal [Bha87, Gar87b]. Moyen [Pet88]. Mr. [Fri81a]. Mrozek [Hol83]. Mrs [Beb80]. Muenger [Com87b]. Muhly [Par82]. Mukerji [Gut86]. Muldowny [Kru85]. Mules [Joh87a]. Mulholland [Yat82]. Mulkay [Wil85a]. Muller [Bül88, Web84]. Multhauf [Tur89, Bee85, KP80]. Multi [Wil81, Cur84a]. Multi-disciplinary [Wil81]. Multi-Media [Cur84a]. Multinational [Scr87a]. Multinationals [Wil84a]. Mumford [Bri85, Mol89, Win87, Bri85, Mor89c]. Murdock [Dow84a]. Murphy [Sel88]. Murray [Hof85, Tay81, Mor86, Mor88]. Museo [Bed80a]. Museum [Arn89, Bar85, Bed87, Coo87, Dav87, Doo87, Eve86, Fri81a, Fri80b, Ger81, Hai84, Hal88b, Han80, Hoy88, Hyd88b, Hyd88a, Hyd89, Jac89b, Kel86, Lan80b, Lev88, Lub86c, MD82, Ovi86a, Per85b, Qui89, Sin88b, Smi87b, Tuc82b, Tur89, Vic87, Vol87, War88, Whi83b, Cud88, Gro88b, McM81b, Mei87, Nye88, Per85b, Rey89c, Whi83c, Wil86a, Bed85a]. Museums [Hil85, Koh87]. Music [Cur84b, Kle87, LeM82]. Musician [Kle87]. Musterbücher [Tod87]. Myers [Jac86, Mil82c]. Mystery [Had80b]. Mystique [See84]. Myth [Cha88a, Cru87, For88b, Upt88, How86, Rhe82]. Mythic [For88a]. Myths [Gri81a, Nev82].

N [Bla89, Bra83a, Car80, Car89, Dan82, Dow85, Duf80, EE81, Gar81a, Goe82, Gru86, Jer85, Klu82, Lan86, Mit85, O’B81, Per84a, Sha81b, Smi80b]. NACA [Rol83, Lev87b]. Nagpaul [Beb80]. Nail [Coo85]. Nair [Wei81b]. naissance [Sle81]. Nakayama [Kob80]. Nancy [Man83]. Nannie [Kor86]. Napoleonic [Ste81a]. NASA [Zig85, Fri88a]. Nash [Bur80b, Sti85b]. Nasmyth [Tan89b]. Nathan [Bri81b, Bri81c, Buc88a, Dor80, Ell87, Har84, Kul88, Lin88, Pyc82]. Nation [Dob81a, Leh87, Sal83]. National [All84, Beb80, Com86, Fri80b, Ger81, Han80, Law85, Mal86a, Ove85, Ovi86a, Pur82b, Rab81, Sin86b, Sin86a, The86, Tob85, Tro87, Tur89, Wal82, Yao88, Arn89, Doo87, MD82, McM81b, Tom83b, Vol87, War88]. Nationale [Tar85a]. Nationalisation [Tuc80a]. Nationalised [Mil84b]. Nationalized [Str88]. Nations [De 84c]. Natural [Ale85, Bed85b, Das85, Fri81a, Tom85, Wor81]. Nature [Bij86, Bry80, Bry82, Bur82, De 82a, Jer82, Kra82, Pin84, Pla86b]. Naturwissenschaft [Fis82, Kra81a]. Naval [Cas88b, Elb87, Lan84, Pri87, San80, Wis81, Woo82]. navigatietechniek [Ung88]. Navigation [Abb86, War89b]. Navy [Cas88b, Pit86, Mor80a]. Nederland [Ung88]. Need [Maz86a]. Needed [Web85a]. Needham [Bed89c, Kid87, Mul88a]. Needs [Coo88]. Neglected [Lev87b]. Neighborhoods [Cow83]. Neil [Bry81, Cor87, Isr87, Wil84d]. Nelkin [SF82, Lew88a, Mil87]. Nelles [Tay89]. Nellie [Lan86]. Nelson [Hof84, Gra88, Hab82, Lan86]. Neolithic [Spi81]. Neoprene [Smi85c]. Neptune [KP80]. Nerheim [Wor86b]. Nesson [Bla84]. Nest [Lan88]. Networks [Ber85, Wei87a, Rey84]. Neu [Goo82a]. neue [Mul89b]. neuere [vW83]. Neumann [Ste82a]. Neustadt [Nel89]. Never [Mah82]. Nevitt [Fow82]. Newark [Fol80b, Win87, HBR^{+}80]. Newcastle [Fit82].

Newcomen [Tuc82b]. Newell [Kop82, You87b]. Newer [Kie86]. Newhouse [Bil85a]. News [Kie87, Lay87]. Newspaper [Ros81]. Newton [Bri80b, Mah82]. Nicholas [Gar88, Gri81a, Jac88, Lie81b, Rit88, Swe82]. Nicholls [Rea81]. Nick [Bil80a]. Nicks [Com87a]. Niebur [Mol84a]. Niederhuemer [Web87a]. Niemi [Aga82]. Nigel [Buc82c]. Night [Est85]. Nihilism [Dur83a]. Nilo [Kli87b]. Nineteenth [Ber85, Cla86, Dub88, Fri80a, Goo86, Gra84, Har83b, Kar87a, Kem84a, Lan84, Leh86, Mar87b, Mil82c, Nye89, Och89, Pre87, Qui89, Ran82a, Ska88, Sti82, Swe88a, Tan89a, War86a, Wri86, Bed86, War87a]. Nineteenth-Century [Ber85, Cla86, Dub88, Har83b, Mar87b, Wri86, Bed86, War87a]. Nippon [Ino82]. Nirmal [Mal86a]. Nissan [Rae87]. Nisser [Vic84]. NMAH [Wor86a]. No [Qur86, Atk89, Ino82, Lan83a, Ros89, Smi83e, McD85a, Bez83, Boy86, Cro86]. No. [Qur86]. Nobel [Car84a]. Noble [Les86a]. Nolan [Hic85a]. Non [BM85, Maz83]. Non-Euclidean [BM85]. Non-Proliferation [Maz83]. Nondecimal [Asp84]. None [Scr88b, Tra85]. None-Too-Porous [Scr88b]. Noneconomic [HH87b]. Nonstop [TL88]. Nora [dSP81]. Norberg [GR86]. Norbert [Ste82a]. Nordeen [Gra87a]. Norgas [Wor86b]. Norma [Con81b]. Normal [Sim86]. Norman [Ali89, Bul82, Car82a, Gar82, Hew83, Lau86, Ola85, Pal81, Sch80, Smi88a, Gia81]. Norris [Kru85]. North [Ano88i, Bed80b, Gol87, Hil85, Keh84, Mac88a, Mar83b, Mar88, GF87, Gor89b, Vie83a]. Northern [Buc83a, Hul88, Row80b, War89a, Hic85a]. Norton [Scr87a]. Norwegian [Wor86b]. nos [Hin86b]. Note [Bul81c, Pre81, Wer81]. Notes [Ano84b, Ano85b, Ano86b, Ano88b]. Notions [Sin86b]. Nouveau [Lev88]. Novel [Nei88]. November [Ano88i, Mel87, Mor87a, Mor89c, TOE^{+}85, Tuc82b, Wer81, Wor86a]. Noventa [Cum86]. Novo [Kel88]. Nowhere [Wer81]. Nriagu [Zee86]. NS [Fis82, Kra81a]. NS-Ideologie [Fis82, Kra81a]. Nuala [Sti84]. Nuclear [Bro83, Dow84b, Hac86, Hac87, Hac88, Hea88, How85a, Maz84a, Maz86a, Scl88, Wal89, Web80a, dlB88, Boy89, Dow84a, Huf82, Maz81, Maz83]. Nuclearism [How85a]. Nukespeak [Maz84a]. Number [Ano80i, Ano80h, Ano80j, Ano80k, Ano81h, Ano81g, Ano81i, Ano81j, Ano82g, Ano82f, Ano82h, Ano82i, Ano83j, Ano83i, Ano83k, Ano83l, Ano84h, Ano84g, Ano84i, Ano84j, Ano85i, Ano85h, Ano85j, Ano85k, Ano86k, Ano86j, Ano86l, Ano86m, Ano87i, Ano87h, Ano87j, Ano87k, Ano88l, Ano88k, Ano88m, Ano88n, Ano89k, Ano89j, Ano89l, Ano89m]. Numbers [Ros85]. Numeracy [Jaf84]. Numeration [Asp84]. Nursing [Mar85a]. Nutrition [Lan84]. Nuttall [Dav80]. Nye [Per86, Sin87, Wik80]. Nyhart [Bro86b].

O [Cre83, Gra87a, Ovi87, Zee86]. Oaking [Gol84]. Oberkampf [Jer85]. Obert [EE81]. Oberth [EE81, Neu88]. Objects [Bur82, Koh87, Mei88b]. Observation [Zie89a]. Observed [Wil85a]. Obsolescence [Bro85b]. Occupational [Kem86c, Mil87]. Ocean [Cro87a]. O’Connor [Maz84a]. October [Ano88i, Ano89n, Bri81a, DKP^{+}86, DHW^{+}87, FBW^{+}81, GK81, GS81, Hei86, HBR^{+}80, HRP^{+}82, Kel84c, KT80, LHH^{+}84, McC88, Pfe83, Rec88, Smi87b, SB82, TLK^{+}83, Wes87a, Win87, Ano80k, Ano81j, Ano82i, Ano83l, Ano84j, Ano85k, Ano86m, Ano87k, Ano88n, Ano89m]. Off [Vol88]. Office [Cos85]. Oft [Lev87b]. Oft-Neglected [Lev87b]. Ogilvie [Och89]. Ohio [Bog85, Fre89, McD85a, V88, Seg89a]. Oil [Cla89, Coc81, Con89, Gie81a, Gie85, Hyd86, Kra81b, Sch84, Vie83a, Wil83, Pre83a, Wea83]. Okun [Num88]. Olai [Hal86a]. Old [Cas88b, Hud86, Joh83, Lan80b, Nye89, Spi81, You87b, McD89, McI81, Mol84b]. Oldest [GDNCH88]. Ole [Hal86a]. Oleson [Gre88a, Gre81]. Olien [Pre83a, Pre83a]. Oligopoly [Vie83b]. Oliver [Pur82a]. Olson [Col84, Fri81b]. Omaha [Cas86]. Once [Kid87, Mel85]. One [Ovi85, Spu89]. Ontario [Fri82b, Hul88, New85, You87b, Fri82b]. ontwikkeling [Ung88]. Open [Hac85, Str84]. Open-cut [Str84]. Operating [Hil85]. Opinion [Fee85]. Oppenheimer [Kar81]. Options [Ain89]. Oran [Com87a]. Order [De 84c, Fol80a, Hin85, Sta88b, Sti86b, Tuc80b]. Orders [Bac82b]. Ordinary [Lic85]. Ordóñez [Cum86, GD81]. Ordway [Emm81]. Oregon [Cas85]. Ores [HS81]. Organism [Sta85a]. Organization [Coo85, Gor85a, Gra84, HH87b, Maa81, Mue89, New83, Smi82b, Spi82b, Tay89]. Organizational [LM87]. Organized [Wei87a]. Orientalis [Fle84]. Origin [Car89, DD88, MB89, Muh81, Wer81]. Origins [Bac85b, Cos88, Emm83a, Eva88, Gué81, Isr87, Lau87, Lav84, Lew82b, Lie81a, Pow83, Run81, Sin88a, Smi87a, Gut84]. Orman [Goo86]. Oro [Kel84a]. Oscar [Har86a]. Oscilloscope [Car84a]. Osiris [Rei89]. Ostwald [Ser83]. Otfried [Mul89a]. Other [Asp84, Bli82, Eis80b, Mar85b, Ras80b, Roe85a, Wil86a]. Ottfried [Tod87]. Otto [And82, Han87a, Pur84b]. ouest [Mul83c]. Our [Bin89, Flo84, Gro88a, Hab85, Les85]. Outlook [Ove85]. Outram [Dav89c]. ouvrière [Ger84b]. Ovitt [Dal88]. Owen [Bra81b, Har87a, Bau86]. Owens [Ble83, Pow83]. oWhaling [Mar88]. Own [Bin89]. Oxygen [Got83].

P [Beb80, Bec86, Bee85, Bil82, Bil87c, Buc89, Cai82b, Con80b, De 80b, Fin83, Fla89, For88a, Fos82, Hig83b, Hol86b, Jac88, Jan86, Jor89, Kan85, Kap82, Kas84, Kea89, KP80, Les85, Mei84, Mor80d, Rey84, Rhe88, Rol80, Sah87, Seg81, Sta88a, Tei88, Tri83, Tur89, War87a]. Pacey [Cow86]. Pacific [Hic85a, Mar86d, Sto88, Wit83, O’B87, Nas89, Kid87]. Packing [Vie83b]. Packinghouse [O’C88]. Packingtown [Ska88]. Page [Pas85]. Pain [Mar87b]. Pais [For84]. Palace [Bry81]. Palaeolithic [Spi81]. Palaeolithic-Neolithic [Spi81]. Palmer [Ste88c]. Pamela [Ovi87, Pap84]. Panel [Wei86]. Panorama [Leh86]. Paolo [Hal88a, Mor80a]. Paper [Ber89, Ler88, McG85a, Mor80d, Pan87]. Papers [Bed89b, Bri81b, Cha87, Ell87, Far80, Har89a, Ken86, Lev87c, Mus83, Pyc82, Sch82d]. Paradigmatic [She81a]. Parallels [PS82]. Paret [Hea88]. Paris [Lev88, Sta89, Tar85a, But87, Con81b]. Parker [Hac80b, Don81a]. Parliament [Mal86a]. Parrott [Bal85]. Part [Cha87, Pan87, Seg87, Tre85, San81b, Tre85]. Particular [Ste88b]. partir [Kra85]. Partner [Bla89]. Partnership [Mus83, Sha81b]. Partridge [Wei82]. Parts [Buc82a, How80b]. Pascarella [Gen81]. Paskoff [Gor85a]. Paso [Joh87a]. Passenger [Sla89c]. Passes [Fli89b]. Passing [Kra81c]. Past [Hac84, Hil85, Mil83, Pen82, Tho84]. pastoral [Jan87]. pastorala [Jan87]. Patent [Doo87, Fri85, Hew87, Mac87b, Rob85]. Patented [Seg87]. Patents [Fel87]. Paternalist [Cla86]. Paterson [Arg88, Per85b, Wal87b]. Paths [Smi80b]. Patricia [Hoc86, Jaf84, Lan86, Sho88]. Patrick [Lub86b, McG85b, Sha81a, Leh83]. Patron [Tob85]. Pattern [Dub88, Seg80, Beb80]. Patterns [Buc86d, Gut86, Smi85d, Fri84b]. Patterson [Kid85]. Paty [Sch82c]. Paul [Bat80, Bul81b, Bur80a, Che84, Con81c, Dan81a, De 80a, Fer81, Fra84, Fre80, Gie81a, Gor85a, Hac82, Hin81, Hir85, Hol87a, Jas87, Lin81a, Mar85c, Mul88b, Pre83b, Row80b, Ryd89, Seg86, She84, Smi86a, Sta83a, Wal87a, Whe83, Wil85b]. Paula [Gra88]. Pawnee [Keh84]. Pawtucket [Moh87b]. Pay [Iri89b]. Payson [Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89a]. PCC [Fos82]. Peace [San80]. Peacock [Leh83]. Peale [Fri81a, Kas84, Whi83b]. Peasant [Bra80b, Mor87d, Wei81b]. Peebles [Bud82]. Peel [Jer85]. Pegg [Str88]. Pellegrino [Zig88]. Pelterie [EE81]. Pel’tri [EE81]. Pen [How85b, Hok87]. pendant [Gut87]. Penfold [Buc81b]. Penhallurick [Ste88b]. Penicillin [How86]. Peninsula [Mer81b]. Penn [Bez83, Seg83]. Pennsylvania [Con81a, Con87, DHW^{+}87, Gor85a, Kou81, May83, Mor89c, Mul82, Hyd86, Tri83]. Penny [Twe88]. Penury [Gli89]. People [Fol80a, Ger84a, Gie86, Har87c, Hen86, Jaf84, Joh87b, Lic85, Mit85, Pas85, The87, Upt83b, Pia84, Run84, Lat86]. Pepper [Gar88]. Perception [Mel87]. Perceptions [Fol80a, Pet83, Smu87]. Percy [Tri89]. Pérez [Mar84a]. Perfection [Dal88]. Performance [Fli89b, Han87b, Smi83d, Bra81a]. Performance-Control [Fli89b]. period [San81b]. Perkins [Coc82, Gil85b]. Perlin [Kap80]. Permanence [Fla89]. Permian [Wea83]. Pernick [Mar87b]. Pernoud [Pet88]. Perrow [Sim86]. Perry [Gen81, Rol87]. Personalities [Goo82a]. Perspective [LR85, Mor89d, Pur89b, Rui80, And86, Bro84, Kan85, Pen84, Per80b, She87]. Perspectives [Bat82, Bil81a, Cor88a, Hac82, HH87a, HN89, Hir84, Joh87b, Lic85, Mil87, Mor87d, Sta85b, Tei82, Wil84c, Wil85a, Wor86b, Bri81c, McL85]. Pervasive [Gol81]. Pest [Gil85b]. Pete [Tin87]. Peter [Ame89, Buc88b, Fli89a, Hal82, Hea88, Hod89, Kli86b, Mor87c, Mul84a, Sla84, Smi80b, Smi89b, Web84, Woo86, Zuk80]. Petrochemical [Sch84]. Petrochemicals [Smi89b]. Petroleum [Bos85]. Petroski [Kem86b]. Pfafflin [Pap84]. Pfau [Boy86]. Pharmaceutical [Ada89]. Phil [Kar86]. Philadelphia [Dan81b, Hil81, Jac89b, Lub86a, Mil82c, McG88a, TLK^{+}83, Hut88]. Philibert [San89]. Philip [Bai83, Car89, Coa80, Gli89, Gre84, Kea89, Lub86a, Mor81b, Rey87, Tri89, Wal87b, Wes86a, Wik82, Zuk80, Har87a]. Philippe [Bac86]. Philippines [Bra80b]. Phillip [Dun81]. Phillips [Bri80b, Ait81, Bra81b, Elb88]. Philosopher [Bru85]. Philosophical [Dur83a, Goo89, Sta82a, Whi82c, Wil86b]. Philosophie [Lan80a, Rap81]. Philosophy [Fer81, Kas83, Lau82, Mac85b, Mor80c, Ran86b, The85a, Smi86a]. Phoebe [Mar87c]. Phoenix [Gre83]. Photographic [Lan86, Leh86]. Photography [Hic85a, Lan87d, Baa87, Vac84]. Physical [Bed85b, Hoc88, The86]. Physician [Mar85a]. Physicist [Bru85]. Physicists [Sie81]. Physick [Eve86]. Physics [Huf82, Vin82, Wis81, Rab81]. Pianos [Roe85a]. Piasecki [Kro86]. Pickstone [Feu88]. Pickup [TL88]. Pictoral [Sal84]. Pictorial [Wea83]. Picture [Wri86, Mul84b]. Pictures [Wri86]. Pierre [Bel84b, Cha82a, Lan87b, Mar88, Mul83c]. Pierre-Marie-Maurice [Cha82a]. Pieterson [vB83]. Pietro [Gil82]. Piggott [Bul86a]. Pilcher [Tri89]. Pile [Boy85]. Pilot [Sla89c]. Pilots [Hig83b, Mor85]. Pin [Lub87b]. Piñero [Mau82]. Pioneer [Kap82, Ste81b, Tro84a, Bud84]. Pioneering [Cro83]. Pioneers [EE81, Kel89, Lev87a, Smi87b]. Pionery [EE81]. Pipeline [Nev80]. Pisani [Jac86]. Pitcairn [Smi83c]. Pittsburgh [Dob81b, Lau82, DHW^{+}87]. Place [Ber83, Cos85, Wil87a]. Places [Pur89a]. Plains [Soc83]. Plan [She81a]. Planet [Gru86]. planétaire [de 82b]. Planetary [Swe82, Web80b]. planètes [de 82b]. Planisferologio [Bru88]. Planispheric [Ovi86a]. Planned [Seg89b]. Planning [Beb80, Con88b, Fri83, Kir80, tZS82, Vie83a, Sti81]. Plans [Seg89a, Beb80]. Plant [Cas88a, Hil83]. Plastic [Bud84]. Platinum [Mau84]. Platt [Sti85a]. Play [Num88]. Playfair [Mor87e]. Plough [Min81]. Plugs [Sch86b]. Pluviculture [Rot82]. Pnina [Dav89c]. Poets [Sla89b]. Point [Cas88b]. Poisoning [Zee86]. Poisonous [Hol88]. Policies [Don81a]. Policy [Ber80, Bil80a, Bra81a, Che84, Cut87, DR83, Fie84, Har86a, Hir84, Hol87a, Jam89, Ove85, Pur84a, Qur86, Reu89, Sta83b, Tei82, Tei88, Mal86a, Maz83, Pur82b, Rei82, Rot85b, Smi85e, Tei88, McL85]. Political [Fri88a, Gie81b, Hal87, Hod89, How85a, McL84, Nic81, Sch85a]. Politicians [Mil84b]. Politics [Bro83, Coo86b, De 84d, Don81b, Dow84b, Fla89, Fos80, Fri86c, Hea89, Jor89, Lub87a, Meh87, Moh87c, Mor83, Pap82, Pia84, Pre87, Reu85, tZS82, Tom84, Whe84, Hac88, Ali89, Per84b, Bal85, Cut85]. Pollak [SF82]. Pollution [Fee85, Fli89a, Wal89, Per81b]. Polmar [Hew83]. Polyhistor [Hal86a]. Polytechnic [Hob85]. Polytechnique [Kra88a]. Pomona [Hob85]. Pont [Hoc86, Ola85, Smi85c]. Ponting [Jer85, Tan82b, Jer85]. Ponts [Smi86d, Tar85a]. Pool [Fri86a, Kem84b, Shu84b]. Poorly [Dan81b]. Popper [She84]. Poppin [Soc89]. Popular [Cur84b, Fri81a, Hun84, Mel87]. Population [Kru85, McN83]. Porfirian [Bee83]. Porous [Scr88b]. Port [Dan81b, Har89b, Kir80, Kou81, May83, Kon89]. Porter [Coa83, Pur82b]. Portfolio [Sei80, Slo81]. Porticello [Gui89]. Portland [Cas85]. Portrait [Mic84b]. Portraits [Baa87]. Ports [Buc86b]. Portsmouth [Coo84b]. portuaria [Kel84a]. Portuguese [Law87]. Positive [Lin88]. Post [Bol88, Bra81a, Hig80c, Hir86, Joh88, Pas85, Roe89, Smi85e, Lub85, Hok83, Pur84b, Sin88a]. Post-Industrial [Joh88, Roe89]. Post-Reconstruction [Pas85]. Post-Typographic [Bol88]. Post-War [Hig80c, Hir86, Smi85e]. Postan [Mul89c]. Poster [Car86]. Posterity [Law89]. Postglacial [Row80b]. Posthumous [Lev87c]. Postindustrial [Nev82, Sla82]. Postwar [War87b]. Pot [Web87b]. Potash [Mil80b]. Poulle [de 82b]. Poulter [Cla88]. pour [Ger84b, Pet88]. Poverty [Gut85, Lew80]. Power [Bla88, Ble83, Bri81a, Bro83, Car80, Car83b, Con81f, DR83, Fit83, Fro88, Gie81b, Gre83, Gui84, Hal84b, Hal84a, Kid88, Mal86b, Maz81, Maz82, Maz83, Maz86b, Min82, Rey84, Rey89b, Saw88, Sta87, Wal89, Wil87a, Maz82, Bin83, Bri80a, Fit83, Gie81b, Hol86b, Saw88, Spu89, Whe87, Gli89, Maj81]. Powered [Hyd86, Gle84]. Powers [Coo84c, Web80b]. Practical [Pur89b]. Practice [Cha82c, Cha88b, Dur83a, McG85a, Mul81, Pre87, Rhe86, Wei86, Sta88b]. Practices [Maa81]. Practised [Ros87b]. Practitioners [Mor81a]. Praxis [Res81]. Pre [Cas88b, Min82]. Pre-World [Min82]. Precolonial [Lec88]. Precursors [Sha85]. Pred [Blu82]. Prehistoric [Dor83, Lec84, PS82, O’B81, Row80a]. Prehistory [Cro87b, Gor88b, Ran82b, Row86, Wil85b, Spi81]. Prelude [Spi81, Bil87b]. Premier [Pal81]. Prendergast [Lub86b]. Presas [Cum86]. Prescott [Car81]. Present [Kle87, Och89, Rah82, Con88b, Hac84, Han89a, May83, Mil83, Rol89, tZS82]. Presented [Tuc80b]. Preservation [Wal82]. Preserving [Upt83b]. Presidency [Mac87a]. President [Gor87, Hol89]. Presidential [Mac87a]. Press [Bot80, Lew88a]. Pressure [Str85a]. Preston [Cri80]. Prevention [Bul87]. Price [All87, Bed84, Ber85, Cro82]. Pricing [HN89]. Pride [Ska88]. Prime [Wil89a]. Prince [Mor81a]. Prince-Practitioners [Mor81a]. Princeton [Bil81a, Bri81b, Ell87, Pyc82]. Print [Wri86]. Printed [Qui89, Smi83b]. Printers [Jer85]. Printing [Bot80, Min82, Wri86, Pan87]. Printmaking [Wri86, Wri86]. Prints [Hin80b]. Priorities [Pur84a]. Prison [Sha89b]. private [Bra80a]. Prize [Ano80b, Ano80f, Ano81c, Ano81e, Ano82c, Ano82e, Ano83d, Ano83h, Ano84c, Ano84e, Ano85c, Ano85f, Ano86c, Ano86f, Ano86h, Ano87b, Ano87e, Ano88c, Ano88e, Ano88g, Ano89b, Ano89e, Ano89g, Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano86i, Ano87a, Ano87g, Ano88a, Ano88j, Ano89a, Ano89i]. Prizewinner [Car84a]. Probability [Das85]. Problem [Fit81, Mue89, Sim89, Vin86a]. Procedures [Mor81a]. Proceedings [Bri81a, De 84a, Huf82, Mor81b, Ber80, Har86a, Zim84]. Process [Bla88, Eis80b, FRCL83, HH87b, Hir86, Nel87, Zie89b, Mye87]. Processes [See87]. Processing [Sho88, Hym84]. Produced [Hai84, Hol87b]. Production [Bro84, Buc88b, Les86a, Mey86, Mey88, Mul88b, Nuw88, Ros87b, San84]. Productivity [Fel87]. Produites [Bul87]. Prof.ssa [Bed80a]. Profession [Mil86, Kli86c]. Professional [Rey86, Wis80]. Professionalism [Mar87b, Mil84a, Sin86b]. professionalità [Bul81a]. Professionalization [Pri87]. Professions [Och89]. Professor [Jer85]. Program [Ano83f, DR83, Han85, Sch89, Tom83b, Wil85b, Hil86]. Programming [Cer84]. Programs [Rol83]. Progress [Cam89, Cam88, GF87, Hir86, McM81b, McM86, Tod80, Arm86, Cha88a, Kea89]. Progressive [Sti86a]. Progressivism [Seg89a]. Project [Bra81b, Pla86a, Tit89, Wor88, Sha81b, Ste81b, Tre85]. Projections [Sha81c]. Projects [Fri83, Hab85, Kon89]. Proliferation [Maz83]. Prometheus [Rol80, For88a]. Promise [GF87]. Promoter [Bed85b]. Propelled [Als81, Kar83]. Prophecy [Fow82]. Prophets [Sti86b]. Proprietary [Lub86a]. Propulsion [Han85, Per80d]. Proseminar [MD82]. Prospect [Mor81b]. Prospective [Fow80b]. Prospects [Ber84, Mul81]. Protest [McD85a]. Proto [And88]. Proto-Industrialisation [And88]. Proulx [Mar88]. Provenzo [Bol88]. Province [Mar86c, Reu87]. Prude [Hin85]. Pruthi [Beb80]. Pseudo [Ovi85]. Pseudo-Juanelo [Ovi85]. Psychological [How85a]. Psychology [Ber84, Sha84]. Pt [Bed87, Mét86, Mul88a]. Ptolémée [de 82b]. Public [Che84, DR83, Fee85, GF87, Hil81, Hou80b, Hoy85, Kon84, Mar87c, Maz81, Ove85, Rei82, See84, Sti84, Sti85a, Fri88b, Mel83b]. Publications [Man89, Smi83b]. Puddled [Pay85]. puissance [Her88]. Pullman [Bel84a]. Pulos [Whi85]. Pump [Han88, Oer89]. Purdy [Dor83]. Pure [Num88, Tob85]. Purple [Tra89]. Pursell [Hou83]. Pursuit [Gui84, Rei83a, She84]. Pushers [Cla89]. Pyatt [The86]. Pynchon [Sla84, Sla82]. Pytlik [Cut81b].

Qaisar [Mic84a]. Quadrant [McC85]. Quaker [Har82b]. Quantitative [KR80, Kid89, Sin88c]. Quarter [Swe88a]. Quarterly [Haf85a]. Quebec [Hic85b]. Queen [Moh87b]. Queensland [Str84, Str84]. quelques [Kra85]. Quershi [Wei80b]. Quest [Dan87, Gil85b, Tri89, Whe87, Han87b]. Question [Nic81]. Quimby [Sei80, Slo81, Wel87b]. Quincy [Mol84b]. Quinta [Wat80]. quoi [Pet88].

R [All86, Ata82, Beb80, Ber84, Bez85, Bos85, Bot80, Bow80, Bri80a, Bus87, Cam89, Col81, Con81c, Con85b, Con80c, Cos88, Dan81b, Dav89a, Dav80, Dav85, Dob81b, Dow84a, Eis80a, Elb87, Emm81, Emm80a, Fit82, Fit83, Gar81c, Gie81b, Gli88, Gor87, Gor88b, Hac80b, Han85, Han89c, Has86, Hev86, Hok82, Hol87a, Jer85, Joh87a, Kid89, Kop89, Kor86, Lan89a, Leh83, Luk89, Maj81, Mar88, McG84, McI81, Mil80a, Mil83, Min81, Moo86, Mor89b, Mor87e, Nei88, Nel89, O’C88, Pal81, Pay85, Ran82b, Rea81, Rei82, Rei83b, Rot83, SS84, Seg87, Seg89b, SF87, Smi86d, Smi87a, Sti82, Str84, Tod84b, Tro87, Vac81, Vie83a, War87a, Wri83, Zie89a, Zig88, aH80, Beb80]. R&D [Smi87b, Fel87]. R. [Bli82]. Race [McL85]. Rachel [Bij86]. Racine [CL80]. Rad [Bul89]. Radar [All84, Swo89]. Räderuhr [Bed80b]. Radforth [Hul88]. Radiation [Tit89]. Radio [Ait81, All84, Bir87, Hun87b, Wik82, Wis81, Wis84, Pyn84].

Radio-Electronics [Wis81]. Radioactivity [Hac80a]. Rae [Fli89c, Les86b]. Rahman [Beb80, Mal86a, And86, Wei80b]. Rahn [Elb88]. Rail [Kou81, Mar84b, May83]. Railroad [Bez83, Cha89, Con81a, Dub88, Fah89, Gra84, Hic85a, McD86, Sau86, Whi83c, Wil88, Wit83, Fre89, Kle89]. Railroading [Mar83b]. Railroads [Bee83, Con85b, Mar84b, Ste89, Lan85, Dun89]. Rails [McD86, Cud88]. Railway [Cai88, Gra87b, Hof85, Lav84, May89, Spu89, War89a, Whi82a, Whi83a]. Railways [Eva81, San81a, Cal81, Hic85b]. Rainer [Bra82a, Web84]. Rainmakers [Rot82]. Raising [Wil89a]. Raistrick [Har82b]. Raj [Cal81]. raketnoy [EE81]. Raleigh [Ano88i]. Ralph [Leh86, Lin88, Tre85, War89a, Yao88]. Ramo [Flo84]. Ramón [Gli82]. RAND [Zie89a]. Randall [Sto89]. Randolph [Gre88c, Hof80]. Range [Per80c]. Rankine [Cha82c]. Ransom [Jan86]. Rapp [Bra82b, Smi86a]. Rate [Wat81]. Ratings [Bir87]. Raumfahrt [Neu88, Ree81]. Raumfahrt-entwicklung [Neu88]. Ravenscroft [Mac87b]. Ravi [Qur86]. Ray [Cal81, Car84a, DeV85]. Raymond [McD85a, Mer86, Gor88a]. Rays [Zie89b]. RCA [Bro88b]. Re [Fer82]. REA [Whi81]. Reactionary [Meh87]. Reactor [Maz82, Maz86b, McG88c]. Reader [Rey89a, Smi83a, Mol89, Ros81]. Readiness [Swe81]. Readings [Pye84]. Reality [De 82a, Gor88c, Fel84]. Reason [Mar86c, Reu87]. Rebirth [Mar86d]. Receptacles [Sch86b]. recherche [Kra85]. Recht [Web85b]. Rechtlicher [Web85b]. Reckoners [Wil85b]. Reclaiming [Jac83]. Reclamation [Jac81, Jac83, Wor88]. Recollections [Kar81]. Reconciling [She80b]. Reconsidered [Mei88d, Mil81]. Reconstruction [Pas85]. Record [Lan86, Mer81b, Wri85a]. Recorded [LeM82]. Records [Bed89c, Lyt88]. Recreational [Jak87]. Recycling [Gal87]. Red [Gru86, Gre83, Lec88]. Reddy [Ben86]. Redefining [Rot85b]. Redemption [Sha89b]. Redesigning [Mor87c]. Rediscovered [LL82]. Rediscovery [Sin88a]. Redman [Mol84a]. Redmond [Ste81b]. Redondi [Gil82]. Redwood [Rob88]. Reed [Sem84]. Rees [Gol80a]. Reese [Maa81]. Reexhibition [Lan80b]. Reference [Ste88b]. References [Pur83]. Reflections [Sha81c, Sta84, Tro87]. Reflexions [Kid88]. Reform [McS83, Mey85b, Moh87a, Per81b, Mor83]. Reformer [Con88a]. Refrigerator [Wil88]. Refuse [McS83]. Regained [Roe85b]. Regelungen [Web85b]. Regime [Joh83]. Régine [Pet88]. Region [Feu88, Sel88, Ger84a]. Regional [Hit85]. Register [Mye89]. Regulation [Hac87, Tay89, Vie83b, Maz81, Mar80]. Regulatory [Bro86b, Con89]. Reich [Bro87a, Meh87]. Reiches [Fis82, Kra81a]. Reid [Bin87, Hev86]. Reinforced [Bil81a]. Reingold [Bri81b, Buc88a, Ell87, Pyc82, Bri81c]. Reinhard [Bra83b, Web81b]. Reinhold [Cro83]. Reisner [Row89]. rekishi [Ino82]. Related [Bed87, Bil83, Zie89a]. Relation [Jam89]. Relations [Gar81b, Hou80b, Mit85, Pap82, Ran82b, Str84, Bai83]. Relationship [Dub83]. Relationships [Das85]. Relays [Wil85b]. Relevant [Ras80b, Bij86]. Reliable [Bar91, Mol84b, Ste89]. Religion [Das85, Str80, Eng87]. Relph [Han89a]. Remaking [Maz83]. Remarks [Siv80]. Remember [Hoy88]. Renaissance [Mor81a, San89, aH80]. René [Hal82, Hen86]. Renfrew [Tod86]. Rennison [Fit82]. Replication [Sta88b]. Report [Bra81b, Hab85, Hil86, Pla86a]. Representation [GDNCH88]. Representations [Fah89]. Reproduction [McG83]. Reproductive [Rot85b]. Reps [Sti81]. Republic [Buc85, Hin80a, Mit85, Mul83d, Mer86]. République [Day89]. Research [All84, Bec86, Ber80, Bra80b, Bro87a, Com86, Com87b, De 80a, Far80, Han87c, Har86a, Hod81, Maz86a, Mer87, Pel83, Rab81, See84, Sha88, Sin86a, Smi85c, Vey89, Wis80, Wis81, San81b, Bro88b, Fre80, Gia81, Smi87a, Bro86b, Cai82b, De 80a, Fer81, Fre80, Lin81a, Mil82a, The85a]. Reserves [Cen80]. Resettlement [Kru85]. Resistance [Boy85]. Resolution [McG88b]. Resolving [Bro86b]. Resonances [Kid87]. Resource [Gal87, Jak87, Sin88c]. Resources [Gor83, Hab85, Mac88a, Wor81]. Respect [Don81a]. Response [Bar91, BM84b, Emm83b, Gor83, Mic84a]. Responses [DeG87, Wos82]. Responsibilities [Wei82]. Responsibility [Kli86c]. Rest [Mah82]. Restoration [Dal88]. Restraint [Cut85]. Results [Hab85]. Retrospect [Huf82]. Retrospective [Fri86a, Rei80a, TMMY84]. Returns [Bow80]. Reuss [Cal85]. Review [DSF89, McM81b, Reu87, Sha81a, Abb86, Ada89, Aga82, Ain89, Ait81, Ale85, Ale86, Ali89, All86, All84, All87, All88, Ame89, And88, And86, And81, And82, And83, Arg88, Arl84, Arm86, Asp84, Ata82, Atk89, Baa87, Baa89, Bac82a, Bac82b, Bac85a, Bac86, Bah87, Bai83, Bal80, Bal85, Bar88, Bar89a, Bar85, Bat82, Bat80, Bau86, Bay80, Bea81, Beb80, Bec86, Bed80a, Bed80b, Bed83c, Bed83a, Bed83b, Bed85a, Bed85b, Bed86, Bed87, Bed89b, Bed89c, Bee85, Bee83, Bel82, Bel84b, Bel84c, Ben86, Ben88, Ber87, Ber85, Ber89, Ber80, Ber83, Ber84, BM85, Bez83, Bez85, Bha87, Big88, Bij86, Bij88, Bil83, Bil80a, Bil80b, Bil81b, Bil82, Bil85a, Bil85b, Bil87b, Bil87c, Bil88, Bin87, Bin89, Bin83, Bir87, Bir81, Bla81, Bla87, Bla84]. review [Bla88, Bla89, Ble83, Bli82, Blo81, Blu86, Blu82, Bly88, Bog85, Bol88, Bos85, Bot80, Bow87, Bow80, Boy86, Boy88, Boy89, Bra81a, Bra81b, Bra83a, Bra80a, Bra82b, Bra82a, Bra82c, Bra83b, Bra80b, Bra84, Bri80a, Bri89, Bri80b, Bri81a, Bri85, Bri81b, Bri81c, Bro83, Bro85a, Bro87a, Bro88b, Bro86b, Bro80, Bro84, Bro87b, Bro86c, Bru88, Bru85, Bry81, Buc81a, Buc81b, Buc82b, Buc82a, Buc82c, Buc83a, Buc86b, Buc86a, Buc88a, Buc88b, Buc89, Buc86d, Bud82, Bud84, Bul81a, Bul86a, Bul89, Bul81b, Bul82, Bul86b, Bül88, Bur80a, Bur80b, Bur87, Bur83, Bur86, Bus87, Cai82a, Cai82b, Cai88, Cal85, Cal81, Cam89, Cam85, Cam88, Car80, Car81, Car82a, Car82c, Car83a, Car84a, Car89, Car83b, Car84b, Cas88a, Cas85, Cas86, Cas88b, Cen80, Cer84, Cer85, Cha88a]. review [Cha80, Cha82a, Cha82b, Cha84, Cha87, Che84, Chr89, Cla80, Cla86, Cla88, Cla89, Coa80, Coa83, Coc81, Coc82, Coh86, Coh89, Col81, Col84, Com86, Com87a, Com87b, Con80a, Con80b, Con81a, Con81b, Con81c, Con81d, Con82, Con83a, Con85b, Con86b, Con86a, Con87, Con88a, Con88b, Con80c, Con81f, Con85d, Con85c, Coo80, Coo84a, Coo85, Coo86a, Coo89, Coo83, Coo84c, Coo86b, Cor85, Cor87, Cor88a, Cor88b, Cor89, Cos85, Cos87, Cos88, Cou88, Cou86, Cow83, Cow86, Coz89, Cre83, Cri80, Cro82, Cro87a, Cro86, Cro87b, Cro83, Cru87, Cum86, Cum88, Cur84a, Cut81a, Cut81b, Cut85, Cut87, Dal88, Dan83a, Dan85, Dan87, Dan81a, Dan82, Dan83b, Dan80, Dan81b, Das85, Dav89a, Dav89b, Dav80, Dav83, Dav85, Dav89c, Daw89, Day89, De 80a, De 80b, De 81a, De 81b]. review [De 83, De 84a, De 84b, De 84d, De 84c, DeG87, DeV85, DeV88, Dea84, Dea85, Del82, Del83, Dob81a, Dob81b, Don81a, Don81b, Don86a, Dor83, Dor80, Dou84, Dow84a, Dow84b, Dow85, Dre89, Dub83, Dub88, Duf80, Dun80, Dun81, Dun89, Dur83b, Dur83a, Dur85, Eat88, Edg84, Eis80b, Eis80a, Elb87, Elb88, Ell87, Emm81, Emm80a, Emm82, Emm83a, Eng87, Eng80, Est85, Eva88, Eve86, EE81, Fah89, Far80, Fee85, Fel84, Fel87, Fer82, Fer83, Fer88, Fer80a, Fer80b, Fer81, Feu88, Fie84, Fin83, Fis82, Fit82, Fit83, Fla89, Fle84, Fle88, Fli89a, Fli89b, Flo84, Fol82, Fol80a, Fol80b, Fol85, For88a, For84, Fos80, Fos82, Fos89, Fow80b, Fow80a, Fow82, Fre82, Fre89, Fre80, Fri81a, Fri80a, Fri80b, Fri82a, Fri86a, Fri89b, Fri89a, Fri81b, Fri81c, Fri82b]. review [Fri83, Fri84b, Fri86c, Fri88b, Fro87, Fro89, Ful81, Gal87, GD81, Gar88, Gar87b, Gar81a, Gar81b, Gar81c, Gar82, Gei80, Gen81, Ger84a, Ger84b, Ger81, Gia81, Gie81a, Gie85, Gie86, Gie81b, Gil85a, Gil82, Gil85b, Gle84, Gli82, Gli88, Gli89, Goe82, Gol84, Gol87, Gol80a, Gol80b, Gol81, Gol89, Goo85a, Goo86, Goo82a, Goo89, Goo80a, Gor84, Gor85a, Gor86, Gor87, Gor88a, Gor88b, Gor89a, Gor82, Got83, Gra87a, Gra84, Gra87b, Gra88, Gre88a, Gre80, Gre88b, Gre84, Gre88c, Gre83, Gre81, Gri81a, Gri81b, Gro87a, Gro81, Gro88a, Gru86, Gué81, Gui84, Gui89, Gut84, Gut85, Gut86, Gut87, Hab82, Hab85, Hac80a, Hac82, Hac84, Hac85, Hac86, Hac87, HH87a, Hac88, Hac80b, Had80b, Had80a, Had81, Had82, Haf80, Haf85a, Hal86a, Hal88a, Hal80a]. review [Hal80b, Hal81, Hal82, Hal84b, Hal84a, Hal86b, Hal87, Hal89b, Han89a, Han87a, Han85, Han88, Han80, Han87c, Han87b, Han89c, Har82a, Har80b, Har80a, Har82b, Har83b, Har83a, Har84, Har85, Har87c, Har87b, Har87a, Har89a, Har86a, Har81a, Har86b, Har87d, Har89b, Has86, Hat89, Hea88, Hea89, Hei88, Hen86, Hen87, Her88, Hev84, Hev86, Hev89, Hew83, Hey88, Hic85a, Hic85b, Hig80a, Hig80b, Hig80c, Hig82, Hig83a, Hig83b, Hil88, Hil81, Hil83, Hil84, Hil85, Hil86, Hin80a, Hin80b, Hin81, Hin85, Hin86b, Hin86a, Hir84, Hir85, Hir86, Hit85, Hoc85, Hoc87, Hoc88, Hoc86, Hod89, Hof80, Hof84, Hof85, Hok82, Hok83, Hok87, Hol86a, Hol83, Hol86b, Hol87a, Hol88, HS81, HS88, Hol89, Hor84, Hor86, Hor88, Hou80a, Hou83, How80a, How85a, How87, How84]. review [How86, Hud86, Huf82, Hug88a, Hug88b, Hul88, Hun87a, Hun84, Hun87b, Hut88, Huy82, Hyd86, Hyd87, Hyd88a, Hym84, Ino81, Ino82, Isr87, Jac80, Jac81, Jac83, Jac86, Jac87, Jac88, Jac89a, Jaf84, Jak87, Jan86, Jan87, Jas87, Jen84, Jer85, Jer88, Jer82, Joh87a, Joh89, Joh87b, Joh83, Joh88, Kah85, Kam81, Kan85, Kap86, Kap80, Kap82, Kar81, Kar86, Kar87a, Kar87b, Kas83, Kas84, Kea89, Keh84, Kel80, Kel82, Kel84b, Kel84a, Kel88, Kel89, Kem84b, Kem84a, Kem86a, Kem86b, Ken86, Ket85, Kid85, Kid87, Kid88, Kid89, Kie86, Kie89, Kil88, Kin81, Kir80, Kle89, Kle87, Kli86a, Kli86b, Kli86c, Kli87b, Klu82, Kni81, Kob80, Koe84, Koh80, Koh87, Kon84, Kop82, Kop89, Kor80, Kor86, Kou81, Kra81a, Kra85, Kra88b, Kra88a, Kra89a, KP80]. review [Kra86b, Kro86, Kru85, Kul88, LK80, Lan87a, Lan80a, Lan85, Lan87b, Lan87c, Lan89a, Lan84, Lan83a, Lan83b, Lan86, Lan88, Lan89b, Lat86, Lau86, Lau87, Lav84, Law89, Law85, Law83, LeM86, LGWK80, Lec88, Leh83, Leh86, Leh87, Ler88, Les85, Les86b, Les86a, Lev87c, Lev87a, Lev83, Lew88a, Lew80, Lew82a, LB82, Lew82b, Lew85, Lew88b, Lic85, Lie83, Lie81b, Lie81a, Lif85, Lin81a, Lin85, Lin88, Lin81b, Lov82, Lub85, Lub86b, Lub86a, Lub87a, Lue86, Luk89, Lun88, Lyt88, Maa81, Mac85b, Mac88b, Mac85c, Mac86b, Mac85a, Mac86a, Mac87a, Mah82, Maj81, Mal86a, Mal86b, Man81, Man83, Mar89a, Mar84a, Mar86b, Mar86a, Mar87a, Mar85a, Mar87b, Mar87c, Mar84b, Mar89b, Mar80, Mar83b, Mar83a, Mar85b, Mar86c, Mar86d, Mar88, Mau82, Mau84, May83, May89, May85]. review [Maz81, Maz83, Maz84b, Maz84a, McC85, McD85a, McD86, McD89, McD82, McD84, McG86, McG84, McG88b, McG88c, McG85b, McI81, McI83, McI84, McK86, McK80, McL84, McL85, McM81a, McM86, McN83, McS83, Meh87, Mei81, Mei82, Mei84, Mei88a, Mei88b, Mei88c, Mei89, Mel83a, Mel85, Mel83b, Mer87, Mer86, Mer81a, Mer81b, Mer88, Mét86, Met83, Mey85a, Mey85b, Mey86, Mic84a, Mic84b, Mil80a, Mil82a, Mil82b, Mil84b, Mil84a, Mil86, Mil87, Mil88a, Mil82c, Mil83, Mil88b, Mil89, Min81, Mir82, Mis84, Mit85, Moh87a, Moh87c, Mol89, Mol84a, Mol84b, Mol88, Moo86, Mor82, Mor80a, Mor89a, Mor87b, Mor87c, Mor89b, Mor83, Mor80b, Mor80c, Mor80d, Mor81b, Mor87d, Mor87e, Mor85, Mor86, Mor88, Mor89e, Moy80, Mul81, Mul82, Mul84a, Mul83a, Mul83d, Mul83b, Mul83c]. review [Mul84b, Mul87, Mul88a, Mul88b, Mul89c, Mul89a, Mul89b, Mus83, Mye87, Mye89, Nas80, Nas89, Nee83, Nei88, Nel81, Nel82, Nel87, Nel89, Neu88, Nev82, New83, Nic81, Num88, Nye89, O’B81, O’C88, Och87, Och88, Och89, Ola85, Ols89, Ove85, Ovi85, Ovi86a, Ovi87, Ovi88, Pal80, Pal81, Pan87, Pap84, Pap82, Par82, Pas85, Pau83, Pay85, Pel83, Pen82, Pen84, Per81a, Per80a, Per80b, Per85a, Per80d, Per80c, Per84b, Per84a, Per81b, Per86, Pet87, Pet88, Pet89, PS82, Pia84, Pin84, Pin85, Pit86, Pla86b, Pla86a, Pon80, dSP81, Por80, Por89a, Por89b, Pow83, Pre83a, Pre83b, Pre87, de 82b, Pri82, Pur84a, Pur82a, Pur82b, Pur83, Pur84b, Pur89a, Pur89b, Pyc82, Pye84, Pyn84, Qui89, Qur86, Rab81, Rae80a, Rae82, Rae85, Rae87, Rah82, Ran82a, Ran82b, Ran82c, Ran86a]. review [Rap81, Ras80a, Ras80b, Rav81, Rea81, Ree81, Rei80a, Rei80b, Rei82, Rei83b, Rei87, Rei89, Res81, Reu89, Rey84, Rey87, Rey89a, Rey89b, Rhe82, Rhe86, Rhe88, V88, Rig81, Rit88, Rob88, Rob80, Rob85, Roe85a, Roe89, Roe82, Roh84, Rol80, Rol82, Rol83, Rol87, Rol89, Ros81, Ros89, Ros85, Ros87b, Rot82, Rot84, Rot85a, Rot85b, Rot83, Row80b, Row80a, Row82a, Row83, Row86, Row82b, Row89, Rui80, Rum87, Rum88, Run81, Run84, Rup88, Rut89, Ryd89, Ryd80, Sah87, Sal84, Sal83, San84, San81a, San81b, San80, Sar81, Sau80, Sau86, SS84, Saw88, Sch85a, Sch82a, Sch81b, Sch82b, Sch82c, Sch86a, Sch80, Sch85b, Sch81c, Sch82d, Scl88, Scr87a, Scr87b, See87, Seg80, Seg81, Seg83, Seg84, Seg86, Seg87, Seg89a, Seg89b, Sei80, Sel88]. review [Sem84, Ser83, Sha88, Sha84, Sha89a, Sha81b, Sha81c, Sha83a, Sha83b, She84, SPV87, She80a, She81b, She81a, She80b, She87, Sho88, SF82, SF87, Shu84a, Shu84b, Sie81, Sim86, Sin86a, Sin88a, Sin87, Sin88c, Ska88, Sla84, Sla88, Sla89a, Sla89b, Sla89c, Sle81, Slo81, Sma80, Smi80b, Smi80a, Smi81b, Smi81a, Smi82b, Smi82a, Smi83a, Smi83e, Smi83c, Smi83b, Smi84b, Smi85b, Smi85d, Smi85e, Smi85f, Smi85a, Smi86d, Smi86c, Smi86a, Smi87a, Smi88a, Smi88b, Smi89b, Smi89a, Smu85, Smu87, Soc83, Soc89, Sok81, Spi81, Spi82a, Spi82b, Spr81, Spu89, Sta82b, Sta83a, Sta87, Sta88a, Sta88b, Sta81, Sta83c, Sta83b, Sta85a, Sta85b, Ste88a, Ste88b, Ste88c, Ste83, Ste85, Ste80, Ste81b, Ste82a, Sti81, Sti82, Sti84, Sti85a, Sti85b, Sti86a, Sti86b, Sti87, Sto89, Sto88]. review [Str84, Str88, Str80, tZS82, Swe81, Swe82, Swe88b, Swe88a, Swo89, Tan81, Tan82a, Tan82b, Tan89a, Tan89b, Tar81, Tay81, Tay89, Tei82, Tei88, Tha86, The83, The84a, The84b, The85b, The85a, The86, The87, Thi80, Tho84, Tin87, Tit89, Tob85, Tod80, Tod84a, Tod84b, Tod85, Tod86, Tod87, Tom83a, Tom84, Tom85, Top80, Tra85, Tra89, Tre87, Tre85, Tri83, Tri89, Tro84a, Tro87, Tro84b, Tuc80b, Tuc80a, Tuc82a, Tul85, Tur89, Ung83, Ung88, Upt83a, Upt83b, Upt88, Uss88, Vac81, Vac82, Vac84, Vey89, Vic84, Vie83b, Vie83a, Vig83, Vin86b, Vol88, Wal83, Wal87a, Wal82, Wal87b, Wal81, Wal88, Wan89, War89a, War82, War83, War84, War86a, War87a, War89b, War86b, Wat80, Way89, Wea83, Web80a, Web80b, Web85a, Web81a, Web81b, Web81c, Web84, Web87a, Web87b, Wei80a, Wei82]. review [Wei86, Wei87a, Wei80b, Wei81a, Wei81b, Wei87b, Wel87a, Wel87b, Wel89, Wer82, Wes84, Wes83, Wes86a, Wes87c, Wes87b, Wes86b, Whe83, Whe84, Whe87, Whi81, Whi82b, Whi82c, Whi83a, Whi84a, Whi84b, Whi85, Whi86b, Wik80, Wik81, Wik82, Wil87a, Wil88, Wil89a, Wil84a, Wil83, Wil84c, Wil84d, Wil84b, Wil85b, Wil85a, Wil87b, Wil87c, Wil89b, Wil80, Wil81, Wil86b, Win89, Win82, Wis81, Wis84, Wit83, Wol83, Woo82, Woo86, Wor88, Wor81, Wor86b, Wre89, Wri80, Wri83, Wri85a, Wri86, Yao88, Yat81, Yat82, Yor88, You87b, Zee82, Zee86, Zie89a, Zig85, Zig88, Zim84, Zuk80, Zup88, aH80, dB80, dM85, dlB88, vB83, vB86, vS80, vdL83, vW83, Hig80c, Sha88]. Reviews [Ano86d, Fin89]. Revising [Sta88c]. Revisited [Coo80, How80b]. Revival [Sin88a]. Revolt [Kli86c, Mei88d]. revoluties [vB83]. Revolution [Bel84b, Cos88, Cow83, De 84d, Gor84, Hit85, Jac80, Jan86, Jer88, Joh83, Lev87c, Lif85, Mir82, Per84b, Pre83b, Pre87, Rob85, Seg87, Sta83a, tZS82, Uss88, War88, Gut87, Bry81, And88, Bau86, Bry81, Gut84, Har85, Hev86, Jer82, Lew82b, Pur82a, Sch86a, Wei81b, Woo86, Bur87, Hev84, May85]. Revolution/Posthumous [Lev87c]. Revolutionary [Cor87]. Revolutions [McD89, vB83]. Revolvers [Smi84b, How80a]. Reweaving [Tha86]. Reyner [Big88]. Reynolds [Kor86, Cla86, Con86b, Con85c, McM86, Pla86a, Wil89b]. Rhode [Moh87b, Moh88]. Rhodes [Wan89]. Rice [Bra80b, Far80, McL84, Rut89, Tin87]. Rich [Kul88]. Richard [Ale86, Bec86, Bel84c, Ber84, Bil87c, Boy86, Cer84, Col84, Cro82, Cru87, DeV85, Don81b, Gal87, Goo86, Hac88, Hig83b, Hol86b, How85a, Lew80, Mac86b, Mah82, Maz84a, Nel89, Pal81, Rol80, Tei88, Wan89, War86a, Whi82b, vS80, Bed80b, de 81c, Con81c]. Richards [May89]. Richardson [Kas84, Sar81]. Richter [Per84a, Fis82, Kra81a]. Rickover [Hew83]. Ridgway [Gui89]. Rifkin [Gri81b]. Rifle [Tre87]. Rifles [Joh87a]. Righini [Bed82, Bed80a]. Righini-Bonelli [Bed80a]. Rima [Och89]. Ring [How86]. Rings [Rot85a]. Rink [Smi83b]. Rise [Bat82, Ben86, Coo85, Cro82, Dre89, Gre83, Hab82, Hok80, Lan89a, Mic84b, Mul83b, Ols89, Pas85, Pur84b, Sch85b, Smi89b, Ste80, Sti86b, Wil84a, Bil87c]. Risk [Mil87, Sim86, De 83]. Risky [Maz86b]. Ristow [Kar87a]. Rita [Dow84a, Rot85b]. Rittberger [De 84c]. Ritterbush [Mor81b]. River [Rey87, Wor88, Ada84, Sma80]. Rivers [Abb86, Ste88c, Hun87a, Swe81]. Riveting [Vin84]. Road [Gar81c, Gor84, Gra87b, Kle89, Leh87, Rae80a]. Roads [Eva81, Upt88, See84]. Roadscape [Hug88b]. Roanoke [Pur80, Siv80]. Robbins [Sch85a]. Robert [Ale85, Baa89, Bah87, Bee85, Bog85, Bra81b, Bri81a, Con83a, De 80b, De 84d, Fer88, Fri86c, Gre80, Gre84, Gro81, Hil83, Hoc88, How85a, Isr87, KP80, Mel83a, Mic84b, Mol88, Mye89, Ols89, Pel83, Por89a, Pri82, Pur84b, Rav81, Rhe86, Roe82, Rot85b, Row83, Rum87, Sch81b, Sch86a, Shu84a, Sin88a, Sin88c, Smi82b, Smi85a, Soc89, Sta83c, Sta83b, Ste80, Tur89, Uss88, Web80a, Web80b, Whi84a, Gie86, Rae84a, Baa87, Con80b, Har87a, Kar81]. Roberts [Gor86, Rhe86]. Robertson [Lav84, Nei88]. Robin [Hol83, Whi84b]. Robinson [Ano80f, Ano81e, Ano82e, Ano83h, Ano84e, Ano85f, Ano86f, Ano87e, Ano88g, Ano89g, Arl84, Har87b, Jac81, Mel83b]. Rochester [Bil80a, Bez85]. Rock [Jac80, Lan83c, Hin80a]. Rockefeller [Bla89]. Rocket [Bil87b, Emm81, EE81]. Rocketry [Han89c]. Rockets [Swe81]. Rockland [Bel82]. Rocky [HS81]. Roe [HH87a]. Roegen [Gri81a]. Roger [Ber85, Cor85, Eat88, Huf82, Kra88b, Pre83a, Sau86, Smi85d, Ste88b, Vey89]. Rogers [Mac85a]. Rohm [Hei88]. Roland [Com86, Joh87b]. Role [Beb80, De 80b, De 81a, Haf85b, Kem86b, Lay88, Mal86a, McI83, Pap84, Tod80, Wis80, Zie89a]. Rolf [Web87a]. Roller [DD88]. Rollie [Lic85]. Rolph [Maz81]. Rolt [Emm80b]. Roman [Cou88, Gre88a, Har86b, Har89b]. Romana [Har87d]. Romance [Lew85, McM81b]. Romantic [Smi85b]. Ron [Coa80]. Ronald [Cam88, Coo86a, Coo83, Ket85, Ros85, Tul85, Woo86]. Ronan [Tom83a]. Rönn [San81a]. Ronnberg [McG84]. Roos [Mac86b]. Roots [Seg86, Fol80b]. Rorabaugh [Yor88]. Rory [Maz84a]. Rosalind [Har83b]. Rose [Ano86i, Ano87g, Ano88j, Ano89i, Hac86, Ros89, Fos80, Hir84]. Rosegger [Bro84]. Rosemary [Maa81]. Rosen [Bla88, Wik82]. Rosenberg [Dor80, Kul88, Har84, Lin88]. Rosenzweig [Kra86b]. Roslynn [Bea81]. Rosner [Mel85]. Ross [Lub87a]. Rossiter [Pur83, War84]. Rota [Bra83a]. Roth [Coo84a, Con80a, Con86a]. Rothenberg [Buc88a, Con85d, Eis80b].

Rothschild [Haf85a, Pur89b, Sta85b]. Roubaix [Ger84b]. Rougé [Wil84b]. Roughnecks [Cla89]. Route [Sau86]. Rowhouse [Whi83b]. Rowland [Tro87]. Rows [Koh80]. Roy [Ber84, Chr89, Dan82, Kra86b, Sau80, dB80, Cut87]. Royal [De 81a, Mor89d, Pit86, Tra89, Fre82]. Royden [Tan89a]. Roys [Kap82]. Rubber [Bul81c, Per85a, Tut81, Roe82]. Ruddock [Dan80]. Rudi [Tom84]. Rudolf [Bra82c]. Ruhr [Bra80a, Web81c]. Ruhrgebiets [Tod84a]. Ruiz [Sho88]. Run [Bez83]. Rundell [Wea83]. Runyan [Elb87]. Rupert [Car82a, Mac85c, Sch80]. Rural [Ben88, Ber85, Dan85, Fri83, Hin85, Mul82, Whi81, Wil80, Win82]. Rürup [Bra83b, Web81b]. Rush [Jac88]. Ruskin [Mei88a]. Russell [Bus87, Emm80b, Gie81a, Rit88]. Russia [Coo86b, Wes87c, Esp82, Sch84, Bla87]. Russian [Kam81]. Rust [Cas88a]. Rustum [Cut87]. Ruth [Cro83, Hac85, Mil83, Rup88, Wer82]. Rutherford [Cam85]. Ryder [Arm86].

S [Ale85, Beb80, Bed83b, Bil81b, Bro87a, Buc82b, Con85c, De 84d, Gei80, Gre88c, Gui89, Hac80b, Hal84b, Jac83, Kel84b, Lan84, Maa81, Mah82, Mar87b, May85, Maz81, McM86, Mey85a, Pla86a, Pur82a, Rae84a, Rav81, Run81, Smi81a, Smi85a, Spi82b, Tan82b, Tar81, Tei82, TL88, Wat80, Wil89b, Zig85]. S.J [Col81]. S.J. [Tha86]. S.S [Buc82b]. Sacramento [Gui83]. Sacrifice [Boy86]. Sadi [Gil82, Kid88]. Safety [Maz81, Maz82, Tit89, Whi82a]. Saga [Bra84, Tro87]. Sahal [Fri84b, Mil82a]. Sailcar [Kar83]. Sailing [Wri83]. Sailors [Lan84]. Salholm [San81a]. Saliba [Ovi86a]. Salines [Mul83c]. Salinity [Hab85]. Salmon [O’B87]. Salsbury [Bez83]. Salt [Cla86, Mul88b, Wor88, Mul88b]. Saltman [Kil88]. Salvadori [Con81d]. Sam [Mar86c]. Sample [Wil89b]. Samuel [Bed85b, Dav83, Fla89, Hac87, McD82, Mor89a, Rob80, Ano86i, Ano87g, Ano88j, Ano89i, Lau87]. Samuels [Whi84a]. San [Joh87a, Lev83]. Sánchez [Gli82]. Sand [Roh84]. Sand-Glasses [Bed87]. Sanders [Yao88]. Sands [Roh84]. Santa [McD86, Sho88]. Santíssimo [Gui83]. Sara [Mil83]. Sarah [Bed83b]. Sarjeant [Run81]. Sass [Lan80a]. Satellite [Zie89a]. Satow [Cal81]. Saul [Gei80, Gei80]. Sault [Klu82, Pla86a]. Savage [Bha87, Cum88, Hil88]. Savage-Smith [Hil88]. savants [Day89]. Saward [DeV88]. Sawdust [Smi85a]. Sawmills [Nei88]. Sawpower [Nei88]. Saws [Sim83, Sim85]. Scale [Mue89]. Scaling [Smi86b]. Scandal [Emm80a]. Scandinavia [And88]. Scandinavian [Mor89d]. Scapegoats [McD82]. Scarpino [Rey87]. Scene [Con85b]. Schaffer [Han88]. Schaffhausen [Mis87]. Schallenberg [Bel84c, de 81c]. Schatz [Coo86a]. Schaub [Mil84a, The84a]. Schell [Del83]. Schemnitzer [Web81a]. Schepelern [Hal86a]. Scheutz [LL82, Bül88]. Schivelbusch [Whi83a]. Schlereth [How84, Hug88b, Mer87, Tho84]. Schmiechen [Ran86a]. Schmitt [Kap82]. Schneider [Fos82]. Scholl [Bra80a]. School [Cas88b, Bra81b]. Schooling [Scr86a]. Schoolman [Dur83b]. Schools [Mor87a]. Schrier [Kam81]. Schubert [Goe82]. Schutz [Row86]. Schuurman [Whi82c]. Schuyler [Wos82]. Schwartz [Hac85]. Science [Ada89, Ain89, All86, Ano80l, Bea81, Bro87a, Bro86b, Bro87b, Bul81b, Bul86b, Bur82, Cha82c, Cha88b, Col84, Con83b, Con85d, Con89, Cut87, Dan87, DR83, Das85, Dav89c, De 84c, Eis80b, For88b, For84, Fri81a, Fri80a, Fri81c, Gli89, Goo80a, Gor85b, Hab85, Hac82, Haf85b, Hat89, Hir86, Hoy88, Hug81, Jam89, Jon81, Kid85, Kli86a, Kli87b, Kli87a, Koe84, Lan85, Lat86, Lau82, Lay88, LS84, Les85, Lew88a, LB82, Lyt88, Mac85c, Mac86b, Mac87a, Mal86a, McG88b, Mil88a, Mor81a, Mor87e, Mul81, Mul83d, Och89, Ove85, Ovi88, Pal81, Per80a, Pia84, Por89a, Pur83, Rei83a, Rei80b, Rei83b, Sin88b, Sla89b, Smi82b, Smi86b, Tei88, Tob85, Tom83a, Tuc80b, Tuc82b, Vin84, War87b, Wes89, Whe84, Tei88, Tuc80b, Bur87, Cut85, Gor84, Hac80a, Kop82, Mac85b, Mar84a]. Science [Mar86c, Mic84b, Pin84, Pye84, Tom85, Wil85a, And86, Bij88, Bro83, Col84, De 80b, De 81a, De 84d, De 84c, Gre80, Gre88c, Hac84, Hol89, Kar86, Mac85c, Mét86, Mor81b, Mul88a, Nel87, Ovi87, Pal81, Pan87, Pel83, Pur84a, Pur82b, Pye84, Rhe86, Tei82, Wei80b, Wil85a]. Sciences [Bri81c, FW87, Lay88, Ove85, Pal81, Pin85, Tom85, Tom85, Bed85b, Hac82, Hoc88, Hug88a, Smi81b]. Scientia [Mer81a]. Scientiae [Sch82c]. Scientific [Bed85a, Bed86, Bij86, Con83b, Dav85, For88a, Goo85a, Hab82, Hou80b, Jer82, Kob80, Lau82, Lif85, Mar80, Rav81, Rot85a, See84, Sie81, Sta88b, War87a, Zie89b, FLC89, Buc88a, McG88b, Pap84]. Scientific-Technical [Lif85]. Scientific-Technological [Pap84]. Scientist [Cre83, Woo82, Mor82]. Scientists [Bha87, Bij88, Hod89, Nas89, Och89, Rei82, War84, Wis80, Bri80a]. Scienza [Bed80a]. Scotia [Nei88]. Scotland [Mal86b, Gle84]. Scott [Bar89a, Emm80b, Wat80]. Scottish [Gle84, Lav84]. Scranton [Lub86a, Wal87b]. scuola [Bul81a]. Sea [Bro86b, Gar87a, Kir80, Maj81, Vie83a, Bul86a]. Seaborg [Wal83]. Seafaring [Lan84]. Seal [Dav85]. Sealander [Seg89a]. Sean [Sch85b]. Seaport [Jac88]. Search [Bur82, Coh86, LK80, McD82, McG88c, Nel81, Per81a, Tre87, Vac82]. Searching [Com87b]. Searing [Och89]. Seaweed [Neu89]. Second [Kap86, Gor82]. Secret [Pur89a, Rig81]. Secretariat [Spi82a]. Security [Yao88]. Seek [Pit86]. Seely [Reu89]. Segal [Rhe88]. Segregation [Rup88]. Seidel [Hoc88]. Select [Gre88a]. Selected [Bul86b, Bur82, Edg84, Far80, Mus83, Ove85, EE81]. Selections [Hor88]. Selective [Cha87, Con85d]. Self [Als81, Kap86, Kar83, Per86, Smi84b]. Self-Loaders [Smi84b]. Self-Propelled [Als81, Kar83]. Sellers [Fri81a]. Selling [Bud84, Dan83a, Maz84a, Lew88a]. Selmer [Blo81]. selon [de 82b]. Seltzer [Don86a]. Semiconductor [Hev84]. SENCSA [Lew89]. September [Buc85, DR83]. Serfdom [Esp82]. Serge [Jer85]. Series [Hou80b]. Servant [Mul87, Whi80]. Service [Bot80, Cla80, Gol87, Hal81, Hig80c, Lev87a, Pyn84, Cla88, Wel89]. Services [Bed85b, Qur86, Shu84a, Sti85a]. servitudes [Her88]. ses [Kra85]. Sessions [Car85]. Sesto [Bro83]. Settlement [Row80b]. Seven [Fer83, Bac85a, Sch81b]. Seventeenth [Con88b, Das85, Elb88, Est85, Fri80a, Joh89, Kle87]. Seventeenth-Century [Das85]. Seventy [Bur82]. Seventy-two [Bur82]. Sewage [Ham88]. Sewell [Joh83]. Sewing [Moh87b]. Sex [Aga82, Rup88]. Sexual [Ben88]. Shākir [Kel80]. Shaiken [Rum88]. Shaker [Spr81]. Shakespeare [Ste87]. Shale [Gie81a]. Shape [Les85]. Shapes [Arn89, Spi82a]. Shapin [Han88]. Shaping [Hoc87, Cal85]. Shapiro [Das85]. Shapley [Cut87]. Sharlin [Moy80, Moy80]. Sharma [Wei80b]. Sharon [Ovi86a]. Sharpe [Emm81]. Sharps [Joh87a]. Sharpshooting [Sha89a]. Shaw [Mal86b]. Shearer [Smu85]. Sheehan [How86]. Sheffield [Twe86, Gor89a]. Sheila [Pap84]. Sheldon [DeG87, Hei88]. Shell [Mor80b]. Shelton [Dub88]. Shemilt [De 84b]. Sherman [Sta81]. Sherry [Fri81b, Kap86]. Sherwood [Run84]. Shield [Lew82a]. Shifting [Mei88c]. Shigeru [Kob80]. Shining [Kle89]. Ship [Bau86, Buc82b, Mer88, Whi82b]. Shipbuilders [Gol87, Pri87]. Shipbuilding [Gol87, vS80]. Ships [Ung81, Wri83, vS80, Gol84, Wil84b]. Shipwreck [Gui89, Oer89]. Shipwright [Pri87]. Sholes [Cas88a]. Shop [Hok82, Pen80, Scr86a]. Shop-Floor [Scr86a]. Shore [Cro87a]. Short [Ful81, Mac85c, Mar83a]. Shortal [Per80c]. Shorter [Tom83a]. SHOT [Rae84b, Sin89, Sta84]. Shreve [Abb86]. Shrum [Wei87a]. Shumaker [Kra88b]. Sicherman [McG86]. Side [Wil86a]. sidérurgique [Gut87]. Sidney [Wil87a]. siècle [de 82b]. siècles [Smi86c, Kra85]. Siege [Gar88, Lan87c, Thi80]. Siegelbaum [Coo86b]. Siemens [Tod80]. Siena [Gar88]. Sienese [Ada84]. Sigaut [Cen80]. Sigfrid [Tod80]. Siglo [Kel84a]. siglos [Mau82]. Sign [McC85]. Signaling [Mue89]. Significance [Col84]. Signs [Sla84]. Sigvard [Cut81b, Har81a]. Silicon [Fri89b]. Silk [Arg88, Hyd87, Per85b, Wal87b, Wal87b]. Sills [Maz84b]. Silvio [Bru88, How87, McC85]. Simeon [Gor89b]. Simmen [Pet87]. Simmons [Cai88]. Simon [Flo84, Gar88, Han88, dSP81, Ste81b]. Simple [Cam85]. Simpson [Mor87e]. since [Buc86d, Fli89a, Gui84, Hal82, Kra88b, McD89, Moo86, Mul84a, Tin87, Whi86b]. Sinclair [Cor88a, Row82b, Smi81a]. Singapore [De 84a]. Singer [Bla87]. Sir [Mor87e, dB80, Rob80]. Siracusa [Fol80a]. Siry [Cro87a]. Sisters [McG86]. Site [Hac86, Moh87b, Pen80]. Sites [Cas86, Con86b, Coo84a, Mer81b, Mil88b, Wri85a]. Six [Gar81c, Elb88]. Sixteenth [Est85, Gar88, Kel82]. Sixteenth-Century [Gar88]. Sixth [Ber80]. Skaggs [Wil89a]. Skempton [Bal80]. Sketch [Mul83d]. Sketchbooks [She81b]. Sketches [Hor88, Bri85]. Skidi [Keh84]. Skies [Hig80a]. Skill [Ran82c]. Skilled [Mey88, Tan89a]. Skills [And88, Gro87b, Nuw88]. Sky [Cro83]. Skylab [Mac86a]. Skywatchers [dM85]. Slaski [Dan81b]. Slater [Lau87, Moh87b, Pen80]. Sloniger [Whi84b]. Sloop [Per80d]. Slotkin [Cru87]. Slotta [Bra82a]. Small [Ber83, Sch86b, Bla81, Fri83]. Smelting [HS81]. Smith [Bry81, Bur82, Isr87, Kar81, Ste81b, Tri83, Bin89, Bur82, Car82a, Gué81, HH87a, Hil88, Lan89b, Lau86, Nas89, Pas85, Sch80, Smi83c, Smi88a, Spu89, Tre85, Wor88].

Smithson [Bir83]. Smithsonian [Fri80b, Hou80a, Hok80, Sch89]. Smoke [Fli80, Fli89a]. Smolenice [Buc83c]. Smothered [Ber87]. So-called [Kel84b]. Sociability [Fis88]. Social [Bay80, Ber89, BBM84, Buc88b, Bul82, Bur80a, Coz89, Cum88, Dan85, Emm83a, Fos89, Gre81, Har87c, Hin86a, Hoc87, Kie89, Kli86c, Law87, Les86a, Lev87c, Lic85, May89, McG86, Moh87c, Nel81, New83, Per84a, Pre83a, Qur86, Rae82, Ran82a, Ras80a, Rot85b, Rui80, Sch82a, Seg89a, Smu85, Tay81, The87, Tuc80b, Upt83a, Wal88, War82, War86a, Web81c, Wil85a, vB83, DeG87]. sociale [Mul83c]. sociales [Cen80]. Socialism [Sti86b]. sociedad [Mau82]. Societal [Tar85b]. société [Fro87]. Societies [Bil87b, Rut89]. Society [Atk89, Bac82b, Ben86, Bro86c, DR83, Eng87, Feu88, Flo84, GR86, Gre88c, Gui84, Hin86b, Hob80, Kra89a, Lan83a, LR85, Lie81a, Mar85c, McD85a, Mei88b, Mul87, Och88, Ovi88, Por89b, Rey84, Row82b, Sch81b, Sha81c, Smu85, Sta82a, Tom84, Tuc82b, Wei86, Wil84d, Wil87c, Ano83e, Ano85d, Ano87c, Ano89d, Cut81b, Fra84, Jor89, Sal83, Str85b, Wei86, Bij88, Bro86b, Cos88, Cut81b, De 83, Goo89, Hab85, Ovi87, dSP81]. Socioeconomic [Dow84a]. Sociological [Buc89]. Sociology [Coz89, Hac80b, Pye84, Pin85]. Sociotechnological [Cur84b]. Software [Hen86]. Soil [Sto89]. Sokal [Sha84]. Sokol’skiy [EE81]. Sola [Fri86a, Shu84b]. Solar [Kap80, Kap81, Str85b]. Solidarity [Pre87]. Solla [Bed84]. Solo [Roe82]. Solomon [Bil81a]. Solutions [Vac82]. Solyom [Smi80a, Smi80a]. Some [Tra85]. Someone [Fis88]. Somerville [Kid85]. Something [Lan80b]. Sonar [Pit86]. Song [Mul89b]. Song-Dynastie [Mul89b]. Sons [Kel80, Cum88]. Sorböm [Rae82]. Sorby [Nut81]. Sound [McG83, Bra84]. Source [Sta88a, aH80, Whe84]. Sourcebook [Mor80d]. Sources [HH87b, Hol83, Lev87b, Per80a, Dun81, Smi83b]. South [Dan83a, Gra88, McD89, McG85b, Mol88, Sti85a, Tod86, Tul85]. Southeast [McL84]. Southern [Bro86c, Cri80, Dan83a, GK81, GS81, Gol84, Gra87b, Hic85b, Jac86, Kel84c, KT80, McD89, Pfe83, Ros87b, SB82, Sto88]. Southwestern [Wel89]. Soviet [Bai83, Bal85, Cam88, Coo83, Gre83, Lie81b, Lie81a, McI83, Swe82, Pap82]. Soviet-Western [Bai83]. Soyuz [Sha81b]. Space [Bil87b, Hac82, Hac84, Hal87, Han85, Hir85, Hol87a, Kop82, MD82, Mac86a, Mac89, McD85b, McM81b, Nee83, Swe82, The85b, Web85a, Wos84, Zie89a, Fri88a, Web85a, Hac82]. Space-Age [McD85b]. Spacecraft [Hal80a]. Spacefaring [Joh87b]. Spaceflight [Joh87b, Web80b]. Spain [Elb88, Gli86, Mye89, Rui80, Gli89]. Spanier [Tra89]. Spanish [Joh87a]. spätere [Neu88]. speaking [Bed80b]. special [Smi83e]. Specialty [Twe86]. Speech [Shu84b]. Speed [Luk86, Rol83]. Spence [Rot82]. Spencer [Bri80a]. Spending [Hev89, Wil87c]. Sperry [Hal80b]. Spiegel [Mul89b]. Spinners [Ben88]. Spirit [Har83a, Kap86]. Spirn [Pla86b]. Spitz [Smi89b]. Sponsored [Har81b]. Sporting [Hal80a]. Sporty [Bil85a]. Sprachraum [Bed80b]. Spread [Fos89, Jaf84]. Spring [Jon81]. Sputnik [Hir85]. Spy [Sta88a]. Squires [Mer88]. SRI [Bir81]. SST [McD84]. St [Bar88, Bil88, Eis80a, Fra84, Hal81, Mar85c, She81a, Sma80]. Staatliche [Bra80a]. Stabile [Sti86b]. Staff [Bra81a, Wit83]. Stahlindustrie [Tod84a]. Stairway [Sha89b]. Stalin [Lie81a]. Stan [Hen87]. Stand [Con81d]. Standard [Bar89b]. Stanier [Bly88]. Stanley [Bur82, Jac83, Jer85, Smi88b, Fer80a]. Stansky [Mor87c]. Stapleton [Kra88b, McM81a, Win89]. Stapp [Baa87]. Star [Mil88b]. Starch [Pec86]. Stares [Hol87a]. Starr [Mor81b]. Stars [Keh84]. Starving [Lan84]. State [Cud88, Gia81, Hob85, How85a, Seg83, Whi83c, Wil83, Wos82, Lie81b, Wei87b]. Statecraft [Hat89]. States [Bil81a, Blu82, Bro83, Cal85, Con81f, Con85d, Fla89, Hal84b, Hol83, Mar83b, Mel83b, Rot84, Saw88, Smi85d, Wil89a, Mil80b, TMMY84, Hir86, Kas83, Mar84b, Mey86, Nel82, O’B81, Sla89c, Wal81, War83, Wei86, Con85a]. Statesmen [Bed83b]. Station [Bro80, May83, May89, Fri88a, Con87, Kou81]. Stations [Cha89, Dan87]. Statistical [Gro87a]. Statistics [Mac88b]. Staudenmaier [Tha86]. Ste [Klu82, Pla86a]. Steam [Car83b, Cas88b, Con81f, Cor88b, Fit83, KR80, Mar83a, Mil82b, Min81, Saw88, Sta87, Whi86a, San81b, Duf80, Car80]. Steamboating [Bra84]. Stearns [Smi80b]. Steel [Cas88a, Gor89a, How85b, Lew82a, McD86, Nuw88, San84, Twe86, War86b, Lew82a, Yat81]. Steele [Gol84, Joh89, dlB88]. Steelmaking [Got83, War86b]. Steelmasters [Mey85b]. Steensberg [PS82, Spi82a]. Steffen [Fis82, Kra81a, Pre87]. Steffens [Car81]. Steiner [Mar87a]. Steinman [Sla89b]. Steinmetz [Jor89]. Steinsalzindustrie [Bra82a]. Stentors [Wik82]. Stephen [Bez83, Fos82, Goo89, Kar87b, Mac88b, Maz84a, Sch82a, Smi83b, tZS82, The85b, Vac81, Wei87b, Wil80]. Stephenson [Buc86a]. Stereoscopic [Wes84]. Stereotomy [San89]. Sterling [Kie86]. Stern [Man83]. Steve [Dow84a, Ste82a]. Steven [Bro83, Han88, Jer88, Kas83, Mul83d]. Stevens [Jac89a, Smu87]. Stewart [Cut81a, Sch81b]. Stigler [Mac88b]. Stilgoe [Con85b]. Stobaugh [Bra81b]. Stockbridge [Lan86]. Stockholm [Mor89d]. Stocking [Koh87]. Stockton [Jac88]. Stockyards [Ska88]. Stoff [Zig88]. Stokes [Wis84]. Stokowski [McG83]. Stone [Dor83, Eva88, Fri89b, Mor87d]. Stonehenge [Row82a]. Storage [Sch81a, Bel84c, Dow84a]. Stored [Tom83b, Wil85b]. Stored-Program [Tom83b]. Storia [Bed80a]. Story [Bez83, Bil85b, Cla88, Emm80a, Ger84a, Gro88a, Hac86, How86, Kah85, Kle89, Min81, Pur89a, Rol87, Tre85, vdL83, Kam81, Dav89a, Daw89, Smi83c, Whi84b]. Storytellers [Tha86]. Stover [Fre89, Law85]. Strandh [Cut81b, Har81a]. Strategic [Mul84a, Tei88]. Strategies [Fri82b, Kro86, War84, Coa83, Gil85b]. Strategy [Hea88, Isr87, Lin88, Yao88, Mor86]. Stratemeyer [Hor84]. Strauss [Boy86]. Streamliners [Eis80a]. Streamlining [Hyd88b, Hyd88a]. Street [Iri89b, Mul85]. Streets [Whi86a]. Strength [Con81d, McD85a]. Strike [Arg88, Pit86]. Strikes [Gar82]. Strindberg [Jan87]. Strohmeyer [Cas88a]. Stromquist [Dub88]. Stronger [Con85c]. Stross [Sto89]. Structural [MB89, Mei84]. Structure [BM84a, Bur82, For88a, Gor85a, Isr87, LM87, Och87, Con83a]. Structures [Ada84, Ber85, Con86b, Kem84a, Wri85a]. Struggle [Bin89]. Struggles [Ger84a, War84]. Struik [Mul83d]. Stuart [Bil80a, Bla81, Bry81, Bul86a, Car82c, Hev84, Por89b, Rae85]. Stubborn [Sto89]. studier [San81b]. Studies [Bla87, Cou86, Cum88, Haf85a, Hob85, Hoc88, How84, McG88b, Mul83b, Nel87, See87, Seg86, Wal87b, Had81, Had82, Por80, Wil89b, Eng80, Sch82d]. Studley [Coo84a]. Study [Buc89, Coc81, Con80c, Coo85, Coo86b, Das85, Fri82b, Gro87b, Hev89, Jer85, Mac85b, Maz81, McN83, Ovi88, Pal81, PS82, She81a, Smi85e, Str84, Tan89b, Tuc80a, Twe86, Wes84, Wil85a, San81b]. Stuewer [Huf82]. Style [Lev88, Gol89]. Styles [Spi82b]. Subject [Maz86a]. Submarine [Pit86, Rei80a, Rol87]. Subordination [HH87b]. Subtle [For84]. Suburbia [Ols89]. Subways [Vig83]. Successful [Kem86b]. Successor [Hok87]. Suction [Oer89]. Sudan [Mul88b]. Sudnik [Hoc86]. Suellen [Kon84, Mel83b]. Sueno [Ino81]. Suffering [Mar87b]. Sugar [Mar89a, Wil87a]. Sugarcane [DD88]. Suggestion [Bir83]. Suicide [Fer80b]. Sum [Lin88]. Summary [Muh81, Pur80]. Summers [Dan85]. Summit [Smi88b]. Sun [Coc81, Gie85, Hac86, Mul88b, Qur86, Whe87]. Sunbelt [Swe81]. Superior [Lan83c]. Superpower [Pap82, Cor88b]. Supplement [Hol83]. Supplies [De 84b]. Supply [Ble83, Cla88, Mil84b, Tuc80a, Bla84]. Survey [Con85a, Dan83b, Wri85a, Jer85]. Surveying [War89b]. Surveyors [Bed83b]. Survival [Bry80, Fer80a, Fer80b, Per81a]. Survive [Cas88a]. Surviving [Kid88]. Survivor [Whi84a]. survivre [Ger84b]. Susan [Fah89, Fol80b, Fri89a, Mil82c, Moh87a, Och89]. Susantha [Bro87b]. Susskind [Car84a]. Sutcliffe [Blu86]. Svante [Dan83b, Sta87]. Svenska [San81a]. Sweat [Tan82a]. Sweated [Ran86a, Ran86a]. Sweden [GR86, Qur86, San81a, Sta87]. Swedish [Lun88, San84, San81a]. Sweetness [Wil87a]. Swenarton [Mor87b]. Swetz [Coh89]. Swift(R) [Hor84]. Switchboard [Mue89]. Switching [Kli86a, Mue89]. Switzerland [Mis87]. Sydenham [Fin83]. Syllabi [Wil89b]. Symbol [BM84a, Bry81, Upt88, Cre83]. Symbolic [Mei89]. Symbols [Con86b]. Symposium [Bri81a, Buc83c, Buc85, FW87, Hin84, Huf82, Mor89c, Mor89d, Sch82d, Sch89, Wer82, Wil84c, Zim84, Mor87e, Web85a, Mor81b, Web85a, Wil81]. Symptoms [Sla84]. Synthetic [Roe82, Tut81]. System [Coo84b, Doo87, Fri88a, Hod81, Hok80, Isr87, Kli86a, Kou81, Lav84, May83, Mey86, Pur84b, Reu89, Rob85, San81a, Del82, Fli89b]. systematische [Web87b]. Systems [Blu82, Bry80, Coz89, Jas87, Lec84, Ste80, Wal81, Zie89a, San81a, Wei87a]. Szasz [Hac86]. Széchenyi [Bro88a].

T [Aga82, Bil85b, Bul81b, Don81a, Emm80b, EE81, Fer83, Fer81, Hac87, Hev86, Hok87, Hol89, Hor84, Jer85, Kem84a, Kli86c, Lan86, Mil88b, Ovi88, Pap82, Smi86a, The85a, Wal83, Wes83, Wes87b, Whe87, Wik82, Wil84a].

Tail [Tit89]. Takeoff [Bil80a]. Takes [Vol88]. Taking [Pos82]. Talbert [Chr89]. Talbot [Vac84]. talet [San81a]. Talking [Ste83]. Tall [Gol84]. Tamara [Dub83, Hof80]. Tancook [Sin88a]. Tang [Pet89]. Tann [Mus83]. Tarr [Nas80, Ros89]. Tartakow [Ano86i, Ano87g, Ano88j, Ano89i]. Tatsuya [Ino82]. Tavern [Dav87]. Tawney [Bli82]. Taxicab [Whi84a]. Taylerson [How80a]. Taylor [Hab82, Hoc86, Hig83a]. Teaching [Dan83b, Mor87e, Tho84, Pur89b]. Team [Emm81]. Teatro [Kel88]. Tech [Mol88, Tei88]. Technical [Buc88b, Eva81, FW87, GK81, GS81, Kel84c, KT80, Kra81b, Law87, Lev87b, Lie81b, Lie81a, Lif85, Mac85b, Pfe83, Por80, Smi83b, SB82, Sta88a, Wei87a, Smi83b, Cam88]. Technics [Dur85, Res81]. Technih [Tod85]. Technik [Bed80b, Fis82, Kra81a, Lan80a, Web85b, Web85b, Sch81c, Web81b, vW83]. Technik-Geschichte [vW83]. Technik-Philosophie [Lan80a]. Techniker [Bra80a]. Technikgeschichte [Bul89, Web87a]. Technikphilosophie [Bra82b]. Technikverständnis [Rap81]. Technique [Smi81b, Smi83e, Mul83c]. Techniques [Cen80, Hug88a, Ros87b, Tho84, Wes86b, Wri86, Ger84b, Smi81b, Smi85f, Coo85]. technisch [vB83]. Technische [Mul89a, Bra82a]. Technischen [Bra83b, Web87b]. Technocracy [Gun82, Sti86b]. Technocratic [Gun82]. Technological [Bij86, Boy85, Con83b, Coo80, Cor87, Cos88, Coz89, Cut81a, DR83, De 83, DeG87, Del82, Dor80, Emm83a, Fah89, Fer80a, Flo84, Fri84b, Fro88, Gol89, Gor83, Gor89a, HH87a, HH87b, Haf85a, Hob80, Kel84b, LM87, Lau82, Lub87b, McG85a, McM81b, McN83, Mer88, Mil82a, Mor87d, Mul83b, Pec86, Roe85a, Ros84, Rum87, Smi82a, Ste88a, Tan89a, Twe86, Vin84, Wol83, vB83, Bur80a, Gol89, Kar87b, McI84, Mer86, Pap84, Rhe88, Sha81a]. Technologie [Kra85, Gil82, Pal81, Sle81]. Technologies [De 82a, Hin81, Kie86, Mir82, Nye89, Rot85b, Smu85, Ste82a, Win89, Sim86, Shu84b]. technologisch [Web84]. technologisch-historischen [Web84]. Technology [And86, Bai83, Bal85, Bee85, Ber87, Ber84, Bez85, Bin83, BBM84, Bry82, Buc82a, Buc85, Bul81c, Bul81b, Bul86b, Bul87, Bus83, Cam89, Car82a, Car85, Cas88a, CL80, Coa83, Con85d, Con89, Coo85, Coo87, Coo89, Cor88a, CRA^{+}89, CRM84, CRM85, CRA86, CRA87, Cut87, CRA^{+}88, Dal88, DR83, Dan83b, De 80b, De 84a, De 84c, DSF89, Esp82, Fol82, FLC89, FW87, Fri86a, Fri81c, Ful80, GK81, GS81, Gli88, Gli89, Goo80b, Goo80a, Goo81, Goo82b, Goo83, Gre88a, Gre81, Gro87b, Gui84, Hac85, Haf85b, Haf85a, Har84, Har81b, Hea89, Hei86, Hin83, Hir86, Hob85, Hol89, Hug81, Hun87b, Hyd86, Ino82, Jam89, Jan87, Jon81, Kar86, Kel84c, KT80, Kie87, Kil88, Kli86a, Kra81b, Kra89a, Kra86a, Lau86, Law85, LS84, Ler88, Les85, Lev87c, Lew80, Lif85]. Technology [Lig87, Lin88, Lue86, Lyt88, Mac86b, Mac87a, Mal86a, Mar86b, Maz84a, McC88, McD82, McI83, Meh87, Mei88c, Mét86, Mey86, Mey88, Mic84a, Mil88a, Mil80b, Min82, Mit85, Moh88, Moh87c, Mor81a, Mor87a, Mor89d, Mor81b, Muh81, Mul83d, Mul88a, Nee83, Nel87, Nev82, Nuw88, Och89, Ove85, Ovi86b, Ovi87, Pal81, Pan87, Per80b, Pfe83, Pur84a, Pur80, Pur82b, Pur89b, Qur86, Rae82, Rae87, Rah82, Rec88, Rei80a, Rei83a, Rei80b, Reu85, Roe85b, Roe82, Rut89, Sar81, Scr88a, Scr88b, See87, SF87, Sla89a, Sla89b, Sla89c, Smi82b, Smi85e, Smi88a, SB82, Sta82a, Sta87, Sta84, Sti82, Str80, TMMY84, Tar85a, Tei82, Tei88, Thr82, Tom85, Tuc80b, Vol88, Wei80a, Wei87a, Wei80b, Wei81a, Wer81, Wes87a, Wes89, Wil89b, Win87, Wos82, You87a, Zie89b, Zim84, Bed80b, EE81]. Technology [Ano83e, Ano85d, Ano87c, Ano89d, Bro86b, Cut81b, Han80, Lan80b, Bow80, Bra80b, Coa80, Cos87, Cow86, Coz89, Dor83, Dur83a, Fer81, Fri80b, Ger81, Gol81, Har86b, Hoc85, Hoc87, Huy82, Kap80, Lat86, Law83, Lew88a, Mac85c, McG88b, McG88c, Mor80c, Per80a, Ros89, Sch80, Shu84a, Smi86a, Spi82b, Sta85b, The85a, Tuc80b, War86b, Wes87c, Zie89a, Ali89, Ano80e, Ano81d, Ano82d, Ano83g, Ano84d, Ano85e, Ano86e, Ano87d, Ano88f, Ano89f, Cut81b, Dan85, Dan81a, Dur83a, Fel84, Gen81, Hol86a, Hou83, Lie81a, McI83, Per84a, Por89b, Qur86, Rol89, SPV87, Smi83b, Sta87, Sta81, Tha86, Tom84, Whi82c, Wil86b, Yao88, You87b]. Technology/Historical [Mor89d]. Technorivalry [McL85]. técnica [Mau82]. tecnología [Gli82, Kel84a]. Ted [Dan80]. Teeth [Hal84b]. Tekhelet [Tra89]. tekhniki [EE81]. Teknik [Jan87]. Tel [Kel86]. Telecommunciations [Gar87b]. Telecommunications [Ber80, Har86a]. Telegraph [Kie87]. Telegrapher [Kie89]. Telegraphy [Rei80a]. Telephone [Fis88, Fri86a, Hod81, Isr87, McG83, Mue89]. Telescopes [Rot85a, Swe88b]. Television [Bir87, Hir85, Fow80a]. Telford [Buc81b]. Tellico [Joh88]. Ten [Smi85c, Smi88b]. Ten-Year [Smi85c]. Tender [Mer88]. Tennessee [Seg89b]. Tensile [Zuk80]. Tension [Smu87]. Tent [Zuk80]. Tentacles [Cam89]. Tenth [Har86a, Smi88a]. Terence [Gol80a]. Teresa [Smi82a]. Term [Dow84a, McN83]. terme [Cen80]. Terminal [Kou81, May83, Seg84]. Terminology [Ryd80]. Territory [Dun80]. Terror [How85a]. Terry [Con85c, Hud86, McM86, Pla86a, Wil89b]. Tesla [DAK89, Car83a]. Test [Per80c, Sha81b, Wal83, Web80a, Rot85b, Hig83b]. Test-Tube [Rot85b]. Testability [Con83b]. Testing [Hac88, Hac88]. Texas [Pre83a, Smi85a, Sti85a, Tom83b, Wea83, Con89, Mil88a]. Text [Owe86]. Textielhistorische [vB86]. Textile [Bar85, Ben86, Jer85, Kor80, Lau87, Lub86a, Mir82, Moh88, Qui89, Scr86b, Spr81, Ger84b, Tan82b]. Textiles [Pon80, Scr87b]. Thackray [Rei87, Rei89]. Theatre [Gol80a]. Their [Dan80, Hil88, Smu85, Swe88b, Tan82a, War87a]. Thematic [Arl84]. Themes [Hug81]. Theo [Fos89]. Theobald [Sta83b]. Theodore [Hok82, Par82, Smi84a]. Theology [Lue86]. théorie [de 82b]. Theorien [Web87b]. Theories [Smi82a]. Theorists [Jam89]. Theory [Bry80, Cha82c, Cha88b, Con83b, DSF89, Dur83b, Goo80a, Gun82, Kem84a, Kli87a, Mye87, Och87, Per81c, SF87, Ste80, Swe88a, Seg86, Ser83]. Thépot [Fro87]. Thermal [Wal89]. Thermodynamics [Cha82b]. Things [Sch86b, Seg84]. Thinking [De 81b, Don86b, Gro87a, Nye89, Vin82, For88a, Nel89]. Third [Bro87b, Law85, Meh87, Mye89, Vic84, Scr86b, Wil84a]. Thirty [Cla89]. Thirty-Three [Cla89]. Tholander [San81b]. Thomas [Bed85b, Bow87, Bur83, Cas88a, Cha87, Eng87, Haf80, Hew83, How84, Hug88b, Kra89a, Mar85b, Mer87, Per86, Pur82b, Rei82, Rey84, Rui80, Seg86, SF87, Sla84, Soc83, Ste81b, Kni81, Sla82, Buc81b, Cha87, How87, Kap82, Ras80b, Rhe82]. Thompson [Kah85]. Thorkild [Hab85]. Those [Cai82a]. Thought [McG88a, Seg81, Wal87a, Bea81, Dav83]. Thoughts [Wei80a]. Thousand [Wei80a]. Thread [Kap80]. Three [Bac82b, Cla89, Gre84, Maz84b, vB83, Kor80]. Threshold [Roe82]. throughout [Ste88b]. Thurber [Abr80, Abr81a]. Tiby [Moy80]. Tichi [Mei88c]. Tilley [Kor86]. Tilmann [Kli86b]. Tim [Och88, Ste88c]. Timber [Gol84]. Time [Bar89b, Bed87, Dav83, Dub83, Hin83, May85, Nel89, Ovi88, Ste89, The85b, Bar91, Ste87, Ste89, Bed80b, Bed83c, Car83a, Ovi88, Ung83]. Times [Hal80b, Mar88, McC85, Per85b, Sla88, Fli89a, Lan87a]. Timothy [Elb87]. Tin [Bur80b, LK80, Ste88b]. Tisdell [Pur84a]. Titus [Cla86, Hac88]. Tobacco [Kor86, Tin87]. Toby [Ger81]. Tolstoy [The84b]. Tom [Gie81b, Gil85a, Rol82, Tro87, Hor84]. Tomorrow [Sla89a, Tro87, Haf85a]. Too [Boy86, Scr88b]. Tool [Cla89, Hev89]. Tools [Fin83, Fit83, Har80b, Hok82, Spi82a, Sti82, Bat80, Hin81]. Topeka [McD86]. Topicality [Buc85, Hol86a]. Tordoff [Mul87]. Toronto [FBW^{+}81]. Torpedo [Whi82a]. Touch [Fis88]. Tour [Twe88]. Tower [Mei84]. Town [Ber83, Cai88, Gra88, Hin85, Lan86, Web81c]. Town-Builders [Gra88]. Towns [Har87c, Pre83a]. Townsend [Hal89a]. Toxic [Kro86]. Toyota [Rae87]. Traces [Pen82]. Traction [Con81a]. Tractor [Soc89]. Trade [Ben86, Gar81c, Har87c, Mar85c, McL85, Mul88b, Mul89c, Ste88b, Har89b, Pap82, Wri86, Bai83]. Tradeoffs [Tei88]. Trades [Ran86a]. Tradesmen [Hin80a]. Tradition [Cas88b, Seg83]. Traditional [Hob80, Ros87b]. Tragicomical [Cha82b]. Trains [Whi83a]. Trans [Goe82, Kid87]. Trans-Mississippi [Goe82]. Trans-Pacific [Kid87]. Transactions [Gor82]. Transatlantic [Rei80a, Mir82]. Transcontinental [Kle89]. Transdisciplinary [Bur80a]. Transfer [Ame89, Cam89, Hea89, Kra81b, McI83, Mor89d, Sta81, Vol88, Win89, Ino82, Sar81]. Transformation [Ame89, Bri89, Haf80, Hit85, Kul88, Tin87, Win82]. Transformations [For88b, Coo89]. Transformed [Rum88]. Transit [Che84, Hil81, McK86]. Transition [Kar87b, Spi81]. Transits [Lan87d]. Translation [Bed80b]. Transmission [Isr87]. Transport [Bul86a, DR83, Hir84, Jan86, Bau86, Rae82]. Transportation [McK86, Whi84a, Sal83]. Transports [Bac85b]. travail [Ger84b]. Travel [Jak87, Whi83a]. Travelers [Com87a]. Traveller [Sta88a]. Travelling [Gar81c]. Treadmill [Sha89b]. Treatment [Ham88]. Trebilcock [Coo84c]. Tree [Rot85a]. Trends [Kie86, Qur86, Smi85d, McN83]. Trescott [Coo80, Mul83b]. Treue [Bul89]. Trevor [Buc82a, Fri89b, Lie83]. Trial [Sta87, Scl88]. Trials [Hew87]. Tribulations [Ste82b]. Tribute [Kra89b, McG84]. Trimble [Dav89b]. Trinder [Bry81]. Trinity [Hac86, Tre85, Tre85]. Tripp [Arg88]. Tristram [McG88b]. Triumph [Tro87]. Triumphs [Ste82b, Leh86]. Troitzsch [vW83]. Trolley [Hil85, Hil85]. Trouble [Hil83]. trudiy [EE81]. True [Daw89, Moh87a].

Truesdell [Cha82b]. Truslow [Tre85]. Truss [Con81e, Mil81]. Trust [Wal82]. Truth [She84, Tra85]. Tsangadas [Bac82a]. Tsatsawassa [Wes86a]. tseins [Bra82c]. Tsien [Pan87]. Tsuen [Pan87]. Tsuen-Hsuin [Pan87]. Tube [Rot85b]. Tubes [Wis84]. Tucker [Swe88b, Lau87, Swe88b]. Tufte [Kid89]. Tule [Kah85]. Tunnel [Dav89a]. Tunnelling [Dre89]. Tunstall [LeM86]. Tunzelmann [Car80]. turbine [San81b]. Turbines [Con83b]. Turbojet [Lew82b]. Turing [Cer85]. Turkle [Kap86]. Turn [Isr87]. Turned [Gor88c]. Turner [Bed86, Bed87, Mor89b]. Turriano [HS88, Ovi85]. tussen [Ung88]. TVA [Seg86, Chr89, Joh88, Kru85]. Tweedale [Gor89a]. Twelfth [Fra84, Kel89]. Twelfth-Century [Kel89]. Twentieth [Ber83, Bra83a, Buc82a, Lin85, Pal80, Qui89, Ros87b, Tuc82a]. Twentieth-Century [Ber83]. Twenty [Fin89, Man81, Ovi85, Spi82a, Sta84, Tro87, Sha81c]. Twenty-five [Spi82a, Sta84, Tro87, Sha81c]. Twenty-One [Ovi85]. Twenty-Year [Man81]. Twice [Hac86]. Two [Doo87, Gil85a, Hev86, Jac84, Kar86, Mar85c, Bur82, McG86, Har82b]. Tylecote [Gor88b]. Tyneside [Fit82, Maj81]. Typewriter [Abr80, Cos85, Hok87]. Typewriters [Hok83]. Typewritten [Abr81a]. Typographic [Bol88].

U [Hol87a]. U-2 [Hol87a]. U.S [All87, Big88, Cai82a, Cas85, Cas86, Cut87, De 80b, Dob81b, Got83, Hal81, Hal89b, Hig80c, Hol87a, Lev87a, Maz83, McK86, Sch85a, Sma80, Smi87a, Sti84, Zie89a]. U.S. [Kra81b, Bil82, Cla80, Dun81, Hug88b]. Übersicht [Web87b]. Ulrich [Bra80a, vW83]. Ultimate [Tre87]. Uncertainty [Mac88b, Vin86a]. Underdeveloping [Wei87b]. Underground [LM87]. Understanding [Fow82, Har81b, Hou80b, Hir86]. Uneasy [Dav89c]. Unemployment [Cos87]. Unequal [Wei87b]. Unesco [Fer80b]. Ungarn [Web81a]. Unger [Whi82b, vS80]. Unified [Och87]. Union [McD85a, Bal85, Coo83]. Unionism [Gar81c]. Unionization [Sho88]. Unions [Cro82]. Unique [Wos84]. unité [Kra85]. United [Bil81a, Blu82, Bro83, Cal85, Con85a, Con81f, Con85d, De 84c, Fla89, Hal84b, Hir86, Hol83, Kas83, Mar84b, Mar83b, Mel83b, Mey86, Nel82, O’B81, Rot84, Saw88, Sla89c, Smi85d, Vie83a, Wal81, War83, Wei86, Wil89a, Mil80b, TMMY84]. Unity [Roe85b]. UNIVAC [Man83, Wre89]. Universe [And82, DeV85, Fri84a]. Universelles [Ryd89]. Universität [Bra83b]. Universities [Bur86, Vey89, De 80b]. University [Bil81a, DR83, Hob85, Lau82, LS84, Lew89, Mac88a, Mor87e, War87b, Wei84, Wil89b, Wos82, Mor89c]. Unknown [LB82]. Unmanned [Web80b]. Unsystematic [Tra85]. Untold [Bez83, How86]. upon [Lan84]. Upper [Mer81b, Rey87, LM87]. Upset [Mul82]. Urban [Cha89, Che84, Con88a, Cou88, Han89a, Hil81, Hoy88, Les82, McK86, Nas80, Pla86b, Ros84, Ros89, Sti81, Whi83b, Whi84a, Blu82, Sch86a]. Urbanization [Lew89]. USA [Dan83b]. Use [Thr82, Hil88, Smi89a]. Uselding [Cai82b, De 80a, Fre80, Lin81a]. User [Tar85b]. Uses [Bin83, Lub86b, Nel89]. Usher [Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89a]. USSR [McI83, Pel83]. Utilities [Tay89]. Utopia [Roe85b]. Utopian [Chr89, Seg81]. Utopianism [Rhe88]. Utopians [Roe89]. Utopias [Ols89]. Uwe [Pet87].

V [Bla87, Car89, Cro87a, EE81, Feu88, HS88, McS83, Per81b, Por89a, Rea81, Rey87, Rol83, Ste83, Sti87, Tan82a, Tay89, Wes87c, Lin81a, Ovi88]. val [San81a]. Valley [Bar85, Ger84a, Kem84b, Pow83, Seg89a, Ble83, Sho88]. Valley-Los [Pow83]. Value [Lec84]. Values [Eng87, Gre81, Jon81, Smi83b, Tar85b, Rei83b, Wei80a]. Valves [Wis84]. Vámos [Ain89]. Vandelvira [San89]. VandenBroeck [Lew80]. Vanderburg [Och87]. Vanguard [Dan83c, Goe83, Goe82]. Vannevar [Owe86]. Vary [Rei80a]. Vauban [Lan87c]. Vehicle [Als81, Kar83]. Vehicles [Jak87, Fos89]. Venereal [Bul81d]. Ventilation [Fit81]. Venus [Lan87d]. Verdigris [Ben84]. Verein [Web85b]. Vereins [Tod85]. Verkstadsbyggnader [San81a]. Vermont [Coo87]. Vernon [Spu89]. versus [Emm80b, Rhe86]. Vertical [Con85c]. vertikalen [Pet87]. Very [Wil87b, Maz86b]. Vessel [Gui89]. Veszprém [Haf85b]. Vicki [Sho88]. Victor [Gia81]. Victoria [Eve86]. Victorian [Car84b, Cla88, Gor86, Gre80, Moh87c, Rhe86, Tan82a, Hai84, Twe88, Moy80]. vid [San81b]. VideoDisc [Bro88b]. Videohistory [Sch89]. Vietnam [Tre87]. Vietor [Don81b]. View [BM84b, Bry81, Con84, Dow85, Han89b, Hor88, Gri81b, Man81, Sok81]. Views [Con86b, Hic85a]. Vig [Ali89]. Vignoles [Smi85b, Smi85b]. Vigueras [Kel84a]. Village [Ben88, Bra80b, Lan83b, McL84]. Village-Level [McL84]. Villainy [Pow83]. Villard [Hey88]. Vincenti [Dor80]. Vinci [Ano80g, Ano81f, Ano84f, Ano85g, Ano86g, Ano87f, Ano88h, Ano89h, Car82b, Hal88b, Hal88a]. Virgil [Wit83]. Virginia [Gol84, Pur80, Hyd86, Bed83b]. Virgins [Coo80]. Vision [Les85, Pow83]. Visionary [Cha80]. Visions [Lew80, Nas89, Seg84]. Vistas [Ryd89]. Visual [Kid89, Per80a, Wil81]. Vittorio [Kel88]. Vivian [Ait81]. Vladimir [Zee82]. Vol [All87, Atk89, Bed87, Bed89b, Bos85, Bra81a, Bra82a, Bri81b, Car89, Con81f, Cou88, Ell87, Goo82a, Hal81, Har85, Har87c, Har89a, Hig80c, Hin86b, Hug88a, Hun84, Ken86, Lan87c, Lan83a, Maa81, Mét86, Mul88a, Mul89c, Mus83, Pan87, Ros89, Saw88, Seg87, Str88, War86b, Zim84, vB86, Pyc82]. Volker [De 84c, Tod84b]. Vols [Smi85f, Buc82a, Mil82b, Nel87]. Volti [Tom84]. Volume [Ano80e, Ano80i, Ano80h, Ano80j, Ano80k, Ano81d, Ano81h, Ano81g, Ano81i, Ano81j, Ano82d, Ano82g, Ano82f, Ano82h, Ano82i, Ano83g, Ano83j, Ano83i, Ano83k, Ano83l, Ano84d, Ano84h, Ano84g, Ano84i, Ano84j, Ano85e, Ano85i, Ano85h, Ano85j, Ano85k, Ano86e, Ano86k, Ano86j, Ano86l, Ano86m, Ano87d, Ano87i, Ano87h, Ano87j, Ano87k, Ano88f, Ano88l, Ano88k, Ano88m, Ano88n, Ano89f, Ano89k, Ano89j, Ano89l, Ano89m, Lau86, Smi88a, Vin82, Car82a]. Voluntary [Lew80]. VW [Whi84b].

W [All84, Bal80, Bos85, Bri80b, Bro88b, Buc83a, Bud82, Cai82a, Coa83, Coc81, Com87a, Coo86a, Cos85, Dan81a, Dav89b, Dav85, De 84b, Fit82, Fos82, Gor86, Hab82, Had80b, Har85, Hic85a, Hil81, Hoc88, Hou83, How80a, Hyd86, Iri89a, Isr87, Kar87a, Ken86, Ket85, Kir80, Koh87, Kou81, Kra89b, Lan84, Lec88, Lew82b, Mac86a, Mac87a, Maj81, Man89, May83, Mor87e, Per81a, Pur83, Rab81, Rae85, Rea81, Rei80b, Rey89a, Rob80, Roh84, Scl88, Scr87a, Seg84, Sin82, Sin86a, Smi83a, Sti81, Tro87, Upt83b, War89a, War84, Wes83, Whi82b, Wil87a, Wil80, Wis84, Yor88, Zee82, vS80]. W. [Cha82c]. Wachhorst [Rhe82]. Wachsmann [Hor86]. Wächtler [Wer82, Mul89a]. Wade [Bin83, Ska88]. Wagar [Seg84]. Wage [War83]. Wage-earning [War83]. Wagen [Bul89]. Wagenbreth [Mul89a]. Wagner [Ros87b]. Wagon [Wik81]. Wajcman [Hoc87]. Wakeman [Wri86]. Wakens [Sta88c]. Wales [Buc82c]. Walker [Hac87, McI81, Sta83a]. Wallace [Swe88b]. Wallingford [Bed80b]. Wallops [Per80c]. Walter [Bil81b, Dor80, Gra84, Hal87, Hig82, Hor86, Hyd86, Kar87a, She81a, Smi81a, Wea83]. Walzer [Mor83]. War [Boy89, Buc86d, Cas88b, Cla80, Coo86b, Dea84, Eva88, Hal81, Hal84b, Hal84a, Hig80c, Hir86, Kra81b, Min82, Nel82, Neu89, Rol89, Rot82, Rup88, San80, Smi85e, Swo89, Tre87, Tut81, Vin86b, Hol88, McD89, Mor85, Bac86]. War-Industries [Coo86b]. Ward [Qur86, See87, Tuc80b, Bed85a, Dun89, Seg87]. Wardens [Dan81b]. Warfare [All87, Bri81a, Gar88, Gra87a, Hol88, Lan87c, Pit86, Thi80, Rol87]. Warner [Mar86c]. Warp [Hin83]. Warren [Rae80a, Seg84]. Warriors [Fle88]. Wars [Gar81a, Kel84b, LeM82, Smu87]. Warships [Ung81]. Was [Gra88, Kar83, She80a]. Washington [Cre83, Dob81a, Doo87, Mor87a, Sch89, Ste87, War88, Wor86a, Wri85b, LHH^{+}84]. Wasserman [Isr87]. Waste [Dow84a, Gal87, dlB88, Dow84b]. Wastes [Hoy88, TMMY84, Tar85b]. Wastewater [TMMY84]. Watch [Har80b]. Watches [And81]. Watchmaking [And83, Ste85]. Water [Bla84, Ble83, Car84b, Con88b, Con83b, Con85c, Fit82, Gle84, Hab85, Hof84, Hun87a, Hyd86, Pow83, Sim83, Sim85, TMMY84, Tar85b, McL84, Row89, SS84, Ble83, Fit82, Jac81, Mal86b, Rey89b, Smi89a]. Water-Bound [Con88b]. Water-Clocks [SS84]. Water-driven [Sim83, Sim85]. Water-Powered [Hyd86, Gle84]. WaterClocks [Bed87]. WaterClocks/Sand [Bed87]. WaterClocks/Sand-Glasses/Fire-Clocks [Bed87]. Watercolors [Hor88]. Waterfront [Con88b]. Waterpower [Con81f, Gor85b]. Waters [Bla84, Fee85]. Waterway [Tro84b]. Waterways [Bro80, Abb86]. Waterworks [Jac89b]. Watkins [Mil82b]. Watrous [Wri86]. Watson [Mil83, Wes86b]. Watt [Lan84, Mus83]. Wave [Ait81, Hun87b]. Waves [O’B87]. Way [Bez83, GF87, Jor89, Ros81]. Wayne [Joh87a, Wil80]. Wayward [Hoc85]. Wayzata [Jon81]. Weapon [Dea84, Tre87, Arn89]. Weapons [How85a, Hal86b]. Weart [Bri80a]. Weather [Fle88]. Weavers [Ben88, LGWK80, Moh87a, Moh88]. Weaving [Moh88]. Webb [Rot85a]. Wedgwood [Mei88b]. Weible [Mel83a, Bah87, Uss88]. Weick [Luk89]. Weiher [Tod80]. Weimar [Meh87]. Weinberg [Lub85]. Weiner [Nel82, Kar81]. Weingartner [Hyd88a]. Weinstein [Per85a, Tar81]. Weiss [Emm83a, Ste85]. Weisz [Bur86, Smi82b]. Weitzman [Cor88b, Pen82]. Welcome [Kap82]. Weld [Mul82]. Welding [Wor86b]. Weldon [Bir81]. Welfare [Har82b]. Wells [Bea81, Wil84a]. Welsh [Jac87]. Welt [Tod87]. Wendorff [Bra82c]. Wendy [Ber87]. Wenk [Tei88]. Were [Nye89]. wereldbeeld [Mul84b]. Werk [Neu88]. Werneke [Cos85]. Werner [Gor82, Ree81]. Werrell [Sah87]. Wertime [Par82, Smi84a]. Wesley [Bri80b, Wei87a]. Wessell [For88a]. West [Cas88b, Goe82, Gra88, Hyd86, Jac81, Jac83, Jac86, Kil88, Kul88, Leh87, Ovi88, Pap82, Rah82, Row89, Sti81, Wea83, Dre89, Iri89b, Muh81, Ran86b, Sch84, Wer81, Ale86, Goo86, Hun87a]. Western [Bai83, Cer85, Col84, Con85a, Dav83, Gie81a, Gra88, Hof85, Isr87, Jac80, Mar88, McI83, Rey84, Sau86, Seg81, Wit83, Ovi86b]. Westfall [Mah82]. Westing [Coo86a]. Westing-house [Coo86a]. Westminster [McI81]. Westralian [You87a]. Westwood [Duf80]. wetenschap [Ung88]. Wexelblat [Cer84]. Whale [McG88c]. Whalecraft [Mar85b]. Whalers [Sin88a]. Whales [Mar88]. Whaling [Mar88, Kap82]. Whalley [Buc88b]. Wheel [FRCL83, Con85c]. Wheeled [Bul86a]. Wheeler [Joh88]. Wheels [Bel82]. Whence [Wer81]. Where [Arn89, Mar85c, Hyd86]. Whiffen [Con82]. Whirlwind [Ste81b]. White [Con86a, For88b, Hal89a, Sch84, Sta82b, Str80, Har86b, Mar83a, Wil88]. Whitfield [Gre88c]. Whither [Whi83b, Bin83]. Whitmore [Str84]. Whitney [Smi87a, Vic87]. Whittle [Daw89]. Who [Gie86, Gor88c]. Wider [Han89b]. Wiebe [Coz89]. Wiener [Ste82a, Har83a]. Wilcox [Wor81]. Wilderness [Rey89c, Rey89b]. Wildes [Kli87b]. Wildlife [Run84]. Wilentz [Sch85b]. Wilfley [Mol84a]. Wilfried [Tod84a]. Wilgus [Kon89]. Wilhelm [Bul89, Ser83]. Wilkening [Dan85]. Wilkinson [Ola85, Pen80]. Will [Kra86b, Les85]. Willem [Och87, Pit86]. William [Baa87, Ben86, Ble83, Cas85, Cou88, Dav89b, Dow84b, Gor84, Gui84, Hal80b, Hin81, Hol89, Jac86, Joh83, Joh88, Law85, Lev87c, Lev87a, McL85, Mei81, Mis84, Mor87c, Mul83a, O’B81, Pap82, Ras80a, Rey89b, Run81, Sch85a, Sok81, Str88, Zie89a, Bac85b, Ham88, Kon89, Ste89, Bal80, McG84]. Williams [Buc82a, Fri89b, Gre84, Har83b, Kle89, Lie83, Soc89, Tro87, Wes87c]. Williamsburg [Wal82]. Williamson [Mor86, Mor88]. Willingham [Cas85, Rey89b]. Willis [Smi87a]. Willson [Fri81a, Kas84]. Wilmington [Lin85, Smi87b, Ano89n]. Wilson [Bed83c, Cam85, Jer85, Pon80]. Wind [Web80a]. Windmill [Kel89]. Windmills [Wes87b]. Windows [Vic87]. Wines [Gal87]. Wing [Vin86a]. Winged [Lew85, Bil85b]. Wings [McD84, Rol82, Smi83c, Smi81a, Bil85b]. Wingspread [CL80]. Winkelmann [Web84]. Winner [McG88c]. Winning [Dea84, McL85]. Winston [Lub86b, All88]. Winter [Bil87b, Bli82, Mor85]. Winterthur [Sei80, Slo81]. Wires [Jac86]. Wisconsin [CL80]. Wise [Smi87a].

Wissenschaft [Bra83b]. Wissenschaftsgeschichte [Fis82, Kra81a]. Witchcraft [Est85]. within [Nel81]. without [McD85a, Vin84]. Witness [Dur83b, Leh86]. Witney [Bat80]. Witold [Zup88]. Wittfoht [Mar86b]. Wizard [Rae85]. Wizards [Dea85]. Wohlauf [vW83]. Wolfe [LB82, Seg89b]. Wolfgang [Tod85, Whi83a]. Wolfs [Mil83]. Woman [Haf85a, Moh87a, Cos85]. Women [Ber87, Coo80, Cos85, Dav89c, Eat88, HH87b, Haf85b, Haf85a, Jer82, Ler88, Nel82, Och89, Pap84, Pur81, Rot85b, Sho88, Tan82a, War83, Och89, Ain89, Coo89, Haf80, Haf85a, Nel82, Och89, Sho88, War82, War84]. Won [Gra88]. Wonderful [Ber89]. Wonders [Bil85b]. Wood [Mil80b, Gol84]. Woodbury [Rae84a]. Wooden [Hin81, Sch82b]. Woodward [Had80b, Nev82]. Woodworking [Bat80]. Wool [Gro87b, Jer85]. Wooldridge [Bil85b]. Woolen [Ran86b]. Woollen [Hit85]. Woolrich [Sta88a, Sta88a]. Word [Lan85]. Work [Buc88b, Cas88a, Coo89, Cos85, Cro82, Dub83, Fri86c, Gie81b, Gra84, Hac85, HH87b, Haf80, Hof80, Kel86, Kul88, Lub87a, Mar85a, Mul82, Nel82, New85, Nuw88, Rup88, Sch82d, Sho88, Sta85a, Ung83, War83, Cos87, Gie81b, Joh83, O’C88, Rum88]. Work-Force [Nuw88]. Worker [Pal80, Nye88, McD85a]. Workers [Coo86a, Cos85, Kra86b, McD85a, McD86, Mey85a, Mey88, Nel82, O’C88, Pre87, Ran86b, Rum87, Sch86a, Sho88, Tan82a, Tan89a, Fre82, Lub87a, Rob88]. Working [Fol80b, Ger84a, Mac86a, Nel81, Ran82c, Sch85b, Scr86b, Gra84]. Workplace [Mey88, Cum88]. Works [GF87, Kon84, Mel83b, Min81, Seg87, Sem84, Vic87, EE81, Bil83]. Workshop [De 84a, Har81b, War87b]. workshops [San81a]. World [Baa89, Car83b, Cas88b, Cla80, Coh86, Con80c, De 84b, Fri82a, Gri81b, Hal81, Hal84b, Hal84a, Hig80c, Hol88, Hud86, Koe84, Lan89a, Law85, Mac88a, McD85a, Mei82, Mor87c, Mor85, Nel82, Rot82, Rup88, Ryd89, Smi80a, Smi84b, Spi81, Sta89, Ste88b, Swo89, Tan81, Thi80, Tre87, Uss88, Vig83, Vin86b, Wes86b, Wil84a, Mul84b, Min82, Neu89, Tei88, Tut81, Bro87b, Gen81, Gol80b, Kas84, Kni81, Kul88, Leh83, May85, Pur89a, Seg81, Sla84, Har80a]. Worlds [Har83b, Mar85c, Sin87]. Worms [Kel82]. Woronoff [Gut87]. Worse [Dan81b]. Worst [Lan88]. Worster [Hun87a]. Wright [Cro86, Rig81, Rol80, Bar85, Coo89, Hut88, McD89, Och88, Upt83a]. Wrights [Cor85]. Writing [SB80]. Writings [Bed80b, Ras80b]. Wrong [Flo84]. Wrought [Kra84, Rae84b, Sta84]. Wulff [Mar80]. Wyatt [Hal80b]. Wyke [Har80b]. Wyman [Jac80]. Wyn [Rhe82].

X [De 83, DeV85]. X-ray [DeV85]. XIII [de 82b]. XIX [Kra85]. XP [Hol87b]. XP-75 [Hol87b]. Xuechen [Qur86]. XVI [Mau82, de 82b]. XVII [Mau82]. XVIII [Kra85, Smi86c]. XVIth [Ano80l]. XX [Smi86c].

Yago [McK86]. Yamamura [McL85]. Yankee [Pur84b]. Yates [Hil84]. Yeager [Vie83b]. Year [Fer80b, Man81, Smi85c]. Years [All87, Bil80a, Bil81b, Bos85, Cas88b, Cla88, Cla89, Doo87, Ell87, Fin89, Gor84, Kap80, Kar86, Kop82, Maj81, Mar83b, McM86, Mil84b, Mil88a, Pyc82, Seg86, Sha81c, Smi81a, Spi82a, Spu89, Sta82b, Sta84, Tob85, Tro87, Web85a, Wel89, Wis84, Wri85b, Bir81, Bri85, Bri81b, Per80c, Str84]. Yehuda [Wil84c]. Yellow [Wil88]. Yergin [Bra81b]. York [Con86b, Cud88, Hil81, Hin80a, Hin84, Kou81, LS84, May83, Maz86b, Mei82, Mel85, Sch85b, Sma80, Wos82, Cor87, Dav87, Iri89b, Kon89, Lev88, Col81]. Yorkshire [Hit85]. Young [Ful81, Hil85, Nel87]. Yvonne [Lew80].

Z [Swe88a]. Zalinski [Han84, Sha85]. Zavella [Sho88]. Zealand [Tro87]. Zeewezen [Ung88]. Zeit [Bra82c]. Zeitbewuß [Bra82c]. Zeitlin [Tan89a]. Zeitmessers [Bed80b]. Zero [vdL83]. Zimmerman [Haf85a]. Zonca [Kel88]. Zsámboki [Web81a]. Zuckerman [Vin86b]. Zur [Bed80b, Web85b, Bra83b, Fis82, Kra81a, Wer82]. Zussman [Fri86c]. Zvi [Fel87].

References

Abbott:1986:CRH

[Abb86] Carl Abbott. Conquering the Rivers: Henry Miller Shreve and the Navigation of America’s Inland Waterways by Edith McCall (review). Technology and Culture, 27(1):161–162, January 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

           URL https://muse.jhu.edu/pub/1/article/889660/pdf.

Abraham:1980:CTT

[Abr80] Terry Abraham. Charles Thurber: Typewriter inventor. Technology and Culture, 21(3):430–434, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

           muse.jhu.edu/pub/1/article/890705/pdf.

Abraham:1981:TTL

[Abr81a] Terry Abraham. Thurber’s typewritten letter. Technology and Culture, 22(4):851, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

           edu/pub/1/article/890978/pdf.

Abrams:1981:HMM

[Abr81b] John W. Abrams. Herbert Marshall McLuhan (1911–1980). Technology and Culture, 22(4):843–844, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

           https://muse.jhu.edu/pub/1/article/890973/pdf.

Adams:1984:AFS

[Ada84] Nicholas Adams. Architecture for fish: The Sienese Dam on the Bruna River — structures and designs, 1468–ca. 1530. Technology and Culture, 25(4):768–797, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889946/pdf.

Adams:1989:MSM

[Ada89] David P. Adams. Medical Science and Medical Industry: The Formation of the American Pharmaceutical Industry by Jonathan Liebenau (review). Technology and Culture, 30(3):711–712, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889077/pdf.

Agassi:1982:ESD

[Aga82] Judith Buber Agassi. The Economies of Sex Differentials by Cynthia B. Lloyd, Beth T. Niemi (review). Technology and Culture, 23(4):698–702, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890278/pdf.

al-Hassan:1980:GAC

[aH80] Ahmad Y. al Hassan. The Genius of Arab Civilization, Source of Renaissance ed. by J. R. Hayes (review). Technology and Culture, 21(3):477–478, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890712/pdf.

Ainley:1989:WSO

[Ain89] Marianne G. Ainley. Women in Science: Options and Access ed. by Éva Vámos (review). Technology and Culture, 30(4):1085–1086, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889012/pdf.

Aitken:1981:ERW

[Ait81] Hugh G. J. Aitken. Early Radio Wave Detectors by Vivian J. Phillips (review). Technology and Culture, 22(3):643–644, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890914/pdf.

Alexander:1985:BDN

[Ale85] J. T. Alexander. The Black Death: Natural and Human Disaster in Medieval Europe by Robert S. Gottfried (review). Technology and Culture, 26(4):831–833, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889684/pdf.

Alexander:1986:NDW

[Ale86] Thomas G. Alexander. The New Deal and the West by Richard Lowitt (review). Technology and Culture, 27(1):164–165, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889662/pdf.

Alic:1989:TPE

[Ali89] John A. Alic. Technology and Politics ed. by Michael E. Kraft, Norman J. Vig (review). Technology and Culture, 30(4):1082–1083, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889008/pdf.

Allison:1984:RDC

[All84] David K. Allison. Radar Development in Canada: The Radio Branch of the National Research Council of Canada 1939–1946 by W. E. Knowles Middleton (review). Technology and Culture, 25(1):120–121, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890090/pdf.

Allen:1986:BAS

[All86] Garland Allen. Black Apollo of Science: The Life of Ernest Everett Just by Kenneth R. Manning (review). Technology and Culture, 27(4):900–901, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889562/pdf.

Allison:1987:HUE

[All87] David K. Allison. The History of U.S. Electronic Warfare. Vol. 1: The Years of Innovation — Beginnings to 1946 by Alfred Price (review). Technology and Culture, 28(1):172–173, January 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889482/pdf.

Allison:1988:MMB

[All88] David K. Allison. Misunderstanding Media by Brian Winston (review). Technology and Culture, 29(2):328–329, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889295/pdf.

Alsop:1981:ISP

[Als81] James D. Alsop. The invention of a self-propelled vehicle in the 16th century. Technology and Culture, 22(4):753–756, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890938/pdf.

Ames:1989:TTI

[Ame89] Kenneth L. Ames. The Transfer and Transformation of Ideas and Material Culture ed. by Peter J. Hugill, D. Bruce Dickson (review). Technology and Culture, 30(4):1094–1097, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889024/pdf.

Andrewes:1981:HCW

[And81] William Andrewes. The History of Clocks and Watches by Eric Bruton (review). Technology and Culture, 22(2):302–303, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890849/pdf.

Andrewes:1982:CUG

[And82] William Andrewes. The Clockwork Universe: German Clocks and Automata 1550–1650 ed. by Klaus Maurice, Otto Mayr (review). Technology and Culture, 23(2):227–228, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891150/pdf.

Andrewes:1983:WGD

[And83] William Andrewes. Watchmaking by George Daniels (review). Technology and Culture, 24(3):545–546, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891424/pdf.

Anderson:1986:STI

[And86] R. G. W. Anderson. Science and Technology in Indian Culture — a Historical Perspective ed. by A. Rahman (review). Technology and Culture, 27(4):857–858, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889516/pdf.

Andersen:1988:PIS

[And88] Håkon With Andersen. Proto-Industrialisation in Scandinavia: Craft Skills in the Industrial Revolution by Maths Isacson, Lars Magnusson (review). Technology and Culture, 29(4):959–960, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889193/pdf.

Anonymous:1980:APU

[Ano80a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 21(3):435, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890706/pdf.

Anonymous:1980:DP

[Ano80b] Anonymous. The Dexter prize. Technology and Culture, 21(3):436– 437, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890707/pdf.

Anonymous:1980:E

[Ano80c] Anonymous. Erratum. Technology and Culture, 21(1):159, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890649/pdf.

Anonymous:1980:IEC

[Ano80d] Anonymous. IEEE to establish center for the history of electrical engineering. Technology and Culture, 21(1):160, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890985/
pdf.

Anonymous:1980:IVT

[Ano80e] Anonymous. Index to volume 21 (1980) of Technology and Culture. Technology and Culture, 21(4):719–738, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890795/pdf.

Anonymous:1980:JCR

[Ano80f] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 21(3):551, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890987/pdf.

Anonymous:1980:LVM

[Ano80g] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 21(3):437–445, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890708/pdf.

Anonymous:1980:VNJa

[Ano80h] Anonymous. Volume 21, number 1, January 1980. Technology and Culture, 21(1):??, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1980:VNA

[Ano80i] Anonymous. Volume 21, number 2, April 1980. Technology and Culture, 21(2):??, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1980:VNJb

[Ano80j] Anonymous. Volume 21, number 3, July 1980. Technology and Culture, 21(3):??, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1980:VNO

[Ano80k] Anonymous. Volume 21, number 4, October 1980. Technology and Culture, 21(4):??, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1980:XIC

[Ano80l] Anonymous. XVIth international Congress of the history of science. Technology and Culture, 21(2):356, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890986/pdf.

Anonymous:1981:APU

[Ano81a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 22(3):560, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890889/pdf.

Anonymous:1981:A

[Ano81b] Anonymous. Announcement. Technology and Culture, 22(4):854, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890981/pdf.

Anonymous:1981:DP

[Ano81c] Anonymous. The Dexter prize. Technology and Culture, 22(3):561– 562, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890890/pdf.

Anonymous:1981:IVT

[Ano81d] Anonymous. Index to volume 22 (1981) of Technology and Culture. Technology and Culture, 22(4):855–874, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890982/pdf.

Anonymous:1981:JCR

[Ano81e] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 22(3):559, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890888/pdf.

Anonymous:1981:LVM

[Ano81f] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 22(3):563–569, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890891/pdf.

Anonymous:1981:VNJa

[Ano81g] Anonymous. Volume 22, number 1, January 1981. Technology and Culture, 22(1):??, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1981:VNA

[Ano81h] Anonymous. Volume 22, number 2, April 1981. Technology and Culture, 22(2):??, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1981:VNJb

[Ano81i] Anonymous. Volume 22, number 3, July 1981. Technology and Culture, 22(3):??, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1981:VNO

[Ano81j] Anonymous. Volume 22, number 4, October 1981. Technology and Culture, 22(4):??, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1982:APU

[Ano82a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 23(3):440, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891186/pdf.

Anonymous:1982:A

[Ano82b] Anonymous. Announcement. Technology and Culture, 23(4):704, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890280/pdf.

Anonymous:1982:DP

[Ano82c] Anonymous. The Dexter prize. Technology and Culture, 23(3):441– 442, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891187/pdf.

Anonymous:1982:IVT

[Ano82d] Anonymous. Index to volume 23 (1982) of Technology and Culture. Technology and Culture, 23(4):705–715, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890281/pdf.

Anonymous:1982:JCR

[Ano82e] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 23(3):439, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891185/pdf.

Anonymous:1982:VNJa

[Ano82f] Anonymous. Volume 23, number 1, January 1982. Technology and Culture, 23(1):??, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1982:VNA

[Ano82g] Anonymous. Volume 23, number 2, April 1982. Technology and Culture, 23(2):??, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1982:VNJb

[Ano82h] Anonymous. Volume 23, number 3, July 1982. Technology and Culture, 23(3):??, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1982:VNO

[Ano82i] Anonymous. Volume 23, number 4, October 1982. Technology and Culture, 23(4):??, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1983:APU

[Ano83a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 24(3):466, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891389/pdf.

Anonymous:1983:Aa

[Ano83b] Anonymous. Announcement. Technology and Culture, 24(2):315, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891293/pdf.

Anonymous:1983:Ab

[Ano83c] Anonymous. Announcements. Technology and Culture, 24(3):554– 557, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891429/pdf.

Anonymous:1983:DP

[Ano83d] Anonymous. The Dexter prize. Technology and Culture, 24(3):467– 468, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891390/pdf.

Anonymous:1983:DIM

[Ano83e] Anonymous. Directory of individual members: The Society for the History of Technology. Technology and Culture, 24(1):157–181, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891253/pdf.

Anonymous:1983:GPH

[Ano83f] Anonymous. Graduate program in the history of industrial America. Technology and Culture, 24(1):156, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891252/pdf.

Anonymous:1983:IVT

[Ano83g] Anonymous. Index to volume 24 (1983) of Technology and Culture. Technology and Culture, 24(4):725–737, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890119/pdf.

Anonymous:1983:JCR

[Ano83h] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 24(3):465, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891388/pdf.

Anonymous:1983:VNJa

[Ano83i] Anonymous. Volume 24, number 1, January 1983. Technology and Culture, 24(1):??, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1983:VNA

[Ano83j] Anonymous. Volume 24, number 2, April 1983. Technology and Culture, 24(2):??, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1983:VNJb

[Ano83k] Anonymous. Volume 24, number 3, July 1983. Technology and Culture, 24(3):??, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1983:VNO

[Ano83l] Anonymous. Volume 24, number 4, October 1983. Technology and Culture, 24(4):??, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1984:APU

[Ano84a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 25(3):604–605, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889997/pdf.

Anonymous:1984:BN

[Ano84b] Anonymous. Book notes. Technology and Culture, 25(1):167–169, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890118/pdf.

Anonymous:1984:DP

[Ano84c] Anonymous. The Dexter prize. Technology and Culture, 25(3):606– 607, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889998/pdf.

Anonymous:1984:IVT

[Ano84d] Anonymous. Index to volume 25 (1984) of Technology and Culture. Technology and Culture, 25(4):913–929, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889989/pdf.

Anonymous:1984:JCR

[Ano84e] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 25(3):602–603, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889996/pdf.

Anonymous:1984:LVM

[Ano84f] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 25(3):608–614, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889999/pdf.

Anonymous:1984:VNJa

[Ano84g] Anonymous. Volume 25, number 1, January 1984. Technology and Culture, 25(1):??, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1984:VNA

[Ano84h] Anonymous. Volume 25, number 2, April 1984. Technology and Culture, 25(2):??, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1984:VNJb

[Ano84i] Anonymous. Volume 25, number 3, July 1984. Technology and Culture, 25(3):??, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1984:VNO

[Ano84j] Anonymous. Volume 25, number 4, October 1984. Technology and Culture, 25(4):??, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1985:APU

[Ano85a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 26(3):581, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889720/pdf.

Anonymous:1985:BN

[Ano85b] Anonymous. Book notes. Technology and Culture, 26(1):142–148, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889843/pdf.

Anonymous:1985:DP

[Ano85c] Anonymous. The Dexter prize. Technology and Culture, 26(3):582– 583, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889721/pdf.

Anonymous:1985:DIM

[Ano85d] Anonymous. Directory of individual members: The Society for the History of Technology. Technology and Culture, 26(1):149–178, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889844/pdf.

Anonymous:1985:IVT

[Ano85e] Anonymous. Index to volume 26 (1985) of Technology and Culture. Technology and Culture, 26(4):891–910, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889715/pdf.

Anonymous:1985:JCR

[Ano85f] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 26(3):579–580, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889719/pdf.

Anonymous:1985:LVM

[Ano85g] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 26(3):584–589, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889722/pdf.

Anonymous:1985:VNJa

[Ano85h] Anonymous. Volume 26, number 1, January 1985. Technology and Culture, 26(1):??, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1985:VNA

[Ano85i] Anonymous. Volume 26, number 2, April 1985. Technology and Culture, 26(2):??, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1985:VNJb

[Ano85j] Anonymous. Volume 26, number 3, July 1985. Technology and Culture, 26(3):??, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1985:VNO

[Ano85k] Anonymous. Volume 26, number 4, October 1985. Technology and Culture, 26(4):??, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1986:APU

[Ano86a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 27(3):563–564, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889537/pdf.

Anonymous:1986:BN

[Ano86b] Anonymous. Book notes. Technology and Culture, 27(4):912–920, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889576/pdf.

Anonymous:1986:DP

[Ano86c] Anonymous. The Dexter prize. Technology and Culture, 27(3):565– 567, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889539/pdf.

Anonymous:1986:ER

[Ano86d] Anonymous. Exhibit reviews. Technology and Culture, 27(3):668, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889609/pdf.

Anonymous:1986:IVT

[Ano86e] Anonymous. Index to volume 27 (1986) of Technology and Culture. Technology and Culture, 27(4):921–939, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889579/pdf.

Anonymous:1986:JCR

[Ano86f] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 27(3):562, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889535/pdf.

Anonymous:1986:LVM

[Ano86g] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 27(3):568–576, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889540/pdf.

Anonymous:1986:LP

[Ano86h] Anonymous. The Levinson prize. Technology and Culture, 27 (1):188, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889673/pdf.

Anonymous:1986:SER

[Ano86i] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 27(3):561, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889533/pdf.

Anonymous:1986:VNJa

[Ano86j] Anonymous. Volume 27, number 1, January 1986. Technology and Culture, 27(1):??, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1986:VNA

[Ano86k] Anonymous. Volume 27, number 2, April 1986. Technology and Culture, 27(2):??, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1986:VNJb

[Ano86l] Anonymous. Volume 27, number 3, July 1986. Technology and Culture, 27(3):??, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1986:VNO

[Ano86m] Anonymous. Volume 27, number 4, October 1986. Technology and Culture, 27(4):??, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1987:APU

[Ano87a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 28(3):611, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889350/pdf.

Anonymous:1987:DP

[Ano87b] Anonymous. The Dexter prize. Technology and Culture, 28(3):612– 614, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889351/pdf.

Anonymous:1987:DIM

[Ano87c] Anonymous. Directory of individual members: The Society for the History of Technology. Technology and Culture, 28(1):189–226, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889491/pdf.

Anonymous:1987:IVT

[Ano87d] Anonymous. Index to volume 28 (1987) of Technology and Culture. Technology and Culture, 28(4):905–922, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889349/pdf.

Anonymous:1987:JCR

[Ano87e] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 28(3):609–610, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889348/pdf.

Anonymous:1987:LVM

[Ano87f] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 28(3):615–618, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889352/pdf.

Anonymous:1987:SER

[Ano87g] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 28(3):608, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889346/pdf.

Anonymous:1987:VNJa

[Ano87h] Anonymous. Volume 28, number 1, January 1987. Technology and Culture, 28(1):??, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1987:VNA

[Ano87i] Anonymous. Volume 28, number 2, April 1987. Technology and Culture, 28(2):??, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1987:VNJb

[Ano87j] Anonymous. Volume 28, number 3, July 1987. Technology and Culture, 28(3):??, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1987:VNO

[Ano87k] Anonymous. Volume 28, number 4, October 1987. Technology and Culture, 28(4):??, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1988:APU

[Ano88a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 29(3):639–640, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889226/pdf.

Anonymous:1988:BN

[Ano88b] Anonymous. Book notes. Technology and Culture, 29(3):713–718, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889272/pdf.

Anonymous:1988:DP

[Ano88c] Anonymous. The Dexter prize. Technology and Culture, 29(3):641– 643, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889227/pdf.

Anonymous:1988:DA

[Ano88d] Anonymous. The dibner award. Technology and Culture, 29 (3):633–634, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889218/pdf.

Anonymous:1988:ILM

[Ano88e] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 29(3):635, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889221/pdf.

Anonymous:1988:IVT

[Ano88f] Anonymous. Index to volume 29 (1988) of Technology and Culture. Technology and Culture, 29(4):995–1013, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889224/pdf.

Anonymous:1988:JCR

[Ano88g] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 29(3):637–638, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889225/pdf.

Anonymous:1988:LVM

[Ano88h] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 29(3):644–652, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889228/pdf.

Anonymous:1988:RNC

[Ano88i] Anonymous. The Raleigh, North Carolina, Meeting, October 29–November 1, 1987. Technology and Culture, 29(3):653–660, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889229/pdf.

Anonymous:1988:SER

[Ano88j] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 29(3):636, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889222/pdf.

Anonymous:1988:VNJa

[Ano88k] Anonymous. Volume 29, number 1, January 1988. Technology and Culture, 29(1):??, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1988:VNA

[Ano88l] Anonymous. Volume 29, number 2, April 1988. Technology and Culture, 29(2):??, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1988:VNJb

[Ano88m] Anonymous. Volume 29, number 3, July 1988. Technology and Culture, 29(3):??, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1988:VNO

[Ano88n] Anonymous. Volume 29, number 4, October 1988. Technology and Culture, 29(4):??, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1989:APU

[Ano89a] Anonymous. The Abbott Payson Usher Prize. Technology and Culture, 30(3):608, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889023/pdf.

Anonymous:1989:DP

[Ano89b] Anonymous. The Dexter prize. Technology and Culture, 30(3):609– 611, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889025/pdf.

Anonymous:1989:DA

[Ano89c] Anonymous. The dibner award. Technology and Culture, 30 (3):601–603, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889016/pdf.

Anonymous:1989:DIM

[Ano89d] Anonymous. Directory of individual members: The Society for the History of Technology. Technology and Culture, 30(1):215–247, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889154/pdf.

Anonymous:1989:ILM

[Ano89e] Anonymous. The IEEE life members prize in electrical history. Technology and Culture, 30(3):604, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889018/pdf.

Anonymous:1989:IVT

[Ano89f] Anonymous. Index to volume 30 (1989) of Technology and Culture. Technology and Culture, 30(4):1101–1116, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889028/pdf.

Anonymous:1989:JCR

[Ano89g] Anonymous. The joan cahalin Robinson prize. Technology and Culture, 30(3):606–607, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889022/pdf.

Anonymous:1989:LVM

[Ano89h] Anonymous. The Leonardo da Vinci Medal. Technology and Culture, 30(3):612–620, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889027/pdf.

Anonymous:1989:SER

[Ano89i] Anonymous. The Samuel Eleazar and Rose Tartakow Levinson Prize. Technology and Culture, 30(3):605, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889020/pdf.

Anonymous:1989:VNJa

[Ano89j] Anonymous. Volume 30, number 1, January 1989. Technology and Culture, 30(1):??, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1989:VNA

[Ano89k] Anonymous. Volume 30, number 2, April 1989. Technology and Culture, 30(2):??, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1989:VNJb

[Ano89l] Anonymous. Volume 30, number 3, July 1989. Technology and Culture, 30(3):??, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1989:VNO

[Ano89m] Anonymous. Volume 30, number 4, October 1989. Technology and Culture, 30(4):??, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:1989:WDM

[Ano89n] Anonymous. The Wilmington, Delaware, Meeting, October 20–23, 1988. Technology and Culture, 30(3):621–627, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889029/pdf.

Argersinger:1988:WWP

[Arg88] Jo Ann E. Argersinger. The I.W.W. and the Paterson Silk Strike of 1913 by Anne H. Tripp (review). Technology and Culture, 29(4):928–929, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889174/pdf.

Arlinghaus:1984:ETM

[Arl84] Sandra Arlinghaus. Early Thematic Mapping in the History of Cartography by Arthur H. Robinson (review). Technology and Culture, 25(4):850–852, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889953/pdf.

Armstrong:1986:EEC

[Arm86] Christopher Armstrong. Engineers & Electrons: A Century of Electrical Progress by John D. Ryder, Donald G. Fink (review). Technology and Culture, 27(4):842–843, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889507/pdf.

Arnold:1989:NAM

[Arn89] Kenneth Arnold. The National Atomic Museum, Albuquerque, New Mexico: Where “Weapon shapes” are not enough. Technology and Culture, 30(3):640–642, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889032/pdf.

Aspray:1984:HBO

[Asp84] William Aspray. History of Binary and Other Nondecimal Numeration by Anton Glaser (review). Technology and Culture, 25 (1):142–144, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890102/pdf.

Atack:1982:EIA

[Ata82] Jeremy Atack. An Emerging Independent American Economy, 1815–1875 ed. by Joseph R. Frese, Jacob Judd (review). Technology and Culture, 23(1):122–123, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891129/pdf.

Atkinson:1989:IJS

[Atk89] D. Scott Atkinson. IA: The Journal of the Society for Industrial Archeology. Vol. 12, no. 2: IA in Art ed. by Betsy Fahlman (review). Technology and Culture, 30(2):460–464, April 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889080/pdf.

Baatz:1987:RCP

[Baa87] Simon Baatz. Robert Cornelius: Portraits from the Dawn of Photography by William F. Stapp (review). Technology and Culture, 28(1):161–162, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889476/pdf.

Baatz:1989:EEB

[Baa89] Simon Baatz. Enterprising Elite: The Boston Associates and the World They Made by Robert F. Dalzell, Jr. (review). Technology and Culture, 30(3):694–695, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889053/pdf.

Bachrach:1982:FDC

[Bac82a] Bernard S. Bachrach. The Fortifications and Defense of Constantinople by Byron Tsangadas (review). Technology and Culture, 23(3):477–479, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891193/pdf.

Bachrach:1982:TOF

[Bac82b] Bernard S. Bachrach. The Three Orders: Feudal Society Imagined by Georges Duby (review). Technology and Culture, 23(4):647–649, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890251/pdf.

Bachrach:1985:SDB

[Bac85a] Bernard S. Bachrach. Seven Decisive Battles of the Middle Ages by Joseph Dahmus (review). Technology and Culture, 26(4):830–831, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889683/pdf.

Bachrach:1985:OWC

[Bac85b] Bernard S. Bachrach. On the origins of William the Conqueror’s horse transports. Technology and Culture, 26(3):505–531, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889713/pdf.

Bachrach:1986:WMA

[Bac86] Bernard S. Bachrach. War in the Middle Ages by Philippe Contamine (review). Technology and Culture, 27(3):609–610, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889551/pdf.

Bahr:1987:ELC

[Bah87] Betsy W. Bahr. Essays from the Lowell Conference on Industrial History 1982 and 1983: The Arts and Industrialism/The Industrial City ed. by Robert Weible (review). Technology and Culture, 28(2):369–371, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889423/pdf.

Bailes:1983:TTS

[Bai83] Kendall E. Bailes. Trade and Technology in Soviet-Western Relations by Philip Hanson (review). Technology and Culture, 24 (1):138–140, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891242/pdf.

Baldwin:1980:WJE

[Bal80] Mark Baldwin. William Jessop, Engineer by Charles Hadfield, A. W. Skempton (review). Technology and Culture, 21(2):246–250, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890684/pdf.

Balzer:1985:PTS

[Bal85] Harley D. Balzer. Politics and Technology in the Soviet Union by Bruce Parrott (review). Technology and Culture, 26(2):301–303, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889779/pdf.

Bartoshesky:1985:MVT

[Bar85] Florence Bartoshesky. The Merrimack Valley Textile Museum: A Guide to the Manuscript Collections by Helena E. Wright (review). Technology and Culture, 26(3):665–666, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889748/pdf.

Barrett:1988:MAC

[Bar88] Paul Barrett. The Motorization of American Cities by David J. St. Clair (review). Technology and Culture, 29(1):169–170, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888526/pdf.

Barrett:1989:AAM

[Bar89a] Paul Barrett. Los Angeles and the Automobile: The Making of the Modern City by Scott L. Bottles (review). Technology and Culture, 30(4):1068–1070, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888998/pdf.

Bartky:1989:AST

[Bar89b] Ian R. Bartky. The adoption of standard time. Technology and Culture, 30(1):25–56, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889112/pdf.

Bartky:1991:CRM

[Bar91] Ian R. Bartky. Comment and response on “The Most Reliable Time”. Technology and Culture, 32(1):183–185, January 1991. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888825/pdf. See

[Ste89].

Battison:1980:AWT

[Bat80] Edwin A. Battison. American Woodworking Tools by Paul B. Kebabian, Dudley Witney (review). Technology and Culture, 21(1): 108–110, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890622/pdf.

Bateman:1982:MHC

[Bat82] Fred Bateman. Managerial Hierarchies: Corporate Perspectives on the Rise of the Modern Industrial Enterprise ed. by Alfred D. Chandler, Herman Daems (review). Technology and Culture, 23(1):140–142, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891140/pdf.

Bauer:1986:TIR

[Bau86] K. Jack Bauer. Transport in the Industrial Revolution ed. by Derek H. Aldcroft, Michael J. Freeman, and: The Manchester Ship Canal by David Owen (review). Technology and Culture, 27(3):623–624, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889567/pdf.

Bayor:1980:SHH

[Bay80] Ronald H. Bayor. A Social History of Housing, 1815–1970 by John Burnett (review). Technology and Culture, 21(2):252–253, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890686/pdf.

Bose:1984:HTS

[BBM84] Christine E. Bose, Philip L. Bereano, and Mary Malloy. Household technology and the social construction of housework. Technology and Culture, 25(1):53–82, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890080/pdf.

Beauchamp:1981:HGW

[Bea81] Gorman Beauchamp. H. G. Wells: Discoverer of the Future/The Influence of Science on His Thought by Roslynn D. Haynes (review). Technology and Culture, 22(2):332–335, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890865/pdf.

Bebarta:1980:RDP

[Beb80] Prafulla C. Bebarta. R & D Planning in the Framework of National Plans by A. Rahman, et al., and: Pattern and Role of Informal Communication in R & D by (Mrs.) S. Pruthi, P. S. Nagpaul (review). Technology and Culture, 21(4):712–714, October 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890792/pdf.

Becker:1986:FFR

[Bec86] John V. Becker. On the Frontier: Flight Research at Dryden, 1946–1981 by Richard P. Hallion (review). Technology and Culture, 27(3):659–661, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889602/pdf.

Bedini:1980:ADM

[Bed80a] Silvio A. Bedini. Annali dell’Istituto e Museo di Storia della Scienza di Firenze ed. by Prof.ssa Dott.ssa Maria Luisa Righini-Bonelli (review). Technology and Culture, 21(1):92–93, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890612/pdf.

Bedini:1980:CTG

[Bed80b] Silvio A. Bedini. Cultures and Time by L. Gardet, et al., and: Richard of Wallingford, an Edition of His Writings with Introductions, English Translation and Commentary by J. D. North, and: Die Deutsche Räderuhr. Zur Kunst und Technik des mechanischen Zeitmessers im deutschen Sprachraum [German clocks. The art and technology of mechanical clocks in German-speaking countries] by Klaus Maurice (review). Technology and Culture, 21(1):87–92, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890611/pdf.

Bedini:1982:MLR

[Bed82] Silvio A. Bedini. Maria Luisa Righini Bonelli (1917–1981). Technology and Culture, 23(3):531–533, July 1982. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891220/pdf.

Bedini:1983:GIB

[Bed83a] Silvio A. Bedini. Geomagnetic Instruments before 1900 by Anita McConnell (review). Technology and Culture, 24(2):267, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891265/pdf.

Bedini:1983:SSL

[Bed83b] Silvio A. Bedini. Surveyors and Statesmen: Land Measuring in Colonial Virginia by Sarah S. Hughes (review). Technology and Culture, 24(2):267–269, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891266/pdf.

Bedini:1983:BTE

[Bed83c] Silvio A. Bedini. The Book of Time ed. by John Grant, Colin Wilson (review). Technology and Culture, 24(1):117, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891229/
pdf.

Bedini:1984:DJS

[Bed84] Silvio A. Bedini. Derek J. de Solla Price (1922–1983). Technology and Culture, 25(3):701–705, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890032/pdf.

Bedini:1985:CES

[Bed85a] Silvio A. Bedini. A Catalogue of European Scientific Instruments in the Department of Medieval and Later Antiquities of the British Museum by F. A. B. Ward (review). Technology and Culture, 26(3):636–637, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889729/pdf.

Bedini:1985:DCS

[Bed85b] Silvio A. Bedini. A Discourse on the Character and Services of Thomas Jefferson, More Especially as a Promoter of Natural and Physical Sciences by Samuel Latham Mitchill (review). Technology and Culture, 26(3):637–638, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889730/pdf.

Bedini:1986:NCS

[Bed86] Silvio A. Bedini. Nineteenth-Century Scientific Instruments by Gerard L’E. Turner (review). Technology and Culture, 27(1): 146–148, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889653/pdf.

Bedini:1987:TMC

[Bed87] Silvio A. Bedini. The Time Museum: Catalogue of the Collection. Vol. 1: Time Measuring Instruments. Pt. 1: Astrolabes/ Astrolabe Related Instruments; pt. 3: WaterClocks/Sand-Glasses/ Fire-Clocks by Anthony J. Turner (review). Technology and Culture, 28(1):159–161, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889475/pdf.

Bedini:1989:BD

[Bed89a] Silvio A. Bedini. Bern Dibner (1897–1988). Technology and Culture, 30(1):189–193, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889152/pdf.

Bedini:1989:CMP

[Bed89b] Silvio A. Bedini. The Correspondence and Miscellaneous Papers of Benjamin Henry Latrobe. Vol. 2: 1805–1810 ed. by John C. Van Horne, et al. (review). Technology and Culture, 30(3):687–689, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889049/pdf.

Bedini:1989:HHR

[Bed89c] Silvio A. Bedini. The Hall of Heavenly Records: Korean Astronomical Instruments and Clocks 1380–1780 by Joseph Needham, et al. (review). Technology and Culture, 30(3):678–680, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889044/pdf.

Beezley:1983:GAD

[Bee83] William H. Beezley. Growth against Development: The Economic Impact of Railroads in Porfirian Mexico by John H. Coatsworth (review). Technology and Culture, 24(1):135–137, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891240/
pdf.

Beer:1985:HCT

[Bee85] John J. Beer. The History of Chemical Technology: An Annotated Bibliography by Robert P. Multhauf (review). Technology and Culture, 26(3):661–662, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889745/pdf.

Belasco:1982:HWM

[Bel82] Warren Belasco. Homes on Wheels by Michael Aaron Rockland (review). Technology and Culture, 23(4):682–684, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890271/
pdf.

Belanger:1984:CP

[Bel84a] Dian O. Belanger. The Corliss at Pullman. Technology and Culture, 25(1):83–90, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890081/pdf.

Belasco:1984:ARI

[Bel84b] Warren Belasco. The Automobile Revolution: The Impact of an Industry by Jean-Pierre Bardou, et al. (review). Technology and Culture, 25(4):881–882, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889972/pdf.

Belmore:1984:BEE

[Bel84c] Karen Belmore. Bottled Energy: Electrical Engineering and the Evolution of Chemical Energy Storage by Richard H. Schallenberg (review). Technology and Culture, 25(4):885–887, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889975/
pdf.

Benhamou:1984:VE

[Ben84] Reed Benhamou. Verdigris and the entrepreneuse. Technology and Culture, 25(2):171–181, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890034/pdf.

Benhamou:1986:RMC

[Ben86] Reed Benhamou. The Rise of Market Culture: The Textile Trade and French Society, 1750–1900 by William M. Reddy (review). Technology and Culture, 27(2):298–299, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889601/pdf.

Benhamou:1988:SWA

[Ben88] Reed Benhamou. Spinners and Weavers of Auffay: Rural Industry and the Sexual Division of Labor in a French Village, 1750– 1850 by Gay L. Gullickson (review). Technology and Culture, 29(4):934–935, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889178/pdf.

Berger:1980:PSA

[Ber80] Arthur Asa Berger. Proceedings of the Sixth Annual Telecommunications Policy Research Conference ed. by Herbert Dordick (review). Technology and Culture, 21(4):708–709, October 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890790/pdf.

Berger:1983:AST

[Ber83] Michael L. Berger. The American Small Town: Twentieth-Century Place Images by John A. Jakle (review). Technology and Culture, 24(2):279–280, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891274/pdf.

Berger:1984:ITP

[Ber84] Michael L. Berger. Information Technology and Psychology: Prospects for the Future ed. by Richard A. Kasschau, Roy Lachman, and Kenneth R. Laughery (review). Technology and Culture, 25(4):849–850, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889952/pdf.

Berenson:1985:MRF

[Ber85] Edward Berenson. The Modernization of Rural France: Communications Networks and Agricultural Market Structures in Nineteenth-Century France by Roger Price (review). Technology and Culture, 26(3):645–647, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889735/pdf.

Berch:1987:SIT

[Ber87] Bettina Berch. Smothered by Invention: Technology in Women’s Lives ed. by Wendy Faulkner, Erik Arnold (review). Technology and Culture, 28(1):135–136, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889460/pdf.

Berg:1989:MWM

[Ber89] Maxine Berg. Most Wonderful Machine: Mechanization and Social Change in Berkshire Paper Making 1801–1885 by Judith A. Mc-Gaw (review). Technology and Culture, 30(3):692–693, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889052/
pdf.

Bezilla:1983:NWR

[Bez83] Michael Bezilla. No Way to Run a Railroad: The Untold, Story of the Penn Central Crisis by Stephen Salsbury (review). Technology and Culture, 24(2):291–293, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891280/pdf.

Bezilla:1985:RIT

[Bez85] Michael Bezilla. Rochester Institute of Technology: Industrial Development and Educational Innovation in an American City by Dane R. Gordon (review). Technology and Culture, 26(4):859–860, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889698/pdf.

Bhathal:1987:BSL

[Bha87] Ragbir Bhathal. “A Bright and Savage Land”: Scientists in Colonial Australia by Ann Moyal (review). Technology and Culture, 28(4):889–891, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889335/pdf.

Biggs:1988:CAU

[Big88] Lindy Biggs. A Concrete Atlantis: U.S. Industrial Building and European Modern Architecture by Reyner Banham (review). Technology and Culture, 29(2):316–317, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889287/pdf.

Bijker:1986:NTK

[Bij86] Wiebe E. Bijker. The Nature of Technological Knowledge: Are Models of Scientific Change Relevant ed. by Rachel Laudan (review). Technology and Culture, 27(2):336–338, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889628/pdf.

Bijker:1988:SAH

[Bij88] Wiebe E. Bijker. Science in Action: How to Follow Scientists and Engineers through Society by Bruno Latour (review). Technology and Culture, 29(4):982–983, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889213/pdf.

Bilstein:1980:BBF

[Bil80a] Roger E. Bilstein. Bonfires to Beacons: Federal Civil Aviation Policy under the Air Commerce Act, 1926–1938 by Nick A. Komons, and: Takeoff at Mid-Century: Federal Civil Aviation Policy in the Eisenhower Years, 1953–1961 by Stuart I. Rochester (review). Technology and Culture, 21(3):525–528, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890736/pdf.

Bilstein:1980:IFE

[Bil80b] Roger E. Bilstein. Introduction to Flight: Its Engineering and History by John D. Anderson, Jr. (review). Technology and Culture, 21(3):524–525, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890735/pdf.

Billington:1981:SCH

[Bil81a] David P. Billington. Solomon Cady Hollister Colloquium: Perspectives on the History of Reinforced Concrete in the United States, 1904–1941: Princeton University, June 2, 1980. Technology and Culture, 22(4):757–759, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890939/pdf.

Bilstein:1981:JAF

[Bil81b] Roger E. Bilstein. The Jet Age: Forty Years of Jet Aviation ed. by Walter J. Boyne, Donald S. Lopez (review). Technology and Culture, 22(1):200–201, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890828/pdf.

Bilstein:1982:UCA

[Bil82] Roger E. Bilstein. U.S. Civil Aircraft by Joseph P. Juptner (review). Technology and Culture, 23(2):272–274, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891174/pdf.

Billington:1983:BDH

[Bil83] David P. Billington. Bridges, Docks and Harbours with Related Works by Julia Elton (review). Technology and Culture, 24(4): 710–712, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890149/pdf.

Bilstein:1985:SGJ

[Bil85a] Roger E. Bilstein. The Sporty Game by John Newhouse (review). Technology and Culture, 26(3):669–671, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889751/pdf.

Bilstein:1985:WWS

[Bil85b] Roger E. Bilstein. Winged Wonders: The Story of the Flying Wings by E. T. Wooldridge (review). Technology and Culture, 26(4):869–870, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889704/pdf.

Bilstein:1987:AHA

[Bil87a] Roger E. Bilstein. Aerospace historians, aerospace enthusiasts. Technology and Culture, 28(1):124–125, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889456/pdf.

Bilstein:1987:PSA

[Bil87b] Roger E. Bilstein. Prelude to the Space Age: The Rocket Societies, 1924–1940 by Frank H. Winter (review). Technology and Culture, 28(2):381–382, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889431/pdf.

Bilstein:1987:RFA

[Bil87c] Roger E. Bilstein. Rise of the Fighter Aircraft 1914–1918 by Richard P. Hallion (review). Technology and Culture, 28(3):737– 738, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889397/pdf.

Bilstein:1988:CFH

[Bil88] Roger E. Bilstein. City of Flight: The History of Aviation in St. Louis by James J. Horgan (review). Technology and Culture, 29(3):707–708, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889265/pdf.

Bindon:1983:LCU

[Bin83] George Bindon. Liberation and Control: The Uses of Knowledge and Power ed. by David Wade Chambers, and: Whither Technology? ed. by Lyndsay Farrall (review). Technology and Culture, 24(3):548–550, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891427/pdf.

Bindocci:1987:MDG

[Bin87] Cynthia Gay Bindocci. The Miners of Decazeville: A Genealogy of Deindustrialization by Donald Reid (review). Technology and Culture, 28(2):362–363, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889418/pdf.

Bindocci:1989:DOO

[Bin89] Gay Bindocci. Digging Our Own Graves: Coal Miners and the Struggle over Black Lung Disease by Barbara Ellen Smith (review). Technology and Culture, 30(3):700–701, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889061/pdf.

Birr:1981:SFY

[Bir81] Kendall Birr. SRI: The Founding Years by Weldon B. Gibson (review). Technology and Culture, 22(2):347–349, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890871/pdf.

Bird:1983:SCJ

[Bir83] William L. Bird, Jr. A suggestion concerning James Smithson’s concept of “Increase and Diffusion”. Technology and Culture, 24 (2):246–255, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891257/pdf.

Bird:1987:ARR

[Bir87] William L. Bird, Jr. Audience Ratings: Radio, Television, Cable by Hugh Malcolm Beville, Jr. (review). Technology and Culture, 28(3):704–706, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889373/pdf.

Blackford:1981:SBA

[Bla81] Mansel Blackford. Small Business in American Life ed. by Stuart Bruchey (review). Technology and Culture, 22(4):800–802, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890955/pdf.

Blake:1984:GWH

[Bla84] Nelson M. Blake. Great Waters: A History of Boston’s Water Supply by Fern L. Nesson (review). Technology and Culture, 25(2):358– 360, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890068/pdf.

Blackford:1987:AEF

[Bla87] Mansel Blackford. American Enterprise in Foreign Markets: Studies of Singer and International Harvester in Imperial Russia by Fred V. Carstensen (review). Technology and Culture, 28(4): 872–874, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889324/pdf.

Blake:1988:LPG

[Bla88] Nelson M. Blake. The Limits of Power: Great Fires and the Process of City Growth in America by Christine M. Rosen (review). Technology and Culture, 29(1):167–168, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888525/pdf.

Blake:1989:FRP

[Bla89] Nelson M. Blake. Flagler: Rockefeller Partner and Florida Baron by Edward N. Akin (review). Technology and Culture, 30(2):467– 468, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889086/pdf.

Blew:1983:WPC

[Ble83] Robert W. Blew. Water and Power: The Conflict over Los Angeles’ Water Supply in the Owens Valley by William L. Kahrl (review). Technology and Culture, 24(3):538–539, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891419/pdf.

Blicksilver:1982:RHT

[Bli82] Jack Blicksilver. R. H. Tawney: The American Labour Movement and Other Essays ed. by J. M. Winter (review). Technology and Culture, 23(3):493–495, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891201/pdf.

Bloombecker:1981:DCL

[Blo81] Jay Bloombecker. A Decade of Computers and Law ed. by Jon Bing, Knut Selmer (review). Technology and Culture, 22(4): 820–822, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890964/pdf.

Blustein:1982:UGC

[Blu82] Bonnie Ellen Blustein. Urban Growth and City-Systems in the United States, 1840–1860 by Allan Pred (review). Technology and Culture, 23(3):503–504, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891208/pdf.

Bluestone:1986:MEA

[Blu86] Daniel M. Bluestone. Metropolis 1890–1940 ed. by Anthony Sutcliffe (review). Technology and Culture, 27(2):329–330, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889624/
pdf.

Blyth:1988:GSJ

[Bly88] John D. Blyth. Gresley and Stanier by John Bellwood, David Jenkinson (review). Technology and Culture, 29(4):954–955, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889190/pdf.

Billington:1984:CBS

[BM84a] David P. Billington and Robert Mark. The cathedral and the bridge: Structure and symbol. Technology and Culture, 25(1): 37–52, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890079/pdf.

Billington:1984:RAV

[BM84b] David P. Billington and Robert Mark. In response to “Another view of ‘the cathedral and the bridge”’. Technology and Culture, 25(3):595–601, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889995/pdf.

Berrett:1985:FDN

[BM85] Joshua Berrett and Virginia Marquardt. The Fourth Dimension and Non-Euclidean Geometry in Modern Art by Linda Dalrymple Henderson (review). Technology and Culture, 26(4):879–882, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889709/pdf.

Bogue:1985:HAO

[Bog85] Allan G. Bogue. History of Agriculture in Ohio to 1880 by Robert Leslie Jones (review). Technology and Culture, 26(1):100–101, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889822/pdf.

Bolter:1988:BGG

[Bol88] Jay David Bolter. Beyond the Gutenberg Galaxy: Microcomputers and the Emergence of Post-Typographic Culture by Eugene F. Provenzo, Jr. (review). Technology and Culture, 29(2):323–324, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889292/pdf.

Bostock:1985:HBP

[Bos85] Frances Bostock. The History of the British Petroleum Company. Vol. 1: The Developing Years 1901–1932 by R. W. Ferrier (review). Technology and Culture, 26(4):846–848, October 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889691/pdf.

Botein:1980:CGF

[Bot80] Stephen Botein. Chronicles of Genius and Folly: R. Hoe & Company and the Printing Press as a Service to Democracy by Frank E. Comparato (review). Technology and Culture, 21(4):675–677, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890774/pdf.

Bowers:1980:BEI

[Bow80] Brian Bowers. The British Electrical Industry, 1875–1914: The Economic Returns of a New Technology by I. C. R. Byatt (review). Technology and Culture, 21(4):664–666, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890769/pdf.

Bowditch:1987:FHA

[Bow87] John Bowditch. Ferracute: The History of an American Enterprise by Arthur J. Cox, Thomas Malim (review). Technology and Culture, 28(4):869–871, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889322/pdf.

Boyer:1985:RTI

[Boy85] Marjorie Nice Boyer. Resistance to technological innovation: The history of the pile driver through the 18th century. Technology and Culture, 26(1):56–68, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889808/pdf.

Boyer:1986:NST

[Boy86] Paul Boyer. No Sacrifice Too Great: The Life of Lewis L. Strauss by Richard Pfau (review). Technology and Culture, 27(4):902–903, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889564/pdf.

Boyer:1988:BCB

[Boy88] Marjorie Nice Boyer. Building Construction before Mechanization by John Fitchen (review). Technology and Culture, 29(3):677–679, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889238/pdf.

Boyer:1989:NHA

[Boy89] Paul Boyer. Nuclear Holocausts: Atomic War in Fiction, 1895– 1984 by Paul Brians (review). Technology and Culture, 30(1): 131–133, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889121/pdf.

Braun:1980:IFS

[Bra80a] Hans-Joachim Braun. Ingenieure in der Frühindustrialisierung. Staatliche und private Techniker im Königreich Hannover und an der Ruhr (1815–1873) by Lars Ulrich Scholl (review). Technology and Culture, 21(2):250–252, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890685/pdf.

Bray:1980:ECN

[Bra80b] Francesca Bray. Economic Consequences of the New Rice Technology by the International Rice Research Institute, and: Anatomy of a Peasant Economy: A Rice Village in the Philippines by Y. Hayami, et al. (review). Technology and Culture, 21(3):547–548, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890746/pdf.

Brainerd:1981:CEP

[Bra81a] John G. Brainerd. Chinese Economy — Post Mao. Vol. 1: Policy and Performance by the Staff of the Joint Economic Committee (review). Technology and Culture, 22(3):686–688, July 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890931/pdf.

Brainerd:1981:EFR

[Bra81b] John G. Brainerd. Energy Future: Report of the Energy Project at the Harvard Business School ed. by Robert Stobaugh, Daniel Yergin, and: The Last Chance Energy Book by Owen Phillips (review). Technology and Culture, 22(1):217–219, January 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890836/pdf.

Braun:1982:TDB

[Bra82a] Hans J. Braun. Technische Denkmäler in der Bundesrepublik Deutschland. Vol. 3: Die Kali-und Steinsalzindustrie by Rainer Slotta (review). Technology and Culture, 23(4):659–660, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890257/pdf.

Braun:1982:ATF

[Bra82b] Hans-Joachim Braun. Analytische Technikphilosophie by Friedrich Rapp (review). Technology and Culture, 23(1):88–89, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891109/pdf.

Braun:1982:ZKG

[Bra82c] Hans-Joachim Braun. Zeit und Kultur. Geschichte des Zeitbewuß tseins in Europa by Rudolf Wendorff (review). Technology and Culture, 23(2):229–230, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891151/pdf.

Brainerd:1983:HCT

[Bra83a] John G. Brainerd. A History of Computing in the Twentieth Century ed. by N. Metropolis; J. Howlett; Gian-Carlo Rota (review). Technology and Culture, 24(1):147–148, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891247/pdf.

Braun:1983:WGB

[Bra83b] Hans-Joachim Braun. Wissenschaft und Gesellschaft: Beiträge zur Geschichte der Technischen Universität Berlin, 1879–1979 ed. by Reinhard Rürup (review). Technology and Culture, 24(4):717–719, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890154/pdf.

Braynard:1984:EJL

[Bra84] Frank O. Braynard. The Era of the Joy Line: A Saga of Steamboating on Long Island Sound by Edwin L. Dunbaugh (review). Technology and Culture, 25(4):880, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889971/pdf.

Brickman:1980:SPS

[Bri80a] Ronald Brickman. Scientists in Power by Spencer R. Weart (review). Technology and Culture, 21(2):268–271, April 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890695/pdf.

Bright:1980:DHA

[Bri80b] Charles D. Bright. Delta: The History of an Airline by W. David Lewis, Wesley Phillips Newton (review). Technology and Culture, 21(3):528–529, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890737/pdf.

Bright:1981:APW

[Bri81a] Charles D. Bright. Air Power and Warfare: The Proceedings of the Eighth Military History Symposium, 18–20 October 1978 ed. by Alfred F. Hurley, Robert C. Ehrhart (review). Technology and Culture, 22(1):198–199, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890827/pdf.

Brittain:1981:PJH

[Bri81b] James E. Brittain. The Papers of Joseph Henry. Vol. 3: January 1836–December 1837: The Princeton Years ed. by Nathan Reingold, et al. (review). Technology and Culture, 22(2):327–329, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890863/pdf.

Brittain:1981:SAC

[Bri81c] James E. Brittain. The Sciences in the American Context: New Perspectives ed. by Nathan Reingold (review). Technology and Culture, 22(2):325–327, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890862/pdf.

Brindley:1985:SLA

[Bri85] Thomas Brindley. Sketches from Life: The Autobiography of Lewis Mumford, the Early Years by Lewis Mumford (review). Technology and Culture, 26(2):317–319, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889787/pdf.

Briggs:1989:TBG

[Bri89] Asa Briggs. The Transformation of Britain 1830–1939 by G. E. Mingay (review). Technology and Culture, 30(3):677–678, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889043/pdf.

Brown:1980:HWE

[Bro80] D. Clayton Brown. A History of the Waterways Experiment Station, 1929–1979 by Gordon Cotton (review). Technology and Culture, 21(3):519–521, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890732/pdf.

Bromberg:1983:SPC

[Bro83] Joan Bromberg. Science, Politics, and Controversy: Civilian Nuclear Power in the United States, 1946–1974 by Steven L. Del Sesto (review). Technology and Culture, 24(3):527–528, July 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891412/pdf.

Brown:1984:EPI

[Bro84] Shannon R. Brown. The Economics of Production and Innovation: An Industrial Perspective by Gerhard Rosegger (review). Technology and Culture, 25(1):149–150, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890106/pdf.

Bromberg:1985:MLH

[Bro85a] Joan Bromberg. Masers and Lasers: An Historical Approach by Mario Bertolotti (review). Technology and Culture, 26(1):134–135, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889839/pdf.

Brose:1985:COG

[Bro85b] Eric Dorn Brose. Competitiveness and obsolescence in the German charcoal iron industry. Technology and Culture, 26(3):532–559, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889716/pdf.

Bromberg:1986:EKL

[Bro86a] Joan Lisa Bromberg. Engineering knowledge in the laser field. Technology and Culture, 27(4):798–818, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889499/pdf.

Bronstein:1986:RIL

[Bro86b] Daniel A. Bronstein. Research, International Law and the Sea in Man’s Future. Impact of Science on Society, and: Law and Science in Collaboration: Resolving Regulatory Issues of Science and Technology ed. by J. D. Nyhart, Milton M. Carrow (review). Technology and Culture, 27(1):185–187, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889672/pdf.

Brownell:1986:ISS

[Bro86c] Blaine A. Brownell. Industrialization and Southern Society 1877– 1984 by James C. Cobb (review). Technology and Culture, 27 (2):322–324, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889621/pdf.

Bromberg:1987:MAI

[Bro87a] Joan Bromberg. The Making of American Industrial Research: Science and Business at GE and Bell, 1876–1926 by Leonard S. Reich (review). Technology and Culture, 28(4):874–875, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889325/pdf.

Brown:1987:ADS

[Bro87b] John Brown. Aborted Discovery: Science and Creativity in the Third World by Susantha Goonatilake (review). Technology and Culture, 28(4):899–901, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889344/pdf.

Brody:1988:SCB

[Bro88a] Judit Brody. The Széchenyi chain bridge at Budapest. Technology and Culture, 29(1):104–117, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888500/pdf.

Bromberg:1988:RVB

[Bro88b] Joan Bromberg. RCA and the VideoDisc: The Business of Research by Margaret B. W. Graham (review). Technology and Culture, 29(2):324–326, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889293/pdf.

Brush:1985:LBM

[Bru85] Stephen G. Brush. Ludwig Boltzmann: Man — Physicist — Philosopher by Engelbert Broda (review). Technology and Culture, 26(3):659–661, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889744/pdf.

Brusa:1988:CCB

[Bru88] Giuseppe Brusa. Clockwork Cosmos: Bernardo Facini and the Farnese Planisferologio by Silvio A. Bedini (review). Technology and Culture, 29(2):289–291, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889266/pdf.

Bryant:1980:STS

[Bry80] John Bryant. Systems theory, survival, and the back-to-nature movement. Technology and Culture, 21(2):227–230, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890676/
pdf.

Bryden:1981:IBS

[Bry81] D. J. Bryden. The Iron Bridge, Symbol of the Industrial Revolution by Neil Cossons, Barrie Trinder, and: A View from the Iron Bridge by Stuart Smith, and: Iron Bridge to Crystal Palace: Impact and Images of the Industrial Revolution by Asa Briggs (review). Technology and Culture, 22(2):306–310, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890852/pdf.

Bryant:1982:BNM

[Bry82] John Bryant. The back-to-nature movement, modern technology, and the inversion of Dr. De gregori. Technology and Culture, 23(2):218–222, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891148/pdf.

Buchanan:1981:CJB

[Buc81a] R. A. Buchanan. Connections by James Burke (review). Technology and Culture, 22(3):660–663, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890922/pdf.

Buchanan:1981:TTE

[Buc81b] R. A. Buchanan. Thomas Telford: Engineer ed. by Alastair Penfold (review). Technology and Culture, 22(4):792–794, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890951/pdf.

Buchanan:1982:HTT

[Buc82a] R. A. Buchanan. A History of Technology: The Twentieth Century, c. 1900 to c. 1950, Parts I and II, vols. 6 and 7 ed. by Trevor I. Williams (review). Technology and Culture, 23(4):631–638, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890244/pdf.

Buchanan:1982:GIS

[Buc82b] R. A. Buchanan. The Greatest Iron Ship: S.S. Great Eastern by George S. Emmerson (review). Technology and Culture, 23(3):510– 512, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/

891212/pdf. See comments [Buc83b, Emm83b].

Buchanan:1982:IAF

[Buc82c] R. A. Buchanan. The Industrial Archaeology of Farming in England and Wales by Nigel Harvey (review). Technology and Culture, 23(2):242–244, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891158/pdf.

Buchanan:1983:IAN

[Buc83a] R. A. Buchanan. The Industrial Archaeology of Northern Ireland by W. A. McCutcheon (review). Technology and Culture, 24(4):673–675, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890128/pdf.

Buchanan:1983:GEC

[Buc83b] R. A. Buchanan. The Great Eastern controversy: A comment. Technology and Culture, 24(1):98–106, January 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891225/pdf. See

[Buc82b].

Buchanan:1983:ISS

[Buc83c] R. A. Buchanan. ICOHTEC Symposium: Smolenice, Czechoslovakia, June 7–12, 1982. Technology and Culture, 24(3):495–497, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891393/pdf.

Buchanan:1985:EHT

[Buc85] R. A. Buchanan. Energy in history: The topicality of the history of technology — ICOHTEC Symposium, Lerbach, Federal Republic of Germany, September 2–7, 1984. Technology and Culture, 26(3):629–630, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889726/pdf.

Buchanan:1986:SBD

[Buc86a] R. A. Buchanan. Stephensons’ Britain by Derrick Beckett (review). Technology and Culture, 27(2):300–301, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889603/pdf.

Buchanan:1986:HAP

[Buc86b] R. A. Buchanan. The History and Archaeology of Ports by Gordon Jackson (review). Technology and Culture, 27(1):130–131, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889644/pdf.

Buchanan:1986:DBE

[Buc86c] R. A. Buchanan. The diaspora of British engineering. Technology and Culture, 27(3):501–524, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889526/pdf.

Bucholz:1986:PWS

[Buc86d] Arden Bucholz. The Patterns of War since the Eighteenth Century by Larry H. Addington (review). Technology and Culture, 27(3):611–612, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889552/pdf.

Buchanan:1988:SCC

[Buc88a] R. A. Buchanan. Scientific Colonialism: A Cross-Cultural Comparison ed. by Nathan Reingold, Marc Rothenberg (review). Technology and Culture, 29(3):685–686, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889242/pdf.

Buchanan:1988:SPT

[Buc88b] R. Angus Buchanan. The Social Production of Technical Work: The Case of British Engineers by Peter Whalley (review). Technology and Culture, 29(1):179–180, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888533/pdf.

Buchanan:1989:EBS

[Buc89] R. Angus Buchanan. Engineers in Britain: A Sociological Study of the Engineering Dimension by Ian A. Glover, Michael P. Kelly (review). Technology and Culture, 30(3):704–705, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889066/pdf.

Bud:1982:EGI

[Bud82] Robert Bud. Evolution of the Gas Industry by Malcolm W. H. Peebles (review). Technology and Culture, 23(4):660–661, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890258/pdf.

Bud:1984:PPM

[Bud84] Robert Bud. Pioneer Plastic: The Making and Selling of Celluloid by Robert Friedel (review). Technology and Culture, 25(3):674, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890018/pdf.

Bulferetti:1981:MSI

[Bul81a] Luigi Bulferetti. Macchine scuola industria dal mestiere alla professionalità ed. by the Comune di Bologna (review). Technology and Culture, 22(3):616–618, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890899/pdf.

Bullough:1981:GCS

[Bul81b] Vern L. Bullough. A Guide to the Culture of Science, Technology, and Medicine ed. by Paul T. Durbin (review). Technology and Culture, 22(3):611–612, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890897/pdf.

Bullough:1981:BNR

[Bul81c] Vern L. Bullough. A brief note on rubber technology and contraception: The diaphragm and the condom. Technology and Culture, 22(1):104–111, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890799/pdf.

Bullough:1981:IVD

[Bul81d] Vern L. Bullough. Infection and venereal disease. Technology and Culture, 22(4):853, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890980/pdf.

Bullough:1982:MMS

[Bul82] Vern L. Bullough. Medieval Medicus: A Social History of Anglo-Norman Medicine by Edward J. Kealey (review). Technology and Culture, 23(4):646–647, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890250/pdf.

Bulliet:1986:EWT

[Bul86a] Richard W. Bulliet. The Earliest Wheeled Transport: From the Atlantic Coast to the Caspian Sea by Stuart Piggott (review). Technology and Culture, 27(4):853–854, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889513/pdf.

Bullough:1986:MST

[Bul86b] Vern L. Bullough. Medieval Science and Technology: A Selected, Annotated Bibliography by Claudia Kren (review). Technology and Culture, 27(2):295, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889595/pdf.

Bullough:1987:TPM

[Bul87] Vern L. Bullough. Technology for the prevention of “les maladies produites par la masturbation”. Technology and Culture, 28(4):828–832, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889304/pdf.

Bulow:1988:GFD

[Bül88] Ralf Bülow. Glory and Failure: The Difference Engines of Johann Muller, Charles Babbage and Georg and Edvard Scheutz by Michael Lindgren (review). Technology and Culture, 29(4):956– 958, October 1988. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL http://www.jstor.org/stable/
3105078;   https://muse.jhu.edu/pub/1/article/889191/pdf.

Bulliet:1989:ARW

[Bul89] Richard W. Bulliet. Achse, Rad und Wagen: Fünftausend Jahre Kultur- und Technikgeschichte ed. by Wilhelm Treue (review). Technology and Culture, 30(4):1039–1041, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888979/pdf.

Burke:1980:TSC

[Bur80a] John G. Burke. Technological and Social Change: A Transdisciplinary Model by Jacob Fried, Paul Molnar (review). Technology and Culture, 21(3):540–543, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890742/pdf.

Burke:1980:WEM

[Bur80b] Melvin Burke. We Eat the Mines and the Mines Eat Us: Dependency and Exploitation in Bolivian Tin Mines by June Nash (review). Technology and Culture, 21(3):545–546, July 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890745/pdf.

Burke:1982:ASS

[Bur82] John G. Burke. From Art to Science: Seventy-two Objects Illustrating the Nature of Discovery by Cyril Stanley Smith, and: A Search for Structure: Selected Essays on Science, Art, and History by Cyril Stanley Smith (breview). Technology and Culture, 23(3):470–473, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891190/pdf.

Burlock:1983:FCE

[Bur83] Joseph Burlock. Frontiers of Change: Early Industrialization in America by Thomas C. Cochran (review). Technology and Culture, 24(2):270–272, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891268/pdf.

Burney:1986:EMU

[Bur86] John M. Burney. The Emergence of Modern Universities in France, 1863–1914 by George Weisz (review). Technology and Culture, 27(4):860–861, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889518/pdf.

Burke:1987:RSB

[Bur87] John G. Burke. Revolution in Science by I. Bernard Gohen (review). Technology and Culture, 28(4):901–903, October 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889347/pdf.

Busch:1983:CCT

[Bus83] Jane Busch. Cooking competition: Technology on the domestic market in the 1930s. Technology and Culture, 24(2):222–245, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891256/pdf.

Bushman:1987:EBA

[Bus87] Richard L. Bushman. The Economy of British America, 1607– 1789 by John J. McCusker, Russell R. Menard (review). Technology and Culture, 28(2):357–359, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889415/pdf.

Butrica:1987:FIC

[But87] Andrew J. Butrica. The first international colloquium on the history of electricity — Paris, April 15–17, 1986. Technology and Culture, 28(2):337–339, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889404/pdf.

Cain:1982:AEH

[Cai82a] Louis P. Cain. Those Army Engineers: A History of the Chicago District, U.S. Army Corps of Engineers by John W. Larson (review). Technology and Culture, 23(1):126, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891132/pdf.

Cain:1982:REH

[Cai82b] P. J. Cain. Research in Economic History ed. by P. J. Uselding (review). Technology and Culture, 23(1):142–144, January 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891141/pdf.

Cain:1988:RTC

[Cai88] P. J. Cain. The Railway in Town and Country 1830–1914 by Jack Simmons (review). Technology and Culture, 29(2):308–310, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889282/pdf.

Callahan:1981:RRM

[Cal81] Raymond Callahan. Railways of the Raj by Michael Satow, Ray Desmond (review). Technology and Culture, 22(4):811–812, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890961/pdf.

Caldwell:1985:SEA

[Cal85] Lynton K. Caldwell. Shaping Environmental Awareness: The United States Army Corps of Engineers Environmental Advisory Board 1970–1980 by Martin Reuss (review). Technology and Culture, 26(2):332–334, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889795/pdf.

Campbell:1985:RSG

[Cam85] John Campbell. Rutherford: Simple Genius by David Wilson (review). Technology and Culture, 26(1):129–130, January 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889836/pdf.

Campbell:1988:TPS

[Cam88] Robert Campbell. Technical Progress and Soviet Economic Development ed. by Ronald Amann, Julian Cooper (review). Technology and Culture, 29(1):181–183, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888535/pdf.

Cameron:1989:TPT

[Cam89] Rondo Cameron. The Tentacles of Progress: Technology Transfer in the Age of Imperialism, 1850–1940 by Daniel R. Headrick (review). Technology and Culture, 30(1):166–167, January 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889143/pdf.

Cardwell:1980:SPB

[Car80] Donald S. L. Cardwell. Steam Power and British Industrialisation to 1860 by G. N. von Tunzelmann (review). Technology and Culture, 21(3):497–499, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890721/pdf.

Cardwell:1981:JPJ

[Car81] Donald S. L. Cardwell. James Prescott Joule and the Concept of Energy by Henry John Steffens (review). Technology and Culture, 22(2):321–322, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890860/pdf.

Cardwell:1982:HTF

[Car82a] Donald Cardwell. History of Technology, Fifth Annual Volume ed. by A. Rupert Hall, Norman Smith (review). Technology and Culture, 23(2):234–235, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891154/pdf.

Cardwell:1982:LVM

[Car82b] Donald S. L. Cardwell. The Leonardo da Vinci Medal. Technology and Culture, 23(3):443–448, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891188/pdf.

Carlson:1982:HCE

[Car82c] W. Bernard Carlson. A History of Control Engineering, 1800– 1930 by Stuart Bennett (review). Technology and Culture, 23(4): 657–658, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890256/pdf.

Carlson:1983:TMT

[Car83a] W. Bernard Carlson. Tesla: Man out of Time by Margaret Cheney (review). Technology and Culture, 24(1):140–142, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891243/
pdf.

Carr:1983:PSI

[Car83b] R. J. M. Carr. The Power of Steam: An Illustrated History of the World’s Steam Age by Asa Briggs (review). Technology and Culture, 24(3):514, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891403/pdf.

Carlson:1984:FBL

[Car84a] W. Bernard Carlson. Ferdinand Braun: A Life of the Nobel Prizewinner and Inventor of the Cathode-Ray Oscilloscope by Friedrich Kurylo, Charles Susskind (review). Technology and Culture, 25(1):116–118, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890088/pdf.

Carr:1984:EVW

[Car84b] R. J. M. Carr. Early Victorian Water Engineers by Geoffrey Morse Binnie (review). Technology and Culture, 25(4):895–896, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889980/pdf.

Carlson:1985:HTS

[Car85] W. Bernard Carlson. The history of technology sessions at the American Historical Association Centennial Meeting, Chicago, December 27–30, 1984. Technology and Culture, 26(4):815–823, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889679/pdf.

Carlson:1986:CEP

[Car86] W. Bernard Carlson. “A century of electricals” — a poster exhibit celebrating the IEEE’s centennial. Technology and Culture, 27(4):828–831, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889501/pdf.

Carlson:1988:AEE

[Car88] W. Bernard Carlson. Academic entrepreneurship and engineering education: Dugald C. Jackson and the MIT-GE cooperative engineering course, 1907–1932. Technology and Culture, 29(3):536–567, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889204/pdf.

Carlson:1989:HVP

[Car89] W. Bernard Carlson. The History of N. V. Philips’ Gloeilampenfabrieken. Vol. 1: The Origin of the Dutch Incandescent Lamp Industry by A. Heerding (review). Technology and Culture, 30 (1):162–164, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889141/pdf.

Cass:1985:AED

[Cas85] Edward Cass. Army Engineers and the Development of Oregon: A History of the Portland District U.S. Army Corps of Engineers by William F. Willingham (review). Technology and Culture, 26 (2):334–335, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889796/pdf.

Cass:1986:FED

[Cas86] Ed Cass. The Federal Engineer/Damsites to Missile Sites: A History of the Omaha District U.S. Army Corps of Engineers (review). Technology and Culture, 27(4):889–890, October 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889548/pdf.

Casey:1988:CBB

[Cas88a] Robert Casey. Crisis in Bethlehem: Big Steel’s Battle to Survive by John Strohmeyer, and: From Fire to Rust: Business, Technology and Work at the Lackawanna Steel Plant, 1899–1983 by Thomas E. Leary, Elizabeth C. Sholes (review). Technology and Culture, 29(4):929–931, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889175/pdf.

Cass:1988:BSW

[Cas88b] Ed Cass. “The Best School in the World”: West Point, The Pre — Civil War Years, 1833–1866 by James L. Morrison, Jr., and: Captains of the Old Steam Navy: Makers of the American Naval Tradition 1840–1880 ed. by James C. Bradford (review). Technology and Culture, 29(4):953–954, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889189/pdf.

Censer:1980:RGL

[Cen80] Jack R. Censer. Les Reserves de grains à long terme: Techniques de conservation et fonctions sociales dans l’histoire by François Sigaut (review). Technology and Culture, 21(1):114–116, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890626/pdf.

Ceruzzi:1984:HPL

[Cer84] Paul Ceruzzi. History of Programming Languages ed. by Richard L. Wexelblat (review). Technology and Culture, 25(1):144–146, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890103/pdf.

Ceruzzi:1985:TMW

[Cer85] Paul E. Ceruzzi. Turing’s Man: Western Culture in the Computer Age by J. David Bolter (review). Technology and Culture, 26(2):341–343, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889800/pdf.

Channell:1980:VEJ

[Cha80] David F. Channell. The Visionary Eye by Jacob Bronowski (review). Technology and Culture, 21(1):139–140, January 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890639/pdf.

Channell:1982:EMP

[Cha82a] David F. Channell. The Evolution of Mechanics by Pierre-Marie-Maurice Duhem (review). Technology and Culture, 23(1):106–107, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891119/pdf.

Channell:1982:THT

[Cha82b] David F. Channell. The Tragicomical History of Thermodynamics, 1822–1854 by C. Truesdell (review). Technology and Culture, 23(1):104–106, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891118/pdf.

Channell:1982:HTP

[Cha82c] David F. Channell. The harmony of theory and practice: The engineering science of W. J. M. Rankine. Technology and Culture, 23(1):39–52, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891103/pdf.

Channell:1984:BMA

[Cha84] David F. Channell. Before Mickey: The Animated Film 1898– 1928 by Donald Crafton (review). Technology and Culture, 25 (1):123–124, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890092/pdf.

Channell:1987:TEP

[Cha87] David F. Channell. Thomas A. Edison Papers. A Selective Microfilm Edition: Part 1 (1850–1878) ed. by Thomas E. Jeffrey, et al. (review). Technology and Culture, 28(1):126–129, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889457/
pdf.

Chambers:1988:MIP

[Cha88a] David Wade Chambers. The Myth of Inevitable Progress by Franco Ferrarotti (review). Technology and Culture, 29(3):709– 710, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889269/pdf.

Channell:1988:EST

[Cha88b] David F. Channell. Engineering science as theory and practice. Technology and Culture, 29(1):98–103, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888499/pdf.

Chappell:1989:UID

[Cha89] Sally A. Kitt Chappell. Urban ideals and the design of railroad stations. Technology and Culture, 30(2):354–375, April 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889067/pdf.

Cheape:1984:AUT

[Che84] Charles Cheape. The Automobile and Urban Transit: The Formation of Public Policy in Chicago, 1900–1930 by Paul Barrett (review). Technology and Culture, 25(2):360–362, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890069/pdf.

Christie:1989:FUA

[Chr89] Jean Christie. FDR’s Utopian: Arthur Morgan of the TVA by Roy Talbert (review). Technology and Culture, 30(1):139–141, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889126/pdf.

Channell:1980:THW

[CL80] David F. Channell and W. David Lewis. Technology and the humanities — the wingspread conference: Racine, Wisconsin, February 1979. Technology and Culture, 21(1):77–81, January 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890608/pdf.

Clark:1980:UAS

[Cla80] Paul W. Clark. U.S. Air Service in World War I ed. by Maurer Maurer (review). Technology and Culture, 21(2):265–267, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890693/pdf.

Clark:1986:GPT

[Cla86] Christopher Clark. The Great Paternalist: Titus Salt and the Growth of Nineteenth-Century Bradford by Jack Reynolds (review). Technology and Culture, 27(2):303–304, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889607/pdf.

Clark:1988:CEY

[Cla88] Christopher Clark. Cornwall Electric: 100 Years of Service by Karen Carter-Edwards, and: An Early History of Electricity Supply: The Story of the Electric Light in Victorian Leeds by J. D. Poulter (review). Technology and Culture, 29(3):694–695, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889249/pdf.

Clark:1989:RDT

[Cla89] Norman H. Clark. Roughnecks, Drillers, and Tool Pushers: Thirty-Three Years in the Oil Fields by Gerald Lynch (review). Technology and Culture, 30(3):698–700, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889059/pdf.

Coates:1980:DTE

[Coa80] Joseph F. Coates. Directing Technology ed. by Ron Johnston, Philip Gummett (review). Technology and Culture, 21(1):151–153, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890646/pdf.

Coates:1983:GTA

[Coa83] Joseph F. Coates. A Guidebook for Technology Assessment and Impact Analysis by Alan Porter, et al., and: Strategies for Conducting Technology Assessments by Joe E. Armstrong, W. H. Harman (review). Technology and Culture, 24(1):151–153, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891249/pdf.

Cochran:1981:BGO

[Coc81] Thomas C. Cochran. Business and Government in the Oil Industry: A Case Study of Sun Oil, 1876–1945 by August W. Giebelhaus (review). Technology and Culture, 22(3):651–653, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890918/
pdf.

Cochran:1982:ECA

[Coc82] Thomas C. Cochran. The Economy of Colonial America by Edwin J. Perkins (review). Technology and Culture, 23(1):117–118, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891125/pdf.

Cohen:1986:DHM

[Coh86] I. Bernard Cohen. The Discoverers: A History of Man’s Search to Know His World and Himself by Daniel J. Boorstin (review). Technology and Culture, 27(3):662–664, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889606/pdf.

Cohen:1989:CAN

[Coh89] Patricia Cline Cohen. Capitalism and Arithmetic: The New Math of the 15th Century by Frank J. Swetz (review). Technology and Culture, 30(3):670–671, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889038/pdf.

Coleman:1981:BEE

[Col81] Peter J. Coleman. Business Enterprise in Early New York ed. by Joseph R. Frese, S.J., Jacob Judd (review). Technology and Culture, 22(1):189–190, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890821/pdf.

Colodny:1984:SDS

[Col84] Robert G. Colodny. Science Deified and Science Defied: The Historical Significance of Science in Western Culture from the Bronze Age to the Beginnings of the Modern Era ca. 3500 B.C. to ca. A.D. 1640 by Richard Olson (review). Technology and Culture, 25(2):322–323, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890050/pdf.

Compton:1986:MRN

[Com86] Wm. David Compton. Model Research: The National Advisory Committee for Aeronautics 1915–1958 by Alex Roland (review). Technology and Culture, 27(3):653–658, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889598/pdf.

Compton:1987:FTE

[Com87a] Wm. David Compton. Far Travelers: The Exploring Machines by Oran W. Nicks (review). Technology and Culture, 28(3):740–741, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889403/pdf.

Compton:1987:SHH

[Com87b] Wm. David Compton. Searching the Horizon: A History of Ames Research Center 1940–1976 by Elizabeth A. Muenger (review). Technology and Culture, 28(2):390–391, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889437/pdf.

Condit:1980:CHA

[Con80a] Carl W. Condit. A Concise History of American Architecture by Leland M. Roth (review). Technology and Culture, 21(3):513–516, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890729/pdf.

Condit:1980:RMB

[Con80b] Carl W. Condit. Robert Maillart’s Bridges: The Art of Engineering by David P. Billington (review). Technology and Culture, 21 (2):243–246, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890683/pdf.

Connolly:1980:LMW

[Con80c] Terry Connolly. Limits of a Modern World: A Study of the “Limits to Growth” Debate by R. McCutcheon (review). Technology and Culture, 21(2):275–276, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890698/pdf.

Condit:1981:ETP

[Con81a] Carl W. Condit. Electric Traction on the Pennsylvania Railroad, 1895–1968 by Michael Bezilla (review). Technology and Culture, 22(3):637–639, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890911/pdf.

Condit:1981:PCC

[Con81b] Carl W. Condit. Paris: A Century of Change, 1878–1978 by Norma Evenson (review). Technology and Culture, 22(1):183–185, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890818/pdf.

Condit:1981:RMH

[Con81c] Carl W. Condit. Richard Morris Hunt by Paul R. Baker (review). Technology and Culture, 22(2):318–320, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890858/pdf.

Condit:1981:WBS

[Con81d] Carl W. Condit. Why Buildings Stand Up: The Strength of Architecture by Mario Salvadori (review). Technology and Culture, 22(4):790–792, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890950/pdf.

Condit:1981:ITB

[Con81e] Carl W. Condit. Iron truss bridge failures. Technology and Culture, 22(4):846, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890975/pdf.

Constant:1981:HIP

[Con81f] Edward W. Constant II. A History of Industrial Power in the United States. Vol. 1: Waterpower in the Century of the Steam Engine by Louis C. Hunter (review). Technology and Culture, 22 (2):371–373, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890882/pdf.

Condit:1982:AAM

[Con82] Carl W. Condit. American Architecture, 1607–1976 by Marcus Whiffen, Frederick Koeper (review). Technology and Culture, 23 (4):684–686, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890272/pdf.

Condit:1983:EGS

[Con83a] Carl W. Condit. Experiments in Gothic Structure by Robert Mark (review). Technology and Culture, 24(2):261–263, April 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891261/pdf.

Constant:1983:STT

[Con83b] Edward W. Constant. Scientific theory and technological testability: Science, dynamometers, and water turbines in the 19th century. Technology and Culture, 24(2):183–198, April 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891254/pdf.

Condit:1984:AVC

[Con84] Carl W. Condit. Another view of “The cathedral and the bridge”. Technology and Culture, 25(3):589–594, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889994/pdf.

Condit:1985:GSG

[Con85a] Carl W. Condit. The Geological Survey’s Guidebook of the Western United States. Technology and Culture, 26(2):277–278, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889769/
pdf.

Condit:1985:MCR

[Con85b] Carl W. Condit. Metropolitan Corridor: Railroads and the American Scene by John R. Stilgoe (review). Technology and Culture, 26(2):319–321, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889788/pdf.

Constant:1985:STH

[Con85c] Edward W. Constant. Stronger than a Hundred Men: A History of the Vertical Water Wheel by Terry S. Reynolds (review). Technology and Culture, 26(4):826–828, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889681/pdf.

Constant:1985:HST

[Con85d] Edward W. Constant. The History of Science and Technology in the United States: A Critical and Selective Bibliography by Marc Rothenberg (review). Technology and Culture, 26(4):850–851, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889693/pdf.

Condit:1986:MMW

[Con86a] Carl W. Condit. McKim, Mead & White, Architects by Leland M. Roth (review). Technology and Culture, 27(1):169–172, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889665/pdf.

Condit:1986:ANY

[Con86b] Carl W. Condit. The Architecture of New York City: Histories and Views of Important Structures, Sites, and Symbols by Donald Martin Reynolds (review). Technology and Culture, 27(3):645–646, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889588/pdf.

Condit:1987:LGP

[Con87] Carl W. Condit. The Late, Great Pennsylvania Station by Lorraine B. Diehl (review). Technology and Culture, 28(4):885–886, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889332/pdf.

Condit:1988:EFB

[Con88a] Carl W. Condit. Ernest Flagg: Beaux-Arts Architect and Urban Reformer by Mardges Bacon (review). Technology and Culture, 29(1):165–167, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888524/pdf.

Condit:1988:MWB

[Con88b] Carl W. Condit. Manhattan Water-Bound: Planning and Developing Manhattan’s Waterfront from the Seventeenth Century to the Present by Ann L. Buttenwieser (review). Technology and Culture, 29(3):700–702, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889257/pdf.

Constant:1989:CCS

[Con89] Edward W. Constant II. Cause or consequence: Science, technology, and regulatory change in the oil business in Texas, 1930– 1975. Technology and Culture, 30(2):426–455, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889073/pdf.

Cooper:1980:DVR

[Coo80] Carolyn C. Cooper. Dynamos and Virgins Revisited: Women and Technological Change in History ed. by Martha Moore Trescott (review). Technology and Culture, 21(4):640–644, October 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890758/pdf.

Coopersmith:1983:IIS

[Coo83] Jonathan Coopersmith. Industrial Innovation in the Soviet Union ed. by Ronald Amann, Julian Cooper (review). Technology and Culture, 24(4):701–703, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890144/pdf.

Cooper:1984:CIH

[Coo84a] Carolyn C. Cooper. Connecticut: An Inventory of Historic Engineering and Industrial Sites by Matthew Roth, and: Connecticut the Industrial Incubator by Gerard L. Studley (review). Technology and Culture, 25(2):345–347, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890062/pdf.

Cooper:1984:PSM

[Coo84b] Carolyn C. Cooper. The portsmouth system of manufacture. Technology and Culture, 25(2):182–225, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890035/pdf.

Coopersmith:1984:ICP

[Coo84c] Jonathan G. Coopersmith. The Industrialization of the Continental Powers 1780–1914 by Clive Trebilcock (review). Technology and Culture, 25(2):341–342, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890060/pdf.

Cooper:1985:RDA

[Coo85] Carolyn C. Cooper. The Rise and Decline of the American Cut Nail Industry: A Study of the Interrelationships of Technology, Business Organization, and Management Techniques by Amos J. Loveday, Jr. (review). Technology and Culture, 26(2):309–310, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889783/pdf.

Cooper:1986:EWH

[Coo86a] Patricia Cooper. The Electrical Workers: A History of Labor at General Electric and Westing-house, 1923–60 by Ronald W. Schatz (review). Technology and Culture, 27(4):875–877, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889530/pdf.

Coopersmith:1986:PIM

[Coo86b] Jonathan Coopersmith. The Politics of Industrial Mobilization in Russia, 1914–17: A Study of the War-Industries Committees by Lewis H. Siegelbaum (review). Technology and Culture, 27(2):311– 313, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889614/pdf.

Cooper:1987:GKH

[Coo87] Carolyn C. Cooper. The ghost in the kitchen: Household technology at the Brattleboro Museum, Vermont. Technology and Culture, 28(2):328–332, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889400/pdf.

Cooper:1988:WCN

[Coo88] Patricia A. Cooper. “What this country needs is a good five-cent cigar”. Technology and Culture, 29(4):779–807, October 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889158/pdf.

Cooper:1989:WWT

[Coo89] Carolyn C. Cooper. Women, Work, and Technology: Transformations ed. by Barbara Wright (review). Technology and Culture, 30(4):1086–1087, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889013/pdf.

Corn:1985:FAW

[Cor85] Joseph J. Corn. Flight in America 1900–1983: From the Wrights to the Astronauts by Roger E. Bilstein (review). Technology and Culture, 26(4):871–873, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889706/pdf.

Corn:1987:MMT

[Cor87] Joseph J. Corn. Mechanical Metamorphosis: Technological Change in Revolutionary America by Neil Longley York (review). Technology and Culture, 28(1):153–155, January 1987. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889472/pdf.

Corn:1988:NPT

[Cor88a] Joseph J. Corn. New Perspectives on Technology and American Culture ed. by Bruce Sinclair (review). Technology and Culture, 29(4):968–970, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889198/pdf.

Corn:1988:SMS

[Cor88b] Joseph J. Corn. Superpower: The Making of a Steam Locomotive by David Weitzman (review). Technology and Culture, 29 (4):917–918, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889167/pdf.

Corn:1989:AAJ

[Cor89] Joseph J. Corn. The Automobile Age by James J. Flink (review). Technology and Culture, 30(4):1066–1067, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888996/pdf.

Costello:1985:WPT

[Cos85] Cynthia Costello. Woman’s Place Is at the Typewriter: Office Work and Office Workers 1870–1930 by Margery W. Davies, and: Microelectronics and Office Jobs: The Impact of the Chip on Women’s Employment by Diane Werneke (review). Technology and Culture, 26(1):104–106, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889824/pdf.

Costello:1987:WUN

[Cos87] Cynthia B. Costello. Work, Unemployment and the New Technology by Colin Gill (review). Technology and Culture, 28(3):707–709, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889375/pdf.

Costello:1988:CRT

[Cos88] Cynthia B. Costello. The Control Revolution: Technological and Economic Origins of the Information Society by James R. Beniger (review). Technology and Culture, 29(2):326–328, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889294/pdf.

Cousins:1986:HES

[Cou86] Peter H. Cousins. History and Ecology: Studies of the Grassland by James C. Malin (review). Technology and Culture, 27(2):318–319, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889618/pdf.

Courtenay:1988:ARE

[Cou88] Lynn T. Courtenay. The Architecture of the Roman Empire. Vol. 2: An Urban Appraisal by William MacDonald (review). Technology and Culture, 29(3):675–677, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889237/pdf.

Cowan:1983:GDR

[Cow83] Ruth Schwartz Cowan. The Grand Domestic Revolution: A History of Feminist Designs for American Homes, Neighborhoods and Cities by Dolores Hayden (review). Technology and Culture, 24 (2):285–288, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891277/pdf.

Cowan:1986:CTA

[Cow86] Ruth Schwartz Cowan. The Culture of Technology by Arnold Pacey (review). Technology and Culture, 27(2):291–293, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889592/pdf.

Cozzens:1989:SCT

[Coz89] Susan E. Cozzens. The Social Construction of Technological Systems: New Directions in the Sociology and History of Technology ed. by Wiebe E. Bijker, et al. (review). Technology and Culture, 30(3):705–707, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889068/pdf.

Cutcliffe:1986:CBH

[CRA86] Stephen H. Cutcliffe, Christine M. Roysdon, and Judith A. Adams. Current bibliography in the history of technology (1984). Technology and Culture, 27(2):343–476, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889631/pdf.

Cutcliffe:1987:CBH

[CRA87] Stephen H. Cutcliffe, Christine M. Roysdon, and Judith A. Adams. Current bibliography in the history of technology (1985). Technology and Culture, 28(2):401–538, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889454/pdf.

Cutcliffe:1988:CBH

[CRA^{+}88] Stephen H. Cutcliffe, Christine M. Roysdon, Judith A. Adams, Jane Morley, Louis Rodriquez, and Ian Winship. Current bibliography in the history of technology (1986). Technology and Culture, 29(2):338–492, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889300/pdf.

Cutclifee:1989:CBH

[CRA^{+}89] Stephen H. Cutclifee, Christine M. Roysdon, Judith A. Adams, Victoria Dow, Jane Morley, Louis Rodriquez, and Ian Winship. Current bibliography in the history of technology (1987). Technology and Culture, 30(3):723–877, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889089/pdf.

Crew:1983:GWC

[Cre83] Spencer R. Crew. George Washington Carver: Scientist and Symbol by Linda O. McMurry (review). Technology and Culture, 24 (2):275–277, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891271/pdf.

Crimmins:1980:AAA

[Cri80] Timothy J. Crimmins. Automobile Age Atlanta: The Making of a Southern Metropolis, 1900–1935 by Howard L. Preston (review). Technology and Culture, 21(1):122–124, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890630/pdf.

Cutcliffe:1984:CBH

[CRM84] Stephen H. Cutcliffe, Christine M. Roysdon, and Judith A. Mistichelli. Current bibliography in the history of technology (1982). Technology and Culture, 25(2):382–472, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890033/pdf.

Cutcliffe:1985:CBH

[CRM85] Stephen H. Cutcliffe, Christine M. Roysdon, and Judith A. Mistichelli. Current bibliography in the history of technology (1983). Technology and Culture, 26(2):353–470, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889805/pdf.

Cronin:1982:MUM

[Cro82] James E. Cronin. Masters, Unions and Men: Work Control in Building and the Rise of Labour, 1830–1914 by Richard Price (review). Technology and Culture, 23(2):256–258, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891165/pdf.

Crough:1983:SPA

[Cro83] Tom D. Crough. Sky Pioneering: Arizona in Aviation History by Ruth M. Reinhold (review). Technology and Culture, 24(2):296– 297, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891283/pdf.

Crouch:1986:NLI

[Cro86] Tom D. Crouch. No Longer an Island: Britain and the Wright Brothers, 1902–1909 by Alfred Gollin (review). Technology and Culture, 27(3):651–652, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889594/pdf.

Cronin:1987:MOS

[Cro87a] L. Eugene Cronin. Marshes of the Ocean Shore: Development of an Ecological Ethic by Joseph V. Siry (review). Technology and Culture, 28(4):860–861, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889316/pdf.

Crouch:1987:PFC

[Cro87b] Tom D. Crouch. The Prehistory of Flight by Clive Hart (review). Technology and Culture, 28(3):733–734, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889393/pdf.

Crunden:1987:FEM

[Cru87] Robert M. Crunden. The Fatal Environment: The Myth of the Frontier in the Age of Industrialization 1800–1890 by Richard Slotkin (review). Technology and Culture, 28(4):859–860, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889315/pdf.

Cudahy:1988:RMN

[Cud88] Brian J. Cudahy. “Rails to the Metropolis” at the New York State Museum. Technology and Culture, 29(2):276–280, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889256/pdf.

Cumiford:1986:CNP

[Cum86] William L. Cumiford. Catálogo de Noventa Presas y Azudes Españoles Anteriores a 1900 by José A. Fernández Ordóñez (review). Technology and Culture, 27(4):858–860, October 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889517/pdf.

Cumiford:1988:SMC

[Cum88] William L. Cumiford. Sons of the Machine: Case Studies of Social Change in the Workplace by Charles H. Savage, Jr., George F. F. Lombard (review). Technology and Culture, 29(1):177–179, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888532/pdf.

Curtis:1984:MMC

[Cur84a] James M. Curtis. Multi-Media Communications ed. by May Katzen (review). Technology and Culture, 25(3):687–688, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890025/pdf.

Curtis:1984:TSI

[Cur84b] James M. Curtis. Toward a sociotechnological interpretation of popular music in the Electronic Age. Technology and Culture, 25 (1):91–102, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890082/pdf.

Cutcliffe:1981:BHT

[Cut81a] Stephen H. Cutcliffe. Being Human in a Technological Age ed. by Donald M. Borchet, David Stewart (review). Technology and Culture, 22(1):204–206, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890831/pdf.

Cutcliffe:1981:TCS

[Cut81b] Stephen H. Cutcliffe. Technology, Change and Society by Edward C. Pytlik, et al., and: Technology and Change ed. by John G. Burke, Marshall C. Eakin, and: Technology and Its Impact on Society ed. by Sigvard Strandh (review). Technology and Culture, 22(3):666–670, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890924/pdf.

Cutcliffe:1985:PRS

[Cut85] Stephen H. Cutcliffe. Politics and the Restraint of Science by Leonard A. Cole (review). Technology and Culture, 26(1):120–121, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889831/pdf.

Cutcliffe:1987:LFU

[Cut87] Stephen H. Cutcliffe. Lost at the Frontier: U.S. Science and Technology Policy Adrift by Deborah Shapley, Rustum Roy (review). Technology and Culture, 28(1):184–186, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889489/pdf.

Dood:1989:TIM

[DAK89] Kendall J. Dood, Leland I. Anderson, and Ronald R. Kline. Tesla and the induction motor. Technology and Culture, 30(4): 1013–1023, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888974/pdf.

Daley:1988:RPL

[Dal88] Brian E. Daley, S.J. The Restoration of Perfection: Labor and Technology in Medieval Culture by George Ovitt, Jr. (review). Technology and Culture, 29(4):913–915, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889165/pdf.

Danko:1980:ABT

[Dan80] George M. Danko. Arch Bridges and Their Builders, 1735–1835 by Ted Ruddock (review). Technology and Culture, 21(2):241–243, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890682/pdf.

Daniels:1981:TIP

[Dan81a] George H. Daniels. Technology: An Introduction by Paul W. De-Vore (review). Technology and Culture, 22(3):608–611, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890896/
pdf.

Danko:1981:PMW

[Dan81b] George Danko. Poorly Marked and Worse Lighted: Being a History of the Port Wardens of Philadelphia, 1766–1907 by Eugene R. Slaski (review). Technology and Culture, 22(2):311–312, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890853/pdf.

Daniels:1982:MDF

[Dan82] George H. Daniels. Meet Dr. Franklin ed. by Roy N. Lokken (review). Technology and Culture, 23(1):120–121, January 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891127/pdf.

Daniel:1983:SSS

[Dan83a] Pete Daniel. The Selling of the South: The Southern Crusade for Industrial Development, 1936–1980 by James C. Cobb (review). Technology and Culture, 24(3):521–522, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891408/pdf.

Daniels:1983:THT

[Dan83b] George H. Daniels. The Teaching of History of Technology in USA: A Critical Survey in 1978 by Svante Lindqvist (review). Technology and Culture, 24(1):122–123, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891233/pdf.

Daniels:1983:CVE

[Dan83c] Roger Daniels. Concerning Vanguard of expansion. Technology and Culture, 24(2):258–259, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891259/pdf.

Daniel:1985:TSC

[Dan85] Pete Daniel. Technology and Social Change in Rural Areas: A Festschrift for Eugene A. Wilkening ed. by Gene F. Summers (review). Technology and Culture, 26(2):300–301, April 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889778/pdf.

Daniel:1987:ASQ

[Dan87] Pete Daniel. Agricultural Science and the Quest for Legitimacy: Farmers, Agricultural Colleges, and Experiment Stations, 1870– 1890 by Alan I. Marcus (review). Technology and Culture, 28 (2):371–372, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889424/pdf.

Daston:1985:PCS

[Das85] Lorraine J. Daston. Probability and Certainty in Seventeenth-Century England: A Study of the Relationships between Natural Science, Religion, History, Law, and Literature by Barbara J. Shapiro (review). Technology and Culture, 26(4):835–836, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889686/pdf.

Daumas:1981:BG

[Dau81] Maurice Daumas. Bertrand Gille (1920–1980). Technology and Culture, 22(4):839–842, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890972/pdf.

Davis:1980:MFC

[Dav80] Audrey B. Davis. Microscopes from the Frank Collection by R. H. Nuttall (review). Technology and Culture, 21(1):93–95, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890613/pdf.

Davis:1983:CCT

[Dav83] Edward B. Davis. Clocks and the Cosmos: Time in Western Life and Thought by Samuel L. Macey (review). Technology and Culture, 24(2):264–265, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891263/pdf.

Davis:1985:CAS

[Dav85] Audrey B. Davis. Cholera: The American Scientific Experience, 1947–1980 by W. E. van Heyningen and John R. Seal (review). Technology and Culture, 26(1):124–126, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889833/pdf.

Davis:1987:HAE

[Dav87] Audrey B. Davis. “The Healing Arts in Early America” — Fraunces Tavern Museum, New York. Technology and Culture, 28(2):333–336, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889402/pdf.

Davidson:1989:CTS

[Dav89a] Frank P. Davidson. The Channel Tunnel Story by Michael R. Bonavia (review). Technology and Culture, 30(3):701–702, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889062/pdf.

Davies:1989:AEH

[Dav89b] R. E. G. Davies. The Airway to Everywhere: A History of All American Aviation, 1937–1953 by W. David Lewis, William F. Trimble (review). Technology and Culture, 30(4):1076–1078, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889003/pdf.

Davis:1989:UCI

[Dav89c] Audrey B. Davis. Uneasy Careers and Intimate Lives: Women in Science, 1789–1979 ed. by Pnina G. Abir-Am, Dorinda Outram (review). Technology and Culture, 30(4):1084–1085, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889010/pdf.

Dawson:1989:WTS

[Daw89] Virginia P. Dawson. Whittle: The True Story by John Golley (review). Technology and Culture, 30(4):1074, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889001/pdf.

Day:1989:RAB

[Day89] C. R. Day. La République avait besoin de savants by Janis Langins (review). Technology and Culture, 30(3):676–677, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889042/pdf.

deLotterBright:1980:JLS

[dB80] Charles deLotter Bright. By Jupiter! The Life of Sir Roy Fedden by Bill Gunston (review). Technology and Culture, 21(1):134–135, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890636/pdf.

Daniels:1988:OSR

[DD88] John Daniels and Christian Daniels. The origin of the sugarcane roller mill. Technology and Culture, 29(3):493–535, July 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889202/pdf.

DeGregori:1980:REH

[De 80a] Thomas R. De Gregori. Research in Economic History: A Research Annual ed. by Paul Uselding (review). Technology and Culture, 21(3):521–522, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890733/pdf.

DeGregori:1980:STD

[De 80b] Thomas R. De Gregori. Science and Technology for Development: The Role of U.S. Universities by Robert P. Morgan, et al. (review). Technology and Culture, 21(3):544–545, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890744/pdf.

DeGregori:1981:SCE

[De 81a] Thomas R. De Gregori. Science and Colonial Expansion: The Role of the British Royal Botanic Gardens by Lucile H. Brockway (review). Technology and Culture, 22(4):809–811, October 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890960/pdf.

DeGregori:1981:TTG

[De 81b] Thomas R. De Gregori. Through the ’80’s: Thinking Globally, Acting Locally ed. by Frank Feather (review). Technology and Culture, 22(3):691–692, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890933/pdf.

deSollaPrice:1981:RHS

[de 81c] Derek de Solla Price. Richard H. Schallenberg (1943–1980). Technology and Culture, 22(4):845, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890974/pdf.

DeGregori:1982:BNM

[De 82a] Thomas R. De Gregori. The back-to-nature movement: Alternative technologies and the inversion of reality. Technology and Culture, 23(2):214–217, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891147/pdf.

Price:1982:ITP

[de 82b] Derek de Solla Price. Les Instruments de la théorie des planètes selon Ptolémée: Équatoires et horlogerie planétaire du XIII^{e} au XVI^{e} siècle by Emmanuel Poulle (review). Technology and Culture, 23(4):651–653, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890253/pdf.

DeGregori:1983:RTS

[De 83] Thomas R. De Gregori. Risk in the Technological Society ed. by Christoph Hohenemser, Jeanne X. Kasperson (review). Technology and Culture, 24(2):311–313, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891291/pdf.

DeGregori:1984:ADI

[De 84a] Thomas R. De Gregori. Absorption and Diffusion of Imported Technology: Proceedings of a Workshop Held in Singapore, 26–30 January 1981 (review). Technology and Culture, 25(4):871, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889965/pdf.

DeGregori:1984:CWF

[De 84b] Thomas R. De Gregori. Chemistry and World Food Supplies: The New Frontiers ed. by L. W. Shemilt (review). Technology and Culture, 25(3):675–676, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890019/pdf.

DeGregori:1984:STC

[De 84c] Thomas R. De Gregori. Science and Technology in a Changing International Order: The United Nations Conference on Science and Technology for Development ed. by Volker Rittberger (review). Technology and Culture, 25(2):315–316, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890045/pdf.

DeGregori:1984:SPA

[De 84d] Thomas R. De Gregori. Science, Politics, and the Agricultural Revolution in Asia ed. by Robert S. Anderson, et al. (review). Technology and Culture, 25(3):676–679, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890020/pdf.

Deasey:1984:WWA

[Dea84] John Deasey. The Winning Weapon: The Atomic Bomb in the Cold War 1945–1950 by Gregg Herken (review). Technology and Culture, 25(3):666–667, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890014/pdf.

Deasey:1985:WAF

[Dea85] John Deasey. The Wizards of Armageddon by Fred Kaplan (review). Technology and Culture, 26(3):674–676, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889754/pdf.

DeGregori:1987:SRT

[DeG87] Thomas R. DeGregori. Social Responses to Technological Change ed. by Augustine Brannigan, Sheldon Goldenberg (review). Technology and Culture, 28(1):138–139, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889462/pdf.

DelSesto:1982:TSJ

[Del82] Steven L. Del Sesto. The Technological System by Jacques Ellul (review). Technology and Culture, 23(1):81–85, January 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891107/pdf.

DelSesto:1983:FEJ

[Del83] Steven L. Del Sesto. The Fate of the Earth by Jonathan Schell (review). Technology and Culture, 24(2):305–309, April 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891289/pdf.

DeVorkin:1985:GIU

[DeV85] David H. DeVorkin. Glimpsing an Invisible Universe: The Emergence of X-ray Astronomy by Richard F. Hirsh (review). Technology and Culture, 26(1):132–134, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889838/pdf.

DeVorkin:1988:BLB

[DeV88] David H. DeVorkin. Bernard Lovell: A Biography by Dudley Saward (review). Technology and Culture, 29(1):162–165, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888523/pdf.

Durbin:1987:PPM

[DHW^{+}87] Paul Durbin, Brooke Hindle, Thomas R. Winpenny, Edward W. Constant II, Joan Rothschild, Susan Smulyan, Daniel Headrick, Colleen A. Dunlavy, W. Bernard Carlson, A. Hunter Dupree, William McGucken, Bernard S. Finn, Kathleen Ochs, James E. Tomayko, Miriam R. Levin, Merritt Roe Smith, Marjorie N.

Boyer, and Jeanne Allen. The Pittsburgh, Pennsylvania, Meeting, October 23–26, 1986. Technology and Culture, 28(3):619–647, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889353/pdf.

Dibner:1981:EB

[Dib81] Bern Dibner. Elmer Belt (1893–1980). Technology and Culture, 22(4):837–838, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890971/pdf.

Douglas:1986:DMM

[DKP^{+}86] Susan J. Douglas, Kathleen Anderson Kraus, William S. Pretzer, George Ovitt, Jr., Larry Owens, O. Allan Gianniny, Jr., W. David Lewis, Sheldon Hochheiser, Edmund N. Todd, and Kathleen Ochs. The Dearborn, Michigan, Meeting, October 17–20, 1985. Technology and Culture, 27(3):577–593, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889542/pdf.

delaBruheze:1988:FNW

[dlB88] Adri A. Albert de la Bruhèze. Forevermore: Nuclear Waste in America by Donald L. Barlett, James B. Steele (review). Technology and Culture, 29(4):979–980, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889209/pdf.

deMontellano:1985:SAM

[dM85] Bernard R. Ortiz de Montellano. Skywatchers of Ancient Mexico by Anthony F. Aveni (review). Technology and Culture, 26(2):343– 345, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889801/pdf.

Dobney:1981:CNA

[Dob81a] Fredrick J. Dobney. A City for the Nation: The Army Engineers and the Building of Washington, D.C., 1790–1967 by Albert E. Cowdrey (review). Technology and Culture, 22(1):186–187, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890819/pdf.

Dobney:1981:HDH

[Dob81b] Fredrick J. Dobney. The Headwaters District: A History of the Pittsburgh District, U.S. Army Corps of Engineers by Leland R. Johnson (review). Technology and Culture, 22(2):313–315, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890855/pdf.

Donaghay:1981:BCP

[Don81a] Marie Donaghay. The Bureau of Commerce in 1781 and Its Policies with Respect to French Industry by Harold T. Parker (review). Technology and Culture, 22(3):625–626, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890904/pdf.

Donovan:1981:EPC

[Don81b] Arthur L. Donovan. Environmental Politics and the Coal Coalition by Richard H. K. Vietor (review). Technology and Culture, 22(4):831–833, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890968/pdf.

Donovan:1986:FHM

[Don86a] Arthur Donovan. Fire in the Hole: Miners and Managers in the American Coal Industry by Curtis Seltzer (review). Technology and Culture, 27(2):321–322, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889620/pdf.

Donovan:1986:TAE

[Don86b] Arthur Donovan. Thinking about engineering. Technology and Culture, 27(4):674–679, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889493/pdf.

Dood:1987:MIE

[Doo87] Kendall J. Dood. Models of invention and enterprise: 150 years of the American patent system — two exhibits at the National Museum of American History, Washington, D.C. Technology and Culture, 28(3):670–674, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889357/pdf.

Dorn:1980:BBG

[Dor80] Harold Dorn. The Britannia Bridge: The Generation and Diffusion of Technological Knowledge by Nathan Rosenberg, Walter G. Vincenti (review). Technology and Culture, 21(1):100–101, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890616/pdf.

Doran:1983:FPS

[Dor83] Glen H. Doran. Florida’s Prehistoric Stone Technology by Barbara A. Purdy (review). Technology and Culture, 24(3):509–510, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891399/pdf.

Douglas:1984:MAM

[Dou84] Susan J. Douglas. Media and the American Mind: From Morse to McLuhan by Daniel J. Czitrom (review). Technology and Culture, 25(3):685–687, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890024/pdf.

Downey:1984:NWS

[Dow84a] Gary L. Downey. Nuclear Waste: Socioeconomic Dimensions of Long-Term Storage ed. by Steve H. Murdock, F. Larry Leistritz, and Rita R. Hamm (review). Technology and Culture, 25(4): 891–893, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889978/pdf.

Downey:1984:PNW

[Dow84b] Gary L. Downey. The Politics of Nuclear Waste ed. by E. William Colglazier, Jr. (review). Technology and Culture, 25(1):158–159, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890112/pdf.

Downey:1985:EDD

[Dow85] Gary L. Downey. Energy for Developed and Developing Countries ed. by Behram N. Kursunoglu, et al., and: A Global View of Energy ed. by Behram N. Kursunoglu, et al. (review). Technology and Culture, 26(1):110–112, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889827/pdf.

Daniels:1983:PTP

[DR83] George H. Daniels and Mark H. Rose. Power, transport, and public policy in modern America: Program in science, technology, and society, Michigan Technological University, September 25–27, 1981. Technology and Culture, 24(1):91–97, January 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891224/pdf.

Dreicer:1989:IRT

[Dre89] Gregory K. Dreicer. Innovation and the Rise of the Tunnelling Industry by Graham West (review). Technology and Culture, 30 (4):1059–1061, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888992/pdf.

DeGregori:1989:RTT

[DSF89] Thomas R. DeGregori and Kristin Shrader-Frechette. The review of a theory of technology. Technology and Culture, 30(4): 1027–1030, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888976/pdf.

Pool:1981:CSS

[dSP81] Ithiel de Sola Pool. The Computerization of Society by Simon Nora, Alain Minc (review). Technology and Culture, 22(2):352– 354, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890873/pdf.

Dublin:1983:FTI

[Dub83] Thomas Dublin. Family Time and Industrial Time: The Relationship between the Family and Work in a New England Industrial Community by Tamara K. Hareven (review). Technology and Culture, 24(4):683–686, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890133/pdf.

Dubofsky:1988:GBP

[Dub88] Melvyn Dubofsky. A Generation of Boomers: The Pattern of Railroad Labor Conflict in Nineteenth-Century America by Shelton Stromquist (review). Technology and Culture, 29(4):918–919, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889168/pdf.

Duffy:1980:LDA

[Duf80] M. C. Duffy. Locomotive Designers in the Age of Steam by J. N. Westwood (review). Technology and Culture, 21(3):499–502, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890722/pdf.

Duncan:1980:THL

[Dun80] Ralph Duncan. The Territory of the Historian by E. LeRoy Ladurie (review). Technology and Culture, 21(2):237–239, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890680/pdf.

Dunlap:1981:UCH

[Dun81] Thomas R. Dunlap. U.S. Cultural History: A Guide to Information Sources ed. by Phillip I. Mitterling (review). Technology and Culture, 22(2):320, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890859/pdf.

Dunlavy:1989:RCA

[Dun89] Colleen Dunlavy. Railroads and the Character of America, 1820– 1887 by James A. Ward (review). Technology and Culture, 30 (2):466–467, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889085/pdf.

Durbin:1983:THN

[Dur83a] Paul T. Durbin. Technology — Humanism or Nihilism: A Critical Analysis of the Philosophical Basis and Practice of Modern Technology by Gregory H. Davis (review). Technology and Culture, 24(1):153, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891250/pdf.

Durbin:1983:IWC

[Dur83b] Paul T. Durbin. The Imaginary Witness: The Critical Theory of Herbert Marcuse by Morton Schoolman (review). Technology and Culture, 24(3):550–553, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891428/pdf.

Durbin:1985:ETD

[Dur85] Paul T. Durbin. Existential Technics by Don Ihde (review). Technology and Culture, 26(4):884–886, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889711/pdf.

Eaton:1988:AMW

[Eat88] Leonard K. Eaton. Architecture, Men, Women and Money in America 1600–1860 by Roger G. Kennedy (review). Technology and Culture, 29(3):695–697, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889250/pdf.

Edge:1984:HMA

[Edg84] David Edge. The History of Modern Astronomy and Astrophysics: A Selected, Annotated Bibliography by David H. DeVorkin (review). Technology and Culture, 25(1):135–136, January 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890098/pdf.

Ezell:1981:PRT

[EE81] Edward C. Ezell and Linda Neuman Ezell. Pionery raketnoy tekhniki: Gansvindt, Goddard, Esno-Pel’tri, Obert, Goman — Izbranniye trudiy (1891–1938) [Pioneers of rocket technology: Ganswindt, Goddard, Esnault-Pelterie, Oberth, and Hohmann — selected works (1891–1938)] ed. by T. M. Mell’kumov, V. N. Sokol’skiy (review). Technology and Culture, 22(1):202–203, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890829/pdf.

Eisenman:1980:HSI

[Eis80a] H. J. Eisenman. From Horsecars to Streamliners: An Illustrated History of the St. Louis Car Company by Alan R. Lind (review). Technology and Culture, 21(1):118–120, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890628/pdf.

Eisenman:1980:EGC

[Eis80b] H. J. Eisenman. The Emerging Goddess: The Creative Process in Art, Science and Other Fields by Albert Rothenberg (review). Technology and Culture, 21(4):689–691, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890781/pdf.

Elbl:1987:ENM

[Elb87] Martin M. Elbl. European Naval and Maritime History, 300–1500 by Archibald R. Lewis, Timothy J. Runyan (review). Technology and Culture, 28(4):852–853, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889310/pdf.

Elbl:1988:SGK

[Elb88] Martin M. Elbl. Six Galleons for the King of Spain: Imperial Defense in the Early Seventeenth Century by Carla Rahn Phillips (review). Technology and Culture, 29(4):947–949, October 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889185/pdf.

Elliott:1987:PJH

[Ell87] Clark A. Elliott. The Papers of Joseph Henry. Vol. 5: The Princeton Years: January 1841–December 1843 ed. by Nathan Reingold, et al. (review). Technology and Culture, 28(1):163–164, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889477/pdf.

Emmerson:1980:SSH

[Emm80a] G. S. Emmerson. The Story and Scandal of HMS Megaera by R. Hayward (review). Technology and Culture, 21(1):124–126, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890631/pdf.

Emmerson:1980:CRG

[Emm80b] George S. Emmerson. L. T. C. Rolt and the Great Eastern Affair of Brunel versus Scott Russell. Technology and Culture, 21 (4):553–569, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890748/pdf.

Emme:1981:RTF

[Emm81] Eugene M. Emme. The Rocket Team by Frederick I. Ordway III, Mitchell R. Sharpe (review). Technology and Culture, 22(2):343– 345, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890869/pdf.

Emmerson:1982:BBD

[Emm82] George S. Emmerson. Brunel’s Britain by Derrick Beckett (review). Technology and Culture, 23(3):508–510, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891211/pdf.

Emmerson:1983:MTM

[Emm83a] George S. Emmerson. The Making of Technological Man: The Social Origins of French Engineering Education by John Hubbel Weiss (review). Technology and Culture, 24(3):498–500, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891394/
pdf.

Emmerson:1983:GEC

[Emm83b] George S. Emmerson. The Great Eastern controversy: In response to Dr. Buchanan. Technology and Culture, 24(1):107–113, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/

891226/pdf. See [Buc82b].

Engle:1980:JGS

[Eng80] Anita Engle. Journal of Glass Studies (review). Technology and Culture, 21(3):481–484, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890714/pdf.

Engelbourg:1987:CAV

[Eng87] Saul Engelbourg. Challenges to American Values: Society, Business, and Religion by Thomas C. Cochran (review). Technology and Culture, 28(1):186–187, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889490/pdf.

Esper:1982:ISM

[Esp82] Thomas Esper. Industrial serfdom and metallurgical technology in 19th-Century Russia. Technology and Culture, 23(4):583–608, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890241/pdf.

Estes:1985:NBW

[Est85] Leland L. Estes. The Night Battles: Witchcraft and Agrarian Cults in the Sixteenth and Seventeenth Centuries by Carlo Ginzburg (review). Technology and Culture, 26(4):833–834, October 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889685/pdf.

Evans:1981:RRC

[Eva81] Francis T. Evans. Roads, railways, and canals: Technical choices in 19th-Century Britain. Technology and Culture, 22(1):1–34, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890796/pdf.

Evans:1988:OWS

[Eva88] John Karl Evans. The Origins of War: From the Stone Age to Alexander the Great by Arther Ferrill (review). Technology and Culture, 29(1):136–137, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/888506/pdf.

Evelyn:1986:VAM

[Eve86] Douglas E. Evelyn. The Victoria and Albert Museum: The History of Its Building by John Physick (review). Technology and Culture, 27(3):627–629, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889572/pdf.

Fahlman:1989:RAA

[Fah89] Betsy Fahlman. The Railroad in American Art: Representations of Technological Change ed. by Susan Danly, Leo Marx (review). Technology and Culture, 30(4):1052–1054, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888987/pdf.

Farmer:1980:IAS

[Far80] B. H. Farmer. Interpretive Analysis of Selected Papers from Changes in Rice Farming in Selected Areas of Asia by International Rice Research Institute (review). Technology and Culture, 21(2):278–280, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890700/pdf.

Friedel:1981:TMO

[FBW^{+}81] Robert Friedel, John G. Brainerd, Vivian Weil, Daniel J. Kevles, Reese V. Jenkins, Bert Hansen, Deborah Warner, Fredrick J. Dobney, Margaret W. Rossiter, Daniel J. Perkins, Arnold E. Roos, John G. Burke, Mark H. Rose, and P. Thomas Carroll. The Toronto Meeting, October 16–19, 1980. Technology and Culture, 22(3):570–597, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890892/pdf.

Fee:1985:CWP

[Fee85] Elizabeth Fee. Chesapeake Waters: Pollution, Public Health, and Public Opinion, 1607–1972 by John Capper, et al. (review). Technology and Culture, 26(2):329–331, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889794/pdf.

Feldman:1984:TRJ

[Fel84] Stanley C. Feldman. Technology and Reality by James K. Feibleman (review). Technology and Culture, 25(3):692–693, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890028/
pdf.

Feller:1987:RPP

[Fel87] Irwin Feller. R&D, Patents, and Productivity ed. by Zvi Griliches (review). Technology and Culture, 28(1):183–184, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889488/
pdf.

Ferkiss:1980:TCS

[Fer80a] Victor Ferkiss. The Technological Conscience: Survival and Dignity in an Age of Expertise by Manfred Stanley (review). Technology and Culture, 21(1):144–146, January 1980. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890642/pdf.

Ferre:1980:SSC

[Fer80b] Frederick Ferré. Suicide or Survival? The Challenge of the Year 2000 by Unesco (review). Technology and Culture, 21(1):146–148, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890643/pdf.

Ferre:1981:RPT

[Fer81] Frederick Ferré. Research in Philosophy and Technology ed. by Paul T. Durbin (review). Technology and Culture, 22(4):833–834, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890969/pdf.

Ferguson:1982:HCH

[Fer82] Eugene S. Ferguson. The Herbert Clark Hoover Collection of Mining and Metallurgy. Bibliotheca De Re Metallica by David Kuhner (review). Technology and Culture, 23(3):473–475, July 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891191/pdf.

Ferguson:1983:FHS

[Fer83] Eugene S. Ferguson. Five Hundred and Seven Mechanical Movements by Henry T. Brown (review). Technology and Culture, 24 (4):719–720, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890155/pdf.

Ferguson:1988:EEL

[Fer88] Eugene S. Ferguson. Edison’s Electric Light: Biography of an Invention by Robert Friedel, et al. (review). Technology and Culture, 29(1):153–155, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888517/pdf.

Feuerwerker:1988:MIS

[Feu88] Elie Feuerwerker. Medicine and Industrial Society: A History of Hospital Development in Manchester and Its Region, 1752–1946 by John V. Pickstone (review). Technology and Culture, 29(4): 958–959, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889192/pdf.

Field:1984:CMM

[Fie84] Mark G. Field. Chinese Medical Modernization: Comparative Policy Continuities, 1930s–1980s by AnElissa Lucas (review). Technology and Culture, 25(3):679–681, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890021/pdf.

Finn:1983:MIT

[Fin83] Bernard S. Finn. Measuring Instruments: Tools of Knowledge and Control by P. H. Sydenham (review). Technology and Culture, 24(4):722, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890158/pdf.

Finn:1989:ERT

[Fin89] Bernard S. Finn. Exhibit reviews — twenty years after. Technology and Culture, 30(4):993–1003, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888972/pdf.

Fischer:1982:NTN

[Fis82] William B. Fischer. Naturwissenschaft, Technik und NS-Ideologie: Beiträge zur Wissenschaftsgeschichte des Dritten Reiches ed. by Herbert Mehrtens, Steffen Richter (review). Technology and Culture, 23(3):485–488, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891197/pdf.

Fischer:1988:TST

[Fis88] Claude S. Fischer. “Touch someone”: The telephone industry discovers sociability. Technology and Culture, 29(1):32–61, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888496/pdf.

Fitchen:1981:PVT

[Fit81] John F. Fitchen III. The problem of ventilation through the ages. Technology and Culture, 22(3):485–511, July 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890884/pdf.

FitzSimons:1982:WTH

[Fit82] Neal FitzSimons. Water to Tyneside: A History of the Newcastle and Gateshead Water Company by R. W. Rennison (review). Technology and Culture, 23(2):249–250, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891162/pdf.

Fitzgerald:1983:AFT

[Fit83] Deborah Fitzgerald. American Farm Tools: From Hand Power to Steam Power by R. Douglas Hurt (review). Technology and Culture, 24(3):510–511, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891400/pdf.

Flader:1989:BHP

[Fla89] Susan Flader. Beauty, Health, and Permanence: Environmental Politics in the United States, 1955–1985 by Samuel P. Hays (review). Technology and Culture, 30(2):494–496, April 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889105/pdf.

Fores:1989:ST

[FLC89] Michael Fores, Edwin T. Layton, Jr., and David Channell. “Scientific technology”. Technology and Culture, 30(4):1024–1026, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888975/pdf.

Fleming:1984:AOA

[Fle84] Stuart Fleming. Ars Orientalis: The Arts of Islam, and the East (review). Technology and Culture, 25(2):325–326, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890052/pdf.

Fleming:1988:AWW

[Fle88] James R. Fleming. America’s Weather Warriors 1814–1985 by Charles C. Bates, John F. Fuller (review). Technology and Culture, 29(1):161–162, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888522/pdf.

Flick:1980:MSA

[Fli80] Carlos Flick. The movement for smoke abatement in 19th-Century Britain. Technology and Culture, 21(1):29–50, January 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890603/pdf.

Flick:1989:BSH

[Fli89a] Carlos Flick. The Big Smoke: A History of Air Pollution in London since Medieval Times by Peter Brimblecombe (review). Technology and Culture, 30(1):155–156, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889136/pdf.

Flink:1989:GPF

[Fli89b] James J. Flink. GM Passes Ford, 1918–1938: Designing the General Motors Performance-Control System by Arthur J. Kuhn (review). Technology and Culture, 30(4):1067–1068, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/888997/
pdf.

Flink:1989:JBR

[Fli89c] James J. Flink. John Bell Rae (1911–1988). Technology and Culture, 30(3):718–722, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889087/pdf.

Florman:1984:WWO

[Flo84] Samuel C. Florman. What’s Wrong with Our Technological Society — and How to Fix It by Simon Ramo (review). Technology and Culture, 25(2):319–320, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890048/pdf.

Florman:1986:EC

[Flo86] Samuel C. Florman. An engineer’s comment. Technology and Culture, 27(4):680–682, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889494/pdf.

Folsom:1980:MPP

[Fol80a] Michael Brewster Folsom. A Mechanical People: Perceptions of the Industrial Order in Massachusetts, 1815–1880 by Carl Siracusa (review). Technology and Culture, 21(3):516–517, July 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890730/pdf.

Folsom:1980:RAW

[Fol80b] Michael Brewster Folsom. Roots of the American Working Class: The Industrialization of Crafts in Newark, 1800–1860 by Susan E. Hirsch (review). Technology and Culture, 21(2):259–262, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890690/pdf.

Foley:1982:FCA

[Fol82] Vernard Foley. Firearms in Colonial America: The Impact on History and Technology, 1492–1792 by M. L. Brown (review). Technology and Culture, 23(1):118–120, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891126/pdf.

Folsom:1985:MDM

[Fol85] Michael Brewster Folsom. Mill by David Macaulay (review). Technology and Culture, 26(4):851–854, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889694/pdf.

Forman:1984:LSL

[For84] Paul Forman. “Subtle Is the Lord … ”: The Science and the Life of Albert Einstein by Abraham Pais (review). Technology and Culture, 25(3):697–698, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890030/pdf.

Fores:1988:PBM

[For88a] Michael Fores. Prometheus Bound: The Mythic Structure of Karl Marx’s Scientific Thinking by Leonard P. Wessell, Jr. (review). Technology and Culture, 29(2):298–299, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889275/pdf.

Fores:1988:TME

[For88b] Michael Fores. Transformations and the myth of “Engineering science”: Magic in a white coat. Technology and Culture, 29(1):62–81, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888497/pdf.

Foster:1980:IEH

[Fos80] Mark S. Foster. Interstate: Express Highway Politics, 1941–1956 by Mark H. Rose (review). Technology and Culture, 21(3):522–524, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890734/pdf.

Foster:1982:PCF

[Fos82] Mark S. Foster. PCC: The Car That Fought Back by Stephen P. Carlson, Fred W. Schneider III (review). Technology and Culture, 23(3):506–507, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891210/pdf.

Foster:1989:ESE

[Fos89] Mark S. Foster. The Economic and Social Effects of the Spread of Motor Vehicles ed. by Theo Barker (review). Technology and Culture, 30(2):472–474, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889091/pdf.

Fowles:1980:THB

[Fow80a] Jib Fowles. Television and Human Behavior by George Comstock, et al. (review). Technology and Culture, 21(1):142–144, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890641/pdf.

Fowles:1980:CFE

[Fow80b] Jib Fowles. The Crises in Forecasting and the Emergence of the “Prospective” Approach by Michel Godet (review). Technology and Culture, 21(4):699–700, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890786/pdf.

Fowles:1982:APU

[Fow82] Jib Fowles. ABC of Prophecy: Understanding the Environment by Barrington Nevitt (review). Technology and Culture, 23(1):96, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891113/pdf.

Frame:1984:SIA

[Fra84] Robert M. Frame III. Society for Industrial Archeology: Twelfth Annual Conference, Minneapolis–St. Paul, May 12–15, 1983. Technology and Culture, 25(3):639–643, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890001/pdf.

Foley:1983:LWL

[FRCL83] Vernard Foley, Steven Rowley, David F. Cassidy, and F. Charles Logan. Leonardo, the wheel lock, and the milling process. Technology and Culture, 24(3):399–427, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891385/pdf.

Freyer:1980:REH

[Fre80] Tony Freyer. Research in Economic History: An Annual Compilation of Research ed. by Paul Uselding (review). Technology and Culture, 21(1):102–104, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890618/pdf.

Freilich:1982:RBC

[Fre82] Morris Freilich. Royal Blue: The Culture of Construction Workers by Herbert A. Applebaum (review). Technology and Culture, 23(3):500–501, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891205/pdf.

Frey:1989:HBO

[Fre89] Robert L. Frey. History of the Baltimore and Ohio Railroad by John F. Stover (review). Technology and Culture, 30(4):1054–1055, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888988/pdf.

Friedel:1980:BDA

[Fri80a] Robert Friedel. Biographical Dictionary of American Science: The Seventeenth through the Nineteenth Centuries by Clark A. Elliott (review). Technology and Culture, 21(3):512–513, July 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890728/pdf.

Friedel:1980:GSA

[Fri80b] Robert Friedel. Guide to the Smithsonian Archives, and: Guide to Manuscript Collections in the National Museum of History and Technology (review). Technology and Culture, 21(3):509–512, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890727/pdf.

Frick:1981:MPM

[Fri81a] George F. Frick. Mr. Peale’s Museum: Charles Willson Peale and the First Popular Museum of Natural Science and Art by Charles Coleman Sellers (review). Technology and Culture, 22(2):317–318, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890857/pdf.

Fries:1981:BBA

[Fri81b] Sylvia Doughty Fries. Baltimore: The Building of an American City by Sherry H. Olson (review). Technology and Culture, 22(3):628–629, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890906/pdf.

Fries:1981:TEF

[Fri81c] Sylvia Doughty Fries. Toward the Endless Frontier: History of the Committee, on Science and Technology, 1959–79 by Ken Hechler (review). Technology and Culture, 22(1):209–211, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890833/
pdf.

Friedel:1982:AIW

[Fri82a] Robert D. Friedel. An Age of Innovation: The World of Electronics, 1930–2000 by the editors of Electronics (review). Technology and Culture, 23(1):128–129, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891134/pdf.

Fries:1982:OHO

[Fri82b] Russell I. Fries. Ontario Hydro, Ontario Heritage: A Study of Strategies for the Conservation of the Heritage of Ontario Hydro by Mark Fram (review). Technology and Culture, 23(4):661–662, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890259/pdf.

Fries:1983:SHP

[Fri83] Russell I. Fries. Small Hydroelectric Projects for Rural Development: Planning and Management ed. by Louis J. Goodman, et al. (review). Technology and Culture, 24(3):526–527, July 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891411/pdf.

Friedman:1984:CU

[Fri84a] Alan J. Friedman. The clockwork universe. Technology and Culture, 25(2):280–286, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890038/pdf.

Fries:1984:PTI

[Fri84b] Russell I. Fries. Patterns of Technological Innovation by Devendra Sahal (review). Technology and Culture, 25(2):317–318, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890047/
pdf.

Friedel:1985:AEC

[Fri85] Robert Friedel. “American enterprise” at the Cooper-hewitt: The patent models again. Technology and Culture, 26(1):69–73, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889809/pdf.

Friedel:1986:FTR

[Fri86a] Robert Friedel. Forecasting the Telephone: A Retrospective Technology Assessment by Ithiel de Sola Pool (review). Technology and Culture, 27(1):137–139, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889648/pdf.

Friedel:1986:EC

[Fri86b] Robert Friedel. Engineering in the 20th century. Technology and Culture, 27(4):669–673, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889492/pdf.

Fries:1986:MMC

[Fri86c] Sylvia D. Fries. Mechanics of the Middle Class: Work and Politics among American Engineers by Robert Zussman (review). Technology and Culture, 27(4):836–838, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889504/pdf.

Fries:1986:EME

[Fri86d] Sylvia D. Fries. Eugene M. Emme (1919–1985). Technology and Culture, 27(3):665–667, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889608/pdf.

Fries:1988:PED

[Fri88a] Sylvia D. Fries. 2001 to 1994: Political environment and the design of NASA’s Space Station system. Technology and Culture, 29(3):568–593, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889206/pdf.

Fries:1988:PHI

[Fri88b] Sylvia Doughty Fries. Public History: An Introduction ed. by Barbara J. Howe, Emory L. Kemp (review). Technology and Culture, 29(2):335–337, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889299/pdf.

Friedel:1989:IAB

[Fri89a] Robert Friedel. Inventing American Broadcasting 1899–1922 by Susan J. Douglas (review). Technology and Culture, 30(2):484–486, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889099/pdf.

Friedel:1989:HIS

[Fri89b] Robert Friedel. The History of Invention: From Stone Axes to Silicon Chips by Trevor I. Williams (review). Technology and Culture, 30(3):666–667, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889035/pdf.

Frost:1987:LSF

[Fro87] Robert L. Frost. L’Ingénieur dans la société française ed. by André Thépot (review). Technology and Culture, 28(4):856–857, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889313/pdf.

Frost:1988:LTI

[Fro88] Robert L. Frost. Labor and technological innovation in French electrical power. Technology and Culture, 29(4):865–887, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889161/pdf.

Frost:1989:FEE

[Fro89] Robert L. Frost. La France des électriciens 1880–1980 ed. by Fabienne Cardot (review). Technology and Culture, 30(1):160–162, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889140/pdf.

Fullmer:1980:TCL

[Ful80] June Z. Fullmer. Technology, chemistry, and the law in Early 19th-Century England. Technology and Culture, 21(1):1–28, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890602/pdf.

Fullmer:1981:SHA

[Ful81] J. Z. Fullmer. A Short History of Ale by Jimmy Young (review). Technology and Culture, 22(1):193–194, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890824/pdf.

Freeze:1987:TTS

[FW87] Karen J. Freeze and James C. Williams. Technology and technical sciences in history — ICOHTEC Symposium, Dresden, August 25–29, 1986. Technology and Culture, 28(4):842–849, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889307/pdf.

Galishoff:1987:FAW

[Gal87] Stuart Galishoff. Fertilizer in America: From Waste Recycling to Resource Exploitation by Richard A. Wines (review). Technology and Culture, 28(2):372–373, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889425/pdf.

Garside:1981:BIB

[Gar81a] W. R. Garside. British Industry between the Wars: Instability and Industrial Development, 1919–1939 by N. K. Buxton, D. H. Aldcroft (review). Technology and Culture, 22(2):370–371, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890881/
pdf.

Garside:1981:IRC

[Gar81b] W. R. Garside. Industrial Relations in the Coal Industry by B. J. McCormick (review). Technology and Culture, 22(1):192–193, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890823/pdf.

Garside:1981:TBS

[Gar81c] W. R. Garside. Travelling Brothers: The Six Centuries’ Road from Craft Fellowship to Trade Unionism by R. Leeson (review). Technology and Culture, 22(3):621–622, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890902/pdf.

Garside:1982:SNM

[Gar82] W. R. Garside. Strikes by Norman McCord (review). Technology and Culture, 23(3):502–503, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891207/pdf.

GarciayRobertson:1987:FHG

[Gar87a] Rodrigo Garcia y Robertson. Failure of the heavy gun at sea, 1898– 1922. Technology and Culture, 28(3):539–557, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889338/pdf.

Garland:1987:CAA

[Gar87b] Roger Garland. Clear across Australia: A History of Telecommunciations by Ann Moyal (review). Technology and Culture, 28(1):175–177, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889484/pdf.

GarciayRobertson:1988:FFM

[Gar88] Rodrigo Garcia y Robertson. Firearms and Fortifications: Military Architecture and Siege Warfare in Sixteenth-Century Siena by Simon Pepper, Nicholas Adams (review). Technology and Culture, 29(1):138–139, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888507/pdf.

Garcia-Diego:1981:EFJ

[GD81] José A. Garcia-Diego. Eugène Freyssinet by José A. Fernández Ordóñez (review). Technology and Culture, 22(4):788–790, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890949/pdf.

Gwei-Djen:1988:ORB

[GDNCH88] Lu Gwei-Djen, Joseph Needham, and Phan Chi-Hsing. The oldest representation of a bombard. Technology and Culture, 29(3):594– 605, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889208/pdf.

Geiger:1980:EDC

[Gei80] Reed Geiger. The Economic Development of Continental Europe, 1780–1870 by Alan S. Milward, S. B. Saul, and: The Development of the Economies of Continental Europe, 1850–1914 by Alan S. Milward, S. B. Saul (review). Technology and Culture, 21(3):491– 493, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890717/pdf.

Gendron:1981:TFD

[Gen81] Bernard Gendron. Technology: Fire in a Dark World by Perry Pascarella (review). Technology and Culture, 22(3):674–677, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890927/pdf.

Gerstner:1981:BIN

[Ger81] Patsy A. Gerstner. Bloodletting Instruments in the National Museum of History and Technology by Audrey Davis, Toby Appel (review). Technology and Culture, 22(1):231–232, January 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890842/pdf.

Gerstle:1984:BVS

[Ger84a] Gary Gerstle. Brass Valley: The Story of Working People’s Lives and Struggles in an American Industrial Region ed. by Jeremy Brecher, et al. (review). Technology and Culture, 25(2):348–351, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890064/pdf.

Gerstle:1984:CPS

[Ger84b] Gary Gerstle. Chanter pour survivre: Culture ouvrière, travail et techniques dans le textile Roubaix 1850–1914 by Laurent Marty (review). Technology and Culture, 25(4):863–865, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889960/
pdf.

Goldfield:1987:PPP

[GF87] David R. Goldfield and Neil Fulghum. “The promise of progress: Public works and the way we live in North Carolina”. Technology and Culture, 28(1):103–105, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889451/pdf.

Gianniny:1981:ISJ

[Gia81] O. Allan Gianniny, Jr. Iowa State Journal of Research ed. by Victor Ferkiss, Buford Norman (review). Technology and Culture, 22(3):613–615, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890898/pdf.

Giebelhaus:1981:HWO

[Gie81a] August W. Giebelhaus. History of Western Oil Shale by Paul L. Russell (review). Technology and Culture, 22(3):653–655, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890919/pdf.

Gies:1981:WPL

[Gie81b] Joseph Gies. Work and Power: The Liberation of Work and the Control of Political Power ed. by Tom R. Burns, et al. (review). Technology and Culture, 22(3):622–624, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890903/pdf.

Giebelhaus:1985:CCS

[Gie85] August W. Giebelhaus. The Challenge of Change: The Sun Oil Company, 1945–1977 by Arthur M. Johnson (review). Technology and Culture, 26(1):113–115, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889828/pdf.

Giebelhaus:1986:EHC

[Gie86] August W. Giebelhaus. Ethyl: A History of the Corporation and the People Who Made It by Joseph C. Robert (review). Technology and Culture, 27(3):637–639, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889581/pdf.

Gillmor:1982:LIS

[Gil82] C. Stewart Gillmor. L’Accueil des idées de Sadi Carnot: Et la technologie française de 1820 à 1860–de la legende à l’histoire by Pietro Redondi (review). Technology and Culture, 23(1):102–104, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891117/pdf.

Gillispie:1985:EAT

[Gil85a] Charles C. Gillispie. The Eagle Aloft: Two Centuries of the Balloon in America by Tom D. Crouch (review). Technology and Culture, 26(1):84–85, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889813/pdf.

Gilmore:1985:IEI

[Gil85b] J. E. Gilmore. Insects, Experts, and the Insecticide Crisis: The Quest for New Pest Management Strategies by John H. Perkins (review). Technology and Culture, 26(1):121–124, January 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889832/pdf.

Gates:1981:IHTa

[GK81] Roberta D. Gates and George E. Kennedy II. Interface ’79 — Humanities and Technology: Southern Technical Institute, Marietta, Georgia, October 25–26, 1979. Technology and Culture, 22(1):149–156, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890803/pdf.

Glen:1984:SCM

[Gle84] Robert Glen. The Scottish Country Miller 1700–1900: A History of Water-powered Meal Milling in Scotland by Enid Gauldie (review). Technology and Culture, 25(2):335–337, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890057/pdf.

Glick:1982:HTI

[Gli82] Thomas F. Glick. Historia de la tecnología y la invención en México by Ramón Sánchez Flores (review). Technology and Culture, 23(1):115–117, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891124/pdf.

Glick:1986:CFS

[Gli86] Thomas F. Glick. “Catalonia, Factory of Spain,” an Exhibition in Barcelona. Technology and Culture, 27(3):597–603, July 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889547/pdf.

Glick:1988:ITI

[Gli88] Thomas F. Glick. Islamic Technology: An Illustrated History by Ahmad Y. al-Hassan, Donald R. Hill (review). Technology and Culture, 29(3):668–670, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889233/pdf.

Glick:1989:PPG

[Gli89] Thomas F. Glick. Power and Penury: Government, Technology and Science in Philip II’s Spain by David C. Goodman (review). Technology and Culture, 30(3):673–674, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889040/pdf.

Goetzmann:1982:VEA

[Goe82] William H. Goetzmann. Vanguard of Expansion, Army Engineers in the Trans-Mississippi West, 1819–1879 by Frank N. Schubert (review). Technology and Culture, 23(1):125, January 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891131/pdf.

Goetzmann:1983:CVE

[Goe83] William H. Goetzmann. Concerning Vanguard of expansion. Technology and Culture, 24(2):260, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891260/pdf.

Goldfarb:1980:TLA

[Gol80a] Stephen J. Goldfarb. Theatre Lighting in the Age of Gas by Terence Rees (review). Technology and Culture, 21(1):101–102, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890617/pdf.

Goldman:1980:AAH

[Gol80b] Steven Louis Goldman. Apollo Agonistes: The Humanities in a Computerized World ed. by M. E. Grenander (review). Technology and Culture, 21(4):704–708, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890789/pdf.

Goldman:1981:EAP

[Gol81] Steven Louis Goldman. Ethics in an Age of Pervasive Technology ed. by Melvin Kranzberg (review). Technology and Culture, 22(4):829–831, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890967/pdf.

Goldenberg:1984:LOS

[Gol84] Joseph A. Goldenberg. Live Oaking: Southern Timber for Tall Ships by Virginia Steele Wood (review). Technology and Culture, 25(3):672–673, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890017/pdf.

Goldenberg:1987:CSE

[Gol87] Joseph A. Goldenberg. A Century of Service to Engineering and Shipbuilding: A Centenary History of the North East Coast Institution of Engineers and Shipbuilders 1884–1984 by J. F. Clarke (review). Technology and Culture, 28(1):151–152, January 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889470/pdf.

Goldman:1989:TET

[Gol89] Steven L. Goldman. Technological Education — Technological Style ed. by Melvin Kranzberg (review). Technology and Culture, 30(1):158–159, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889138/pdf.

Goodwin:1980:CIT

[Goo80a] Jack Goodwin. Citation Indexing — Its Theory and Application in Science, Technology, and Humanities by Eugene Garfield (review). Technology and Culture, 21(4):714–715, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890793/pdf.

Goodwin:1980:CBH

[Goo80b] Jack Goodwin. Current bibliography in the history of technology (1978). Technology and Culture, 21(2):281–355, April 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890701/pdf.

Goodwin:1981:CBH

[Goo81] Jack Goodwin. Current bibliography in the history of technology (1979). Technology and Culture, 22(2):374–484, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890883/pdf.

Goodstein:1982:ICB

[Goo82a] Judith R. Goodstein. ISIS Cumulative Bibliography, 1966–1975. Vol. 1: Personalities and Institutions ed. by John Neu (review). Technology and Culture, 23(4):702–703, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890279/pdf.

Goodwin:1982:CBH

[Goo82b] Jack Goodwin. Current bibliography in the history of technology (1980). Technology and Culture, 23(2):282–340, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891178/pdf.

Goodwin:1983:CBH

[Goo83] Jack Goodwin. Current bibliography in the history of technology (1981). Technology and Culture, 24(2):316–398, April 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891294/pdf.

Good:1985:EAS

[Goo85a] Gregory Good. Encyclopedia of Antique Scientific Instruments by John Fitz Maurice Mills (review). Technology and Culture, 26(3):634–636, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889728/pdf.

Goodway:1985:HM

[Goo85b] Martha Goodway. Historical materials at MIT. Technology and Culture, 26(2):282, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889771/pdf.

Good:1986:ENC

[Goo86] Gregory Good. The Explorers: Nineteenth Century Expeditions in Africa and the American West by Richard A. Van Orman (review). Technology and Culture, 27(1):162–164, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889661/pdf.

Goodstein:1989:NGC

[Goo89] Judith Goodstein. A New Guide to the Collections in the Library of the American Philosophical Society ed. by J. Stephen Catlett (review). Technology and Culture, 30(3):716–717, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889084/pdf.

Gorman:1982:SIC

[Gor82] Mel Gorman. Second International Congress on the Conservation of Industrial Monuments: Transactions ed. by Werner Kroker (review). Technology and Culture, 23(2):238–240, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891156/pdf.

Gordon:1983:MMR

[Gor83] Robert B. Gordon. Materials for manufacturing: The response of the Connecticut iron industry to technological change and limited resources. Technology and Culture, 24(4):602–634, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890121/
pdf.

Gordon:1984:RJC

[Gor84] Robert B. Gordon. The Road to Jaramillo: Critical Years of the Revolution in Earth Science by William Glen (review). Technology and Culture, 25(2):371–373, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890075/pdf.

Gordon:1985:IEO

[Gor85a] Robert B. Gordon. Industrial Evolution: Organization, Structure, and Growth of the Pennsylvania Iron Industry, 1750–1860 by Paul E. Paskoff (review). Technology and Culture, 26(2):306–308, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889782/pdf.

Gordon:1985:HSD

[Gor85b] Robert B. Gordon. Hydrological science and the development of waterpower for manufacturing. Technology and Culture, 26(2):204– 235, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889762/pdf.

Gordon:1986:LVE

[Gor86] Robert B. Gordon. A Legacy from Victorian Enterprise: The Briton Ferry Ironworks and the Daughter Companies by C. W. Roberts (review). Technology and Culture, 27(3):626–627, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889570/pdf.

Gordon:1987:ECP

[Gor87] Dane R. Gordon. The Education of a College President: A Memoir by James R. Killian, Jr. (review). Technology and Culture, 28(1):170–171, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889481/pdf.

Gordon:1988:FFI

[Gor88a] Robert B. Gordon. Out of the Fiery Furnace: The Impact of Metals on the History of Mankind by Robert Raymond (review). Technology and Culture, 29(3):673–675, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889236/pdf.

Gordon:1988:PMB

[Gor88b] Robert B. Gordon. The Prehistory of Metallurgy in the British Isles by R. F. Tylecote (review). Technology and Culture, 29(1): 133–134, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888504/pdf.

Gordon:1988:WTM

[Gor88c] Robert B. Gordon. Who turned the mechanical ideal into mechanical reality? Technology and Culture, 29(4):744–778, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889157/pdf.

Gordon:1989:SSA

[Gor89a] Robert B. Gordon. Sheffield Steel and America: A Century of Commercial and Technological Interdependence, 1830–1930 by Geoffrey Tweedale (review). Technology and Culture, 30(1):164–165, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889142/pdf.

Gordon:1989:SNJ

[Gor89b] Robert B. Gordon. Simeon North, John Hall, and mechanized manufacturing. Technology and Culture, 30(1):179–188, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889151/pdf.

Goto:1983:HJI

[Got83] Akira Goto. How Japan Innovates: A Comparison with the U.S. in the Case of Oxygen Steelmaking by Leonard H. Lynn (review). Technology and Culture, 24(4):700–701, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890143/pdf.

Gordon:1986:MIS

[GR86] Robert B. Gordon and Terry S. Reynolds. Medieval iron in society — Norberg, Sweden, May 6–10, 1985. Technology and Culture, 27(1):110–117, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889638/pdf.

Grant:1984:WRO

[Gra84] H. Roger Grant. Working for the Railroad: The Organization of Work in the Nineteenth Century by Walter Licht (review). Technology and Culture, 25(4):865–866, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889961/pdf.

Graham:1987:AWM

[Gra87a] Dominick Graham. Air Warfare in the Missile Age by Lon O. Nordeen, Jr. (review). Technology and Culture, 28(2):383–384, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889432/pdf.

Grant:1987:SRR

[Gra87b] H. Roger Grant. The Southern Railway: Road of the Innovators by Burke Davis (review). Technology and Culture, 28(3):723–724, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889384/pdf.

Grant:1988:AWW

[Gra88] H. Roger Grant. After the West Was Won: Homesteaders and Town-Builders in Western South Dakota, 1900–1917 by Paula M. Nelson (review). Technology and Culture, 29(3):697–698, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889252/pdf.

Greenaway:1980:SVM

[Gre80] Frank Greenaway. Science in Victorian Manchester by Robert H. Kargon (review). Technology and Culture, 21(4):660–662, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890767/pdf.

Gregori:1981:ATS

[Gre81] Thomas R. De. Gregori. Appropriate Technology and Social Values: A Critical Appraisal ed. by Franklin A. Long, Alexandra Oleson (review). Technology and Culture, 22(3):680–683, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890929/
pdf.

Greenwood:1983:RPR

[Gre83] John T. Greenwood. Red Phoenix: The Rise of Soviet Air Power, 1941–1945 by Von Hardesty (review). Technology and Culture, 24(4):703–704, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890145/pdf.

Greenberg:1984:CCD

[Gre84] Dolores Greenberg. Crisis Contained: The Department of Energy at Three Mile Island by Philip L. Cantelon and Robert C. Williams (review). Technology and Culture, 25(3):667–669, July 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890015/pdf.

Green:1988:BAG

[Gre88a] Peter Green. Bronze Age, Greek and Roman Technology: A Select, Annotated Bibliography by John Peter Oleson (review). Technology and Culture, 29(1):135–136, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888505/pdf.

Greenaway:1988:NFH

[Gre88b] Frank Greenaway. New Flame: How Gas Changed the Commercial, Domestic and Industrial Life of Britain between 1813 and 1984 by Hugh Barty-King (review). Technology and Culture, 29 (2):295–296, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889273/pdf.

Greene:1988:SSE

[Gre88c] John C. Greene. Science and Society in Early America: Essays in Honor of Whitfield J. Bell, Jr. ed. by Randolph S. Klein (review). Technology and Culture, 29(1):149–150, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888515/pdf.

Grinevald:1981:EEM

[Gri81a] Jacques Grinevald. Energy and Economic Myths: Institutional and Analytical Economic Essays by Nicholas Georgescu-Roegen (review). Technology and Culture, 22(3):655–658, July 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890920/pdf.

Grinevald:1981:ENW

[Gri81b] Jacques Grinevald. Entropy: A New World View by Jeremy Rifkin (review). Technology and Culture, 22(4):834–836, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890970/
pdf.

Gross:1981:CCC

[Gro81] Laurence F. Gross. The Cape Cod Canal by Robert H. Farson (review). Technology and Culture, 22(2):312–313, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890854/pdf.

Grob:1987:AMS

[Gro87a] Gerald N. Grob. American Medicine and Statistical Thinking, 1800–1860 by James H. Cassedy (review). Technology and Culture, 28(3):730–731, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889389/pdf.

Gross:1987:WCS

[Gro87b] Laurence F. Gross. Wool carding: A study of skills and technology. Technology and Culture, 28(4):804–827, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889303/pdf.

Gross:1988:OHS

[Gro88a] Laurence F. Gross. With Our Hands: The Story of Carpenters in Massachusetts by Mark Erlich (review). Technology and Culture, 29(4):920–921, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889169/pdf.

Gross:1988:LMC

[Gro88b] Laurence F. Gross. “Lobstering and the Maine coast” at the Maine Maritime Museum. Technology and Culture, 29(4):904–908, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889163/pdf.

Gruen:1986:MER

[Gru86] Adam L. Gruen. On Mars: Exploration of the Red Planet 1958– 1978 by Edward C. Ezell, Linda N. Ezell (review). Technology and Culture, 27(3):661–662, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889604/pdf.

Gates:1981:IHTb

[GS81] Roberta Dixon Gates and Amos St. Germain. Interface ’80 — Humanities and Technology: Southern Technical Institute, Marietta, Georgia, October 23–24, 1980. Technology and Culture, 22(4):763–770, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890941/pdf.

Guedon:1981:OED

[Gué81] Jean-Claude Guédon. The Origins and Early Development of Heavy Chemical Industry in France by John Graham Smith (review). Technology and Culture, 22(2):366–370, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890880/pdf.

Guilmartin:1983:GSS

[Gui83] John F. Guilmartin, Jr. The guns of the Santíssimo Sacramento. Technology and Culture, 24(4):559–601, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890120/pdf.

Guilmartin:1984:PPT

[Gui84] John F. Guilmartin, Jr. The Pursuit of Power: Technology, Armed Force, and Society since A.D. 1000 by William H. McNeill (review). Technology and Culture, 25(2):308–313, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890043/pdf.

Guilmartin:1989:PSM

[Gui89] John F. Guilmartin. The Porticello Shipwreck: A Mediterranean Merchant Vessel of 415–385 B.C. by Cynthia J. Eiseman, Brunilde S. Ridgway (review). Technology and Culture, 30(3):668–669, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889037/pdf.

Gunnell:1982:TIT

[Gun82] John G. Gunnell. The technocratic image and the theory of technocracy. Technology and Culture, 23(3):392–416, July 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891181/pdf.

Guthrie:1984:OEB

[Gut84] Christopher E. Guthrie. Origins of Enterprise: Business Leadership in the Industrial Revolution by Katrina Honeyman (review). Technology and Culture, 25(2):340–341, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890059/pdf.

Guthrie:1985:IPE

[Gut85] Christopher E. Guthrie. The Idea of Poverty: England in the Early Industrial Age by Gertrude Himmelfarb (review). Technology and Culture, 26(3):643–645, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889734/pdf.

Guthrie:1986:GIP

[Gut86] Christopher E. Guthrie. From Graven Images: Patterns of Modern Materialism by Chandra Mukerji (review). Technology and Culture, 27(4):908–910, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889573/pdf.

Guthrie:1987:LSF

[Gut87] Christopher E. Guthrie. L’Industrie sidérurgique en France pendant la Révolution et l’Empire by Denis Woronoff (review). Technology and Culture, 28(2):352–353, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889412/pdf.

Haber:1982:FWT

[Hab82] Samuel Haber. Frederick W. Taylor and the Rise of Scientific Management by Daniel Nelson (review). Technology and Culture, 23(2):266–267, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891170/pdf.

Habib:1985:SIA

[Hab85] Kamal Muhammad Habib. Salinity and Irrigation Agriculture in Antiquity; Diyala Basin Archaeological Projects: Report on Essential Results, 1957–58 by Thorkild Jacobsen, and: Managing Our Fresh-Water Resources. Impact of Science on Society (review). Technology and Culture, 26(1):137–139, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889841/pdf.

Hacker:1980:RAG

[Hac80a] Barton C. Hacker. Radioactivity in America: Growth and Decay of a Science by Lawrence Badash (review). Technology and Culture, 21(4):695–697, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890784/pdf.

Hackett:1980:SIR

[Hac80b] Edward J. Hackett. The Sociology of Industry by S. R. Parker, et al. (review). Technology and Culture, 21(4):666–668, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890770/pdf.

Hacker:1982:SSC

[Hac82] Barton C. Hacker. Space Science Comes of Age: Perspectives in the History of the Space Sciences ed. by Paul A. Hanle, Von Del Chamberlain (review). Technology and Culture, 23(3):518–520, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891215/pdf.

Hacker:1984:SFS

[Hac84] Barton C. Hacker. Science Fiction and Space Futures: Past and Present ed. by Eugene M. Emme (review). Technology and Culture, 25(3):690–691, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890027/pdf.

Hacker:1985:MWM

[Hac85] Sally Hacker. More Work for Mother: The Ironies of Household Technology from the Open Hearth to the Microwave by Ruth Schwartz Cowan (review). Technology and Culture, 26(2):291–293, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889774/pdf.

Hacker:1986:DSR

[Hac86] Barton Hacker. The Day the Sun Rose Twice: The Story of the Trinity Site Nuclear Explosion, July 16, 1945 by Ferenc M. Szasz (review). Technology and Culture, 27(3):614–615, July 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889556/pdf.

Hacker:1987:CAB

[Hac87] Barton G. Hacker. Controlling the Atom: The Beginnings of Nuclear Regulation 1946–1962 by George T. Mazuzan, J. Samuel Walker (review). Technology and Culture, 28(2):391–392, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889439/pdf.

Hacker:1988:UCD

[Hac88] Barton C. Hacker. Under the Cloud: The Decades of Nuclear Testing by Richard L. Miller, and: Bombs in the Backyard: Atomic Testing and American Politics by A. Costandina Titus (review).

Technology and Culture, 29(2):331–334, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889297/pdf.

Haden:1980:AGC

[Had80a] H. J. Haden. Ancient Glass in Its Context by Anita Engle (review). Technology and Culture, 21(2):235–237, April 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890679/pdf.

Haden:1980:AFM

[Had80b] H. J. Haden. Art, Feat and Mystery by H. W. Woodward (review). Technology and Culture, 21(2):253–254, April 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890687/pdf.

Haden:1981:JGS

[Had81] H. J. Haden. Journal of Glass Studies ed. by John H. Martin (review). Technology and Culture, 22(1):162–164, January 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890806/pdf.

Haden:1982:JGS

[Had82] H. J. Haden. Journal of Glass Studies ed. by John H. Martin (review). Technology and Culture, 23(2):269–271, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891172/pdf.

Hafter:1980:WWT

[Haf80] Daryl M. Hafter. Women at Work: The Transformation of Work and Community in Lowell, Massachusetts, 1826–1860 by Thomas Dublin (review). Technology and Culture, 21(4):677–679, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890775/pdf.

Hafter:1985:WTI

[Haf85a] Daryl M. Hafter. Women, Technology and Innovation. Women’s Studies International Quarterly ed. by Joan Rothschild, and: The Technological Woman: Interfacing with Tomorrow ed. by Jan Zimmerman (review). Technology and Culture, 26(2):287–290, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889773/pdf.

Hafter:1985:ICR

[Haf85b] Daryl M. Hafter. International conference on the role of women in the history of science, technology, and medicine in the 19th and 20th centuries — Veszprém, Hungary, August 15–19, 1983. Technology and Culture, 26(2):262–267, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889766/pdf.

Haines:1984:MVA

[Hai84] Charles Haines. The making of Victorian America: Henry Ford Museum’s “Mass-Produced Elegance”. Technology and Culture, 25(4):832–838, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889948/pdf.

Hallion:1980:CAH

[Hal80a] Richard P. Hallion. Chariots for Apollo: A History of Manned Lunar Spacecraft by Courtney G. Brooks, et al., and: We Have a Sporting Chance: The Decision to Go to the Moon by Leonard C. Bruno (review). Technology and Culture, 21(3):531–533, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890739/
pdf.

Hallion:1980:GLT

[Hal80b] Richard P. Hallion. Gyro! The Life and Times of Lawrence Sperry by William Wyatt Davenport (review). Technology and Culture, 21(1):132–133, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890635/pdf.

Hallion:1981:UAS

[Hal81] Richard P. Hallion. The U.S. Air Service in World War I. Vol. 3: The Battle of St. Mihiel ed. by Maurer Maurer (review). Technology and Culture, 22(2):341–343, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890868/pdf.

Hallion:1982:CAP

[Hal82] Richard P. Hallion. Curtiss Aircraft, 1907–1947 by Peter M. Bowers, and: McDonnell Douglas Aircraft since 1920 by René J. Francillon (review). Technology and Culture, 23(1):133–136, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891137/pdf.

Hallion:1984:GAP

[Hal84a] Richard P. Hallion. German Air Power in World War I by John H. Morrow, Jr. (review). Technology and Culture, 25(3):663–664, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890012/pdf.

Hallion:1984:DTC

[Hal84b] Richard P. Hallion. The Dragon’s Teeth? The Creation of United States Air Power for World War II by Benjamin S. Kelsey (review). Technology and Culture, 25(3):664–665, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890013/pdf.

Hall:1986:OBI

[Hal86a] Marie Boas Hall. Olai Borrichii Itinerarium 1660–1665: The Journal of the Danish Polyhistor Ole Borch ed. by H. D. Schepelern (review). Technology and Culture, 27(2):297–298, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889599/pdf.

Hallion:1986:IWB

[Hal86b] Richard P. Hallion. Ideas and Weapons by I. B. Holley, Jr. (review). Technology and Culture, 27(4):895–897, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889557/pdf.

Hallion:1987:LHE

[Hal87] Richard P. Hallion. … the Heavens and the Earth: A Political History of the Space Age by Walter A. McDougall (review). Technology and Culture, 28(1):130–132, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889458/pdf.

Hall:1988:LVEb

[Hal88a] Bert S. Hall. Leonardo da Vinci: Engineer and Architect ed. by Paolo Galluzzi (review). Technology and Culture, 29(3):663–666, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889231/pdf.

Hall:1988:LVEa

[Hal88b] Bert S. Hall. “Leonardo da Vinci: Engineer and Architect,” an Exhibition at the Montreal Museum of Fine Arts. Technology and Culture, 29(3):606–612, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889210/pdf.

Hall:1989:LTW

[Hal89a] Bert S. Hall. Lynn Townsend White, Jr. (1907–1987). Technology and Culture, 30(1):194–213, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889153/pdf.

Hallion:1989:AUA

[Hal89b] Richard P. Hallion. Aviation in the U.S. Army, 1919–1939 by Maurer Maurer (review). Technology and Culture, 30(4): 1075–1076, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889002/pdf.

Hamlin:1988:WDI

[Ham88] Christopher Hamlin. William dibdin and the idea of biological sewage treatment. Technology and Culture, 29(2):189–218, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889247/pdf.

Hansen:1980:ECE

[Han80] James R. Hansen. Einstein: A Centenary Exhibition ed. by the National Museum of History and Technology (review). Technology and Culture, 21(4):693–695, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890783/pdf.

Hansen:1984:ZDG

[Han84] David M. Hansen. Zalinski’s dynamite gun. Technology and Culture, 25(2):264–279, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890037/pdf.

Hanle:1985:JAS

[Han85] Paul A. Hanle. JPL and the American Space Program: A History of the Jet Propulsion Laboratory by Clayton R. Koppes (review). Technology and Culture, 26(4):877–879, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889708/pdf.

Hankins:1987:ALA

[Han87a] Thomas Hankins. Authority, Liberty, and Automatic Machinery in Early Modern Europe by Otto Mayr (review). Technology and Culture, 28(3):685–687, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889361/pdf.

Hansen:1987:QPE

[Han87b] James R. Hansen. Quest for Performance: The Evolution of Modern Aircraft by Laurence K. Loftin, Jr. (review). Technology and Culture, 28(3):734–736, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889395/pdf.

Hansen:1987:MRI

[Han87c] James R. Hansen. The Management of Research Institutions: A Look at Government Laboratories by Hans Mark, Arnold Levine (review). Technology and Culture, 28(2):387–390, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889436/pdf.

Hannaway:1988:LAP

[Han88] Owen Hannaway. Leviathan and the Air-Pump: Hobbes, Boyle and the Experimental Life by Steven Shapin, Simon Schaffer (review). Technology and Culture, 29(2):291–293, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889268/pdf.

Hancock:1989:MUL

[Han89a] John Hancock. The Modern Urban Landscape: 1880 to the Present by Edward C. Relph (review). Technology and Culture, 30(2):456– 457, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889074/pdf.

Hansen:1989:AHW

[Han89b] James R. Hansen. Aviation history in the wider view. Technology and Culture, 30(3):643–656, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889033/pdf.

Hansen:1989:HRA

[Han89c] James R. Hansen. History of Rocketry and Astronautics ed. by R. Cargill Hall (review). Technology and Culture, 30(1):170–172, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889146/pdf.

Harris:1980:WIA

[Har80a] J. R. Harris. World Industrial Archaeology by Kenneth Hudson (review). Technology and Culture, 21(2):239–241, April 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890681/pdf.

Harris:1980:CTW

[Har80b] John R. Harris. A Catalogue of Tools for Watch and Clock Makers by John Wyke (review). Technology and Culture, 21(3):495–497, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890720/pdf.

Hartenberg:1981:MIH

[Har81a] Richard S. Hartenberg. Machines: An Illustrated History by Sigvard Strandh (review). Technology and Culture, 22(3):663–666, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890923/pdf.

Hartman:1981:CWT

[Har81b] J. Paul Hartman. Curriculum Workshop on Technology for the Liberal Arts Major sponsored by the Council for the Understanding of Technology in Human Affairs: Massachusetts Institute of Technology, June 25–27, 1980. Technology and Culture, 22(4): 760–762, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890940/pdf.

Harmond:1982:AAC

[Har82a] Richard P. Harmond. The Automobile and American Culture ed. by David L. Lewis (review). Technology and Culture, 23(3):505– 506, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891209/pdf.

Harris:1982:TCI

[Har82b] J. R. Harris. Two Centuries of Industrial Welfare: The London (Quaker) Lead Company, 1692–1905 by Arthur Raistrick (review). Technology and Culture, 23(2):250–252, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891163/pdf.

Harris:1983:ECD

[Har83a] J. R. Harris. English Culture and the Decline of the Industrial Spirit, 1850–1980 by Martin J. Wiener (review). Technology and Culture, 24(1):130–132, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891237/pdf.

Harris:1983:DWM

[Har83b] Neil Harris. Dream Worlds: Mass Consumption in Late Nineteenth-Century France by Rosalind H. Williams (review). Technology and Culture, 24(3):500–502, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891395/pdf.

Harris:1984:IBB

[Har84] John R. Harris. Inside the Black Box: Technology and Economics by Nathan Rosenberg (review). Technology and Culture, 25(4): 846–847, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889950/pdf.

Harris:1985:HBC

[Har85] John Harris. The History of the British Coal Industry. Vol. 2: 1700–1830: The Industrial Revolution by Michael W. Flinn (review). Technology and Culture, 26(4):844–846, October 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889690/pdf.

Hart:1986:PTA

[Har86a] Christopher A. Hart. Proceedings from the Tenth Annual Telecommunications Policy Research Conference ed. by Oscar H. Gandy, Jr., et al. (review). Technology and Culture, 27(1):139–140, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889649/pdf.

Hartmann:1986:GRT

[Har86b] Nicholas Hartmann. Greek and Roman Technology by K. D. White (review). Technology and Culture, 27(4):854–855, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889514/
pdf.

Harris:1987:RFB

[Har87a] C. M. Harris. Robert Fulton: A Biography by Cynthia Owen Philip (review). Technology and Culture, 28(3):712–713, July 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889377/pdf.

Harris:1987:LME

[Har87b] Elizabeth Harris. The Look of Maps: An Examination of Cartographic Design by Arthur H. Robinson (review). Technology and Culture, 28(4):863–865, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889318/pdf.

Harris:1987:EES

[Har87c] J. R. Harris. Economic Expansion and Social Change: England 1500–1700. Vol. 1: People, Land and Towns; vol. 2: Industry, Trade and Government by Christopher G. A. Clay (review). Technology and Culture, 28(2):355–357, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889414/pdf.

Hartmann:1987:IHR

[Har87d] Nicholas Hartmann. Ingeniería Hidráulica Romana by Carlos Fernández Casado (review). Technology and Culture, 28(4): 851–852, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889309/pdf.

Harris:1989:CMP

[Har89a] C. M. Harris. The Correspondence and Miscellaneous Papers of Benjamin Henry Latrobe. Vol. 3: 1811–1820 ed. by John C. Van Horne, et al. (review). Technology and Culture, 30(3):689–690, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889050/pdf.

Hartmann:1989:RPF

[Har89b] Nicholas Hartmann. The Roman Port and Fishery of Cosa: A Center of Ancient Trade by Anna Marguerite McCann, et al. (review). Technology and Culture, 30(1):149–150, January 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889132/pdf.

Hassan:1986:EDB

[Has86] John A. Hassan. Economic Development of the British Coal Industry 1800–1914 by Brian R. Mitchell (review). Technology and Culture, 27(4):884–885, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889541/pdf.

Hatch:1989:CFS

[Hat89] Robert A. Hatch. Cartography in France 1660–1848: Science, Engineering, and Statecraft by Josef Konvitz (review). Technology and Culture, 30(4):1042–1043, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888981/pdf.

Hindlr:1980:NMO

[HBR^{+}80] Brooke Hindlr, James E. Brittain, Leonard S. Reich, Carroll Pursell, Eugene S. Ferguson, Arthur L. Norberg, Merritt Roe Smith, David F. Channell, Patrick M. Malone, Sylvia Doughty Fries, Rondo Cameron, Carol E. Hoffecker, Susan J. Kleinberg, John J. Beer, and Robert E. Schofield. The Newark Meeting, October 17–20, 1979. Technology and Culture, 21(3):446–471, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890709/pdf.

Headrick:1988:MMS

[Hea88] Daniel R. Headrick. Makers of Modem Strategy: From. Machiavelli to the Nuclear Age ed. by Peter Paret (review). Technology and Culture, 29(1):139–141, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/888508/pdf.

Headrick:1989:MMA

[Hea89] Daniel R. Headrick. Mechanization and Maize: Agriculture and the Politics of Technology Transfer in East Africa by Constance G. Anthony (review). Technology and Culture, 30(3):712–714, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889079/pdf.

Hein:1986:IHT

[Hei86] Virginia H. Hein. Interface ’84: Humanities and technology — Atlanta, Georgia, October 24–25, 1984. Technology and Culture, 27(1):121–122, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889640/pdf.

Heitmann:1988:RHH

[Hei88] John A. Heitmann. Rohm and Haas: History of a Chemical Company by Sheldon Hochheiser (review). Technology and Culture, 29(1):156–157, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888519/pdf.

Hensley:1986:CCA

[Hen86] John Hensley. The Computer Comes of Age: The People, the Hardware, and the Software by René Moreau (review). Technology and Culture, 27(4):868–870, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889523/pdf.

Hensley:1987:BBI

[Hen87] John Hensley. Bit by Bit: An Illustrated History of Computers by Stan Augarten (review). Technology and Culture, 28(4):893–894, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889337/pdf.

Herlea:1985:MD

[Her85] Alexandre Herlea. Maurice Daumas (1910–1984). Technology and Culture, 26(3):698–702, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889760/pdf.

Herlea:1988:SPH

[Her88] Alexandre Herlea. Les servitudes de la puissance: Une histoire de l’énergie by Jean-Claude Debeir, et al. (review). Technology and Culture, 29(1):184–186, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/888537/pdf.

Hevly:1984:RMH

[Hev84] Bruce Hevly. Revolution in Miniature: The History and Impact of Semiconductor Electronics by Ernest Braun and Stuart Macdonald (review). Technology and Culture, 25(4):882–883, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889973/pdf.

Hevly:1986:CHT

[Hev86] Bruce Hevly. The Chip: How Two Americans Invented the Microchip and Launched a Revolution by T. R. Reid (review). Technology and Culture, 27(4):873–875, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889528/pdf.

Hevly:1989:MSI

[Hev89] Bruce Hevly. Military Spending and Industrial Decline: A Study of the American Machine Tool Industry by Anthony DiFilippo (review). Technology and Culture, 30(3):709–710, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889072/pdf.

Hewlett:1983:RCG

[Hew83] Richard G. Hewlett. Rickover: Controversy and Genius by Norman Polmar, Thomas B. Allen (review). Technology and Culture, 24(2):300–302, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891286/pdf.

Hewish:1987:CCL

[Hew87] John Hewish. From cromford to chancery lane: New light on the arkwright patent trials. Technology and Culture, 28(1):80–86, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889444/pdf.

Heyman:1988:CVH

[Hey88] Jacques Heyman. Carnet de Villard de Honnecourt by Alain Erlande-Brandenburg, et al. (review). Technology and Culture, 29(1):142–144, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888510/pdf.

Hacker:1987:MET

[HH87a] Barton C. Hacker and Sally L. Hacker. Military Enterprise and Technological Change: Perspectives on the, American Experience ed. by Merritt Roe Smith (review). Technology and Culture, 28 (3):709–712, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889376/pdf.

Hacker:1987:MIL

[HH87b] Barton C. Hacker and Sally L. Hacker. Military institutions and the labor process: Noneconomic sources of technological change, women’s subordination, and the organization of work. Technology and Culture, 28(4):743–775, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889301/pdf.

Hickcox:1985:NPV

[Hic85a] David H. Hickcox. Northern Pacific Views: The Railroad Photography of F. Jay Haynes, 1876–1905 by Edward W. Nolan (review). Technology and Culture, 26(2):321–322, April 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889789/pdf.

Hickcox:1985:RSQ

[Hic85b] David H. Hickcox. Railways of Southern Quebec by J. Derek Booth (review). Technology and Culture, 26(3):647–648, July 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889736/pdf.

Higham:1980:HHS

[Hig80a] Robin Higham. HERK: Hero of the Skies by Joseph Earl Dabney (review). Technology and Culture, 21(2):264–265, April 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890692/pdf.

Higham:1980:JMH

[Hig80b] Robin Higham. Journal of Modern History (review). Technology and Culture, 21(4):635–637, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890756/pdf.

Higham:1980:UAS

[Hig80c] Robin Higham. The U.S. Air Service in World War I. Vol. 4: Post-War Review ed. by Maurer Maurer (review). Technology and Culture, 21(4):688–689, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890780/pdf.

Higham:1982:MMA

[Hig82] Robin Higham. The Messerschmitt Me-262: Arrow to the Future by Walter J. Boyne (review). Technology and Culture, 23 (1):136–138, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891138/pdf.

Higham:1983:JEA

[Hig83a] Robin Higham. Jane’s Encyclopedia of Aviation ed. by Michael J. H. Taylor (review). Technology and Culture, 24(2):297–298, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891284/pdf.

Higham:1983:TPF

[Hig83b] Robin Higham. Test Pilots — the Frontiersmen of Flight: An Illustrated History by Richard P. Hallion (review). Technology and Culture, 24(1):146–147, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891246/pdf.

Hilton:1981:MMU

[Hil81] George W. Hilton. Moving the Masses: Urban Public Transit in New York, Boston and Philadelphia, 1880–1912 by Charles W. Cheape (review). Technology and Culture, 22(1):180–181, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890816/pdf.

Hilton:1983:ITR

[Hil83] George W. Hilton. Industries in Trouble by Robert Plant (review). Technology and Culture, 24(3):517–519, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891406/pdf.

Hilton:1984:DFA

[Hil84] George W. Hilton. The Decline and Fall of the American Automobile Industry by Brock Yates (review). Technology and Culture, 25(2):363–365, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890071/pdf.

Hilton:1985:TPB

[Hil85] George W. Hilton. Trolley to the Past: A Brief History and Companion to the Operating Trolley Museums of North America by Andrew D. Young (review). Technology and Culture, 26(3):667– 668, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889749/pdf.

Hilton:1986:FAR

[Hil86] George W. Hilton. The Future of the Automobile: The Report of MIT’s International Automobile Program by Alan Altshuler, et al. (review). Technology and Culture, 27(4):894–895, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889554/
pdf.

Hill:1988:ICG

[Hil88] Donald R. Hill. Islamicate Celestial Globes: Their History, Construction, and Use by Emilie Savage-Smith (review). Technology and Culture, 29(3):670–671, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889234/pdf.

Hindle:1980:ANR

[Hin80a] Brooke Hindle. Artisans of the New Republic: The Tradesmen of New York City in the Age of Jefferson by Howard B. Rock (review). Technology and Culture, 21(3):507–508, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890725/pdf.

Hindle:1980:ECP

[Hin80b] Brooke Hindle. Eighteenth-Century Prints of Colonial America: To Educate and Decorate ed. by Joan D. Dolmetsch (review). Technology and Culture, 21(1):110–111, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890623/pdf.

Hindle:1981:TTA

[Hin81] Brooke Hindle. Tools and Technologies: Americas Wooden Age ed. by Paul B. Kebabian, William C. Lipke (review). Technology and Culture, 22(1):166–168, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890808/pdf.

Hindle:1983:TTD

[Hin83] Brooke Hindle. Technology through the 3-d time warp. Technology and Culture, 24(3):450–464, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891387/pdf.

Hindle:1984:BFS

[Hin84] Brooke Hindle. Bridge to the Future: A Symposium Commemorating the Centennial of the Brooklyn Bridge — New York City, May 18–20, 1983. Technology and Culture, 25(2):294–298, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890041/
pdf.

Hindle:1985:CIO

[Hin85] Brooke Hindle. The Coming of Industrial Order: Town and Factory Life in Rural Massachusetts, 1810–1860 by Jonathan Prude (review). Technology and Culture, 26(4):854–855, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889695/
pdf.

Hindle:1986:SHE

[Hin86a] Brooke Hindle. A Social History of England by Asa Briggs (review). Technology and Culture, 27(1):132–133, January 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889645/pdf.

Hindle:1986:MSV

[Hin86b] Brooke Hindle. Materials and Society. Vol. 7, nos. 3 and 4: Energy in American History ed. by Arthur L. Donovan (review). Technology and Culture, 27(4):861–863, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889519/pdf.

Hirsh:1984:ETH

[Hir84] Richard F. Hirsh. Energy and Transport: Historical Perspectives on Policy Issues ed. by George H. Daniels, Mark H. Rose (review). Technology and Culture, 25(2):362–363, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890070/pdf.

Hirsh:1985:ASS

[Hir85] Richard F. Hirsh. Astronomy from Space: Sputnik to Space Television ed. by James Cornell, Paul Gorenstein (review). Technology and Culture, 26(2):345–346, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889802/pdf.

Hirtle:1986:UPP

[Hir86] Peter B. Hirtle. Understanding Progress as Process: Documentation of the History of Post-War Science and Technology in the United States ed. by Clark A. Elliott (review). Technology and Culture, 27(4):899–900, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889560/pdf.

Hitz:1985:RTI

[Hit85] Elizabeth Hitz. Regional Transformation and Industrial Revolution: A Geography of the Yorkshire Woollen Industry by Derek Gregory (review). Technology and Culture, 26(1):89–90, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889816/pdf.

Hausman:1989:EEH

[HN89] William J. Hausman and John L. Neufeld. Engineers and economists: Historical perspectives on the pricing of electricity. Technology and Culture, 30(1):83–104, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889114/pdf.

Hoberman:1980:TCT

[Hob80] Louisa Schell Hoberman. Technological change in a traditional society: The case of the Desagüe in colonial Mexico. Technology and Culture, 21(3):386–407, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890703/pdf.

Hobson:1985:TCC

[Hob85] Wayne K. Hobson. Technology in culture; culture in technology: California American Studies Association Conference — California State Polytechnic University, Pomona, April 27–29, 1984. Technology and Culture, 26(1):77–81, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889811/pdf.

Hoch:1985:WTE

[Hoc85] Paul K. Hoch. Wayward Technology by Ernst Braun (review). Technology and Culture, 26(4):888–890, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889714/pdf.

Hochheiser:1986:DPI

[Hoc86] Sheldon Hochheiser. Du Pont and the International Chemical Industry by Graham D. Taylor, Patricia E. Sudnik (review). Technology and Culture, 27(3):636–637, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889580/pdf.

Hoch:1987:SST

[Hoc87] Paul K. Hoch. The Social Shaping of Technology ed. by Donald MacKenzie, Judy Wajcman (review). Technology and Culture, 28(1):132–135, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889459/pdf.

Hoch:1988:HSP

[Hoc88] Paul K. Hoch. Historical Studies in the Physical and Biological Sciences ed. by Robert W. Seidel (review). Technology and Culture, 29(4):991–993, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889223/pdf.

Hoddeson:1981:EBR

[Hod81] Lillian Hoddeson. The emergence of basic research in the Bell telephone system, 1875–1915. Technology and Culture, 22(3):512– 544, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890885/pdf.

Hodes:1989:BLS

[Hod89] Elizabeth Hodes. Beyond the Laboratory: Scientists as Political Activists in 1930s America by Peter J. Kuznick (review). Technology and Culture, 30(1):135–136, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889123/pdf.

Hoffecker:1980:ALW

[Hof80] Carol E. Hoffecker. Amoskeag: Life and Work in an American Factory City by Tamara K. Hareven, Randolph Langenbach (review). Technology and Culture, 21(4):679–681, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890776/pdf.

Hoffman:1984:LWW

[Hof84] Abraham Hoffman. Land, into Water — Water into Land: A History of Water Management in Florida by Nelson Manfred Blake (review). Technology and Culture, 25(1):160–161, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890113/
pdf.

Hofsommer:1985:BEJ

[Hof85] Don L. Hofsommer. Bonds of Enterprise: John Murray Forbes and Western Development in America’s Railway Age by John Lauritz Larson (review). Technology and Culture, 26(3):648–649, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889737/pdf.

Hoke:1980:RAS

[Hok80] Donald Hoke. The rise of the American system of manufactures, 1800–1870: Smithsonian institution, March 1978. Technology and Culture, 21(1):67–70, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890606/pdf.

Hoke:1982:HST

[Hok82] Donald Hoke. Horological Shop Tools, 1700 to 1900 by Theodore R. Crom (review). Technology and Culture, 23(2):271–272, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891173/pdf.

Hoke:1983:CGA

[Hok83] Donald Hoke. Collector’s Guide to Antique Typewriters ed. by Dan Post (review). Technology and Culture, 24(4):721, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890156/
pdf.

Hoke:1987:HTS

[Hok87] Donald Hoke. The History of the Typewriter, Successor to the Pen by George Carl Mares, and: The Typewriter Legend ed. by Frank T. Masi (review). Technology and Culture, 28(3):701–702, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889371/pdf.

Holley:1983:GSU

[Hol83] I. B. Holley, Jr. A Guide to the Sources of United States Military History: Supplement I ed. by Robin Higham, Donald J. Mrozek (review). Technology and Culture, 24(4):709–710, October 1983.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890148/
pdf.

Holl:1986:EGE

[Hol86a] Jack M. Holl. Energie in der Geschichte/Energy in History: The Topicality of the History of Technology (review). Technology and Culture, 27(3):618–619, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889561/pdf.

Holley:1986:LAA

[Hol86b] I. B. Holley, Jr. The Literature of Aeronautics, Astronautics, and Air Power by Richard P. Hallion (review). Technology and Culture, 27(4):897–899, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889559/pdf.

Holley:1987:MEK

[Hol87a] I. B. Holley, Jr. Mayday: Eisenhower, Khrushchev and the U-2 Affair by Michael R. Beschloss, and: The Militarization of Space: U.S. Policy, 1945–1984 by Paul B. Stares (review). Technology and Culture, 28(4):877–879, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889327/pdf.

Holley:1987:DDM

[Hol87b] I. B. Holley, Jr. A Detroit dream of mass-produced fighter aircraft: The XP-75 fiasco. Technology and Culture, 28(3):578–593, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889343/pdf.

Holley:1988:PCC

[Hol88] I. B. Holley, Jr. The Poisonous Cloud: Chemical Warfare in the First World War by L. F. Haber (review). Technology and Culture, 29(1):158–159, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888520/pdf.

Holmfeld:1989:STA

[Hol89] John D. Holmfeld. Science and Technology Advice to the President, Congress, and Judiciary ed. by William T. Golden (review). Technology and Culture, 30(2):501–503, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889110/pdf.

Horrigan:1984:TSC

[Hor84] Brian Horrigan. Tom Swift(R) and Company: “Boys’ Books” by Stratemeyer and Others by John T. Dizer (review). Technology and Culture, 25(1):124–126, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890093/pdf.

Horrigan:1986:DFM

[Hor86] Brian Horrigan. The Dream of the Factory-made House: Walter Gropius and Konrad Wachsmann by Gilbert Herbert (review). Technology and Culture, 27(3):643–644, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889586/pdf.

Horrigan:1988:LVA

[Hor88] Brian Horrigan. Latrobe’s View of America, 1795–1820: Selections from the Watercolors and Sketches ed. by Edward C. Carter II, et al. (review). Technology and Culture, 29(2):300–302, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889277/
pdf.

Hounshell:1980:SBI

[Hou80a] David Hounshell. The Smithsonian Book of Invention (review). Technology and Culture, 21(3):508–509, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890726/pdf.

Hounshell:1980:PRP

[Hou80b] David A. Hounshell. Public relations or public understanding?: The American industries series in scientific American. Technology and Culture, 21(4):589–593, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890750/pdf.

Hounshell:1983:TAH

[Hou83] David A. Hounshell. Technology in America: A History of Individuals and Ideas ed. by Carroll W. Pursell, Jr. (review). Technology and Culture, 24(1):120–122, January 1983. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891232/pdf.

Howard:1980:ARW

[How80a] Robert A. Howard. Adams’ Revolvers by W. H. J. Chamberlain, A. W. F. Taylerson (review). Technology and Culture, 21(3):518–519, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890731/pdf.

Howard:1980:IPR

[How80b] Robert A. Howard. Interchangeable parts revisited. Technology and Culture, 21(3):549–550, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890747/pdf.

Howe:1984:MCS

[How84] Barbara Howe. Material Culture Studies in America by Thomas J. Schlereth (review). Technology and Culture, 25(2):353–356, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890066/pdf.

Howard:1985:IWP

[How85a] Michael W. Howard. Indefensible Weapons: The Political and Psychological Case against Nuclearism by Robert Jay Lifton, Richard Falk, and: Against the State of Nuclear Terror by Joel Kovel (review). Technology and Culture, 26(2):349–352, April 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889804/pdf.

Howard:1985:SPM

[How85b] Seymour Howard. The steel pen and the modern line of beauty. Technology and Culture, 26(4):785–798, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889677/pdf.

Howell:1986:AFM

[How86] Joel D. Howell. Alexander Fleming: The Man and the Myth by Gwyn Macfarlane, and: The Enchanted Ring: The Untold Story of Penicillin by John C. Sheehan (review). Technology and Culture, 27(2):309–311, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889613/pdf.

Howard:1987:TJH

[How87] Seymour Howard. Thomas Jefferson and His Copying Machines by Silvio A. Bedini (review). Technology and Culture, 28(1):157–158, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889474/pdf.

Hoy:1985:GDP

[Hoy85] Suellen Hoy. The garbage disposer, the public health, and the good life. Technology and Culture, 26(4):758–784, October 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889676/pdf.

Hoy:1988:RMM

[Hoy88] Suellen Hoy. Remember the mobro — “Managing Urban Wastes” at the Chicago Museum of Science and Industry. Technology and Culture, 29(2):271–275, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889254/pdf.

Hitz:1982:MMO

[HRP^{+}82] Elizabeth Hitz, Margaret W. Rossiter, Theodore Z. Penn, Jeffrey L. Sturchio, Carroll Pursell, Raymond H. Merritt, Bernard S. Finn, Thomas P. Hughes, Laurence F. Gross, Bernard S. Finn, Hans J. Braun, Reese V. Jenkins, Edwin T. Layton, and Terry S. Reynolds. The Milwaukee Meeting, October 14–17, 1981. Technology and Culture, 23(3):449–469, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891189/pdf.

Hollister-Short:1981:OMR

[HS81] Graham Hollister-Short. Ores to Metals: The Rocky Mountain Smelting Industry by James E. Fell, Jr. (review). Technology and Culture, 22(4):797–798, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890953/pdf.

Hollister-Short:1988:JTC

[HS88] Graham Hollister-Short. Juanelo Turriano, Charles V’s Clockmaker: The Man and His Legend by José A. García–Diego (review). Technology and Culture, 29(2):287–289, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889264/pdf.

Hudgins:1986:ALB

[Hud86] Carter L. Hudgins. American Log Buildings: An Old World Heritage by Terry G. Jordan (review). Technology and Culture, 27 (2):316–318, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889617/pdf.

Hufbauer:1982:NPR

[Huf82] Karl Hufbauer. Nuclear Physics in Retrospect: Proceedings of a Symposium on the 1930s ed. by Roger H. Stuewer (review). Technology and Culture, 23(1):107–108, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891120/pdf.

Hughes:1981:CTH

[Hug81] Thomas P. Hughes. Convergent themes in the history of science, medicine, and technology. Technology and Culture, 22(3):550–558, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890887/pdf.

Hughes:1988:HTV

[Hug88a] Thomas P. Hughes. The History of Techniques. Vol. 1: Techniques and Civilizations; vol. 2: Techniques and Sciences ed. by Bertrand Gille (review). Technology and Culture, 29(3):688–690, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889244/pdf.

Hugill:1988:URA

[Hug88b] Peter J. Hugill. U.S. 40: A Roadscape of the American Experience by Thomas J. Schlereth (review). Technology and Culture, 29(1):170–171, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888527/pdf.

Hull:1988:BBL

[Hul88] James P. Hull. Bushworkers and Bosses: Logging in Northern Ontario, 1900–1980 by Ian Radforth (review). Technology and Culture, 29(4):921–922, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889170/pdf.

Hungerford:1984:FPL

[Hun84] Constance Cain Hungerford. French Popular Lithographic Imagery 1815–1870. Vol. 1: Lithographs and Literature by Beatrice Farwell (review). Technology and Culture, 25(1):113–115, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890086/
pdf.

Hundley:1987:REW

[Hun87a] Norris Hundley. Rivers of Empire: Water, Aridity, and the Growth of the American West by Donald Worster (review). Technology and Culture, 28(3):716–718, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889380/pdf.

Hunt:1987:CWT

[Hun87b] Bruce J. Hunt. The Continuous Wave: Technology and American Radio, 1900–1932 by Hugh G. J. Aitken (review). Technology and Culture, 28(1):173–175, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889483/pdf.

Hutson:1988:FPE

[Hut88] James H. Hutson. Franklin of Philadelphia by Esmond Wright (review). Technology and Culture, 29(1):148–149, January 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888514/pdf.

Huyssen:1982:MAT

[Huy82] Andreas Huyssen. Man in the Age of Technology by Arnold Gehlen (review). Technology and Culture, 23(1):85–88, January 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891108/pdf.

Hyde:1986:HFW

[Hyd86] Charles K. Hyde. Hopewell Furnace by W. David Lewis, Walter Hugins, and: Where Industry Failed: Water-Powered Mills at Harpers Ferry, West Virginia by Dave Gilbert, and: The Bethlehem Oil Mill 1745–1934: German Technology in Early Pennsylvania by Carter Litchfield, et al. (review). Technology and Culture, 27(1):148–150, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889654/pdf.

Hyde:1987:ACG

[Hyd87] Charles K. Hyde. Automobile and Culture by Gerald Silk, et al. (review). Technology and Culture, 28(2):378–379, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889429/pdf.

Hyde:1988:SAH

[Hyd88a] Charles K. Hyde. Streamlining America: A Henry Ford Museum Exhibit ed. by Fannia Weingartner (review). Technology and Culture, 29(1):174–175, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888530/pdf.

Hyde:1988:SAE

[Hyd88b] Charles K. Hyde. “Streamlining America,” an Exhibit at the Henry Ford Museum, Dearborn, Michigan. Technology and Culture, 29(1):125–129, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888502/pdf.

Hyde:1989:AAL

[Hyd89] Charles K. Hyde. “The Automobile in American Life,” an Exhibit at Henry Ford Museum, Dearborn, Michigan. Technology and Culture, 30(1):105–111, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889115/pdf.

Hyman:1984:HHF

[Hym84] Anthony Hyman. Herman Hollerith: Forgotten Giant of Information Processing by Geoffrey D. Austrian (review). Technology and Culture, 25(1):141–142, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890101/pdf.

Inohara:1981:EGC

[Ino81] Hideo Inohara. Entrepreneur and Gentleman: A Case History of a Japanese Company by Akira Sueno (review). Technology and Culture, 22(1):196–198, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890826/pdf.

Inohara:1982:GIR

[Ino82] Hideo Inohara. Gijitsu iteu, rekishi kara no kosatsu: America to Nippon [Technology transfer in history: comparison between American and Japanese experiences] by Tatsuya Kobayashi (review). Technology and Culture, 23(4):642–644, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890248/pdf.

Irish:1989:EHC

[Iri89a] Sharon Irish. Essays in honor of Carl W. Condit. Technology and Culture, 30(2):249–254, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889054/pdf.

Irish:1989:MML

[Iri89b] Sharon Irish. A “Machine that makes the land pay”: The West Street Building in New York. Technology and Culture, 30(2):376– 397, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889069/pdf.

Israel:1987:TEE

[Isr87] Paul B. Israel. The Telephone Enterprise: The Evolution of the Bell System’s Horizontal Structure, 1876–1909 by Robert W. Garnet, and: The Anatomy of a Business Strategy: Bell, Western Electric, and the Origins of the American Telephone Industry by George David Smith, and: From Invention to Innovation: Long-Distance Telephone Transmission at the Turn of the Century by Neil H. Wasserman (review). Technology and Culture, 28(1):164–167, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889478/pdf.

Jackson:1980:HRE

[Jac80] W. Turrentine Jackson. The Hard Rock Epic: Western Miners and the Industrial Revolution, 1860–1910 by Mark Wyman (review).

Technology and Culture, 21(4):681–684, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890777/pdf.

Jackson:1981:WWB

[Jac81] Donald Jackson. Water for the West: The Bureau of Reclamation, 1902–1977 by Michael C. Robinson (review). Technology and Culture, 22(3):635–637, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890910/pdf.

Jackson:1983:RAW

[Jac83] Donald C. Jackson. Reclaiming the American West: An Historiography and Guide by Lawrence B. Lee, and: The Leadership of the Reclamation Movement, 1875–1902 by Stanley S. Davison (review). Technology and Culture, 24(4):713–715, October 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890151/pdf.

Jackson:1984:CCT

[Jac84] Donald C. Jackson. Centennial celebrations: Two exhibits commemorating the 100th anniversary of the Brooklyn bridge. Technology and Culture, 25(2):287–291, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890039/pdf.

Jackson:1986:FFA

[Jac86] Donald C. Jackson. From the Family Farm to Agribusiness: The Irrigation Crusade in California and the West, 1850–1931 by Donald J. Pisani, and: Iron Men and Copper Wires: A Centennial History of the Southern California Edison Company by William A. Myers (review). Technology and Culture, 27(4):890–892, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889550/pdf.

Jackson:1987:MDA

[Jac87] Donald C. Jackson. A Mission in the Desert: Albuquerque District 1935–1985 by Michael E. Welsh (review). Technology and Culture, 28(3):720, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889382/pdf.

Jackson:1988:HHH

[Jac88] W. Turrentine Jackson. Harbor of the Heartlands: A History of the Inland Seaport of Stockton, California, from the Gold Rush to 1985 by Nicholas P. Hardeman (review). Technology and Culture, 29(2):322–323, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889291/pdf.

Jackson:1989:HDA

[Jac89a] Donald C. Jackson. Hoover Dam: An American Adventure by Joseph E. Stevens (review). Technology and Culture, 30(4): 1062–1064, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888994/pdf.

Jackson:1989:FWP

[Jac89b] Donald C. Jackson. “The Fairmount Waterworks, 1812–1911,” at the Philadelphia Museum of Art. Technology and Culture, 30(3):635–639, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889031/pdf.

Jaffee:1984:CPS

[Jaf84] David Jaffee. A Calculating People: The Spread of Numeracy in Early America by Patricia Cline Cohen (review). Technology and Culture, 25(4):875–876, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889968/pdf.

Jakle:1987:RVT

[Jak87] John Jakle. Recreational Vehicles and Travel: A Resource Guide by Bernard Mergen (review). Technology and Culture, 28(3):700– 701, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889370/pdf.

Jamison:1989:TTC

[Jam89] Andrew Jamison. Technology’s theorists: Conceptions of innovation in relation to science and technology policy. Technology and Culture, 30(3):505–533, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889005/pdf.

Janson:1986:ATR

[Jan86] Richard Janson. The Archaeology of the Transport Revolution 1750–1850 by P. J. G. Ransom (review). Technology and Culture, 27(3):621–623, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889565/pdf.

Janson:1987:MOI

[Jan87] R. H. Janson. Maskin och idyll: Teknik och pastorala ideal hos Strindberg och Heidenstam [Machine and idyll: Technology and pastoral ideals in Strindberg and Heidenstam] by Martin Kylhammar (review). Technology and Culture, 28(4):857–858, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889314/pdf.

Jason:1987:ESA

[Jas87] Gary Jason. Expert Systems: Artificial Intelligence in Business by Paul Harmon, David King (review). Technology and Culture, 28(4):894–895, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889339/pdf.

Jenkins:1984:EIB

[Jen84] Reese V. Jenkins. Emulation and Invention by Brooke Hindle (review). Technology and Culture, 25(3):660–662, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890010/pdf.

Jervis:1982:DNW

[Jer82] Jane L. Jervis. The Death of Nature: Women, Ecology, and the Scientific Revolution by Carolyn Merchant (review). Technology and Culture, 23(4):653–654, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890254/pdf.

Jeremy:1985:ETP

[Jer85] David J. Jeremy. European Textile Printers in the Eighteenth Century: A Study of Peel and Oberkampf by Stanley D. Chapman, Serge Chassagne, and: Cloth and Clothing in Medieval Europe: Essays in Memory of Professor E. M. Carus-Wilson ed. by N. B. Harte, K. G. Ponting, and: Medieval English Clothmaking: An Economic Survey by A. R. Bridbury, and: The British Wool Textile Industry 1770–1914 by David T. Jenkins, K. G. Ponting (review). Technology and Culture, 26(4):837–840, October 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889687/pdf.

Jeremy:1988:ECA

[Jer88] David J. Jeremy. Engines of Change: The American Industrial Revolution, 1790–1860 by Brooke Hindle, Steven Lubar (review). Technology and Culture, 29(1):151–152, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888516/pdf.

Johnson:1983:WRF

[Joh83] Christopher H. Johnson. Work and Revolution in France: The Language of Labor from the Old Regime to 1848 by William H. Sewell, Jr. (review). Technology and Culture, 24(4):675–677, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890129/pdf.

John:1987:SRS

[Joh87a] Richard R. John, Jr. Sharps Rifles and Spanish Mules: The San Antonio — El Paso Mail, 1851–1881 by Wayne R. Austerman (review). Technology and Culture, 28(3):715–716, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889379/pdf.

Johns:1987:SPP

[Joh87b] Gerald Johns. A Spacefaring People: Perspectives on Early Spaceflight ed. by Alex Roland (review). Technology and Culture, 28 (2):386–387, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889434/pdf.

Johnson:1988:TTD

[Joh88] Leland R. Johnson. TVA and the Tellico Dam 1936–1979: A Bureaucratic Crisis in Post-Industrial America by William Bruce Wheeler, Michael J. McDonald (review). Technology and Culture, 29(2):320–322, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889290/pdf.

John:1989:CMC

[Joh89] Richard R. John, Jr. Coming Over: Migration and Communication between England and New England in the Seventeenth Century by David Cressy, and: English Atlantic 1675–1740: An Exploration of Communication and Community by Ian K. Steele (review). Technology and Culture, 30(3):682–685, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889047/pdf.

Jones:1981:STH

[Jon81] Thomas B. Jones. Science, technology, and human values: Spring Hill Center, Wayzata, Minnesota, April 13–16, 1980. Technology and Culture, 22(3):598–604, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890893/pdf.

Jordan:1989:SIW

[Jor89] John M. Jordan. “Society improved the way you can improve a dynamo”: Charles P. Steinmetz and the politics of efficiency. Technology and Culture, 30(1):57–82, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889113/pdf.

Kahrl:1985:TBS

[Kah85] William Kahrl. The Tule Breakers: The Story of the California Dredge by John Thompson, Edward A. Dutra (review). Technology and Culture, 26(2):336–337, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889797/pdf.

Kammen:1981:RLA

[Kam81] Michael Kammen. A Russian Looks at America: The Journey of Aleksandr Borisovich Lakier in 1857 ed. by Arnold Schrier, Joyce Story (review). Technology and Culture, 22(1):168–169, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890809/pdf.

Kantrowitz:1985:MEF

[Kan85] Arthur Kantrowitz. Macro-Engineering and the Future: A Management Perspective ed. by Frank P. Davidson and C. Lawrence Meador (review). Technology and Culture, 26(1):117–120, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889830/pdf.

Kapstein:1980:GTY

[Kap80] Ethan B. Kapstein. A Golden Thread: 2500 Years of Solar Architecture and Technology by Ken Butti, John Perlin (review). Technology and Culture, 21(4):650–652, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890762/pdf.

Kapstein:1981:SC

[Kap81] Ethan Barnaby Kapstein. The solar cooker. Technology and Culture, 22(1):112–121, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890800/pdf.

Kapstein:1982:TWR

[Kap82] Ethan B. Kapstein. Thomas Welcome Roys: America’s Pioneer of Modern Whaling by Frederick P. Schmitt, et al. (review). Technology and Culture, 23(1):124–125, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891130/pdf.

Kaplan:1986:SSC

[Kap86] Bonnie Kaplan. The Second Self: Computers and the Human Spirit by Sherry Turkle (review). Technology and Culture, 27(4):870–872, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889525/pdf.

Kargon:1981:ROL

[Kar81] Robert Kargon. Robert Oppenheimer: Letters and Recollections ed. by Alice Kimball Smith, Charles Weiner (review). Technology and Culture, 22(3):649–651, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890917/pdf.

Karrow:1983:WJM

[Kar83] Robert W. Karrow, Jr. Was John Marmin’s “Self-Propelled Vehicle” a sailcar? Technology and Culture, 24(3):493–494, July 1983.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891392/
pdf.

Kargon:1986:STM

[Kar86] Robert Kargon. Science and Technology in Manchester: Two Hundred Years of the Lit. and Phil by Chris E. Makepeace (review). Technology and Culture, 27(2):304–305, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889610/pdf.

Karrow:1987:AMM

[Kar87a] Robert W. Karrow, Jr. American Maps and Mapmakers: Commercial Cartography in the Nineteenth Century by Walter W. Ristow (review). Technology and Culture, 28(4):862–863, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889317/pdf.

Karrow:1987:TTC

[Kar87b] Robert W. Karrow, Jr. Technological Transition in Cartography by Mark Stephen Monmonier (review). Technology and Culture, 28(1):181–183, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889487/pdf.

Kasson:1983:PME

[Kas83] John F. Kasson. The Philosophy of Manufactures: Early Debates over Industrialization in the United States ed. by Michael Brewster Folsom, Steven D. Lubar (review). Technology and Culture, 24(4):680–681, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890131/pdf.

Kasson:1984:CWP

[Kas84] John F. Kasson. Charles Willson Peale and His World by Edgar P. Richardson, Brooke Hindle, and Lillian B. Miller (review). Technology and Culture, 25(4):876–878, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889969/pdf.

Keating:1989:ABM

[Kea89] Ann Durkin Keating. The Ambassador Bridge: A Monument to Progress by Philip P. Mason (review). Technology and Culture, 30(2):475–476, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889093/pdf.

Kehoe:1984:WSC

[Keh84] Alice B. Kehoe. When Stars Came Down to Earth: Cosmology of the Skidi Pawnee Indians of North America by Von Del Chamberlain (review). Technology and Culture, 25(4):873–875, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889967/pdf.

Keller:1980:BID

[Kel80] Alex Keller. The Book of Ingenious Devices (Kitāb al Hiyal) by the Banū (Sons of) Mūsà bin \underline{Sh}ākir (review). Technology and Culture, 21(2):231–233, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890677/pdf.

Keller:1982:CWC

[Kel82] Alex Keller. The Cheese and the Worms: The Cosmos of a Sixteenth Century Miller by Carlo Ginzburg (review). Technology and Culture, 23(4):650–651, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890252/pdf.

Keller:1984:TPS

[Kel84a] Alex Keller. La tecnología portuaria del Siglo de Oro Español by Modesto Vigueras González (review). Technology and Culture, 25 (2):332–334, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890055/pdf.

Keller:1984:TIC

[Kel84b] Alex Keller. The Technological Illustrations of the So-called “Anonymous of the Hussite Wars” by Bert S. Hall (review). Technology and Culture, 25(1):109–111, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890084/pdf.

Kelly:1984:IHT

[Kel84c] Rebecca Kelly. Interface ’83 — Humanities and Technology: Southern Technical Institute, Marietta, Georgia, October 20–21, 1983. Technology and Culture, 25(4):839–845, October 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889949/pdf.

Keller:1986:MHW

[Kel86] Alex Keller. The “Man and His Work” Center at the Ha’aretz Museum, Tel Aviv. Technology and Culture, 27(3):594–596, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889545/
pdf.

Keller:1988:NTM

[Kel88] Alex Keller. Novo Teatro di Machine et Edificii 1607 by Vittorio Zonca (review). Technology and Culture, 29(2):285–287, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889262/pdf.

Keller:1989:HAW

[Kel89] Alex Keller. Harvesting the Air: Windmill Pioneers in Twelfth-Century England by Edward J. Kealey (review). Technology and Culture, 30(1):153–155, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889135/pdf.

Kemp:1984:HTS

[Kem84a] Emory L. Kemp. A History of Theory of Structures in the Nineteenth Century by T. M. Charlton (review). Technology and Culture, 25(4):898–899, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889982/pdf.

Kemp:1984:AVF

[Kem84b] Emory L. Kemp. America’s Valley Forges and Valley Furnaces by J. Lawrence Pool (review). Technology and Culture, 25(2):347–348, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890063/pdf.

Kemp:1986:BAD

[Kem86a] Emory Kemp. Bridges: Aesthetics and Design by Fritz Leonhardt (review). Technology and Culture, 27(2):327–328, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889623/pdf.

Kemp:1986:EHR

[Kem86b] Emory L. Kemp. To Engineer Is Human: The Role of Failure in Successful Design by Henry Petroski (review). Technology and Culture, 27(4):846–848, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889510/pdf.

Kemp:1986:AEO

[Kem86c] Louis Ward Kemp. Aesthetes and engineers: The occupational ideology of highway design. Technology and Culture, 27(4):759–797, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889498/pdf.

Kennedy:1986:CMP

[Ken86] Roger G. Kennedy. The Correspondence and Miscellaneous Papers of Benjamin Henry Latrobe. Vol. 1: 1784–1804 ed. by John C. Van Horne, Lee W. Formwalt (review). Technology and Culture, 27(1):150–153, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889655/pdf.

Ketcham:1985:BFB

[Ket85] Ralph Ketcham. Benjamin Franklin: A Biography by Ronald W. Clark (review). Technology and Culture, 26(3):641–642, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889732/
pdf.

Kidwell:1985:MSC

[Kid85] Peggy Kidwell. Mary Somerville and the Cultivation of Science, 1815–1840 by Elizabeth Patterson (review). Technology and Culture, 26(1):128–129, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889835/pdf.

Kidwell:1987:TPE

[Kid87] Clara Sue Kidwell. Trans-Pacific Echoes and Resonances; Listening Once Again by Joseph Needham, Lu Gwei-Djen (review). Technology and Culture, 28(3):687–688, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889362/pdf.

Kidwell:1988:RMP

[Kid88] Peggy Kidwell. Reflexions on the Motive Power of Fire: A Critical Edition with the Surviving Manuscripts by Sadi Carnot (review). Technology and Culture, 29(1):144–145, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888511/pdf.

Kidwell:1989:VDQ

[Kid89] Peggy Kidwell. The Visual Display of Quantitative Information by Edward R. Tufte (review). Technology and Culture, 30(1):175–176, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889149/pdf.

Kielbowicz:1986:EMG

[Kie86] Richard B. Kielbowicz. Electronic Media: A Guide to Trends in Broadcasting and Newer Technologies 1920–1983 by Christopher H. Sterling (review). Technology and Culture, 27(4):867–868, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889522/pdf.

Kielbowicz:1987:NGM

[Kie87] Richard B. Kielbowicz. News gathering by mail in the age of the telegraph: Adapting to a new technology. Technology and Culture, 28(1):26–41, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889438/pdf.

Kielbowicz:1989:ATS

[Kie89] Richard B. Kielbowicz. The American Telegrapher: A Social History, 1860–1900 by Edwin Gabler (review). Technology and Culture, 30(2):483, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889098/pdf.

Killick:1988:BBT

[Kil88] David Killick. The Blooms of Banjeli: Technology and Gender in West African Ironmaking by Carlyn Saltman, et al. (review). Technology and Culture, 29(4):932–934, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889177/pdf.

King:1981:AAE

[Kin81] David A. King. Astronomy of the Ancients ed. by K. Brecher, M. Feirtag (review). Technology and Culture, 22(2):300–301, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890848/pdf.

Kirchner:1980:CSP

[Kir80] Walther Kirchner. Cities and the Sea: Port City Planning in Early Modern Europe by Josef W. Konvitz (review). Technology and Culture, 21(4):646–647, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890760/pdf.

Klepper:1987:BMA

[Kle87] David L. Klepper. Buildings for Music: The Architect, the Musician, and the Listener from the Seventeenth Century to the Present Day by Michael Forsyth (review). Technology and Culture, 28 (1):143–145, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889465/pdf.

Klein:1989:GSR

[Kle89] Maury Klein. A Great and Shining Road: The Epic Story of the Transcontinental Railroad by John Hoyt Williams (review). Technology and Culture, 30(4):1055–1057, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888989/pdf.

Kline:1986:HES

[Kli86a] Ronald R. Kline. A History of Engineering and Science in the Bell System: Switching Technology (1925–1975) by A. E. Joel, Jr. (review). Technology and Culture, 27(4):865–866, October 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889521/pdf.

Kline:1986:IPB

[Kli86b] Ronald R. Kline. Industriekultur: Peter Behrens and the AEG, 1907–1914 by Tilmann Buddensieg (review). Technology and Culture, 27(1):172–173, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889666/pdf.

Kline:1986:RES

[Kli86c] Ronald R. Kline. The Revolt of the Engineers: Social Responsibility and the American Engineering Profession by Edwin T. Layton, Jr. (review). Technology and Culture, 27(4):835–836, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889503/pdf.

Kline:1987:SET

[Kli87a] Ronald Kline. Science and engineering theory in the invention and development of the induction motor, 1880–1900. Technology and Culture, 28(2):283–313, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889394/pdf.

Kline:1987:CEE

[Kli87b] Ronald R. Kline. A Century of Electrical Engineering and Computer Science at MIT, 1882–1982 by Karl L. Wildes, Nilo A. Lindgren (review). Technology and Culture, 28(1):169–170, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889480/pdf.

Klunder:1982:BCS

[Klu82] Willard C. Klunder. To Build a Canal: Sault Ste. Marie, 1853– 1854 and After by John N. Dickinson (review). Technology and Culture, 23(4):672–673, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890265/pdf.

Knight:1981:HWH

[Kni81] Thomas J. Knight. A History of the World by Hugh Thomas (review). Technology and Culture, 22(1):160–162, January 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890805/pdf.

Kobayashi:1980:CSD

[Kob80] Tatsuya Kobayashi. Characteristics of Scientific Development in Japan by Shigeru Nakayama (review). Technology and Culture, 21(1):156–158, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890648/pdf.

Koenigsberger:1984:MWS

[Koe84] Dorothy Koenigsberger. Machaut’s World: Science and Art in the Fourteenth Century ed. by Madeleine Cosman, Bruce Chandler (review). Technology and Culture, 25(2):328–332, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890054/pdf.

Kohlmeyer:1980:BFR

[Koh80] Fred W. Kohlmeyer. Beyond the Fence Rows: A History of Farmland Industries, Inc., 1929–1978 by Gilbert C. Fite (review). Technology and Culture, 21(1):111–113, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890624/pdf.

Kohlstedt:1987:OEM

[Koh87] Sally Gregory Kohlstedt. Objects and Others: Essays on Museums and Material Culture ed. by George W. Stocking, Jr. (review). Technology and Culture, 28(3):690–692, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889364/pdf.

Konvitz:1984:EPW

[Kon84] Josef W. Konvitz. Essays in Public Works History ed. by Suellen M. Hoy, et al. (review). Technology and Culture, 25(2):356–357, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890067/pdf.

Konvitz:1989:WJW

[Kon89] Josef W. Konvitz. William J. Wilgus and engineering projects to improve the Port of New York, 1900–1930. Technology and Culture, 30(2):398–425, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889071/pdf.

Koppes:1982:BAE

[Kop82] Clayton R. Koppes. Beyond the Atmosphere: Early Years of Space Science by Homer E. Newell (review). Technology and Culture, 23(3):516–518, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891214/pdf.

Koppes:1989:ECH

[Kop89] Clayton R. Koppes. Engineer in Charge: A History of the Langley Aeronautical Laboratory 1917–1958 by James R. Hansen (review). Technology and Culture, 30(2):488–490, April 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889101/pdf.

Kornwolf:1980:TGF

[Kor80] James D. Kornwolf. Three Generations in a Family Textile Firm by Jocelyn Morton (review). Technology and Culture, 21(4):658–660, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890766/pdf.

Korstad:1986:RJR

[Kor86] Robert Korstad. The R. J. Reynolds Tobacco Company by Nannie M. Tilley (review). Technology and Culture, 27(3):639–640, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889583/pdf.

Kouwenhoven:1981:PNY

[Kou81] John A. Kouwenhoven. The Port of New York: A History of the Rail and Terminal System from the Beginnings to Pennsylvania Station by Carl W. Condit (review). Technology and Culture, 22(1):181–183, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890817/pdf.

Kouwenhoven:1982:DEB

[Kou82] John A. Kouwenhoven. The designing of the Eads bridge. Technology and Culture, 23(4):535–568, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890239/pdf.

Kratz:1980:NGR

[KP80] O. Krätz and Claus Priesner. Neptune’s Gift by Robert P. Multhauf (review). Technology and Culture, 21(4):648–650, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890761/pdf.

Kanefsky:1980:SEC

[KR80] John Kanefsky and John Robey. Steam engines in 18th-Century Britain: A quantitative assessment. Technology and Culture, 21 (2):161–186, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890672/pdf.

Krammer:1981:NTN

[Kra81a] Arnold Krammer. Naturwissenschaft, Technik und NS-Ideologie: Beiträge zur Wissenschaftsgeschichte des Dritten Reiches ed. by Herbert Mehrtens, Steffen Richter (review). Technology and Culture, 22(4):825–828, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890966/pdf.

Krammer:1981:TTW

[Kra81b] Arnold Krammer. Technology transfer as war booty: The U.S. Technical Oil Mission to Europe, 1945. Technology and Culture, 22(1):68–103, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890798/pdf.

Kranzberg:1981:PB

[Kra81c] Melvin Kranzberg. Passing the baton. Technology and Culture, 22(4):695–699, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890935/pdf.

Kranakis:1982:FCG

[Kra82] Eda Fowlks Kranakis. The French connection: Giffard’s injector and the nature of heat. Technology and Culture, 23(1):3–38, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891102/pdf.

Kranzberg:1984:LGW

[Kra84] Melvin Kranzberg. Let’s not get wrought up about it. Technology and Culture, 25(4):735–749, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889944/pdf.

Kranakis:1985:TIR

[Kra85] Eda Fowlks Kranakis. La Technologie introuvable: recherche sur la définition et l’unité de la technologie à partir de quelques modèles du XVIII^{e} et XIX^{e} siècles by Jean-Claude Beaune, and: L’Automate et ses mobiles by Jean-Claude Beaune (review). Technology and Culture, 26(3):685–688, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889758/pdf.

Kranzberg:1986:THK

[Kra86a] Melvin Kranzberg. Technology and history: “Kranzberg’s laws”. Technology and Culture, 27(3):544–560, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889531/pdf.

Kraut:1986:EHW

[Kra86b] Alan M. Kraut. Eight Hours for What We Will: Workers and Leisure in an Industrial City, 1870–1920 by Roy Rosenzweig (review). Technology and Culture, 27(4):877–879, October 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889532/pdf.

Kranakis:1988:HEP

[Kra88a] Eda Kranakis. Histoire de l’École Polytechnique by Ambroise Fourcy (review). Technology and Culture, 29(4):952–953, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889188/pdf.

Kranakis:1988:HCE

[Kra88b] Eda Kranakis. The History of Civil Engineering since 1600: An Annotated Bibliography by Darwin H. Stapleton, Roger L. Shumaker (review). Technology and Culture, 29(1):145–146, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888512/pdf.

Kranakis:1989:DTS

[Kra89a] Eda Kranakis. Diesel: Technology and Society in Industrial Germany by Donald E. Thomas, Jr. (review). Technology and Culture, 30(4):1046–1049, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/888984/pdf.

Kranzberg:1989:TCW

[Kra89b] Melvin Kranzberg. A tribute to Carl W. Condit. Technology and Culture, 30(2):255–257, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889056/pdf.

Kronick:1986:BDN

[Kro86] Jane C. Kronick. Beyond Dumping: New Strategies for Controlling Toxic Contamination ed. by Bruce Piasecki (review). Technology and Culture, 27(4):905–907, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889568/pdf.

Krueger:1985:TDR

[Kru85] Thomas A. Krueger. TVA and the Dispossessed: The Resettlement of Population in the Norris Dam Area by Michael J. McDonald, John Muldowny (review). Technology and Culture, 26(3):651–653, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889739/pdf.

Kennedy:1980:IHT

[KT80] George E. Kennedy II and John S. Tumlin. Interface ’78 — Humanities and Technology: Southern Technical Institute, Marietta, Georgia, October 1978. Technology and Culture, 21(1):71–76, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890607/pdf.

Kulik:1988:HWG

[Kul88] Gary Kulik. How the West Grew Rich: The Economic Transformation of the Industrial World by Nathan Rosenberg, L. E. Birdzell, Jr., and: The Age of Manufactures 1700–1820: Industry, Innovation and Work in Britain by Maxine Berg (review). Technology and Culture, 29(4):962–964, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889195/pdf.

Langenbruch:1980:ETP

[Lan80a] Theodor Langenbruch. Einführung in die Technik-Philosophie von Ernst Kapp [Introduction: new edition of Ernst Kapp, Grundlinien einer Philosophie der Technik] by Hans-Martin Sass (review). Technology and Culture, 21(2):276–278, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890699/pdf.

Lankton:1980:SOS

[Lan80b] Larry Lankton. Something old, something new: The reexhibition of the Henry Ford Museum’s Hall of Technology. Technology and Culture, 21(4):594–613, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890751/pdf.

Lankton:1983:IJS

[Lan83a] Larry Lankton. IA: The Journal of the Society for Industrial Archeology. Vol. 8, no. 1 (review). Technology and Culture, 24 (3):516–517, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891405/pdf.

Lankton:1983:NEM

[Lan83b] Larry D. Lankton. The New England Mill Village, 1790–1860 ed. by Gary Kulik, et al. (review). Technology and Culture, 24 (4):682–683, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890132/pdf.

Lankton:1983:MUG

[Lan83c] Larry D. Lankton. The machine under the garden: Rock drills arrive at the Lake Superior copper mines, 1868–1883. Technology and Culture, 24(1):1–37, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891221/pdf.

Langley:1984:DLD

[Lan84] Harold D. Langley. Dogwatch and Liberty Days: Seafaring Life in the Nineteenth Century by Margaret S. Creighton, and: Starving Sailors: The Influence of Nutrition upon Naval and Maritime History ed. by J. Watt, E. J. Freeman, and W. F. Bynum (review). Technology and Culture, 25(4):867–868, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889962/pdf.

Langins:1985:MBI

[Lan85] Janis Langins. The Montgolfıer Brothers and the Invention of Aviation, 1783–1784; with a Word on the Importance of Ballooning for the Science of Heat and the Art of Building Railroads by Charles Coulston Gillispie (review). Technology and Culture, 26 (1):85–87, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889814/pdf.

Lankton:1986:MTP

[Lan86] Larry D. Lankton. Mining Town: The Photographic Record of T. N. Barnard and Nellie Stockbridge from the Coeur d’Alenes by Patricia Hart, Ivar Nelson (review). Technology and Culture, 27(4):886–887, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889544/pdf.

Landels:1987:HEC

[Lan87a] J. G. Landels. A History of Engineering in Classical and Medieval Times by Donald Hill (review). Technology and Culture, 28(4):850–851, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889308/pdf.

Langins:1987:HFM

[Lan87b] Janis Langins. A History of French Military Aviation by Charles Christienne, Pierre Lissarrague (review). Technology and Culture, 28(3):738–739, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889399/pdf.

Langins:1987:SWV

[Lan87c] Janis Langins. Siege Warfare. Vol. 2: The Fortress in the Age of Vauban and Frederick the Great 1660–1789 by Christopher Duffy (review). Technology and Culture, 28(2):350–351, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889411/pdf.

Lankford:1987:PCT

[Lan87d] John Lankford. Photography and the 19th-Century transits of Venus. Technology and Culture, 28(3):648–657, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889354/pdf.

Lankton:1988:HNI

[Lan88] Larry Lankton. The Hawk’s Nest Incident: America’s Worst Industrial Disaster by Martin Cherniack (review). Technology and Culture, 29(4):931–932, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889176/pdf.

Langins:1989:EIW

[Lan89a] Janis Langins. Education for the Industrial World: The Écoles d’Arts et Métiers and the Rise of French Industrial Engineering by Charles R. Day (review). Technology and Culture, 30(1):159–160, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889139/pdf.

Lankton:1989:MAI

[Lan89b] Larry Lankton. Mining America: The Industry and the Environment, 1800–1980 by Duane A. Smith (review). Technology and Culture, 30(2):496–497, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889106/pdf.

Latour:1986:PST

[Lat86] Bruno Latour. People, Science and Technology by Charles Boyle, et al. (review). Technology and Culture, 27(4):910–911, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889575/pdf.

Laudan:1982:MST

[Lau82] Rachel Laudan. Models of scientific and technological change: Center for Philosophy of Science, University of Pittsburgh, April 9–12, 1981. Technology and Culture, 23(1):78–80, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891106/
pdf.

Laudan:1986:HTE

[Lau86] Rachel Laudan. History of Technology: Eighth Annual Volume, 1983 ed. by Norman Smith (review). Technology and Culture, 27(4):851–853, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889512/pdf.

Laurie:1987:SSO

[Lau87] Bruce Laurie. Samuel Slater and the Origins of the American Textile Industry, 1790–1860 by Barbara M. Tucker (review). Technology and Culture, 28(1):155–157, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889473/pdf.

Lavallee:1984:OSR

[Lav84] Omer Lavallée. The Origins of the Scottish Railway System 1722– 1844 by C. J. A. Robertson (review). Technology and Culture, 25(3):653–655, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890006/pdf.

Lawrence:1983:MTA

[Law83] Christopher Lawrence. Medicine and Its Technology by Audrey B. Davis (review). Technology and Culture, 24(3):543–545, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891423/
pdf.

Lawless:1985:ITT

[Law85] Robert Lawless. Information Technology in the Third World: Can I.T. Lead to Humane National Development by William James Stover (review). Technology and Culture, 26(1):106–109, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889825/pdf.

Law:1987:SET

[Law87] John Law. On the social explanation of technical change: The case of the Portuguese maritime expansion. Technology and Culture, 28(2):227–252, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889388/pdf.

Lawler:1989:IMD

[Law89] Nan Lawler. Iron Monuments to Distant Posterity: Indiana’s Metal Bridges, 1870–1930 by James L. Cooper (review). Technology and Culture, 30(2):474–475, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889092/pdf.

Layton:1987:TLG

[Lay87] Edwin T. Layton, Jr. Through the looking glass, or news from lake mirror image. Technology and Culture, 28(3):594–607, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889345/
pdf.

Layton:1988:SFA

[Lay88] Edwin T. Layton. Science as a form of action: The role of the engineering sciences. Technology and Culture, 29(1):82–97, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888498/pdf.

Lewis:1982:KUI

[LB82] W. David Lewis and Guy V. Beckwith. The Known and the Unknown: The Iconography of Science Fiction by Gary K. Wolfe (review). Technology and Culture, 23(4):697–698, October 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890277/pdf.

Lechtman:1984:AVS

[Lec84] Heather Lechtman. Andean value systems and the development of prehistoric metallurgy. Technology and Culture, 25(1):1–36, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890078/pdf.

Lechtman:1988:RGA

[Lec88] Heather Lechtman. Red, Gold of Africa: Copper in Precolonial History and Culture by Eugenia W. Herbert (review). Technology and Culture, 29(1):130–133, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/888503/pdf.

Lehmann:1983:BPL

[Leh83] Fritz Lehmann. Beyer, Peacock: Locomotive Builders to the World by R. L. Hills, D. Patrick (review). Technology and Culture, 24 (4):695–696, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890140/pdf.

Lehmann:1986:EWP

[Leh86] Fritz Lehmann. Engineer’s Witness: A Photographic Panorama of Nineteenth Century Engineering Triumphs by Ralph Greenhill (review). Technology and Culture, 27(4):885–886, October 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889543/pdf.

Lehmann:1987:CWI

[Leh87] Fritz Lehmann. The CPR West: The Iron Road and the Making of a Nation ed. by Hugh A. Dempsey (review). Technology and Culture, 28(1):177–178, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889485/pdf.

LeMahieu:1982:GRM

[LeM82] D. L. LeMahieu. The gramophone: Recorded music and the cultivated mind in Britain between the wars. Technology and Culture, 23(3):372–391, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891180/pdf.

LeMahieu:1986:MBJ

[LeM86] D. L. LeMahieu. The Media in Britain by Jeremy Tunstall (review). Technology and Culture, 27(1):136, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889647/pdf.

Lerman:1988:CCP

[Ler88] Nina Lerman. Computer Chips and Paper Clips: Technology and Women’s Employment ed. by Heidi I. Hartmann, et al. (review). Technology and Culture, 29(4):939–940, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889181/pdf.

Leslie:1982:UHC

[Les82] Stuart W. Leslie. The urban habitat: The city and beyond. Technology and Culture, 23(3):417–429, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891182/pdf.

Leslie:1985:MCV

[Les85] Stuart W. Leslie. Macro: A Clear Vision of How Science and Technology Will Shape Our Future by Frank P. Davidson (review). Technology and Culture, 26(3):656–658, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889742/pdf.

Leslie:1986:FPS

[Les86a] Stuart W. Leslie. Forces of Production: A Social History of Industrial Automation by David F. Noble (review). Technology and Culture, 27(1):158–160, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889658/pdf.

Leslie:1986:AAI

[Les86b] Stuart W. Leslie. The American Automobile Industry by John B. Rae (review). Technology and Culture, 27(4):892–893, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889553/pdf.

Levy:1983:EGG

[Lev83] Richard Michael Levy. Engineers at the Golden Gate: A History of the San Francisco Army Corps of Engineers, 1866–1980 by Joseph J. Hagwood, Jr. (review). Technology and Culture, 24(3):533–534, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891416/pdf.

Levinson:1987:APU

[Lev87a] Mark Levinson. Aerial Pioneers: The U.S. Air Mail Service, 1918–1927 by William M. Leary (review). Technology and Culture, 28(3):739–740, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889401/pdf.

Levinson:1987:GON

[Lev87b] Mark Levinson. Gleanings from oft-neglected sources: Institutional history from technical documents — the case of the Naca. Technology and Culture, 28(2):314–323, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889396/pdf.

Levinson:1987:AAI

[Lev87c] Paul Levinson. The Age of Access: Information Technology and Social Revolution/Posthumous Papers of Colin Cherry ed. by William Edmondson (review). Technology and Culture, 28(4): 895–897, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889340/pdf.

Levin:1988:ANB

[Lev88] Miriam R. Levin. “Art Nouveau Bing: Paris Style 1900” and “Berlin 1900–1933: Architecture and Design” at the Cooper-Hewitt Museum, New York. Technology and Culture, 29(3):622– 632, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889216/pdf.

Lewis:1980:AVT

[Lew80] W. David Lewis. Appropriate Visions: Technology, the Environment, and the Individual ed. by Richard C. Dorf, Yvonne L. Hunter, and: Less Is More: The Art of Voluntary Poverty ed. by Goldian VandenBroeck (review). Technology and Culture, 21 (4):701–704, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890788/pdf.

Lewis:1982:HSB

[Lew82a] W. David Lewis. Henderson Steel: Birmingham’s First Steel by Bernice Shield Hassinger (review). Technology and Culture, 23 (3):498–499, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891204/pdf.

Lewis:1982:OTR

[Lew82b] W. David Lewis. The Origins of the Turbojet, Revolution by Edward W. Constant II (review). Technology and Culture, 23(3):512– 516, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891213/pdf.

Lewis:1985:WGA

[Lew85] W. David Lewis. The Winged Gospel: America’s Romance with Aviation, 1900–1950 by Joseph J. Corn (review). Technology and Culture, 26(4):874–877, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889707/pdf.

Lewenstein:1988:SSH

[Lew88a] Bruce V. Lewenstein. Selling Science: How the Press Covers Science and Technology by Dorothy Nelkin (review). Technology and Culture, 29(4):985–986, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889217/pdf.

Lewis:1988:FMM

[Lew88b] Arthur O. Lewis. The Flying Machine and Modern Literature by Laurence Goldstein (review). Technology and Culture, 29(3):705– 706, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889263/pdf.

Lewis:1989:IUS

[Lew89] W. David Lewis. Industrialization and urbanization — the SENCSA Conference, University of Alabama at Birmingham, April 1987. Technology and Culture, 30(3):628–634, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889030/pdf.

Leavitt:1980:WDB

[LGWK80] Thomas W. Leavitt, Laurence F. Gross, Helena Wright, and Katherine R. Koob. The Weavers Draft Book and Clothiers Assistant by John Hargrove (review). Technology and Culture, 21 (3):504–506, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890724/pdf.

Leslie:1984:WDC

[LHH^{+}84] Stuart W. Leslie, John Heitmann, Bert Hall, Bonnie Kaplan, Edward W. Constant II, Edward C. Ezell, Neal FitzSimons, John Fleckner, James C. Williams, Stephen H. Cutcliffe, James E. Brittain, George Wise, Tom D. Crouch, Wendy Pollock, and Judith McGaw. The Washington, D.C., Meeting, October 20–23, 1983. Technology and Culture, 25(3):615–638, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890000/pdf.

Licht:1985:OPE

[Lic85] Walter Licht. Ordinary People and Everyday Life: Perspectives on the New Social History ed. by James B. Gardner, George Rollie Adams (review). Technology and Culture, 26(2):293–294, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889775/pdf.

Lieberstein:1981:TSU

[Lie81a] Samuel Lieberstein. Technology and Society under Lenin and Stalin: Origins of the Soviet Technical Intelligentsia 1917–1941 by Kendall E. Bailes (review). Technology and Culture, 22(4):823–825, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890965/pdf.

Lieberstein:1981:TIS

[Lie81b] Samuel Lieberstein. The Technical Intelligentsia and the Soviet State by Nicholas Lampert (review). Technology and Culture, 22(1):203–204, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890830/pdf.

Liebenau:1983:HBG

[Lie83] Jonathan Liebenau. A History of the British Gas Industry by Trevor I. Williams (review). Technology and Culture, 24(1): 134–135, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891239/pdf.

Liff:1985:NTI

[Lif85] Sonia Liff. New Technology and Industrial Change: The Impact of the Scientific-Technical Revolution on Labour and Industry by Ian Benson, John Lloyd (review). Technology and Culture, 26(2):298– 300, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889777/pdf.

Light:1987:BTF

[Lig87] John D. Light. Blacksmithing technology and forge construction. Technology and Culture, 28(3):658–665, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889355/pdf.

Lindstrom:1981:REH

[Lin81a] Diane Lindstrom. Research in Economic History, V ed. by Paul Uselding (review). Technology and Culture, 22(2):361–363, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890877/pdf.

Linsley:1981:IAB

[Lin81b] S. M. Linsley. Industrial Archaeology in the British Isles by John Butt, Ian Donnachie (review). Technology and Culture, 22(2):303– 305, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890850/pdf.

Lindstrom:1985:CCW

[Lin85] Diane Lindstrom. Corporate Capital: Wilmington in the Twentieth Century by Carol E. Hoffecker (review). Technology and Culture, 26(4):861, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889699/pdf.

Link:1988:PSS

[Lin88] Albert N. Link. The Positive Sum Strategy: Harnessing Technology for Economic Growth ed. by Ralph Landau, Nathan Rosenberg (review). Technology and Culture, 29(1):183–184, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/888536/
pdf.

Lamberg-Karlovsky:1980:SAT

[LK80] C. C. Lamberg-Karlovsky. The Search for Ancient Tin ed. by Alan D. Franklin, et al. (review). Technology and Culture, 21(3):474– 477, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890711/pdf.

Lindgren:1982:SFD

[LL82] Michael Lindgren and Svante Lindqvist. Scheutz’s first difference engine rediscovered. Technology and Culture, 23(2):207–213, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891146/pdf.

Lankton:1987:TAO

[LM87] Larry D. Lankton and Jack K. Martin. Technological advance, organizational structure, and underground fatalities in the Upper Michigan copper mines, 1860–1929. Technology and Culture, 28 (1):42–66, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889440/pdf.

Lovekin:1982:JEI

[Lov82] David Lovekin. Jacques Ellul: Interpretive Essays ed. by Clifford G. Christians, Jay M. Van Hook (review). Technology and Culture, 23(4):694–696, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890276/pdf.

Lankton:1985:IIH

[LR85] Larry D. Lankton and Terry S. Reynolds. Industrialization and the industrial heritage in international perspective: Lowell Conference on Industrial History; The International Conference on the Conservation of the Industrial Heritage; and Conference of the Society for Industrial Archeology — Lowell, Cambridge, and Boston, Massachusetts, June 7–17, 1984. Technology and Culture, 26(2):268– 274, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889767/pdf.

Lee:1984:STL

[LS84] Judith Yaross Lee and Joseph W. Slade. Science, technology, and literature — Long Island University, Brooklyn, New York, February 23–25, 1983. Technology and Culture, 25(2):299–307, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890042/pdf.

Lubar:1985:AEH

[Lub85] Steven Lubar. America’s Economic Heritage: From a Colonial to a Capitalist Economy, 1634–1900, and America’s Economic Heritage: A Mature Economy, post 1900 by Meyer Weinberg (review). Technology and Culture, 26(2):304–305, April 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889781/pdf.

Lubar:1986:PCT

[Lub86a] Steven Lubar. Proprietary Capitalism: The Textile Manufacture at Philadelphia, 1800–1885 by Philip B. Scranton (review). Technology and Culture, 27(1):153–155, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889656/pdf.

Lubar:1986:ABC

[Lub86b] Steven Lubar. The AI Business: Commercial Uses of Artificial Intelligence ed. by Patrick H. Winston, Karen A. Prendergast (review). Technology and Culture, 27(4):872–873, October 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889527/pdf.

Lubar:1986:CMB

[Lub86c] Steven Lubar. The Computer Museum, Boston, Massachusetts. Technology and Culture, 27(1):96–105, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889636/pdf.

Lubar:1987:WEW

[Lub87a] Steven Lubar. Workers on the Edge: Work, Leisure, and Politics in Industrializing Cincinnati, 1788–1890 by Steven J. Ross (review). Technology and Culture, 28(4):867–868, October 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889320/pdf.

Lubar:1987:CTD

[Lub87b] Steven Lubar. Culture and technological design in the 19th-Century pin industry: John Howe and the Howe Manufacturing Company. Technology and Culture, 28(2):253–282, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889391/pdf.

Luegenbiehl:1986:TTE

[Lue86] Heinz C. Luegenbiehl. Theology and Technology: Essays in Christian Analysis and Exegesis ed. by Carl Mitcham, Jim Grote (review). Technology and Culture, 27(2):340–342, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889630/pdf.

Lukasiewicz:1986:CEH

[Luk86] Julius Lukasiewicz. Canada’s encounter with high-speed aeronautics. Technology and Culture, 27(2):223–261, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889585/pdf.

Lukasiewicz:1989:GAA

[Luk89] Julius Lukasiewicz. From the Ground Up: The Autobiography of an Aeronautical Engineer by Fred E. Weick, James R. Hansen (review). Technology and Culture, 30(2):486–488, April 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889100/pdf.

Lundgren:1988:CLE

[Lun88] Lars J. Lundgren. City Lights: The Establishment of the First Swedish Gasworks by Arne Kaijser (review). Technology and Culture, 29(2):296–297, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889274/pdf.

Lytle:1988:ARM

[Lyt88] Richard Lytle. Appraising the Records of Modern Science and Technology: A Guide by Joan K. Haas, et al. (review). Technology and Culture, 29(3):710–712, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889270/pdf.

Maass:1981:BHO

[Maa81] Eleanor Maass. A Bibliography on Historical Organization Practices. Vol. 4: Documentation of Collections by Rosemary S. Reese (review). Technology and Culture, 22(1):229–231, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890841/
pdf.

MacKenzie:1984:MM

[Mac84] Donald MacKenzie. Marx and the machine. Technology and Culture, 25(3):473–502, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889990/pdf.

Mack:1985:DIE

[Mac85a] Pamela Mack. Diffusion of Innovations by Everett M. Rogers (review). Technology and Culture, 26(1):109–110, January 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889826/pdf.

MacKenzie:1985:ETC

[Mac85b] Donald MacKenzie. Explaining Technical Change: A Case Study in the Philosophy of Science by Jon Elster (review). Technology and Culture, 26(4):886–888, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889712/pdf.

MacLeod:1985:SIS

[Mac85c] Roy MacLeod. Science for Industry: A Short History of the Imperial College of Science and Technology by A. Rupert Hall (review). Technology and Culture, 26(4):848–850, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889692/pdf.

Mack:1986:LWS

[Mac86a] Pamela E. Mack. Living and Working in Space: A History of Skylab by W. David Compton, Charles D. Benson (review). Technology and Culture, 27(3):658–659, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889600/pdf.

MacLeod:1986:CIH

[Mac86b] Roy MacLeod. Critical Issues in the History of Canadian Science, Technology and Medicine ed. by Richard A. Jarrell, Arnold E. Roos (review). Technology and Culture, 27(1):140–143, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889650/pdf.

Mack:1987:PMS

[Mac87a] Pamela E. Mack. Presidential Management of Science and Technology: The Johnson Presidency by W. Henry Lambright (review). Technology and Culture, 28(4):879–881, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889328/pdf.

Macleod:1987:ADG

[Mac87b] Christine Macleod. Accident or design? George Ravenscroft’s patent and the invention of lead-crystal glass. Technology and Culture, 28(4):776–803, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889302/pdf.

Mack:1988:FHR

[Mac88a] Pamela E. Mack. Forests, habitats, and resources: a conference in world environmental history — Duke University, Durham, North Carolina, April 30–May 2, 1987. Technology and Culture, 29 (2):281–282, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889258/pdf.

MacKenzie:1988:HSM

[Mac88b] Donald MacKenzie. The History of Statistics: The Measurement of Uncertainty before 1900 by Stephen M. Stigler (review). Technology and Culture, 29(2):299–300, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889276/pdf.

Mack:1989:SH

[Mac89] Pamela E. Mack. Space history. Technology and Culture, 30 (3):657–665, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889034/pdf.

Mahoney:1982:NRB

[Mah82] Michael S. Mahoney. Never at Rest: A Biography of Isaac Newton by Richard S. Westfall (review). Technology and Culture, 23(4):655–656, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890255/pdf.

Major:1981:PLS

[Maj81] J. Kenneth Major. Power on Land and Sea — 160 Years of Industrial Enterprise on Tyneside — a History of R. and W. Hawthorn Leslie by J. F. Clarke (review). Technology and Culture, 22(1): 190–191, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890822/pdf.

Malik:1986:RPF

[Mal86a] Yogendra K. Malik. The Role of Parliament in the Formulation of National Science and Technology Policy by A. Rahman, Nirmal Haritash (review). Technology and Culture, 27(1):183–185, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889671/pdf.

Malone:1986:WPS

[Mal86b] Patrick M. Malone. Water Power in Scotland 1550–1870 by John Shaw (review). Technology and Culture, 27(3):620–621, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889563/
pdf.

Mander:1981:CAT

[Man81] Mary S. Mander. The Computer Age: A Twenty-Year View ed. by Michael L. Dertouzos, Joel Moses (review). Technology and Culture, 22(1):220–224, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890838/pdf.

Mander:1983:EUA

[Man83] M. S. Mander. From ENIAC to UNIVAC: An Appraisal of the Eckert-Mauchly Computers by Nancy Stern (review). Technology and Culture, 24(1):148–150, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891248/pdf.

Mancoff:1989:CWC

[Man89] Debra N. Mancoff. Carl W. Condit’s publications — a chronological bibliography, 1946–1988. Technology and Culture, 30(2):258– 265, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889058/pdf.

Marrett:1980:RSI

[Mar80] Cora B. Marrett. Regulation of Scientific Inquiry ed. by Keith M. Wulff (review). Technology and Culture, 21(4):697–699, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890785/pdf.

Martin:1983:SHA

[Mar83a] Albro Martin. A Short History of American Locomotive Builders in the Steam Era by John H. White (review). Technology and Culture, 24(4):694–695, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890139/pdf.

Martin:1983:ENA

[Mar83b] Albro Martin. Encyclopedia of North American Railroading: 150 Years of Railroading in the United States and Canada by Freeman Hubbard (review). Technology and Culture, 24(4):692–693, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890138/pdf.

Mark:1984:ACM

[Mar84a] Robert Mark. Architecture and the Crisis of Modern Science by Alberto Pérez–Gómez (review). Technology and Culture, 25(4): 900–902, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889983/pdf.

Marre:1984:DCR

[Mar84b] Louis A. Marre. Doodlebug Country: The Rail Motorcar on the Class I Railroads of the United States by Edmund Keilty (review). Technology and Culture, 25(1):164–165, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890116/pdf.

Marks:1985:PHW

[Mar85a] Harry M. Marks. “The Physician’s Hand”: Work Culture and Conflict in American Nursing by Barbara Melosh (review). Technology and Culture, 26(2):324–326, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889791/pdf.

Martin:1985:HOW

[Mar85b] Kenneth R. Martin. Harpoons and Other Whalecraft by Thomas G. Lytle (review). Technology and Culture, 26(3):668–669, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889750/
pdf.

Martin:1985:WTW

[Mar85c] Patrick Martin. “Where Two Worlds Meet: The Great Lakes Fur Trade,” Minnesota Historical Society, St. Paul, Minnesota. Technology and Culture, 26(3):620–622, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889724/pdf.

Mark:1986:BFC

[Mar86a] Robert Mark. Bridge to the Future: A Centennial Celebration of the Brooklyn Bridge ed. by Margaret Latimer, et al. (review). Technology and Culture, 27(2):324–326, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889622/pdf.

Mark:1986:BBH

[Mar86b] Robert Mark. Building Bridges: History, Technology, Construction by Hans Wittfoht (review). Technology and Culture, 27(4): 887–889, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889546/pdf.

Martin:1986:PRA

[Mar86c] Albro Martin. Province of Reason: American Lives in a New Age of Science by Sam Bass Warner, Jr. (review). Technology and Culture, 27(3):646–648, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889590/pdf.

Martin:1986:RMP

[Mar86d] Albro Martin. The Rebirth of the Missouri Pacific, 1956–1983 by H. Craig Miner (review). Technology and Culture, 27(1):166–167, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889663/pdf.

Mark:1987:FIA

[Mar87a] Robert Mark. French Iron Architecture by Frances H. Steiner (review). Technology and Culture, 28(2):354–355, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889413/pdf.

Marks:1987:CSP

[Mar87b] Harry M. Marks. A Calculus of Suffering: Pain, Professionalism, and Anesthesia in Nineteenth-Century America by Martin S. Pernick (review). Technology and Culture, 28(3):728–729, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889387/
pdf.

Marling:1987:PLN

[Mar87c] Karal Ann Marling. The Public Landscape of the New Deal by Phoebe Cutler (review). Technology and Culture, 28(2):377–378, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889428/pdf.

Martin:1988:WIM

[Mar88] Kenneth R. Martin. Whales, Ice, and Men: The History of Whaling in the Western Arctic by John R. Bockstoce, and: oWhaling in the North Atlantic: From Earliest Times to the Mid-19th Century by Jean-Pierre Proulx (review). Technology and Culture, 29 (2):305–307, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889280/pdf.

Marcus:1989:MLS

[Mar89a] Alan I. Marcus. The Modernization of the Louisiana Sugar Industry 1830–1910 by John Alfred Heitmann (review). Technology and Culture, 30(3):697–698, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889057/pdf.

Marre:1989:LHE

[Mar89b] Louis A. Marre. Lima: The History by Eric Hirsimaki (review). Technology and Culture, 30(2):469–470, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889088/pdf.

Mauskopf:1982:CTS

[Mau82] Seymour H. Mauskopf. Ciencia y técnica en la sociedad espãnola de los siglos XVI y XVII by José María López Piñero (review). Technology and Culture, 23(1):112–114, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891123/pdf.

Mauskopf:1984:HPA

[Mau84] Seymour H. Mauskopf. A History of Platinum and Its Allied Metals by Donald McDonald and Leslie B. Hunt (review). Technology and Culture, 25(4):853–854, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889955/pdf.

Mayer:1983:PNY

[May83] Harold M. Mayer. The Port of New York: A History of the Rail and Terminal System from the Beginnings to Pennsylvania Station, and The Port of New York: A History of the Rail and Terminal System from the Grand Central Electrification to the Present by Carl W. Condit (review). Technology and Culture, 24(2):289– 291, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891279/pdf.

Mayr:1984:FK

[May84] Otto Mayr. Friedrich Klemm (1904–1983). Technology and Culture, 25(2):377–380, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890077/pdf.

Mayr:1985:RTC

[May85] Otto Mayr. Revolution in Time: Clocks and the Making of the Modern World by David S. Landes (review). Technology and Culture, 26(3):631–634, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889727/pdf.

Mayer:1989:RSS

[May89] Harold M. Mayer. The Railway Station: A Social History by Jeffrey Richards, John M. Mackenzie (review). Technology and Culture, 30(2):464–465, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889083/pdf.

Mazuzan:1981:NPP

[Maz81] George T. Mazuzan. Nuclear Power and the Public Safety: A Study in Regulation by Elizabeth S. Rolph (review). Technology and Culture, 22(1):213–217, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890835/pdf.

Mazuzan:1982:APS

[Maz82] George T. Mazuzan. Atomic power safety: The case of the Power Reactor Development Company fast breeder, 1955–1956. Technology and Culture, 23(3):341–371, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891179/pdf.

Mazuzan:1983:NPN

[Maz83] George T. Mazuzan. Nuclear Power and Non-Proliferation: The Remaking of U.S. Policy by Michael J. Brenner (review). Technology and Culture, 24(2):303–305, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891288/pdf.

Mazuzan:1984:NSN

[Maz84a] George Mazuzan. Nukespeak: The Selling of Nuclear Technology in America by Stephen Hilgartner, Richard C. Bell, and Rory O’Connor (review). Technology and Culture, 25(4):889–891, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889977/pdf.

Mazuzan:1984:ATM

[Maz84b] George T. Mazuzan. Accident at Three Mile Island: The Human Dimensions ed. by David L. Sills, et al. (review). Technology and Culture, 25(1):156–158, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890111/pdf.

Mazuzan:1986:NES

[Maz86a] George T. Mazuzan. Nuclear energy — a subject in need of historical research. Technology and Culture, 27(1):123–126, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889641/pdf.

Mazuzan:1986:VRB

[Maz86b] George T. Mazuzan. “Very risky business”: A power reactor for New York City. Technology and Culture, 27(2):262–284, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889587/pdf.

Mark:1989:SIO

[MB89] Robert Mark and David P. Billington. Structural imperative and the origin of new form. Technology and Culture, 30(2):300–329, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889063/pdf.

McCusker:1985:SCQ

[McC85] John J. McCusker. At the Sign of the Compass and Quadrant: The Life and Times of Anthony Lamb by Silvio A. Bedini (review). Technology and Culture, 26(3):639–640, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889731/pdf.

McCoy:1988:IHT

[McC88] Joan D. McCoy. Interface ’86: Humanities and Technology — Atlanta, Georgia, October 23–24, 1986. Technology and Culture, 29 (2):283–284, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889260/pdf.

McDougall:1982:BTI

[McD82] Walter A. McDougall. Blaming Technology: The Irrational Search for Scapegoats by Samuel C. Florman (review). Technology and Culture, 23(3):491–493, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891200/pdf.

McDougall:1984:CWA

[McD84] Walter A. McDougall. Clipped Wings: The American SST Conflict by Mel Horwitch (review). Technology and Culture, 25(2):367–369, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890073/pdf.

McDonnell:1985:WWK

[McD85a] Lawrence T. McDonnell. Workers’ World: Kinship, Community, and Protest in an Industrial Society, 1900–1940 by John Bodnar, and: No Strength without Union: An Illustrated History of Ohio Workers 1803–1980 by Raymond Boryczka and Lorin Lee Cary (review). Technology and Culture, 26(1):101–103, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889823/
pdf.

McDougall:1985:SAE

[McD85b] Walter A. McDougall. Space-age Europe: Gaullism, Eurogaullism, and the American dilemma. Technology and Culture, 26(2):179–203, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889761/pdf.

McDonnell:1986:MSR

[McD86] Lawrence T. McDonnell. Men of the Steel Rails: Workers on the Atchison, Topeka & Santa Fe Railroad, 1869–1900 by James H. Ducker (review). Technology and Culture, 27(2):319–321, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889619/pdf.

McDonnell:1989:OSN

[McD89] Lawrence T. McDonnell. Old South, New South: Revolutions in the Southern Economy since the Civil War by Gavin Wright (review). Technology and Culture, 30(3):695–697, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889055/pdf.

McGinn:1983:SBT

[McG83] Robert E. McGinn. Stokowski and the Bell telephone laboratories: Collaboration in the development of high-fidelity sound reproduction. Technology and Culture, 24(1):38–75, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891222.

McGowan:1984:WAB

[McG84] A. P. McGowan. William Avery Baker: A Tribute from His Friends by Erik A. R. Ronnberg, Jr. (review). Technology and Culture, 25(4):878–880, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889970/pdf.

McGaw:1985:AIT

[McG85a] Judith A. McGaw. Accounting for innovation: Technological change and business practice in the Berkshire County paper industry. Technology and Culture, 26(4):703–725, October 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889674/pdf.

McGuire:1985:IEN

[McG85b] Robert A. McGuire. Independence and Empire: The New South’s Cotton Mill Campaign, 1865–1901 by Patrick J. Hearden (review). Technology and Culture, 26(3):649–651, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889738/pdf.

McGaw:1986:AHL

[McG86] Judith McGaw. Alice Hamilton: A Life in Letters by Barbara Sicherman, and: Two Sisters for Social Justice: A Biography of Grace and Edith Abbott by Lela B. Costin (review). Technology and Culture, 27(3):648–651, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889591/pdf.

McGaw:1988:FTP

[McG88a] Judith A. McGaw. Food for thought in Philadelphia: “The larder invaded”. Technology and Culture, 29(2):266–270, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889253/pdf.

McGucken:1988:SCC

[McG88b] William McGucken. Scientific Controversies: Case Studies in the Resolution and Closure of Disputes in Science and Technology ed. by H. Tristram Englehardt, Jr., Arthur L. Caplan (review). Technology and Culture, 29(4):980–981, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889211/pdf.

McGucken:1988:WRS

[McG88c] William McGucken. The Whale and the Reactor: A Search for Limits in an Age of High Technology by Langdon Winner (review).

Technology and Culture, 29(1):176–177, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888531/pdf.

McInnes:1981:OWB

[McI81] Angus McInnes. Old Westminster Bridge: The Bridge of Fools by R. J. B. Walker (review). Technology and Culture, 22(2):305–306, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890851/pdf.

McIntyre:1983:TTU

[McI83] John R. McIntyre. Technology Transfer to the USSR, 1928–1937 and 1966–1975: The Role of Western Technology in Soviet Economic Development by George D. Holliday (review). Technology and Culture, 24(3):546–547, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891425/pdf.

McIntyre:1984:TDI

[McI84] John R. McIntyre. Technological Diffusion and Industrialisation before 1914 by A. G. Kenwood and A. L. Lougheed (review). Technology and Culture, 25(4):868–869, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889963/pdf.

McKnight:1980:MIM

[McK80] John L. McKnight. Mathematical Instrument-Makers in the Grocers’ Company, 1688–1800 by Joyce Brown (review). Technology and Culture, 21(3):494–495, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890719/pdf.

McKay:1986:DTU

[McK86] John P. McKay. The Decline of Transit: Urban Transportation in German and U.S. Cities, 1900–1970 by Glenn Yago (review). Technology and Culture, 27(2):331–332, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889625/pdf.

McLeod:1984:VLM

[McL84] Alexander McLeod. Village-Level Modernization in Southeast Asia: The Political Economy of Rice and Water ed. by Geoffrey B. Hainsworth (review). Technology and Culture, 25(4):872–873, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889966/pdf.

McLeod:1985:PTI

[McL85] Alexander Dean McLeod. Policy and Trade Issues of the Japanese Economy: American and Japanese Perspectives ed. by Kozo Yamamura, and: The Amazing Race: Winning the Technorivalry with Japan by William H. Davidson (review). Technology and Culture, 26(1):115–117, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889829/pdf.

McMahon:1981:EDB

[McM81a] Michal McMahon. The Engineering Drawings of Benjamin Henry Latrobe ed. by Darwin H. Stapleton (review). Technology and Culture, 22(4):794–797, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890952/pdf.

McMahon:1981:RTP

[McM81b] Michal McMahon. The romance of technological progress: A critical review of the National Air and Space Museum. Technology and Culture, 22(2):281–296, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890846/pdf.

McMahon:1986:YPH

[McM86] A. Michal McMahon. 75 Years of Progress — a History of the American Institute of Chemical Engineers 1908–1983 by Terry S. Reynolds (review). Technology and Culture, 27(4):845–846, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889509/pdf.

McNeill:1983:PTC

[McN83] William H. McNeill. Population and Technological Change: A Study of Long-Term Trends by Ester Boserup (review). Technology and Culture, 24(1):114–115, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891227/pdf.

McShane:1983:GCR

[McS83] Clay McShane. Garbage in the Cities: Refuse, Reform, and the Environment, 1880–1980 by Martin V. Melosi (review). Technology and Culture, 24(3):541–542, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891421/pdf.

Mack:1982:PSH

[MD82] Pamela E. Mack and David H. De Vorkin. Proseminar in space history: The National Air and Space Museum, May 22, 1981. Technology and Culture, 23(2):202–206, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891145/pdf.

Mehrtens:1987:RMT

[Meh87] Herbert Mehrtens. Reactionary Modernism: Technology, Culture, and Politics in Weimar and the Third Reich by Jeffrey Herf (review). Technology and Culture, 28(1):149–150, January 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889468/pdf.

Meikle:1981:DHA

[Mei81] Jeffrey L. Meikle. Dictionary of the History of the American Brewing and Distilling Industries by William L. Downard (review). Technology and Culture, 22(4):803–805, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890957/pdf.

Meikle:1982:DND

[Mei82] Jeffrey L. Meikle. Dawn of a New Day: The New York World’s Fair, 1939/40 ed. by Helen A. Harrison (review). Technology and Culture, 23(1):130–131, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891135/pdf.

Meikle:1984:TBN

[Mei84] Jeffrey L. Meikle. The Tower and the Bridge: The New Art of Structural Engineering by David P. Billington (review). Technology and Culture, 25(4):896–898, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889981/pdf.

Meikle:1987:MAA

[Mei87] Jeffrey L. Meikle. “The Machine Age in America 1918–1941,” an Exhibition at the Brooklyn Museum. Technology and Culture, 28(3):666–669, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889356/pdf.

Meikle:1988:ALR

[Mei88a] Jeffrey L. Meikle. Art and Labor: Ruskin, Morris, and the Craftsman Ideal in America by Eileen Boris (review). Technology and Culture, 29(1):173–174, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/888529/pdf.

Meikle:1988:ODD

[Mei88b] Jeffrey L. Meikle. Objects of Desire: Design and Society from Wedgwood to IBM by Adrian Forty (review). Technology and Culture, 29(2):318–319, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889288/pdf.

Meikle:1988:SGT

[Mei88c] Jeffrey L. Meikle. Shifting Gears: Technology, Literature, Culture in Modernist America by Cecelia Tichi (review). Technology and Culture, 29(3):704–705, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889261/pdf.

Meiksins:1988:RER

[Mei88d] Peter Meiksins. The “Revolt of the engineers” reconsidered. Technology and Culture, 29(2):219–246, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889248/pdf.

Meikle:1989:CCN

[Mei89] Jeffrey L. Meikle. Culture and Consumption: New Approaches to the Symbolic Character of Consumer Goods and Activities by Grant McCracken (review). Technology and Culture, 30(4): 1091–1092, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889019/pdf.

Melder:1983:ELC

[Mel83a] Keith Melder. Essays from the Lowell Conference on Industrial History, 1980 and 1981 ed. by Robert Weible, et al. (review). Technology and Culture, 24(4):687–689, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890135/pdf.

Melosi:1983:PWH

[Mel83b] Martin Melosi. Public Works History in the United States: A Guide to the Literature ed. by Suellen M. Hoy, Michael C. Robinson (review). Technology and Culture, 24(4):712–713, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890150/
pdf.

Melosh:1985:OCE

[Mel85] Barbara Melosh. A Once Charitable Enterprise: Hospitals and Health Care in Brooklyn and New York, 1885–1915 by David Rosner (review). Technology and Culture, 26(2):326–328, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889792/
pdf.

Melder:1987:PPI

[Mel87] Keith Melder. The popular perception of industrial history — Lowell, Massachusetts, November 1–3, 1985. Technology and Culture, 28(1):113–119, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889453/pdf.

Merritt:1981:LSR

[Mer81a] Raymond H. Merritt. Lex et Scientia (review). Technology and Culture, 22(1):206–209, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890832/pdf.

Merritt:1981:UPM

[Mer81b] Raymond H. Merritt. The Upper Peninsula of Michigan: An Inventory of Historic Engineering and Industrial Sites ed. by the Historic American Engineering Record (review). Technology and Culture, 22(3):634–635, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890909/pdf.

Mernitz:1986:TCG

[Mer86] Kenneth S. Mernitz. Technological Change in the German Democratic Republic by Raymond Bentley (review). Technology and Culture, 27(1):177–179, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889668/pdf.

Mergen:1987:MCR

[Mer87] Bernard Mergen. Material Culture: A Research Guide ed. by Thomas J. Schlereth (review). Technology and Culture, 28(3):698– 700, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889369/pdf.

Merritt:1988:TSG

[Mer88] Raymond Merritt. The Tender Ship: Governmental Management of Technological Change by Arthur M. Squires (review). Technology and Culture, 29(4):987, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889219/pdf.

Metz:1983:CBA

[Met83] Lance E. Metz. The Canal Builders by Anthony Burton (review). Technology and Culture, 24(3):512, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891401/pdf.

Metailie:1986:SCC

[Mét86] Georges Métailié. Science and Civilisation in China. Vol. 6: Biology and Biological Technology. Pt. 2: Agriculture by Francesca Bray (review). Technology and Culture, 27(4):855–857, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889515/pdf.

Meyer:1985:IWI

[Mey85a] Stephen Meyer. Immigrant Workers in Industrial France: The Making of a New Laboring Class by Gary S. Cross (review). Technology and Culture, 26(1):92–94, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889818/pdf.

Meyer:1985:SLR

[Mey85b] Stephen Meyer. Steelmasters and Labor Reform, 1886–1923 by Gerald G. Eggert (review). Technology and Culture, 26(2):311–313, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889784/pdf.

Meyer:1986:ASM

[Mey86] Stephen Meyer. From the American System to Mass Production 1800–1932: The Development of Manufacturing Technology in the United States by David A. Hounshell (review). Technology and Culture, 27(1):155–158, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889657/pdf.

Meyer:1988:TWS

[Mey88] Stephen Meyer. Technology and the workplace: Skilled and production workers at Allis-chalmers, 1900–1941. Technology and Culture, 29(4):839–864, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889160/pdf.

Michael:1984:IRE

[Mic84a] Loren Michael. The Indian Response to European Technology and Culture (A.D. 1498–1707) by Ahsan Jan Qaisar (review). Technology and Culture, 25(1):111–113, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890085/pdf.

Michel:1984:RRM

[Mic84b] John L. Michel. The Rise of Robert Millikan: Portrait of a Life in American Science by Robert Kargon (review). Technology and Culture, 25(1):138–141, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890100/pdf.

Millard:1980:HIC

[Mil80a] A. J. Millard. History and Industrial Civilisation by R. A. Buchanan (review). Technology and Culture, 21(4):652–654, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890763/pdf.

Miller:1980:PWA

[Mil80b] Harry Miller. Potash from wood ashes: Frontier technology in Canada and the United States. Technology and Culture, 21(2):187– 208, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890673/pdf.

Miller:1981:TFR

[Mil81] Howard S. Miller. Truss failures reconsidered. Technology and Culture, 22(4):849–850, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890977/pdf.

Millard:1982:RDT

[Mil82a] A. J. Millard. Research, Development and Technological Innovation ed. by Devendra Sahal (review). Technology and Culture, 23 (2):278–279, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891176/pdf.

Millard:1982:SEI

[Mil82b] A. J. Millard. The Steam Engine in Industry, vols 1 and 2 by George Watkins (review). Technology and Culture, 23(4):638–639, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890245/pdf.

Miller:1982:HIP

[Mil82c] George L. Miller. Handcraft to Industry: Philadelphia Ceramics in the First Half of the Nineteenth Century by Susan H. Myers (review). Technology and Culture, 23(4):668–669, October 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890263/pdf.

Miller:1983:BMC

[Mil83] Carol Poh Miller. Bridges of Metropolitan Cleveland: Past and Present by Sara Ruth Watson, John R. Wolfs (review). Technology and Culture, 24(3):539–541, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891420/pdf.

Millard:1984:EPE

[Mil84a] A. J. Millard. Engineering Professionalism and Ethics ed. by James H. Schaub, et al. (review). Technology and Culture, 25 (4):893–895, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890238/pdf.

Millard:1984:EMP

[Mil84b] A. J. Millard. Engineers, Managers, and Politicians: The First Fifteen Years of Nationalised Electricity Supply in Britain by Leslie Hannah (review). Technology and Culture, 25(1):153–154, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890109/pdf.

Millard:1986:MPC

[Mil86] A. J. Millard. The Making of a Profession: A Century of Electrical Engineering in America by A. Michal McMahon (review). Technology and Culture, 27(4):843–845, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889508/pdf.

Millard:1987:LRC

[Mil87] A. J. Millard. The Language of Risk: Conflicting Perspectives on Occupational Health ed. by Dorothy Nelkin (review). Technology and Culture, 28(3):731–732, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889390/pdf.

Millbrooke:1988:YST

[Mil88a] Anne Millbrooke. 100 Years of Science and Technology in Texas ed. by Leo J. Klosterman, et al. (review). Technology and Culture, 29(4):970–971, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889199/pdf.

Miller:1988:BLS

[Mil88b] Carol Poh Miller. Building the Lone Star: An Illustrated Guide to Historic Sites by T. Lindsay Baker (review). Technology and Culture, 29(3):699–700, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889255/pdf.

Miller:1989:GAB

[Mil89] Carol Poh Miller. Great American Bridges and Dams by Donald C. Jackson (review). Technology and Culture, 30(4):1061–1062, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888993/pdf.

Mingay:1981:SSP

[Min81] Gordon Mingay. The Story of the Steam Plough Works: Fowlers of Leeds by Michael R. Lane (review). Technology and Culture, 22(3):618–620, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890900/pdf.

Minami:1982:MPP

[Min82] Ryōshin Minami. Mechanical power and printing technology in pre-World War II Japan. Technology and Culture, 23(4):609–624, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890242/pdf.

Mirowski:1982:TIR

[Mir82] Philip Mirowski. Transatlantic Industrial Revolution: The Diffusion of Textile Technologies between Britain and America, 1790– 1830s by David J. Jeremy (review). Technology and Culture, 23 (4):641–642, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890247/pdf.

Misa:1984:CCE

[Mis84] Thomas J. Misa. A Century of Chemical Engineering ed. by William F. Furter (review). Technology and Culture, 25(2):370– 371, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890074/pdf.

Misa:1987:EBS

[Mis87] Thomas J. Misa. The eisen-bibliothek, schaffhausen, Switzerland. Technology and Culture, 28(2):324–327, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889398/pdf.

Mitchell:1985:PRC

[Mit85] Stephanie J. Mitchell. The People’s Republic of China: Foreign Economic Relations and Technology Acquisition 1972–1981 by Henry N. Geraedts (review). Technology and Culture, 26(2):303– 304, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889780/pdf.

Mohanty:1987:LTW

[Moh87a] Gail F. Mohanty. Labor’s True Woman: Carpet Weavers, Industrialization, and Labor Reform in the Gilded Age by Susan Levine (review). Technology and Culture, 28(2):368–369, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889422/pdf.

Mohanty:1987:SMQ

[Moh87b] Gail Fowler Mohanty. The sewing machine — “The Queen of inventions,” an exhibition at Slater mill historic site, pawtucket, rhode Island. Technology and Culture, 28(4):839–841, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889306/pdf.

Mohl:1987:GMP

[Moh87c] Raymond A. Mohl. Governing the Metropolis — Politics, Technology and Social Change in a Victorian City: Melbourne 1850– 1891 by David Dunstan (review). Technology and Culture, 28(4): 888–889, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889334/pdf.

Mohanty:1988:ETT

[Moh88] Gail Fowler Mohanty. Experimentation in textile technology, 1788–1790, and its impact on handloom weaving and weavers in Rhode Island. Technology and Culture, 29(1):1–31, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/888032/
pdf.

Molloy:1984:ARW

[Mol84a] Peter Molloy. Arthur Redman Wilfley: Miner, Inventor, and Entrepreneur by Jay E. Niebur (review). Technology and Culture, 25(4):861–863, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889959/pdf.

Molloy:1984:ORI

[Mol84b] Peter Molloy. Old Reliable: An Illustrated History of the Quincy Mining Company by Larry D. Lankton, Charles K. Hyde (review). Technology and Culture, 25(1):163, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890115/pdf.

Molloy:1988:ENS

[Mol88] Peter Molloy. Engineering the New South: Georgia Tech, 1885– 1985 by Robert C. McMath, Jr., et al. (review). Technology and Culture, 29(1):155, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888518/pdf.

Molella:1989:LMR

[Mol89] Arthur P. Molella. The Lewis Mumford Reader ed. by Donald L. Miller (review). Technology and Culture, 30(1):128–130, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889119/pdf.

Moon:1986:ALA

[Moo86] Howard Moon. Airlines of Latin America since 1919 by R. E. G. Davies (review). Technology and Culture, 27(3):652–653, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889596/pdf.

Morison:1980:ABF

[Mor80a] Elting E. Morison. Admiral Bradley A. Fiske and the American Navy by Paolo E. Coletta (review). Technology and Culture, 21(1):127–128, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890632/pdf.

Morris:1980:ELM

[Mor80b] Bertram Morris. The Economy of Literature by Marc Shell (review). Technology and Culture, 21(1):140–142, January 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890640/pdf.

Morris:1980:HPT

[Mor80c] Bertram Morris. The History and Philosophy of Technology ed. by George Bugliarello, Dean B. Doner (review). Technology and Culture, 21(3):534–538, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890740/pdf.

Morris:1980:ISP

[Mor80d] Henry Morris. An International Sourcebook of Paper History by Irving P. Leif (review). Technology and Culture, 21(3):493–494, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890718/pdf.

Moran:1981:GPP

[Mor81a] Bruce T. Moran. German Prince-practitioners: Aspects in the development of courtly science, technology, and procedures in the Renaissance. Technology and Culture, 22(2):253–274, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890844/pdf.

Morris:1981:STH

[Mor81b] Bertram Morris. Science, Technology and the Human Prospect: Proceedings of the Edison Centennial Symposium ed. by Chauncey Starr, Philip C. Ritterbush (review). Technology and Culture, 22 (2):335–339, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890866/pdf.

Moravcsik:1982:EIS

[Mor82] Michael J. Moravcsik. Essays of an Information Scientist by Eugene Garfield (review). Technology and Culture, 23(3):524–526, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891218/pdf.

Morman:1983:HCM

[Mor83] Edward T. Morman. The Healthiest City: Milwaukee and the Politics of Health Reform by Judith Walzer Leavitt (review). Technology and Culture, 24(3):542–543, July 1983. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891422/pdf.

Morrow:1985:FFF

[Mor85] John H. Morrow, Jr. The First of the Few: Fighter Pilots of the First World War by Denis Winter (review). Technology and Culture, 26(4):867–868, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889703/pdf.

Morrow:1986:CAG

[Mor86] John H. Morrow, Jr. The Command of the Air by Giulio Douhet, and: Strategy for Defeat: The Luftwaffe 1933–1945 by Williamson Murray (review). Technology and Culture, 27(1):174–177, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889667/pdf.

Morley:1987:ACS

[Mor87a] Jane Morley. The association of Collegiate Schools of Architecture Technology Conference — Washington, D.C., November 11–12, 1985. Technology and Culture, 28(1):120–123, January 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889455/pdf.

Morley:1987:CHJ

[Mor87b] Jane Morley. Construction History: Journal of the Construction History Group ed. by Mark Swenarton (review). Technology and Culture, 28(1):152–153, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889471/pdf.

Morley:1987:RWW

[Mor87c] Jane Morley. Redesigning the World: William Morris, the 1880s, and the Arts and Crafts by Peter Stansky (review). Technology and Culture, 28(3):694–696, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889367/pdf.

Morris:1987:CIB

[Mor87d] Morris David Morris. Canal Irrigation in British India: Perspectives on Technological Change in a Peasant Economy by Ian Stone (review). Technology and Culture, 28(1):179–181, January 1987.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889486/
pdf.

Morris:1987:PCT

[Mor87e] Peter Morris. The Playfair Collection: And the Teaching of Chemistry at the University of Edinburgh 1713–1858 by R. G. W. Anderson, and: Joseph Black 1728–1799: A Commemorative Symposium ed. by A. D. C. Simpson, and: “Martyr of Science”: Sir David Brewster 1781–1868 ed. by A. D. Morrison-Low, J. R. R. Christie (review). Technology and Culture, 28(3):693–694, July 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889366/pdf.

Morrow:1988:LWM

[Mor88] John H. Morrow, Jr. Luftwaffe by Williamson Murray (review). Technology and Culture, 29(1):159–161, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888521/pdf.

Morison:1989:CES

[Mor89a] Elting Morison. The Civilized Engineer by Samuel C. Florman (review). Technology and Culture, 30(1):138–139, January 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889125/pdf.

Morley:1989:CEB

[Mor89b] Jane Morley. Construction Economics and Building Design: A Historical Approach by R. Gregory Turner (review). Technology and Culture, 30(2):477–478, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889094/pdf.

Morley:1989:ISL

[Mor89c] Jane Morley. International symposium on Lewis Mumford — University of Pennsylvania, November 5–7, 1987. Technology and Culture, 30(1):122–127, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889118/pdf.

Morley:1989:TTH

[Mor89d] Jane Morley. Technology transfer in historical perspective — the 1988 Scandinavian Symposium on the History of Technology/ Historical Library of the Royal Institute of Technology, Stockholm. Technology and Culture, 30(4):1004–1012, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888973/pdf.

Morrow:1989:BDH

[Mor89e] John H. Morrow, Jr. De Blériot à Dassault: Histoire de l’Industrie aéronautique en France 1900–1950 by Emmanuel Chadeau (review). Technology and Culture, 30(4):1072–1074, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889000/
pdf.

Moyer:1980:LKD

[Moy80] Donald Franklin Moyer. Lord Kelvin, the Dynamic Victorian by Harold Issadore Sharlin, Tiby Sharlin (review). Technology and Culture, 21(3):502–504, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890723/pdf.

Mueller:1989:SPS

[Mue89] Milton Mueller. The switchboard problem: Scale, signaling, and organization in manual telephone switching, 1877–1897. Technology and Culture, 30(3):534–560, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889007/pdf.

Muhly:1981:ISO

[Muh81] J. D. Muhly. II. Summary: The origin of agriculture and technology — West or East Asia? Technology and Culture, 22(1):125–148, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890802/pdf.

Mulholland:1981:MSE

[Mul81] James A. Mulholland. Materials Science and Engineering: Its Evolution, Practice and Prospects ed. by Morris Cohen (review). Technology and Culture, 22(1):164–166, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890807/pdf.

Mulholland:1982:DUW

[Mul82] James A. Mulholland. To Draw, Upset, and Weld: The Work of the Pennsylvania Rural Blacksmith 1742–1935 by Jeanette Lasansky (review). Technology and Culture, 23(3):497–498, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891203/
pdf.

Multhauf:1983:HCE

[Mul83a] R. P. Multhauf. History of Chemical Engineering ed. by William F. Furter (review). Technology and Culture, 24(4):716–717, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890153/pdf.

Multhauf:1983:RAE

[Mul83b] R. P. Multhauf. The Rise of the American Electro-Chemicals Industry, 1880–1910: Studies in the American Technological Environment by Martha Moore Trescott (review). Technology and Culture, 24(3):519–520, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891407/pdf.

Multhauf:1983:SOL

[Mul83c] Robert P. Multhauf. Les Salines de l’ouest: Logique technique. Logique sociale by Pierre Lemonnier (review). Technology and Culture, 24(1):115–116, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891228/pdf.

Multhauf:1983:LSH

[Mul83d] Robert P. Multhauf. The Land of Steven and Huygens: A Sketch of Science and Technology in the Dutch Republic during the Golden Century by Dirk J. Struik (review). Technology and Culture, 24(2): 263, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891262/pdf.

Mulligan:1984:EIA

[Mul84a] William H. Mulligan, Jr. The Emergence of Industrial America: Strategic Factors in American Economic Growth since 1870 by Peter George (review). Technology and Culture, 25(3):662–663, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890011/pdf.

Multhauf:1984:IBM

[Mul84b] Lettie S. Multhauf. Isaac Beeckman (1588–1637) en de mechanisering van het wereldbeeld [Isaac Beeckman (1588–1637) and the mechanization of the world picture] by Klaas van Berkel (review). Technology and Culture, 25(2):334–335, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890056/pdf.

Multhauf:1985:LLL

[Mul85] Lettie S. Multhauf. The light of lamp-lanterns: Street lighting in 17th-Century Amsterdam. Technology and Culture, 26(2):236–252, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889763/pdf.

Multhauf:1987:SCH

[Mul87] R. P. Multhauf. The Servant of Colour: A History of the Society of Dyers and Colourists 1884–1984 by Maurice Tordoff (review). Technology and Culture, 28(1):150–151, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889469/pdf.

Multhauf:1988:SCC

[Mul88a] R. P. Multhauf. Science and Civilisation in China. Vol. 5: Chemistry and Chemical Technology. Pt. 7: Military Technology; the Gunpowder Epic by Joseph Needham (review). Technology and Culture, 29(4):943–946, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889183/pdf.

Multhauf:1988:SDS

[Mul88b] Robert P. Multhauf. Salt of the Desert Sun: A History of Salt Production and Trade in the Central Sudan by Paul E. Lovejoy (review). Technology and Culture, 29(3):682–683, July 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889240/pdf.

Multhauf:1989:FBT

[Mul89a] R. P. Multhauf. Der Freiberger Bergbau: Technische Denkmale und Geschichte ed. by Otfried Wagenbreth, Eberhard Wächtler (review). Technology and Culture, 30(3):674–675, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889041/pdf.

Multhauf:1989:SDB

[Mul89b] R. P. Multhauf. Die Song-Dynastie (960 bis 1279): Eine neue Gesellschaft im Spiegel ihrer Kultur by Dieter Kuhn (review). Technology and Culture, 30(1):152–153, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889134/pdf.

Multhauf:1989:CEH

[Mul89c] R. P. Multhauf. The Cambridge Economic History of Europe. Vol. 2: Trade and Industry in the Middle Ages ed. by M. M. Postan, Edward Miller (review). Technology and Culture, 30(4):1041–1042, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888980/pdf.

Musson:1983:SPB

[Mus83] A. E. Musson. The Selected Papers of Boulton and Watt. Vol. 1: The Engine Partnership, 1775–1825 ed. by Jennifer Tann (review). Technology and Culture, 24(3):513, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891402/pdf.

Myers:1987:CTC

[Mye87] Emlen Myers. Ceramic Theory and Cultural Process by Dean E. Arnold (review). Technology and Culture, 28(3):689–690, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889363/pdf.

Myers:1989:ACS

[Mye89] Emlen Myers. Andalusian Ceramics in Spain and New Spain: A Cultural Register from the Third Century B.C. to 1700 by Florence C. Lister, Robert H. Lister (review). Technology and Culture, 30(3):671–672, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889039/pdf.

Nash:1980:JUH

[Nas80] Gerald D. Nash. Journal of Urban History ed. by Joel A. Tarr (review). Technology and Culture, 21(4):637–639, October 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890757/pdf.

Nash:1989:PVC

[Nas89] Gerald D. Nash. Pacific Visions: California Scientists and the Environment 1850–1915 by Michael L. Smith (review). Technology and Culture, 30(2):498–499, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889107/pdf.

Needell:1983:IES

[Nee83] Allan A. Needell. The Illustrated Encyclopedia of Space Technology: A Comprehensive History of Space Exploration ed. by Kenneth Gatland (review). Technology and Culture, 24(4):707–709, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890147/pdf.

Neithercut:1988:SML

[Nei88] Mark E. Neithercut. Sawpower: Making Lumber in the Sawmills of Novel Scotia by Barbara R. Robertson (review). Technology and Culture, 29(4):960–961, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889194/pdf.

Nelson:1981:SNW

[Nel81] Daniel Nelson. In Search of the New Working Class: Automation and Social Integration within the Capitalist Enterprise by Duncan Gallie (review). Technology and Culture, 22(2):358–361, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890876/pdf.

Nelson:1982:WWW

[Nel82] Daniel Nelson. Women, War and Work: The Impact of World War I on Women Workers in the United States by Maurine Weiner Greenwald (review). Technology and Culture, 23(2):264– 265, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891169/pdf.

Nelson:1987:STL

[Nel87] Daniel Nelson. Science, Technology and the Labour Process. Marxist Studies, vols. 1 and 2 ed. by Les Levidow, Bob Young (review). Technology and Culture, 28(1):136–137, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889461/pdf.

Nelson:1989:TTU

[Nel89] Anna Kasten Nelson. Thinking in Time: The Uses of History for Decision-Makers by Richard E. Neustadt, Ernest R. May (review). Technology and Culture, 30(2):503–504, April 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889111/pdf.

Neufeld:1988:HOL

[Neu88] Michael J. Neufeld. Hermann Oberth: Leben, Werk und Auswirkung auf die spätere Raumfahrt-entwicklung by Hans Barth (review). Technology and Culture, 29(2):334–335, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889298/pdf.

Neushul:1989:SWC

[Neu89] Peter Neushul. Seaweed for war: California’s World War I kelp industry. Technology and Culture, 30(3):561–583, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889009/pdf.

Nevitt:1980:PGC

[Nev80] Barrington Nevitt. Pipeline or Grapevine: The changing communications environment. Technology and Culture, 21(2):217–226, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890675/pdf.

Nevitt:1982:MIT

[Nev82] Barrington Nevitt. The Myths of Information: Technology and Postindustrial Culture ed. by Kathleen Woodward (review). Technology and Culture, 23(1):93–95, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891112/pdf.

Newell:1983:SOE

[New83] Dianne Newell. The Social Organization of Early Industrial Capitalism by Michael B. Katz, et al. (review). Technology and Culture, 24(4):678–680, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890130/pdf.

Newell:1985:ADW

[New85] Dianne Newell. “All in a day’s work”: Local invention on the Ontario mining frontier. Technology and Culture, 26(4):799–814, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889678/pdf.

Nicholas:1980:AEI

[Nic80] S. J. Nicholas. The American export invasion of Britain: The case of the engineering industry, 1870–1914. Technology and Culture, 21(4):570–588, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890749/pdf.

Nicholas:1981:MQM

[Nic81] S. J. Nicholas. The Machinery Question and the Making of Political Economy, 1815–1848 by Maxine Berg (review). Technology and Culture, 22(2):356–358, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890875/pdf.

Numbers:1988:FPM

[Num88] Ronald L. Numbers. Fair Play in the Marketplace: The First Battle for Pure Food and Drugs by Mitchell Okun (review). Technology and Culture, 29(2):315–316, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889286/pdf.

Nuttall:1981:FMH

[Nut81] R. H. Nuttall. The first microscope of Henry Clifton sorby. Technology and Culture, 22(2):275–280, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890845/pdf.

Nuwer:1988:BFP

[Nuw88] Michael Nuwer. From batch to flow: Production technology and work-force skills in the steel industry, 1880–1920. Technology and Culture, 29(4):808–838, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889159/pdf.

Nye:1988:WMC

[Nye88] David E. Nye. The Workers’ Museum in Copenhagen. Technology and Culture, 29(4):909–912, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889164/pdf.

Nye:1989:WOT

[Nye89] David E. Nye. When Old Technologies Were New: Thinking about Electric Communication in the Late Nineteenth Century by Carolyn Marvin (review). Technology and Culture, 30(4):1045–1046, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888983/pdf.

OBrien:1981:PAE

[O’B81] Michael J. O’Brien. Prehistoric Architecture in the Eastern United States by William N. Morgan (review). Technology and Culture, 22(4):780–782, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890945/pdf.

OBannon:1987:WCM

[O’B87] Patrick W. O’Bannon. Waves of change: Mechanization in the Pacific Coast canned-salmon industry, 1864–1914. Technology and Culture, 28(3):558–577, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889341/pdf.

OConnor:1988:WCJ

[O’C88] Richard O’Connor. Work and Community in the Jungle: Chicago’s Packinghouse Workers, 1894–1922 by James R. Barrett (review). Technology and Culture, 29(4):924–925, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889172/pdf.

Ochs:1987:GMC

[Och87] Kathleen H. Ochs. The Growth of Minds and Cultures: A Unified Theory of the Structure of Human Experience by Willem H. Vanderburg (review). Technology and Culture, 28(2):399–400, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889450/pdf.

Ochs:1988:CMC

[Och88] Kathleen Ochs. Coal Mining in China’s Economy and Society 1895–1937 by Tim Wright (review). Technology and Culture, 29 (3):684–685, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889241/pdf.

Ochs:1989:WSA

[Och89] Kathleen Ochs. Women Scientists from Antiquity to the Present: An Index by Caroline L. Herzenberg, and: Women of Science, Technology, and Medicine: A Bibliography by Else Høyrup, and: Women in Science: Antiquity through the Nineteenth Century — a Biographical Dictionary with Annotated Bibliography by Marilyn Bailey Ogilvie, and: The History of Women and Science, Health, and Technology: A Bibliographic Guide to the Professions and Disciplines ed. by Susan E. Searing, Rima D. Apple (review). Technology and Culture, 30(4):1088–1089, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889015/pdf.

Oertling:1989:SPE

[Oer89] Thomas J. Oertling. A suction pump from an Early-16th-Century shipwreck. Technology and Culture, 30(3):584–595, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889011/pdf.

Olasky:1985:LDP

[Ola85] Marvin Olasky. Lammot du Pont and the American Explosives Industry 1850–1884 by Norman B. Wilkinson (review). Technology and Culture, 26(4):862–863, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889700/pdf.

Olson:1989:BUR

[Ols89] Gordon Olson. Bourgeois Utopias: The Rise and Fall of Suburbia by Robert Fishman (review). Technology and Culture, 30(2):457– 458, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889076/pdf.

Overmann:1985:FST

[Ove85] Ronald Overmann. Frontiers in Science and Technology: A Selected Outlook by the Committee on Science, Engineering, and Public Policy (National Academy of Sciences, National Academy of Engineering, Institute of Medicine) (review). Technology and Culture, 26(3):683–685, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889757/pdf.

Ovitt:1985:PJT

[Ovi85] George Ovitt, Jr. Pseudo-Juanelo Turriano: The Twenty-One Books of Devices and of Machines by José A. García–Diego (review). Technology and Culture, 26(4):828–829, October 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889682/pdf.

Ovitt:1986:PAN

[Ovi86a] George Ovitt. Planispheric Astrolabes from the National Museum of American History by Sharon Gibbs, George Saliba (review). Technology and Culture, 27(2):296–297, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889597/pdf.

Ovitt:1986:CCW

[Ovi86b] George Ovitt, Jr. The cultural context of Western technology: Early Christian attitudes toward manual labor. Technology and Culture, 27(3):477–500, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889524/pdf.

Ovitt:1987:STM

[Ovi87] George Ovitt, Jr. Science and Technology in Medieval Society ed. by Pamela O. Long (review). Technology and Culture, 28(2):349– 350, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889410/pdf.

Ovitt:1988:TSS

[Ovi88] George Ovitt, Jr. Time, Science, and Society in China and the West: The Study of Time V ed. by J. T. Fraser, et al. (review). Technology and Culture, 29(4):946–947, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889184/pdf.

Owens:1986:VBD

[Owe86] Larry Owens. Vannevar Bush and the differential analyzer: The text and context of an early computer. Technology and Culture, 27(1):63–95, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889635/pdf.

Palmer:1980:CWT

[Pal80] Bryan D. Palmer. The Canadian Worker in the Twentieth Century ed. by Irving Abella, David Millar (review). Technology and Culture, 21(2):257–259, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890689/pdf.

Palmer:1981:STC

[Pal81] Bryan D. Palmer. Science, Technology, and Canadian History: The First Conference on the Study of the History of Canadian Science and Technology (Les Sciences, la technologie et l’histoire canadienne: Premier Congrès sur l’histoire des sciences et de la technologie canadiennes) ed. by Richard A. Jarrell, Norman R. Ball (review). Technology and Culture, 22(2):322–325, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890861/pdf.

Pan:1987:SCC

[Pan87] Jixing Pan. Science and Civilisation in China. Vol. 5: Chemistry and Chemical Technology. Part 1: Paper and Printing by Tsien Tsuen-Hsuin (review). Technology and Culture, 28(2):345–346, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889407/pdf.

Papp:1982:SAR

[Pap82] Daniel S. Papp. Soviet–American Relations in the 1980s: Superpower Politics and East-West Trade by Lawrence T. Caldwell, William Diebold, Jr. (review). Technology and Culture, 23(4): 691–692, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890274/pdf.

Papanek:1984:STC

[Pap84] Hanna Papanek. Scientific-Technological Change and the Role of Women in Development ed. by Pamela M. D’Onofrio-Flores, Sheila M. Pfafflin (review). Technology and Culture, 25(1):126–129, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890094/pdf.

Parr:1982:CAI

[Par82] J. Gordon Parr. The Coming of the Age of Iron ed. by Theodore Wertime, James Muhly (review). Technology and Culture, 23 (2):230–232, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891152/pdf.

Paskoff:1985:RIA

[Pas85] Paul Paskoff. The Rise of Industrial America: A People’s History of the Post-Reconstruction Era by Page Smith (review). Technology and Culture, 26(4):857–859, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889697/pdf.

Paul:1983:CDC

[Pau83] Rodman W. Paul. The California Debris Commission: A History by Joseph J. Hagwood, Jr. (review). Technology and Culture, 24(3):534–535, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891417/pdf.

Payne:1985:HCG

[Pay85] Peter L. Payne. Henry Cort: The Great Finer, Creator of Puddled Iron by R. A. Mott (review). Technology and Culture, 26 (4):842–844, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889689/pdf.

Peckham:1986:TCB

[Pec86] Brian W. Peckham. Technological change in the British and French starch industries, 1750–1850. Technology and Culture, 27(1):18–39, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889633/pdf.

Pelzman:1983:SIU

[Pel83] Joseph Pelzman. Science and Industrialisation in the USSR: industrial Research and Development, 1917–1940 by Robert Lewis (review). Technology and Culture, 24(1):137–138, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891241/
pdf.

Penn:1980:SMH

[Pen80] Theodore Z. Penn. The Slater mill historic site and the Wilkinson mill machine shop exhibit. Technology and Culture, 21(1):56–66, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890605/pdf.

Penn:1982:TPF

[Pen82] Theodore Z. Penn. Traces of the Past: A Field Guide to Industrial Archeology by David Weitzman (review). Technology and Culture, 23(2):240–242, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891157/pdf.

Penn:1984:IDN

[Pen84] Theodore Z. Penn. Innovation Diffusion: A New Perspective by Lawrence A. Brown (review). Technology and Culture, 25(2):316– 317, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890046/pdf.

Perkins:1980:DAV

[Per80a] Daniel J. Perkins. Directory of Audio-Visual Sources: History of Science, Medicine, and Technology by Bruce Eastwood (review). Technology and Culture, 21(4):716–718, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890794/pdf.

Perkins:1980:ETH

[Per80b] John H. Perkins. Environment, Technology, and Health: Human Ecology in Historical Perspective by Merrill Eisenbud (review). Technology and Culture, 21(2):272–275, April 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890697/pdf.

Perry:1980:NDW

[Per80c] Robert Perry. A New Dimension, Wallops Island Flight Test Range: The First Fifteen Years by Joseph Adams Shortal (review). Technology and Culture, 21(3):530–531, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890738/pdf.

Perry:1980:LHP

[Per80d] Robert Perry. Liquid Hydrogen as a Propulsion Fuel, 1945–1959 by John L. Sloop (review). Technology and Culture, 21(1):136–138, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890638/pdf.

Perelman:1981:CSS

[Per81a] Lewis J. Perelman. The Comedy of Survival: In Search of an Environmental Ethic by Joseph W. Meeker (review). Technology and Culture, 22(3):692–693, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890934/pdf.

Pershey:1981:PRA

[Per81b] Edward J. Pershey. Pollution and Reform in American Cities, 1870–1930 ed. by Martin V. Melosi (review). Technology and Culture, 22(3):630–633, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890908/pdf.

Pershey:1981:IGT

[Per81c] Edward Jay Pershey. Infection and the germ theory. Technology and Culture, 22(4):852, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890979/pdf.

Perry:1984:TSC

[Per84a] Charles R. Perry. Technology and Social Complexity by Maurice N. Richter, Jr. (review). Technology and Culture, 25(4):847–849, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889951/pdf.

Perry:1984:CRP

[Per84b] Charles R. Perry. The Communications Revolution in Politics ed. by Gerald Benjamin (review). Technology and Culture, 25 (1):122–123, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890091/pdf.

Perkins:1985:ARB

[Per85a] Patricia Perkins. The Amazon Rubber Boom 1850–1920 by Barbara Weinstein (review). Technology and Culture, 26(4):865–867, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889702/pdf.

Pershey:1985:LTS

[Per85b] Edward Jay Pershey. “Life and Times in Silk City,” an Exhibit in Paterson, New Jersey, at the American Labor Museum, the Paterson Museum, and the Lambert Castle Museum. Technology and Culture, 26(3):623–628, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889725/pdf.

Pershey:1986:ISA

[Per86] Edward Jay Pershey. The Invented Self: An Anti-Biography, from Documents of Thomas A. Edison by David E. Nye (review). Technology and Culture, 27(4):863–865, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889520/pdf.

Petroski:1983:CPI

[Pet83] Henry Petroski. On 19th-Century perceptions of iron bridge failures. Technology and Culture, 24(4):655–659, October 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890124/pdf.

Peters:1987:FGV

[Pet87] Tom F. Peters. Der Fahrstuhl: Die Geschichte der vertikalen Eroberung by Jeannot Simmen, Uwe Drepper (review). Technology and Culture, 28(4):883–885, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889331/pdf.

Peters:1988:MAP

[Pet88] Edward Peters. Le Moyen Age pour quoi faire? by Régine Pernoud, et al. (review). Technology and Culture, 29(3):667–668, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889232/pdf.

Peters:1989:ACB

[Pet89] Tom F. Peters. Ancient Chinese Bridges by Tang Huancheng (review). Technology and Culture, 30(1):151–152, January 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889133/pdf.

Pfeiffer:1983:IHT

[Pfe83] William Sanborn Pfeiffer. Interface ’82 — Humanities and Technology: Southern Technical Institute, Marietta, Georgia, October 21–22, 1982. Technology and Culture, 24(4):664–671, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890126/pdf.

Phillips:1983:MAM

[Phi83] Mona Spangler Phillips. The manufacture of ancient Middle Eastern clay envelopes. Technology and Culture, 24(2):256–257, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891258/pdf.

Piasecki:1984:LCS

[Pia84] Bruce Piasecki. Love Canal: Science, Politics, and People by Adeline Gordon Levine (review). Technology and Culture, 25(3):669– 672, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890016/pdf.

Pinch:1984:MKE

[Pin84] Trevor Pinch. The Manufacture of Knowledge: An Essay on the Constructivist and Contextual Nature of Science by Karin D. Knorr-Cetina (review). Technology and Culture, 25(1):130–131, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890096/pdf.

Pinch:1985:SSA

[Pin85] Trevor Pinch. Sociology of Sciences: An Annotated Bibliography on Invisible Colleges, 1972–1981 by Daryl E. Chubin (review). Technology and Culture, 26(3):662–664, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889746/pdf.

Pitts:1986:SSS

[Pit86] John A. S. Pitts. Seek and Strike: Sonar, Anti-Submarine Warfare and the Royal Navy, 1914–1954 by Willem Hackmann (review). Technology and Culture, 27(2):307–309, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889612/pdf.

Platt:1986:SSM

[Pla86a] Harold L. Platt. Sault Ste. Marie: A Project Report by Terry S. Reynolds (review). Technology and Culture, 27(3):634–635, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889578/pdf.

Platt:1986:GGU

[Pla86b] Harold L. Platt. The Granite Garden: Urban Nature and Human Design by Anne Spirn (review). Technology and Culture, 27 (2):332–334, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889626/pdf.

Ponting:1980:HTK

[Pon80] K. G. Ponting. A History of Textiles by Kax Wilson (review). Technology and Culture, 21(4):655–657, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890765/pdf.

Porter:1980:ATD

[Por80] Alan L. Porter. Assessment of Technical Decisions: Case Studies by Ernest Braun, et al. (review). Technology and Culture, 21(3):543–544, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890743/pdf.

Porter:1989:LMA

[Por89a] Charlotte M. Porter. The Launching of Modern American Science 1846–1876 by Robert V. Bruce (review). Technology and Culture, 30(1):136–138, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889124/pdf.

Porter:1989:TES

[Por89b] Glenn Porter. Technology, the Economy, and Society: The American Experience ed. by Joel Colton, Stuart Bruchey (review). Technology and Culture, 30(4):1049–1050, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888985/pdf.

Post:1982:MET

[Pos82] Robert C. Post. A message from the Editor: Taking the baton. Technology and Culture, 23(1):1–2, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891101/pdf.

Powers:1983:VVO

[Pow83] Judith Powers. Vision or Villainy: Origins of the Owens Valley-Los Angeles Water Controversy by Abraham Hoffman (review). Technology and Culture, 24(3):536–538, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891418/pdf.

Prescott:1981:JPJ

[Pre81] John Prescott. James prescott Joule: A biographical note. Technology and Culture, 22(2):297–299, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890847/pdf.

Presley:1983:OBS

[Pre83a] James Presley. Oil Booms: Social Change in Five Texas Towns by Roger M. Olien, Diana Davids Olien (review). Technology and Culture, 24(3):531–532, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891415/pdf.

Pretzer:1983:MME

[Pre83b] William S. Pretzer. Mechanics and Manufacturers in the Early Industrial Revolution: Lynn, Massachusetts, 1780–1860 by Paul G. Faler (review). Technology and Culture, 24(2):272–273, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891269/
pdf.

Pretzer:1987:MBW

[Pre87] William S. Pretzer. The Mechanics of Baltimore: Workers and Politics in the Age of Revolution, 1763–1812 by Charles G. Steffen, and: The Practice of Solidarity: American Hat Finishers in the Nineteenth Century by David Bensman (review). Technology and Culture, 28(2):366–368, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889421/pdf.

Pritchard:1982:EEC

[Pri82] Michael S. Pritchard. Ethics and Engineering Curricula by Robert J. Baum (review). Technology and Culture, 23(3):522–523, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891217/pdf.

Pritchard:1987:SNC

[Pri87] James Pritchard. From shipwright to naval constructor: The professionalization of 18th-Century French naval shipbuilders. Technology and Culture, 28(1):1–25, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889435/pdf.

Pi-Sunyer:1982:NGG

[PS82] Oriol Pi-Sunyer. New Guinea Gardens: A Study of Husbandry with Parallels in Prehistoric Europe by Axel Steensberg (review). Technology and Culture, 23(4):644–646, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890249/pdf.

Pursell:1980:RCC

[Pur80] Carroll W. Pursell, Jr. The Roanoke Conference — critical issues in the history of technology: Roanoke, Virginia, August 14–18, 1978, I. Summary. Technology and Culture, 21(4):617–620, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890753/pdf.

Pursell:1981:WIA

[Pur81] Carroll Pursell. Women inventors in America. Technology and Culture, 22(3):545–549, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890886/pdf.

Pursell:1982:OEI

[Pur82a] Carroll W. Pursell, Jr. Oliver Evans, Inventive Genius of the American Industrial Revolution by Eugene S. Ferguson (review). Technology and Culture, 23(1):121–122, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891128/pdf.

Pursell:1982:STN

[Pur82b] Carroll W. Pursell, Jr. Science, Technology, and National Policy ed. by Thomas J. Kuehn, Alan L. Porter (review). Technology and Culture, 23(3):488–490, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891198/pdf.

Pursell:1983:LRH

[Pur83] Carroll W. Pursell, Jr. A List of References for the History of Agricultural Science in America by Margaret W. Rossiter (review). Technology and Culture, 24(4):715–716, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890152/pdf.

Puro:1984:STP

[Pur84a] Steven Puro. Science and, Technology Policy: Priorities of Governments by F. Tisdell (review). Technology and Culture, 25(1): 148–149, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890105/pdf.

Pursell:1984:YER

[Pur84b] Carroll W. Pursell, Jr. Yankee Enterprise: The Rise of the American System of Manufactures ed. by Otto Mayr, Robert C. Post (review). Technology and Culture, 25(1):150–151, January 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890107/pdf.

Pursell:1989:FHP

[Pur89a] Carroll Pursell. Friends in High Places: The Bechtel Story — the Most Secret Corporation and How It Engineered the World by Laton McCartney (review). Technology and Culture, 30(3):702– 704, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889064/pdf.

Pursell:1989:TTF

[Pur89b] Carroll Pursell. Teaching Technology from a Feminist Perspective: A Practical Guide by Joan Rothschild (review). Technology and Culture, 30(4):1089–1091, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889017/pdf.

Pycior:1982:PJH

[Pyc82] Helena M. Pycior. The Papers of Joseph Henry: The Princeton Years, January 1838–December 1840, vol 4 ed. by Nathan Reingold, et al. (review). Technology and Culture, 23(4):669–671, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890264/pdf.

Pyenson:1984:SCR

[Pye84] Lewis Pyenson. Science in Context: Readings in the Sociology of Science ed. by Barry Barnes, David Edge (review). Technology and Culture, 25(1):132–135, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890097/pdf.

Pyne:1984:RFH

[Pyn84] Stephen J. Pyne. Radio for the Fireline: A History of Electronic Communication in the Forest Service, 1905–1975 by Gary Craven Gray (review). Technology and Culture, 25(1):118–120, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890089/pdf.

Quimby:1989:JNM

[Qui89] Maureen Quimby. Just New from the Mills: Printed Cottons in America, Late Nineteenth and Early Twentieth Centuries, from the Collection of the Museum of American Textile History by Diane L. F. Affleck (review). Technology and Culture, 30(1):147–149, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889131/pdf.

Qureshi:1986:TCN

[Qur86] M. A. Qureshi. Technology & Culture. No. 9: Technology Policy Trends and Information Services in Sweden and China — a Comparison by Sun Xuechen, No. 10: Chicken and Egg: Electronics and Social Change in India by Ward Morehouse, Ravi Chopra, No. 11: Information Infrastructures in India and China by Erik Baark (review). Technology and Culture, 27(1):181–183, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889670/pdf.

Rabkin:1981:PNR

[Rab81] Yakov M. Rabkin. Physics at the National Research Council of Canada, 1929–1952 by W. E. K. Middleton (review). Technology and Culture, 22(1):211–213, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890834/pdf.

Rae:1980:ARA

[Rae80a] John B. Rae. Americans on the Road: From Auto-Camp to Motel, 1911–1945 by Warren James Belasco (review). Technology and Culture, 21(2):262–264, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890691/pdf.

Rae:1980:HHC

[Rae80b] John B. Rae. The Herbert Hoover Collection of Mining and Metallurgy. Technology and Culture, 21(4):614–616, October 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890752/pdf.

Rae:1982:TTS

[Rae82] John B. Rae. Transport Technology and Social Change ed. by Per Sorböm (review). Technology and Culture, 23(1):138–140, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891139/pdf.

Rae:1984:RW

[Rae84a] John B. Rae. Robert S. Woodbury (1907–1983). Technology and Culture, 25(4):911–912, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889988/pdf.

Rae:1984:WDW

[Rae84b] John B. Rae. What did we expect SHOT to wrought? Technology and Culture, 25(4):731–734, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889943/pdf.

Rae:1985:BKW

[Rae85] John B. Rae. Boss Kettering: Wizard of General Motors by Stuart W. Leslie (review). Technology and Culture, 26(3):654–655, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889741/pdf.

Rae:1987:JAI

[Rae87] John B. Rae. The Japanese Automobile Industry: Technology and Management at Nissan and Toyota by Michael A. Cusumano (review). Technology and Culture, 28(2):380–381, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889430/pdf.

Rahman:1982:HFT

[Rah82] A. Rahman. Homo Faber: Technology and Culture in India, China and the West from 1500 to the Present Day by Claude Alvares (review). Technology and Culture, 23(3):479–481, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891194/pdf.

Randall:1982:BCN

[Ran82a] Adrian Randall. British Coalminers in the Nineteenth Century: A Social History by John Benson (review). Technology and Culture, 23(2):260–263, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891167/pdf.

Randall:1982:MJP

[Ran82b] Adrian Randall. Masters and Journeymen: A Prehistory of Industrial Relations, 1717–1800 by C. R. Dobson (review). Technology and Culture, 23(2):252–256, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891164/pdf.

Randall:1982:SEW

[Ran82c] Adrian Randall. Skill and the English Working Class, 1870–1914 by Charles More (review). Technology and Culture, 23(2):258–260, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891166/pdf.

Randall:1986:SIS

[Ran86a] Adrian J. Randall. Sweated Industries and Sweated Labor: The London Clothing Trades, 1860–1914 by James A. Schmiechen (review). Technology and Culture, 27(1):133–135, January 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889646/pdf.

Randall:1986:PLC

[Ran86b] Adrian J. Randall. The philosophy of luddism: The case of the West of England woolen workers, ca. 1790–1809. Technology and Culture, 27(1):1–17, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889632/pdf.

Rapp:1981:TPD

[Rap81] Friedrich Rapp. Das Technikverständnis in der Philosophie der DDR by Bernd M. Lindenberg (review). Technology and Culture, 22(3):670–671, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890925/pdf.

Rasmussen:1980:MGG

[Ras80a] Wayne D. Rasmussen. Manufacturing Green Gold: The Conditions and Social Consequences of Lettuce Harvest Mechanization by William H. Friedland, et al. (review). Technology and Culture, 21(1):113–114, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890625/pdf.

Rasmussen:1980:TJF

[Ras80b] Wayne D. Rasmussen. Thomas Jefferson’s Farm Book, with Commentary and Relevant Extracts from Other Writings ed. by Edwin Morris Betts (review). Technology and Culture, 21(4):668–669, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890771/pdf.

Ravetz:1981:LSI

[Rav81] J. R. Ravetz. Limits of Scientific Inquiry ed. by Gerald Holton, Robert S. Monson (review). Technology and Culture, 22(1): 219–220, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890837/pdf.

Rostoker:1984:CIB

[RBD84] William Rostoker, Bennet Bronson, and James Dvorak. The castiron bells of China. Technology and Culture, 25(4):750–767, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889945/pdf.

Reader:1981:LCL

[Rea81] W. J. Reader. The Lilleshall Company Limited: A History, 1764– 1964 by W. K. V. Gale, C. R. Nicholls (review). Technology and Culture, 22(4):807–809, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890959/pdf.

Recoulley:1988:IHT

[Rec88] Alfred L. Recoulley. Interface ’87: Humanities and Technology — Atlanta, Georgia, October 22–23, 1987. Technology and Culture, 29(3):661–662, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889230/pdf.

Rees:1981:GRW

[Ree81] Eberhard F. M. Rees. Geschichte der Raumfahrt by Werner Buedeler (review). Technology and Culture, 22(4):816–819, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890963/pdf.

Reich:1980:RTA

[Rei80a] Leonard S. Reich. A Retrospective Technology Assessment: Submarine Telegraphy — the Transatlantic Cable of 1866 by Vary Coates, et al. (review). Technology and Culture, 21(4):684–686, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890778/pdf.

Reingold:1980:AST

[Rei80b] Nathan Reingold. American Science and Technology: A Bicentennial Bibliography by George W. Black, Jr. (review). Technology and Culture, 21(1):107–108, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890621/pdf.

Reingold:1982:DSC

[Rei82] Nathan Reingold. DDT: Scientists, Citizens, and Public Policy by Thomas R. Dunlap (review). Technology and Culture, 23(3):490– 491, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891199/pdf.

Reich:1983:ILP

[Rei83a] Leonard S. Reich. Irving Langmuir and the pursuit of science and technology in the corporate environment. Technology and Culture, 24(2):199–221, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891255/pdf.

Reingold:1983:BSV

[Rei83b] Nathan Reingold. Between Science and Values by Loren R. Graham (review). Technology and Culture, 24(2):309–311, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891290/
pdf.

Reingold:1987:CAH

[Rei87] Nathan Reingold. Chemistry in America, 1876–1976: Historical Indicators by Arnold Thackray, et al. (review). Technology and Culture, 28(3):732–733, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889392/pdf.

Reingold:1989:OEA

[Rei89] Nathan Reingold. Osiris ed. by Arnold Thackray (review). Technology and Culture, 30(4):1097–1099, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889026/pdf.

Restivo:1981:TPD

[Res81] Sal Restivo. Technics and Praxis by Don Ihde (review). Technology and Culture, 22(3):672–674, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890926/pdf.

Reuss:1985:AHD

[Reu85] Martin Reuss. Andrew A. Humphreys and the development of hydraulic engineering: Politics and technology in the Army Corps of Engineers, 1850–1950. Technology and Culture, 26(1):1–33, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889806/pdf.

Reuss:1987:RPR

[Reu87] Martin Reuss. The review of Province of Reason. Technology and Culture, 28(3):683–684, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889360/pdf.

Reuss:1989:BAH

[Reu89] Martin Reuss. Building the American Highway System: Engineers as Policy Makers by Bruce E. Seely (review). Technology and Culture, 30(2):470–472, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889090/pdf.

Reynolds:1984:NPE

[Rey84] Terry S. Reynolds. Networks of Power: Electrification in Western Society, 1880–1930 by Thomas P. Hughes (review). Technology and Culture, 25(3):644–647, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890002/pdf.

Reynolds:1986:DPB

[Rey86] Terry S. Reynolds. Defining professional boundaries: Chemical engineering in the early 20th century. Technology and Culture, 27(4):694–716, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889496/pdf.

Reynolds:1987:GRE

[Rey87] Terry S. Reynolds. Great River: An Environmental History of the Upper Mississippi, 1890–1950 by Philip V. Scarpino (review). Technology and Culture, 28(3):718–719, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889381/pdf.

Reynolds:1989:HIE

[Rey89a] Terry S. Reynolds. A History of the Institution of Electrical Engineers 1871–1971 by W. J. Reader (review). Technology and Culture, 30(2):479–481, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889096/pdf.

Reynolds:1989:WPW

[Rey89b] Terry S. Reynolds. Water Power in the “Wilderness”: The History of Bonneville Lock and Dam by William F. Willingham (review). Technology and Culture, 30(1):141–142, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889127/pdf.

Reynolds:1989:IWM

[Rey89c] Terry S. Reynolds. Iron in the wilderness: The Michigan Iron Industry Museum. Technology and Culture, 30(1):112–117, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889116/pdf.

Rhees:1982:TAE

[Rhe82] David Rhees. Thomas Alva Edison: An American Myth by Wyn Wachhorst (review). Technology and Culture, 23(4):676–678, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890267/pdf.

Rhees:1986:SVP

[Rhe86] David J. Rhees. Science versus Practice: Chemistry in Victorian Britain by Robert F. Bud, Gerrylynn K. Roberts (review). Technology and Culture, 27(2):301–302, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889605/pdf.

Rhees:1988:TUA

[Rhe88] David J. Rhees. Technological Utopianism in American Culture by Howard P. Segal (review). Technology and Culture, 29(2):319–320, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889289/pdf.

Rightmyer:1981:KDH

[Rig81] Don Rightmyer. Kill Devil Hill: Discovering the Secret of the Wright Brothers by Harry Combs (review). Technology and Culture, 22(2):339–341, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890867/pdf.

Ritvo:1988:LEL

[Rit88] Harriet Ritvo. Like Engend’ring Like: Heredity and Animal Breeding in Early Modern England by Nicholas Russell (review). Technology and Culture, 29(2):293–295, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889271/pdf.

Robinson:1980:SSM

[Rob80] Eric Robinson. Sir Samuel Morland, Diplomat and Inventor, 1625–1695 by H. W. Dickinson (review). Technology and Culture, 21(4):654–655, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890764/pdf.

Robinson:1985:PSI

[Rob85] Eric Robinson. The Patent System and Inventive Activity during the Industrial Revolution, 1750–1852 by Harold I. Dutton (review). Technology and Culture, 26(4):840–841, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889688/pdf.

Robbins:1988:WDR

[Rob88] William G. Robbins. Workers and Dissent in the Redwood Empire by Daniel A. Cornford (review). Technology and Culture, 29(4):922–924, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889171/pdf.

Roessner:1982:AHT

[Roe82] J. David Roessner. Across the High Technology Threshold: The Case of Synthetic Rubber by Robert Solo (review). Technology and Culture, 23(2):274–277, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891175/pdf.

Roell:1985:GBD

[Roe85a] Craig H. Roell. Giraffes, Black Dragons, and Other Pianos: A Technological History from Cristofori to the Modern Concert Grand by Edwin M. Good (review). Technology and Culture, 26 (2):337–339, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889798/pdf.

Roemer:1985:TCU

[Roe85b] Kenneth M. Roemer. Technology, corporation, and utopia: Gillette’s unity regained. Technology and Culture, 26(3):560–570, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889717/pdf.

Roemer:1989:PIU

[Roe89] Kenneth M. Roemer. The Post-Industrial Utopians by Boris Frankel (review). Technology and Culture, 30(1):174–175, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889148/pdf.

Rohe:1984:BSH

[Roh84] Randall Rohe. Black Sands: A History of the Mineral Sand Mining Industry in Eastern Australia by I. W. Morley (review). Technology and Culture, 25(4):860–861, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889958/pdf.

Roland:1980:WBH

[Rol80] Alex Roland. The Wright Brothers: Heirs of Prometheus ed. by Richard P. Hallion (review). Technology and Culture, 21(1): 130–132, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890634/pdf.

Roland:1982:DWA

[Rol82] Alex Roland. A Dream of Wings: Americans and the Airplane, 1875–1905 by Tom D. Crouch (review). Technology and Culture, 23(1):132–133, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891136/pdf.

Roland:1983:HSF

[Rol83] Alex Roland. The High-Speed Frontier: Case Histories of Four NACA Programs, 1920–1950 by John V. Becker (review). Technology and Culture, 24(1):144–146, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891245/pdf.

Roland:1987:IMS

[Rol87] Alex Roland. Infernal Machines: The Story of Confederate Submarine and Mine Warfare by Milton F. Perry (review). Technology and Culture, 28(3):713–714, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889378/pdf.

Roland:1989:TWB

[Rol89] Alex Roland. Technology and War: From 2000 B.C. to the Present by Martin van Creveld (review). Technology and Culture, 30(4): 1037–1039, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888978/pdf.

Rose:1981:EWW

[Ros81] Mark H. Rose. Energy and the Way We Live: A Course by Newspaper Reader ed. by Melvin Kranzberg, et al. (review). Technology and Culture, 22(3):658–660, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890921/pdf.

Rose:1984:UET

[Ros84] Mark H. Rose. Urban environments and technological innovation: Energy choices in Denver and Kansas city, 1900–1940. Technology and Culture, 25(3):503–539, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889991/pdf.

Rosner:1985:CHI

[Ros85] David Rosner. Compulsory Health Insurance: The Continuing American Debate ed. by Ronald L. Numbers (review). Technology and Culture, 26(2):328–329, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889793/pdf.

Rosen:1987:CCD

[Ros87a] Howard Rosen. “Chips and changes” — the development and impact of the computer chip. Technology and Culture, 28(1):99–102, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889449/pdf.

Rostoker:1987:DTC

[Ros87b] William Rostoker. Dabieshan: Traditional Chinese Iron-Production Techniques Practised in Southern Henan in the Twentieth Century by Donald B. Wagner (review). Technology and Culture, 28(2):346–347, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889408/pdf.

Rosen:1989:JUH

[Ros89] Christine Rosen. Journal of Urban History. Vol. 14, no. 1: The City and Technology ed. by Mark H. Rose, Joel A. Tarr (review). Technology and Culture, 30(4):1070–1072, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888999/pdf.

Rothenberg:1982:RAP

[Rot82] Marc Rothenberg. The Rainmakers: American “Pluviculture” to World War II by Clark C. Spence (review). Technology and Culture, 23(2):248–249, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891161/pdf.

Rothstein:1983:AIE

[Rot83] Morton Rothstein. American Industrialization, Economic Expansion, and the Law ed. by Joseph R. Frese, Jacob Judd (review). Technology and Culture, 24(4):689–690, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890136/pdf.

Rothenberg:1984:BBM

[Rot84] Marc Rothenberg. Bullets and Bureaucrats: The Machine Gun and the United States Army, 1861–1916 by David A. Armstrong (review). Technology and Culture, 25(1):151–153, January 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890108/pdf.

Rothenberg:1985:TRT

[Rot85a] Marc Rothenberg. Tree Rings and Telescopes: The Scientific Career of A. E. Douglass by George Ernest Webb (review). Technology and Culture, 26(1):131–132, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889837/pdf.

Rothschild:1985:RHL

[Rot85b] Joan Rothschild. Redefining Human Life: Reproductive Technologies and Social Policy by Robert H. Blank, and: Test-Tube Women: What Future for Motherhood? ed. by Rita Arditti, et al. (review). Technology and Culture, 26(3):676–680, July 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889755/pdf.

Rowlett:1980:PAA

[Row80a] Elsebet Sander-Jørgensen Rowlett. Prehistoric Avebury by Aubrey Burl (review). Technology and Culture, 21(4):644–646, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890759/pdf.

Rowlett:1980:EPS

[Row80b] Ralph M. Rowlett. The Early Postglacial Settlement of Northern Europe ed. by Paul Mellars (review). Technology and Culture, 21(3):472–474, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890710/pdf.

Rowlett:1982:ASC

[Row82a] Ralph M. Rowlett. The Age of Stonehenge by Colin Burgess (review). Technology and Culture, 23(3):475–477, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891192/pdf.

Rowley:1982:CHA

[Row82b] Louis N. Rowley, Jr. A Centennial History of the American Society of Mechanical Engineers, 1880–1980 by Bruce Sinclair (review). Technology and Culture, 23(4):678–679, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890268/pdf.

Rowlett:1983:HCR

[Row83] Ralph M. Rowlett. Heartland of Cities by Robert McC. Adams (review). Technology and Culture, 24(3):506–508, July 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891398/pdf.

Rowlett:1986:PGE

[Row86] Ralph M. Rowlett. The Prehistory of Germanic Europe by Herbert Schutz (review). Technology and Culture, 27(1):127–129, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889642/pdf.

Rowley:1989:CDA

[Row89] William D. Rowley. Cadillac Desert: The American West and Its Disappearing Water by Marc Reisner (review). Technology and Culture, 30(2):493–494, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889104/pdf.

Ruiz:1980:ICS

[Rui80] Teofilo F. Ruiz. Islamic and Christian Spain in the Early Middle Ages: Comparative Perspective on Social and Cultural Formation by Thomas F. Glick (review). Technology and Culture, 21(3):479– 481, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890713/pdf.

Rumm:1987:CWT

[Rum87] John C. Rumm. Connecticut Workers and Technological Change by Robert Asher (review). Technology and Culture, 28(4):868–869, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889321/pdf.

Rumm:1988:WTA

[Rum88] John Rumm. Work Transformed: Automation and Labor in the Computer Age by Harley Shaiken (review). Technology and Culture, 29(4):941–942, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889182/pdf.

Rundell:1981:GHG

[Run81] Walter Rundell, Jr. Geologists and the History of Geology: An International Bibliography from the Origins to 1978 by William A. S. Sarjeant (review). Technology and Culture, 22(3):648–649, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890916/pdf.

Runte:1984:BGA

[Run84] Alfred Runte. Big Game in Alaska: A History of Wildlife and People by Morgan Sherwood (review). Technology and Culture, 25(4):903–904, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889985/pdf.

Rupp:1988:GWD

[Rup88] Leila J. Rupp. Gender at Work: The Dynamics of Job Segregation by Sex during World War II by Ruth Milkman (review). Technology and Culture, 29(4):936–937, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889179/pdf.

Ruttan:1989:RET

[Rut89] Vernon W. Ruttan. The Rice Economies: Technology and Development in Asian Societies by Francesca Bray (review). Technology and Culture, 30(3):714–715, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889081/pdf.

Ryder:1980:CTH

[Ryd80] Tom Ryder. Carriage Terminology: An Historical Dictionary by Don H. Berkebile (review). Technology and Culture, 21(1):116–118, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890627/pdf.

Rydell:1989:EVE

[Ryd89] Robert W. Rydell. Ephemeral Vistas: The Expositions Universelles, Great Exhibitions and World’s Fairs, 1851–1939 by Paul Greenhalgh (review). Technology and Culture, 30(4):1043–1045, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888982/pdf.

Sahm:1987:ECM

[Sah87] Virginia H. Sahm. The Evolution of the Cruise Missile by Kenneth P. Werrell (review). Technology and Culture, 28(2):384–386, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889433/pdf.

Salsbury:1983:TEN

[Sal83] Stephen Salsbury. Transportation in the Early Nation ed. by the Indiana Historical Society (review). Technology and Culture, 24 (4):690–692, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890137/pdf.

Salay:1984:CBM

[Sal84] David Salay. Coal: British Mining in Art 1680–1980, and: Coal: A Pictoral History of the British Coal Industry by D. Anderson (review). Technology and Culture, 25(4):854–856, October 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889956/pdf.

Sandler:1980:PWI

[San80] S. Sandler. In Peace and War: Interpretations of American Naval History, 1775–1978 ed. by Kenneth Hagan (review). Technology and Culture, 21(1):129–130, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890633/pdf.

Sandblom:1981:EAS

[San81a] Ulrika Sandblom. Elektrifieringen av Svenska Järnvägar. Motiv för val av system [The electrification of Swedish railways: motives for choice of systems] by Göran Rönn, and: Mekaniska Verkstadsbyggnader under 1800-talet [Mechanical workshops in Sweden during the 19th century] by Bo Salholm (review). Technology and Culture, 22(1):194–196, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890825/pdf.

Sandblom:1981:MHK

[San81b] Ulrika Sandblom. Marinturbinens historia, Del 1: Kolefs och ångans ind 1897–1927 [The history of the marine turbine, part 1: the period of coal and steam, 1897–1927] by Ingmar Jung, and: Bergshistoria: Mal, medel och arbetsmetodik vid studier och forskning [The history of mining and metallurgy: goals, means, and methods for study and research] by Erik Tholander (review). Technology and Culture, 22(4):805–807, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890958/pdf.

Sandberg:1984:ISE

[San84] Lars G. Sandberg. Iron and Steel on the European Market in the 17th Century: A Contemporary Account of Production Forms and Marketing ed. by the Historical Metallurgy Group of the Swedish Ironmasters’ Association (review). Technology and Culture, 25(4): 852–853, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889954/pdf.

Sanabria:1989:GRS

[San89] Sergio Luis Sanabria. From Gothic to Renaissance stereotomy: The design methods of philibert de l’Orme and alonso de vandelvira. Technology and Culture, 30(2):266–299, April 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889060/pdf.

Sardar:1981:ITT

[Sar81] Ziauddin Sardar. Integrated Technology Transfer ed. by Jacques Richardson (review). Technology and Culture, 22(3):683–686, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890930/pdf.

Saul:1980:HAB

[Sau80] S. B. Saul. Herbert Austin: The British Motor Car Industry to 1941 by Roy Church (review). Technology and Culture, 21(4): 662–664, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890768/pdf.

Saunders:1986:CBR

[Sau86] Richard Saunders. The Corn Belt Route: A History of the Chicago Great Western Railroad Company by H. Roger Grant (review). Technology and Culture, 27(1):168–169, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889664/pdf.

Sawyer:1988:HIP

[Saw88] William D. Sawyer. A History of Industrial Power in the United States, 1780–1930. Vol. 2: Steam Power by Louis C. Hunter (review). Technology and Culture, 29(3):691–694, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889246/pdf.

Schmandt-Besserat:1980:EBF

[SB80] Denise Schmandt-Besserat. The envelopes that bear the first writing. Technology and Culture, 21(3):357–385, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890702/pdf.

StGermain:1982:IHT

[SB82] Amos St. Germain and Carol M. Barnum. Interface ’81 — Humanities and Technology: Southern Technical Institute, Marietta, Georgia, October 22–23, 1981. Technology and Culture, 23(4): 625–630, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890243/pdf.

Schneider:1980:HTE

[Sch80] Ivo Schneider. History of Technology ed. by A. Rupert Hall, Norman Smith (review). Technology and Culture, 21(4):633–635, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890755/pdf.

Schallenberg:1981:ASB

[Sch81a] Richard H. Schallenberg. The anomalous storage battery: An American lag in early electrical engineering. Technology and Culture, 22(4):725–752, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890937/pdf.

Schlebecker:1981:SDC

[Sch81b] John T. Schlebecker. Seven Decades That Changed America: A History of the American Society of Agricultural Engineers, 1907– 1977 by Robert E. Stewart (review). Technology and Culture, 22(1):176–177, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890813/pdf.

Schoen:1981:GTE

[Sch81c] Lothar Schoen. Geschichte der Technik ed. by Burchard Brentjes, et al. (review). Technology and Culture, 22(3):607–608, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890895/
pdf.

Schatz:1982:FDD

[Sch82a] Ronald Schatz. The Five Dollar Day: Labor Management and Social Control in the Ford Motor Company, 1908–1921 by Stephen Meyer III (review). Technology and Culture, 23(4):679–680, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890269/pdf.

Schlereth:1982:MCW

[Sch82b] Thomas J. Schlereth. Material Culture of the Wooden Age ed. by Brooke Hindle (review). Technology and Culture, 23(4):666–668, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890262/pdf.

Schmaus:1982:FSE

[Sch82c] Warren Schmaus. Fundamenta Scientiae ed. by Baudoin Jurdant, Michel Paty (review). Technology and Culture, 23(1):98–100, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891115/pdf.

Schofield:1982:SCH

[Sch82d] Robert E. Schofield. Studies on Christiaan Huygens: Invited Papers from the Symposium on the Life and Work of Christiaan Huygens, Amsterdam, 22–25 August 1979 ed. by H. J. M. Bos, et al. (review). Technology and Culture, 23(1):100–102, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891116/
pdf.

Schindler:1984:WOP

[Sch84] Hans Schindler. White oil: Petrochemical development in Russia and the West. Technology and Culture, 25(3):577–588, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889993/
pdf.

Schaffer:1985:LLP

[Sch85a] Daniel Schaffer. Lumberjacks and Legislators: Political Economy of the U.S. Lumber Industry, 1890–1941 by William G. Robbins (review). Technology and Culture, 26(1):98–99, January 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889821/pdf.

Schneider:1985:CDN

[Sch85b] Dorothee Schneider. Chants Democratic: New York City and the Rise of the American Working Class, 1788–1850 by Sean Wilentz (review). Technology and Culture, 26(4):855–857, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889696/
pdf.

Schmiechen:1986:UWE

[Sch86a] James A. Schmiechen. Urban Workers in the Early Industrial Revolution by Robert Glen (review). Technology and Culture, 27(4):880–882, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889536/pdf.

Schroeder:1986:MST

[Sch86b] Fred E. H. Schroeder. More “Small things forgotten”: Domestic electrical plugs and receptacles, 1881–1931. Technology and Culture, 27(3):525–543, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889529/pdf.

Schorzman:1989:SVP

[Sch89] Terri A. Schorzman. Smithsonian Videohistory Program Symposium — Washington, D.C., May 5, 1988. Technology and Culture, 30(1):118–121, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889117/pdf.

Sclove:1988:IAD

[Scl88] Richard E. Sclove. Insuring against Disaster: The Nuclear Industry on Trial by John W. Johnson (review). Technology and Culture, 29(1):186–187, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/888538/pdf.

Scranton:1986:LME

[Scr86a] Philip Scranton. Learning manufacture: Education and shop-floor schooling in the family firm. Technology and Culture, 27(1):40–62, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889634/pdf.

Scranton:1986:TBM

[Scr86b] Philip Scranton. The Third Biennial Meeting of the International Working Group on Textile Geography, Barcelona, June 3–7, 1985. Technology and Culture, 27(3):604–608, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889549/pdf.

Scranton:1987:FFM

[Scr87a] Philip Scranton. Family Firm to Modern Multinational: Norton Company, a New England Enterprise by Charles W. Cheape (review). Technology and Culture, 28(4):871–872, October 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889323/pdf.

Scranton:1987:TAF

[Scr87b] Philip B. Scranton. Textiles in America 1650–1870 by Florence M. Montgomery (review). Technology and Culture, 28(2):363–364, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889419/pdf.

Scranton:1988:LT

[Scr88a] Philip Scranton. Labor and technology. Technology and Culture, 29(4):719–721, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889155/pdf.

Scranton:1988:NTP

[Scr88b] Philip Scranton. None-too-porous boundaries: Labor history and the history of technology. Technology and Culture, 29(4):722–743, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889156/pdf.

Seely:1984:SME

[See84] Bruce E. Seely. The scientific mystique in engineering: Highway research at the Bureau of Public Roads, 1918–1940. Technology and Culture, 25(4):798–831, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889947/pdf.

Seely:1987:THA

[See87] Bruce E. Seely. The Technology of Historic American Buildings: Studies of the Materials, Craft Processes, and the Mechanization of Building Construction ed. by H. Ward Jandl (review). Technology and Culture, 28(4):882–883, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889330/pdf.

Segal:1980:PEF

[Seg80] Howard P. Segal. The Pattern of Expectation, 1644–2001 by I. F. Clarke (review). Technology and Culture, 21(2):271–272, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890696/pdf.

Segal:1981:UTW

[Seg81] Howard P. Segal. Utopian Thought in the Western World by Frank E. Manuel, Fritzie P. Manuel (review). Technology and Culture, 22(4):784–786, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890947/pdf.

Segal:1983:EEP

[Seg83] Howard P. Segal. Engineering Education at Penn State: A Century in the Land Grant Tradition by Michael Bezilla (review). Technology and Culture, 24(2):293–294, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891281/pdf.

Segal:1984:TVL

[Seg84] Howard P. Segal. Terminal Visions: The Literature of Last Things by W. Warren Wagar (review). Technology and Culture, 25(4): 905–906, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889986/pdf.

Segal:1986:AES

[Seg86] Howard P. Segal. The Authority of Experts: Studies in History and Theory ed. by Thomas L. Haskell, and: TVA: Fifty Years of Grass-Roots Bureaucracy ed. by Erwin C. Hargrove, Paul K. Conkin (review). Technology and Culture, 27(3):640–642, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889584/pdf.

Segal:1987:IPW

[Seg87] Howard P. Segal. Inventions: The Patented Works of R. Buckminster Fuller by R. Buckminster Fuller, and: The Artifacts of Buckminster Fuller: A Comprehensive Collection of His Designs and Drawings. Vol. 1: The Dymaxion Experiment, 1926–1943; vol. 2: Dymaxion Deployment, 1927–1946; vol. 3: The Geodesic Revolution, Part 1, 1947–1959; vol. 4: The Geodesic Revolution, Part 2, 1960–1983 ed. by James Ward (review). Technology and Culture, 28(3):697–698, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889368/pdf.

Segal:1989:GPB

[Seg89a] Howard P. Segal. Grand Plans: Business Progressivism and Social Change in Ohio’s Miami Valley, 1890–1929 by Judith Sealander (review). Technology and Culture, 30(4):1064–1066, October 1989.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/888995/
pdf.

Segal:1989:KTP

[Seg89b] Howard P. Segal. Kingsport, Tennessee: A Planned American City by Margaret R. Wolfe (review). Technology and Culture, 30(1):143–144, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889128/pdf.

Seidler:1980:WPA

[Sei80] Jan Seidler. Winterthur Portfolio 13: American Furniture and Its Makers ed. by Ian M. G. Quimby (review). Technology and Culture, 21(4):669–673, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890772/pdf.

Seligmann:1988:IBR

[Sel88] Linda J. Seligmann. Irrigation in the Bajío Region of Colonial Mexico by Michael E. Murphy (review). Technology and Culture, 29(3):679–682, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889239/pdf.

Semmens:1984:CLW

[Sem84] P. W. B. Semmens. Crewe Locomotive Works and Its Men by Brian Reed (review). Technology and Culture, 25(3):655–656, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890007/
pdf.

Servos:1983:EHT

[Ser83] John W. Servos. Electrochemistry: History and Theory by Wilhelm Ostwald (review). Technology and Culture, 24(1):125–127, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891235/pdf.

Shrader-Frechette:1982:ABE

[SF82] K. S. Shrader-Frechette. The Atom Besieged: Extraparliamentary Dissent in France and Germany by Dorothy Nelkin, Michael Pollak (review). Technology and Culture, 23(4):662–664, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890260/pdf.

Shrader-Frechette:1987:TTC

[SF87] Kristin Shrader-Frechette. A Theory of Technology: Continuity and Change in Human Development by Thomas R. De Gregori (review). Technology and Culture, 28(4):897–899, October 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889342/pdf.

Sharlin:1981:TIC

[Sha81a] Harold Issadore Sharlin. Technological Innovation: A Critical Review of Current Knowledge ed. by Patrick Kelly, Melvin Kranzberg (review). Technology and Culture, 22(2):349–352, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890872/pdf.

Sharpe:1981:PHA

[Sha81b] Mitchell R. Sharpe. The Partnership, a History of the Apollo-Soyuz Test Project by Edward C. Ezell, Linda N. Ezell (review). Technology and Culture, 22(2):346–347, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890870/pdf.

Sharpe:1981:TFY

[Sha81c] Mitchell R. Sharpe. Twenty-Five Years of the American Astronautical Society: Historical Reflections and Projections, 1954–1979 ed. by Eugene M. Emme (review). Technology and Culture, 22(3):639– 641, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890912/pdf.

Sharrer:1983:AFN

[Sha83a] C. Terry Sharrer. American Farmers: The New Minority by Gilbert C. Fite (review). Technology and Culture, 24(2):277–278, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891272/pdf.

Shayt:1983:HIE

[Sha83b] David H. Shayt. The Hoghead: An Industrial Ethnology of the Locomotive Engineer by Frederick C. Gamst (review). Technology and Culture, 24(3):522–524, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891409/pdf.

Shapiro:1984:EPJ

[Sha84] Henry D. Shapiro. An Education in Psychology: James McKeen Cattell’s Journals and Letters from Germany and England, 1880–1888 ed. by Michael M. Sokal (review). Technology and Culture, 25(2):374–376, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890076/pdf.

Sharpe:1985:PZD

[Sha85] Mitchell R. Sharpe. Precursors to Zalinski’s dynamite gun. Technology and Culture, 26(1):82–83, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889812/pdf.

Shallat:1988:HHR

[Sha88] Todd Shallat. Hydraulics and Hydraulic Research: A Historical Review ed. by Günther Garbrecht (review). Technology and Culture, 29(4):950–952, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889187/pdf.

Shapiro:1989:AFS

[Sha89a] Henry D. Shapiro. Arguments in Favor of Sharpshooting by David Clarke (review). Technology and Culture, 30(4):1092–1094, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889021/pdf.

Shayt:1989:SRA

[Sha89b] David H. Shayt. Stairway to redemption: America’s encounter with the British prison treadmill. Technology and Culture, 30(4): 908–938, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888969/pdf.

Shelby:1980:GWM

[She80a] Lon Shelby. Gold Was the Mortar: The Economics of Cathedral Building by Henry Kraus (review). Technology and Culture, 21 (1):85–87, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890610/pdf.

Sherwood:1980:RME

[She80b] Morgan Sherwood. Reconciling Man with the Environment by Eric Ashby (review). Technology and Culture, 21(1):149–151, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890645/pdf.

Shelby:1981:PSG

[She81a] Lon Shelby. The Plan of St. Gall: A Study of the Architecture and Economy of, and Life in a Paradigmatic Carolingian Monastery by Walter Horn, Ernest Born (review). Technology and Culture, 22(1):157–160, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890804/pdf.

Shelby:1981:ALB

[She81b] Lon R. Shelby. Architector: The Lodge Books and Sketchbooks of Medieval Architects ed. by François Bucher (review). Technology and Culture, 22(4):786–788, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890948/pdf.

Shearmur:1984:PTE

[She84] Jeremy Shearmur. In Pursuit of Truth: Essays in Honour of Karl Popper’s 80th Birthday ed. by Paul Levinson (review). Technology and Culture, 25(3):694–696, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890029/pdf.

Sherwood:1987:EHC

[She87] Morgan Sherwood. Environmental History: Critical Issues in Comparative Perspective ed. by Kendall E. Bailes (review). Technology and Culture, 28(1):139–141, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889463/pdf.

Shopes:1988:WWC

[Sho88] Linda Shopes. Women’s Work and Chicano Families: Cannery Workers of the Santa Clara Valley by Patricia Zavella, and: Cannery Women, Cannery Lives: Mexican Women, Unionization, and the California Food Processing Industry, 1930–1950 by Vicki L. Ruiz (review). Technology and Culture, 29(4):925–927, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889173/pdf.

Shuchman:1984:ISE

[Shu84a] Hedvah L. Shuchman. Information Services: Economics, Management, and Technology ed. by Robert M. Mason, John E. Creps, Jr. (review). Technology and Culture, 25(1):146–148, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890104/
pdf.

Shulman:1984:TFF

[Shu84b] Holly C. Shulman. Technologies of Freedom: On Free Speech in an Electronic Age by Ithiel de Sola Pool (review). Technology and Culture, 25(3):683–685, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890023/pdf.

Siegel:1981:PHS

[Sie81] Daniel M. Siegel. The Physicists: The History of a Scientific Community in Modern America by Daniel J. Kevles (review). Technology and Culture, 22(2):329–332, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890864/pdf.

Simms:1983:WDS

[Sim83] D. L. Simms. Water-driven saws, Ausonius, and the authenticity of the Mosella. Technology and Culture, 24(4):635–643, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890122/pdf.

Simms:1985:WDS

[Sim85] D. L. Simms. Water-driven saws in late antiquity. Technology and Culture, 26(2):275–276, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889768/pdf.

Simms:1986:NAL

[Sim86] D. L. Simms. Normal Accidents: Living with High-Risk Technologies by Charles Perrow (review). Technology and Culture, 27(4):903–905, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889566/pdf.

Simms:1987:AIA

[Sim87] D. L. Simms. Archimedes and the invention of artillery and gunpowder. Technology and Culture, 28(1):67–79, January 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889441/pdf.

Simms:1989:PA

[Sim89] D. L. Simms. “A problem for Archimedes”. Technology and Culture, 30(1):177–178, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889150/pdf.

Sinclair:1982:JWA

[Sin82] Bruce Sinclair. John W. Abrams (1913–1981). Technology and Culture, 23(3):527–530, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891219/pdf.

Sinclair:1986:MEN

[Sin86a] Bruce Sinclair. Mechanical Engineering at the National Research Council of Canada, 1929–1951 by W. E. Knowles Middleton (review). Technology and Culture, 27(2):314–315, April 1986. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889616/pdf.

Sinclair:1986:LHN

[Sin86b] Bruce Sinclair. Local history and national culture: Notions on engineering professionalism in America. Technology and Culture, 27(4):683–693, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889495/pdf.

Singer:1987:IWC

[Sin87] Bayla Singer. Image Worlds: Corporate Identities at General Electric, 1890–1930 by David E. Nye (review). Technology and Culture, 28(4):876–877, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889326/pdf.

Sinclair:1988:TWO

[Sin88a] Bruce Sinclair. The Tancook Whalers: Origins, Rediscovery, and Revival by Robert C. Post (review). Technology and Culture, 29 (2):307–308, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889281/pdf.

Singer:1988:DMC

[Sin88b] Bayla Singer. “The decorated machine” at the Chicago Museum of Science and Industry. Technology and Culture, 29(3):619–621, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889214/pdf.

Singer:1988:TNI

[Sin88c] S. Fred Singer. Toward a New Iron Age? Quantitative Modeling of Resource Exhaustion by Robert B. Gordon, et al. (review). Technology and Culture, 29(4):974–975, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889203/pdf.

Sinclair:1989:AS

[Sin89] Bruce Sinclair. An agenda for SHOT. Technology and Culture, 30(3):596–600, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889014/pdf.

Sivin:1980:RCI

[Siv80] N. Sivin. The Roanoke Conference, II. Concluding remarks on conference. Technology and Culture, 21(4):621–632, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890754/
pdf.

Skaggs:1988:CPS

[Ska88] Jimmy M. Skaggs. Chicago’s Pride: The Stockyards, Packingtown, and Environs in the Nineteenth Century by Louise Carroll Wade (review). Technology and Culture, 29(2):313–314, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889285/pdf.

Stock:1980:JDI

[SL80] John T. Stock and Philip S. Laurie. John Dover, instrument maker (1824–1881). Technology and Culture, 21(1):51–55, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890604/pdf.

Staudenmaier:1989:AH

[SL89] John Staudenmaier, S.J. and Pamela Walker Lurito Laird. Advertising history. Technology and Culture, 30(4):1031–1036, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888977/pdf.

Slade:1982:TPP

[Sla82] Joseph W. Slade. Thomas Pynchon, postindustrial humanist. Technology and Culture, 23(1):53–72, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891104/pdf.

Slade:1984:SST

[Sla84] Joseph W. Slade. Signs and Symptoms: Thomas Pynchon and the Contemporary World by Peter L. Cooper (review). Technology and Culture, 25(4):902–903, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889984/pdf.

Slade:1988:HAB

[Sla88] Joseph W. Slade. Hopes and Ashes: The Birth of Modern Times 1929–1939 by Alice G. Marquis (review). Technology and Culture, 29(1):171–172, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/888528/pdf.

Slade:1989:ITH

[Sla89a] Joseph W. Slade. Imagining Tomorrow: History, Technology, and the American Future ed. by Joseph J. Corn (review). Technology and Culture, 30(2):459–460, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889078/pdf.

Slade:1989:MAS

[Sla89b] Joseph W. Slade. Made in America: Science, Technology, and American Modernist Poets by Lisa M. Steinman (review). Technology and Culture, 30(1):130–131, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889120/pdf.

Slade:1989:PPE

[Sla89c] Joseph W. Slade. The Pilot and the Passenger: Essays on Literature, Technology, and Culture in the United States by Leo Marx (review). Technology and Culture, 30(4):1050–1052, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/888986/
pdf.

Sleeswyk:1981:MGN

[Sle81] André W. Sleeswyk. Les Mechaniciens grecs: La naissance de la technologie by Bertrand Gille (review). Technology and Culture, 22(4):776–778, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890943/pdf.

Sloat:1981:WPE

[Slo81] Caroline Sloat. Winterthur Portfolio 12 ed. by Ian M. G. Quimby (review). Technology and Culture, 22(1):169–172, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890810/
pdf.

Smallwood:1980:REM

[Sma80] J. B. Smallwood. River Engineers on the Middle Mississippi: A History of the St. Louis District, U.S. Army Corps of Engineers by Frederick J. Dobney, and: Cradle of the Corps: A History of the New York District, U.S. Army Corps of Engineers, 1775– 1975 by Marion J. Klawonin (review). Technology and Culture, 21(1):120–122, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890629/pdf.

Smith:1980:WJK

[Smi80a] Cyril Stanley Smith. The World of the Javanese Keris by Garrett Solyom, Bronwen Solyom (review). Technology and Culture, 21(2):233–235, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890678/pdf.

Smith:1980:PAM

[Smi80b] Cecil O. Smith, Jr. Paths to Authority: The Middle Class and the Industrial Labor Force in France, 1820–48 by Peter N. Stearns (review). Technology and Culture, 21(1):104–106, January 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890619/pdf.

Smith:1981:JAF

[Smi81a] Richard K. Smith. The Jet Age: Forty Years of Jet Aviation ed. by Walter J. Boyne, Donald S. Lopez, and: Wings of Gold: How the Aeroplane Developed New Guinea by James Sinclair (review). Technology and Culture, 22(3):641–643, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890913/pdf.

Smith:1981:HTT

[Smi81b] Cecil O. Smith, Jr. Histoire des techniques: Technique et civilisations, technique et sciences by Bertrand Gille (review). Technology and Culture, 22(4):771–776, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890942/pdf.

Smith:1982:TIT

[Smi82a] Carl S. Smith. The Technological Imagination: Theories and Fictions ed. by Teresa de Lauretis, et al. (review). Technology and Culture, 23(1):91–93, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891111/pdf.

Smith:1982:OST

[Smi82b] Cecil O. Smith, Jr. The Organization of Science and Technology in France, 1808–1914 ed. by Robert Fox, George Weisz (review). Technology and Culture, 23(2):235–238, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891155/pdf.

Smith:1983:BHW

[Smi83a] David C. Smith. Bowater: A History by W. J. Reader (review). Technology and Culture, 24(4):697–698, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890141/pdf.

Smith:1983:TVA

[Smi83b] Merritt Roe Smith. Technology and Values in American Civilization: A Guide to Information Sources ed. by Stephen H. Cutcliffe, et al,, and: Technical Americana: A Checklist of Technical Publications Printed before 1831 by Evald Rink (review). Technology and Culture, 24(1):119–120, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891231/pdf.

Smith:1983:LWH

[Smi83c] Richard K. Smith. Legacy of Wings: The Harold F. Pitcairn Story by Frank Kingston Smith (review). Technology and Culture, 24(4):705–706, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890146/pdf.

Smith:1983:IAE

[Smi83d] Richard K. Smith. The intercontinental airliner and the essence of airplane performance, 1929–1939. Technology and Culture, 24 (3):428–449, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891386/pdf.

Smith:1983:ECT

[Smi83e] Cecil O. Smith, Jr. Ethnotechnologie, Culture Technique, special issue no. 4, February 1981, and: Design, Culture Technique, special issue no. 5, April 1981 (review). Technology and Culture, 24(3):504–506, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891397/pdf.

Smith:1984:TAW

[Smi84a] Cyril Stanley Smith. Theodore Allen Wertime (1919–1982). Technology and Culture, 25(4):907–910, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889987/pdf.

Smith:1984:HWM

[Smi84b] Merritt Roe Smith. Handguns of the World: Military Revolvers and Self-Loaders from 1870 to 1945 by Edward C. Ezell (review). Technology and Culture, 25(4):887–889, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889976/pdf.

Smith:1985:SET

[Smi85a] David C. Smith. Sawdust Empire: The Texas Lumber Industry, 1830–1940 by Robert S. Maxwell, Robert D. Baker (review). Technology and Culture, 26(4):863–865, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889701/pdf.

Smith:1985:CBV

[Smi85b] Denis Smith. Charles Blacker Vignoles: Romantic Engineer by K. H. Vignoles (review). Technology and Culture, 26(3):642–643, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889733/pdf.

Smith:1985:TYI

[Smi85c] John K. Smith. The ten-year invention: Neoprene and du pont research, 1930–1939. Technology and Culture, 26(1):34–55, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889807/pdf.

Smith:1985:FPT

[Smi85d] Richard K. Smith. Flight Patterns: Trends of Aeronautical Development in the United States, 1918–1929 by Roger E. Bilstein (review). Technology and Culture, 26(4):870–871, October 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889705/pdf.

Smith:1985:GBC

[Smi85e] Richard K. Smith. Government and British Civil Aerospace: A Case Study in Post-War Technology Policy by Keith Hayward (review). Technology and Culture, 26(1):87–89, January 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889815/pdf.

Smith:1985:HGT

[Smi85f] Cecil O. Smith, Jr. Histoire générale des techniques. Vols. 4 and 5: Les Techniques de la civilisation industrielle ed. by Maurice Daumas (review). Technology and Culture, 26(3):689–697, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889759/
pdf.

Smith:1986:PTE

[Smi86a] J. Merrill Smith. Philosophy and Technology ed. by Paul T. Durbin, Friedrich Rapp (review). Technology and Culture, 27 (2):338–340, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889629/pdf.

Smith:1986:SSE

[Smi86b] John K. Smith. “Scaling up: Science, engineering, and the American chemical industry” — center for history of chemistry. Technology and Culture, 27(4):832–834, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889502/pdf.

Smith:1986:FMH

[Smi86c] Michael S. Smith. De la Forge à la manufacture horlogère (XVIII^{e}– XX^{e} siècles) by François Jequier (review). Technology and Culture, 27(1):143–145, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889651/pdf.

Smith:1986:CPC

[Smi86d] Cecil O. Smith, Jr. Le Corps des Ponts et Chaussées: Histoire de l’administration française by A. Brunot, R. Coquand (review). Technology and Culture, 27(1):145–146, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889652/pdf.

Smith:1987:WRW

[Smi87a] John K. Smith. Willis R. Whitney, General Electric, and the Origins of U.S. Industrial Research by George Wise (review). Technology and Culture, 28(1):167–169, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889479/pdf.

Smith:1987:RPC

[Smi87b] John K. Smith. R&D Pioneers Conference — Hagley Museum and Library, Wilmington, Delaware, October 7, 1985. Technology and Culture, 28(2):340–342, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889405/pdf.

Smith:1988:HTT

[Smi88a] Denis Smith. History of Technology. Tenth Annual Volume, 1985; Eleventh Annual Volume, 1986 ed. by Norman Smith (review). Technology and Culture, 29(4):965–966, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889196/pdf.

Smith:1988:MSC

[Smi88b] Duane A. Smith. Mining the Summit: Colorado’s Ten Mile District, 1860–1960 by Stanley Dempsey, James E. Fell, Jr. (review). Technology and Culture, 29(2):304–305, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889279/pdf.

Smith:1989:WNM

[Smi89a] Karen L. Smith. Water in New Mexico: A History of Its Management and Use by Ira G. Clark (review). Technology and Culture, 30(2):491–493, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889103/pdf.

Smith:1989:PRI

[Smi89b] John Kenly Smith, Jr. Petrochemicals: The Rise of an Industry by Peter H. Spitz (review). Technology and Culture, 30(3):710–711, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889075/pdf.

Smulyan:1985:CSB

[Smu85] Susan Smulyan. Communications and Society: A Bibliography on Communications Technologies and Their Social Impact comp. by Benjamin F. Shearer, Marilyn Huxford (review). Technology and Culture, 26(3):664–665, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889747/pdf.

Smulyan:1987:MMB

[Smu87] Susan Smulyan. Mass Media between the Wars: Perceptions of Cultural Tension, 1918–1941 ed. by Catherine L. Covert, John D. Stevens (review). Technology and Culture, 28(3):702–704, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889372/pdf.

Socolofsky:1983:CCG

[Soc83] Homer E. Socolofsky. Custom Combining on the Great Plains: A History by Thomas D. Isern (review). Technology and Culture, 24(2):278–279, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891273/pdf.

Socolofsky:1989:FFP

[Soc89] Homer E. Socolofsky. Fordson, Farmall, and Poppin’ Johnny: A History of the Farm Tractor and Its Impact on America by Robert C. Williams (review). Technology and Culture, 30(1):146–147, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889130/pdf.

Sokal:1981:HCE

[Sok81] Michael M. Sokal. The Human Condition: An Ecological and Historical View by William H. McNeill (review). Technology and Culture, 22(4):782–783, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890946/pdf.

Spier:1981:MPP

[Spi81] Robert F. G. Spier. Mesolithic Prelude: The Palaeolithic-Neolithic Transition in Old World Prehistory by Grahame Clark (review). Technology and Culture, 22(4):779–780, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890944/pdf.

Spier:1982:ATF

[Spi82a] Robert F. G. Spier. Agricultural Tools and Field Shapes: Twentyfive Years of Activity by the International Secretariat by Grith Lerche, Axel Steensberg (review). Technology and Culture, 23 (2):244–245, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891159/pdf.

Spier:1982:MCS

[Spi82b] Robert F. G. Spier. Material Culture: Styles, Organization, and Dynamics of Technology ed. by Heather Lechtman, Robert S. Merrill (review). Technology and Culture, 23(1):110–112, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891122/pdf.

Sprigg:1981:STA

[Spr81] June Sprigg. Shaker Textile Arts by Beverly Gordon (review). Technology and Culture, 22(1):172–174, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890811/pdf.

Spude:1989:OML

[Spu89] Robert L. Spude. One Man’s Locomotives: 50 Years Experience with Railway Motive Power by Vernon L. Smith (review). Technology and Culture, 30(4):1058–1059, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888991/pdf.

Sheets-Pyenson:1987:TCM

[SPV87] Susan Sheets-Pyenson and Mary Vipond. Technology and the Canadian Mind: Innis/McLuhan/Grant by Arthur Kroker (review). Technology and Culture, 28(2):359–360, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889416/pdf.

Savage-Smith:1984:AWC

[SS84] Emilie Savage-Smith. Arabic Water-Clocks by Donald R. Hill (review). Technology and Culture, 25(2):326–328, April 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890053/pdf.

Starron:1981:TTI

[Sta81] John E. Starron, Jr. Technology Transfer in Industrialized Countries ed. by Sherman Gee (review). Technology and Culture, 22 (1):224–225, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890839/pdf.

Stapleton:1982:APS

[Sta82a] Darwin Stapleton. The American Philosophical Society Library: Holdings in Early American technology. Technology and Culture, 23(3):430–434, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891183/pdf.

Stapleton:1982:JBY

[Sta82b] Darwin H. Stapleton. The John Bull: 150 Years a Locomotive by John H. White, Jr. (review). Technology and Culture, 23(4): 639–640, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890246/pdf.

Stapleton:1983:EID

[Sta83a] Darwin H. Stapleton. Engineers of Independence: A Documentary History of the Army Engineers in the American Revolution, 1775–1783 ed. by Paul K. Walker (review). Technology and Culture, 24(2):269–270, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891267/pdf.

Staudenmaier:1983:BDP

[Sta83b] John M. Staudenmaier, S.J. Beyond Despair: A Policy Guide to the Communications Era by Robert Theobald (review). Technology and Culture, 24(2):313–314, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891292/pdf.

Staudenmaier:1983:ALR

[Sta83c] John M. Staudenmaier, S.J. The Assembly Line by Robert Linhart (review). Technology and Culture, 24(3):502–503, July 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891396/pdf.

Staudenmaier:1984:WSH

[Sta84] John M. Staudenmaier, S.J. What SHOT hath wrought and what SHOT hath not: Reflections on twenty-five years of the history of technology. Technology and Culture, 25(4):707–730, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889942/
pdf.

Staudenmaier:1985:FOL

[Sta85a] John M. Staudenmaier, S.J. A Feeling for the Organism: The Life and Work of Barbara McClintock by Evelyn Fox Keller (review). Technology and Culture, 26(3):681–682, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889756/pdf.

StaudenmaierSJ:1985:MED

[Sta85b] John M. Staudenmaier S.J. Machina ex Dea: Feminist Perspectives on Technology ed. by Joan Rothschild (review). Technology and Culture, 26(2):283–287, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889772/pdf.

Stabile:1986:HHF

[Sta86] Donald R. Stabile. Herbert Hoover, the FAES, and the AF of L. Technology and Culture, 27(4):819–827, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889500/pdf.

Stapleton:1987:TTI

[Sta87] Darwin H. Stapleton. Technology on Trial: The Introduction of Steam Power Technology into Sweden, 1715–1736 by Svante Lindqvist (review). Technology and Culture, 28(1):146–147, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889466/pdf.

Stapleton:1988:MAM

[Sta88a] Darwin H. Stapleton. Mechanical Arts & Merchandise: Industrial Espionage and Travellers’ Accounts as a Source for Technical Historians by A. P. Woolrich, and: Ferrner’s Journal 1759/1760: An Industrial Spy in Bath and Bristol ed. by A. P. Woolrich (review). Technology and Culture, 29(1):146–148, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888513/pdf.

Star:1988:COR

[Sta88b] Susan Leigh Star. Changing Order: Replication and Induction in Scientific Practice by H. M. Collins (review). Technology and Culture, 29(4):983–985, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889215/pdf.

Staudenmaier:1988:GWR

[Sta88c] John M. Staudenmaier, S.J. The giant wakens: Revising Henry Ford’s history book. Technology and Culture, 29(1):118–124, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888501/pdf.

Stamper:1989:GMP

[Sta89] John W. Stamper. The Galerie des Machines of the 1889 Paris World’s Fair. Technology and Culture, 30(2):330–353, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889065/
pdf.

Stern:1980:RST

[Ste80] Nancy Stern. The Rise of Systems Theory: An Ideological Analysis by Robert Lilienfeld (review). Technology and Culture, 21(3):538– 540, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890741/pdf.

Steiner:1981:BIN

[Ste81a] Frances H. Steiner. Building with iron: a Napoleonic controversy. Technology and Culture, 22(4):700–724, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890936/pdf.

Stern:1981:PWH

[Ste81b] Nancy Stern. Project Whirlwind: The History of a Pioneer Computer by Kent C. Redmond, Thomas M. Smith, and: Early British Computers by Simon H. Lavington (review). Technology and Culture, 22(3):644–647, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890915/pdf.

Stern:1982:JNN

[Ste82a] Nancy Stern. John von Neumann and Norbert Wiener: From Mathematics to the Technologies of Life and Death by Steve J. Heims (review). Technology and Culture, 23(1):108–110, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891121/pdf.

Stern:1982:EMC

[Ste82b] Nancy Stern. The Eckert–Mauchly computers: Conceptual triumphs, commercial tribulations. Technology and Culture, 23(4): 569–582, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890240/pdf.

Stephens:1983:TMV

[Ste83] Carlene Stephens. Talking Machines by V. K. Chew (review). Technology and Culture, 24(4):721–722, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890157/pdf.

Stephens:1985:WEL

[Ste85] Carlene Stephens. Watchmaking in England, 1760–1820 by Leonard Weiss (review). Technology and Culture, 26(1):90–92, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889817/pdf.

Stephens:1987:TGI

[Ste87] Carlene Stephens. “Time: The Greatest Innovator,” an Exhibition at the Folger Shakespeare Library, Washington, D.C. Technology and Culture, 28(4):833–838, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889305/pdf.

Stearns:1988:BMD

[Ste88a] Peter N. Stearns. Brothers: Male Dominance and Technological Change by Cynthia Cockburn (review). Technology and Culture, 29(4):937–938, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889180/pdf.

Stech:1988:TAM

[Ste88b] Tamara Stech. Tin in Antiquity: Its Mining and Trade throughout the Ancient World with Particular Reference to Cornwall by Roger David Penhallurick (review). Technology and Culture, 29(3):671– 673, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889235/pdf.

Steinberg:1988:ERC

[Ste88c] Theodore Steinberg. Endangered Rivers and the Conservation Movement by Tim Palmer (review). Technology and Culture, 29 (4):973–974, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889201/pdf.

Stephens:1989:MRT

[Ste89] Carlene Stephens. “The Most Reliable Time”: William Bond, the New England railroads, and time awareness in 19th-Century America. Technology and Culture, 30(1):1–24, January 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889109/pdf.

Stine:1981:CAW

[Sti81] Jeffrey K. Stine. Cities of the American West: A History of Frontier Urban Planning by John W. Reps (review). Technology and Culture, 22(1):178–180, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890815/pdf.

Stine:1982:TET

[Sti82] Jeffrey K. Stine. The Tools of Empire: Technology and European Imperialism in the Nineteenth Century by Daniel R. Headrick (review). Technology and Culture, 23(3):481–483, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891195/pdf.

Stine:1984:EPG

[Sti84] Jeffrey K. Stine. Engineers for the Public Good: A History of the Buffalo District U.S. Army Corps of Engineers by Nuala Mc-Cann Drescher (review). Technology and Culture, 25(1):161–162, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890114/pdf.

Stine:1985:CBN

[Sti85a] Jeffrey K. Stine. City Building in the New South: The Growth of Public Services in Houston, Texas, 1830–1915 by Harold L. Platt (review). Technology and Culture, 26(1):94–95, January 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889819/pdf.

Stine:1985:LHH

[Sti85b] Jeffrey K. Stine. The Life of Herbert Hoover: The Engineer, 1874–1914 by George H. Nash (review). Technology and Culture, 26(3):653–654, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889740/pdf.

Stine:1986:MLC

[Sti86a] Jeffrey K. Stine. Morris Llewellyn Cooke: Progressive Engineer by Jean Christie (review). Technology and Culture, 27(1):160–161, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889659/pdf.

Stine:1986:POR

[Sti86b] Jeffrey K. Stine. Prophets of Order: The Rise of the New Class, Technocracy and Socialism in America by Donald Stabile (review). Technology and Culture, 27(4):838–840, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889505/pdf.

Stine:1987:CAE

[Sti87] Jeffrey K. Stine. Coping with Abundance: Energy and Environment in Industrial America by Martin V. Melosi (review). Technology and Culture, 28(2):374, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889426/pdf.

Stover:1988:SPD

[Sto88] John F. Stover. The Southern Pacific, 1901–1985 by Don L. Hofsommer (review). Technology and Culture, 29(2):310–312, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889283/pdf.

Stone:1989:SEA

[Sto89] Bruce Stone. The Stubborn Earth: American Agriculturalists on Chinese Soil, 1898–1937 by Randall E. Stross (review). Technology and Culture, 30(1):167–169, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889144/pdf.

Strayer:1980:MRT

[Str80] Joseph R. Strayer. Medieval Religion and Technology — Collected Essays by Lynn White, jr (review). Technology and Culture, 21 (1):82–85, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890609/pdf.

Stranges:1984:CQF

[Str84] Anthony N. Stranges. Coal in Queensland: The First Fifty Years by R. L. Whitmore, and: Conflict and Coal: A Case Study of Industrial Relations in the Open-cut Mining Industry of Central Queensland by Kevin Hince (review). Technology and Culture, 25(4):857–859, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889957/pdf.

Stranges:1985:BBH

[Str85a] Anthony N. Stranges. From Birmingham to Billingham: Highpressure coal hydrogenation in Great Britain. Technology and Culture, 26(4):726–757, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889675/pdf.

Strum:1985:AAS

[Str85b] Harvey Strum. The Association for Applied Solar Energy/Solar Energy Society, 1954–1970. Technology and Culture, 26(3):571– 578, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889718/pdf.

Stranges:1988:HBC

[Str88] Anthony N. Stranges. The History of the British Coal industry. Vol. 5: 1946–1982: The Nationalized Industry by William Ashworth, Mark Pegg (review). Technology and Culture, 29(4): 975–977, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889205/pdf.

Swenson:1981:RRR

[Swe81] Loyd S. Swenson, Jr. Rivers, Rockets and Readiness: Army Engineers in the Sunbelt by D. Clayton Brown (review). Technology and Culture, 22(1):187–188, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890820/pdf.

Swenson:1982:HSL

[Swe82] Loyd S. Swenson, Jr. Handbook of Soviet Lunar and Planetary Exploration; and Handbook of Soviet Manned Space Flight by Nicholas L. Johnson (review). Technology and Culture, 23(3):520– 522, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891216/pdf.

Swenson:1988:MMA

[Swe88a] Loyd S. Swenson, Jr. From Maxwell to Microphysics: Aspects of Electromagnetic Theory in the Last Quarter of the Nineteenth Century by Jed Z. Buchwald (review). Technology and Culture, 29(3):686–688, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889243/pdf.

Swenson:1988:CIM

[Swe88b] Loyd S. Swenson, Jr. The Cosmic Inquirers: Modern Telescopes and Their Makers by Wallace Tucker, Karen Tucker (review). Technology and Culture, 29(4):977–978, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889207/pdf.

Swords:1989:RWW

[Swo89] Sean S. Swords. Radar in World War II by Henry E. Guerlac (review). Technology and Culture, 30(3):707–708, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889070/pdf.

Tann:1981:ECI

[Tan81] Jennifer Tann. The English Cotton Industry and the World Market, 1815–1896 by D. A. Farnie (review). Technology and Culture, 22(3):620–621, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890901/pdf.

Tann:1982:STB

[Tan82a] Jennifer Tann. By the Sweat of Their Brow: Women Workers at Victorian Coal Mines by Angela V. John (review). Technology and Culture, 23(2):263–264, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891168/pdf.

Tann:1982:THE

[Tan82b] Jennifer Tann. Textile History ed. by K. G. Ponting, S. D. Chapman (review). Technology and Culture, 23(2):267–269, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891171/
pdf.

Tann:1989:DLS

[Tan89a] Jennifer Tann. Divisions of Labour: Skilled Workers and Technological Change in Nineteenth Century Britain ed. by Royden Harrison, Jonathan Zeitlin (review). Technology and Culture, 30(3):691– 692, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889051/pdf.

Tann:1989:JNB

[Tan89b] Jennifer Tann. James Nasmyth and the Bridgewater Foundry: A Study of Entrepreneurship in the Early Engineering Industry by J. A. Cantrell (review). Technology and Culture, 30(2):478–479, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889095/pdf.

Tarr:1981:HCE

[Tar81] Joel A. Tarr. Health in the City: Environmental and Behavioral Influences by Malcolm S. Weinstein (review). Technology and Culture, 22(3):629–630, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890907/pdf.

Tarr:1985:ICC

[Tar85a] Joel A. Tarr. International conference on the city and technology — École Nationale des Ponts et Chaussées, Paris, December 12–15, 1983. Technology and Culture, 26(1):74–76, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889810/pdf.

Tarr:1985:WWS

[Tar85b] Joel A. Tarr. “Water and wastes” — societal values and user values. Technology and Culture, 26(2):279–281, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889770/pdf.

Taylor:1981:CSC

[Tay81] John G. Taylor, Jr. Computers and Social Change by Murray Laver (review). Technology and Culture, 22(2):354–356, April 1981. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890874/pdf.

Taylor:1989:MMO

[Tay89] John Taylor. Monopoly’s Moment: The Organization and Regulation of Canadian Utilities, 1830–1930 by Christopher Armstrong, H. V. Nelles (review). Technology and Culture, 30(2):481–482, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889097/pdf.

Teich:1982:STP

[Tei82] Albert H. Teich. Science and Technology Policy: Perspectives for the 1980s ed. by Herbert I. Fusfeld, Carmela S. Haklisch (review). Technology and Culture, 23(4):692–694, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890275/pdf.

Teich:1988:SAS

[Tei88] Albert H. Teich. A Strategic Analysis of Science and Technology Policy by Harvey A. Averch, and: Science, Technology, and Policy Decisions by Anne L. Hiskes, Richard P. Hiskes, and: Tradeoffs: Imperatives of Choice in a High-Tech World by Edward Wenk, Jr. (review). Technology and Culture, 29(4):988–991, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889220/
pdf.

Thackray:1986:TSR

[Tha86] Arnold Thackray. Technology’s Storytellers: Reweaving the Human Fabric by John M. Staudenmaier, S.J. (review). Technology and Culture, 27(4):849–851, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889511/pdf.

Theerman:1983:CBD

[The83] Paul Theerman. Cambridge before Darwin: The Ideal of Liberal Education, 1800–1860 by Martha MacMackin Garland (review). Technology and Culture, 24(1):127–129, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891236/pdf.

Theerman:1984:EHE

[The84a] Paul Theerman. Engineering and Humanities ed. by James H. Schaub, et al. (review). Technology and Culture, 25(2):321–322, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890049/pdf.

Theerman:1984:JCM

[The84b] Paul Theerman. James Clerk Maxwell: A Biography by Ivan Tolstoy (review). Technology and Culture, 25(1):136–138, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890099/pdf.

Theerman:1985:RPT

[The85a] Paul Theerman. Research in Philosophy & Technology ed. by Paul T. Durbin (review). Technology and Culture, 26(1):139–141, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889842/pdf.

Theerman:1985:CTS

[The85b] Paul Theerman. The Culture of Time and Space, 1880–1918 by Stephen Kern (review). Technology and Culture, 26(4):882–884, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889710/pdf.

Theerman:1986:NPL

[The86] Paul Theerman. The National Physical Laboratory: A History by Edward Pyatt (review). Technology and Culture, 27(2):305–307, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889611/pdf.

Theerman:1987:CPS

[The87] Paul Theerman. Cities and People: A Social and Architectural History by Mark Girouard (review). Technology and Culture, 28 (4):886–888, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889333/pdf.

Thibaut:1980:SWF

[Thi80] Jacqueline Thibaut. Siege Warfare: The Fortress in the Early Modern World, 1494–1660 by Christopher Duffy (review). Technology and Culture, 21(3):485–488, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890715/pdf.

Thomas:1984:AAP

[Tho84] Selma Thomas. Artifacts and the American Past: Techniques for the Teaching Historian by Thomas J. Schlereth (review). Technology and Culture, 25(2):351–353, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890065/pdf.

Thrall:1982:CUM

[Thr82] Charles A. Thrall. The conservative use of modern household technology. Technology and Culture, 23(2):175–194, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891143/pdf.

Tindall:1987:BLT

[Tin87] George B. Tindall. Breaking the Land: The Transformation of Cotton, Tobacco, and Rice Cultures since 1880 by Pete Daniel (review). Technology and Culture, 28(3):724–725, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889385/pdf.

Titus:1989:DTR

[Tit89] A. Costandina Titus. The Dragon’s Tail: Radiation Safety in the Manhattan Project, 1942–1946 by Barton C. Hacker (review). Technology and Culture, 30(1):133–134, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889122/pdf.

Trimble:1988:LAA

[TL88] William F. Trimble and W. David Lewis. Lytle S. Adams, the apostle of nonstop airmail pickup. Technology and Culture, 29(2):247– 265, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889251/pdf.

Tarr:1983:PMO

[TLK^{+}83] Joel A. Tarr, W. David Lewis, Carole Kitti, Edward W. Constant II, Brooke Hindle, David A. Hounshell, Bradford B. Blaine, Nancy Stern, Helena Wright, Paul Durbin, Darwin Stapleton, Philip Scranton, Raymond H. Merritt, S. Reid Warren, Jr., Stephen H. Cutcliffe, John T. Greenwood, and Jack M. Holl. The Philadelphia Meeting, October 28–31, 1982. Technology and Culture, 24(3):469– 492, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891391/pdf.

Tarr:1984:WWR

[TMMY84] Joel A. Tarr, James McCurley III, Francis C. McMichael, and Terry Yosie. Water and wastes: A retrospective assessment of wastewater technology in the United States, 1800–1932. Technology and Culture, 25(2):226–263, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890036/pdf.

Tobey:1985:PPS

[Tob85] Ronald C. Tobey. A Patron for Pure Science: The National Science Foundation’s Formative Years, 1945–57 by J. Merton England (review). Technology and Culture, 26(1):126–128, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889834/pdf.

Todd:1980:SCH

[Tod80] Edmund N. Todd. The Siemens Company: Its Historical Role in the Progress of Electrical Engineering by Sigfrid von Weiher, Herbert Goetzeler (review). Technology and Culture, 21(1):106–107, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890620/pdf.

Todd:1984:ESR

[Tod84a] Edmund N. Todd. Die Eisen- und Stahlindustrie des Ruhrgebiets 1879–1914 by Wilfried Feldenkirchen (review). Technology and Culture, 25(3):647–648, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890003/pdf.

Todd:1984:MHI

[Tod84b] Edmund N. Todd. Militarism: The History of an International Debate 1861–1979 by Volker R. Berghahn (review). Technology and Culture, 25(2):313–314, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890044/pdf.

Todd:1985:TIG

[Tod85] Edmund N. Todd. Technih, Ingenieure und Gesellschaft: Geschichte des Vereins Deutscher Ingenieure 1856–1981 ed. by Karl-Heinz Ludwig, Wolfgang König (review). Technology and Culture, 26(3):658–659, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889743/pdf.

Todd:1986:EIC

[Tod86] Edmund N. Todd. Electricity, Industry and Class in South Africa by Renfrew Christie (review). Technology and Culture, 27(2):313– 314, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889615/pdf.

Todd:1987:MFW

[Tod87] Edmund N. Todd. “Mein Feld ist die Welt”: Musterbücher und Katologe 1784–1914 ed. by Ottfried Dascher, et al. (review). Technology and Culture, 28(1):147–148, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889467/pdf.

Thackray:1985:CMN

[TOE^{+}85] Arnold Thackray, John E. O’Connor, James M. Edmonson, Eda Kranakis, Leonard S. Reich, Nathan Sivin, Edmund N. Todd, Stephen Victor, Eric Robinson, Arnold Thackray, Robert Friedel, Joseph J. Corn, Cecil O. Smith, Jr., John Kasson, Donald Hoke, P. Thomas Carroll, Julia C. Bonham, Thomas F. Gieryn, Robert C. Post, Alfred D. Chandler, Jr., Mark Wilde, Pamela E. Mack, and Howard Rosen. The Cambridge Meeting, November 1–4, 1984. Technology and Culture, 26(3):590–619, July 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889723/pdf.

Tomayko:1983:SSC

[Tom83a] James E. Tomayko. The Shorter Science and Civilization in China by Colin A. Ronan (review). Technology and Culture, 24(3):547– 548, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891426/pdf.

Tomayko:1983:SPC

[Tom83b] James E. Tomayko. The stored-program concept: National Computer Conference, Houston, Texas, June 9, 1982. Technology and Culture, 24(4):660–663, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890125/pdf.

Tomayko:1984:TPS

[Tom84] James A. Tomayko. Technology, Politics, and Society in China by Rudi Volti (review). Technology and Culture, 25(3):681–683, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890022/pdf.

Tomayko:1985:ACT

[Tom85] James E. Tomayko. Ancient China’s Technology and Science by the Institute of the History of Natural Sciences, Chinese Academy of Sciences (review). Technology and Culture, 26(1):135–136, January 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889840/pdf.

Topper:1980:BAB

[Top80] David R. Topper. Borromini by Anthony Blunt (review). Technology and Culture, 21(1):95–97, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890614/pdf.

Trachtenberg:1985:HTB

[Tra85] Alan Trachtenberg. Half a Truth Is Better than None: Some Unsystematic Conjectures about Art, Disorder, and American Experience by John A. Kouwenhoven (review). Technology and Culture, 26(2):316–317, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889786/pdf.

Travis:1989:RPB

[Tra89] Anthony S. Travis. The Royal Purple and the Biblical Blue: Argaman and Tekhelet ed. by Ehud Spanier (review). Technology and Culture, 30(3):667–668, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889036/pdf.

Trenn:1985:PAS

[Tre85] Thaddeus J. Trenn. Project Y: The Los Alamos Story. Part 1: Toward Trinity by David Hawkins, and: Part 2: Beyond Trinity by Edith C. Truslow, Ralph C. Smith (review). Technology and Culture, 26(3):672–674, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889753/pdf.

Treadwell:1987:GRC

[Tre87] James Treadwell. The Great Rifle Controversy: Search for the Ultimate Infantry Weapon from World War II through Vietnam and, Beyond by Edward C. Ezell (review). Technology and Culture, 28(4):881–882, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889329/pdf.

Trimble:1983:APF

[Tri83] William F. Trimble. Aviation and Pennsylvania by Frank Kingston Smith, James P. Harrington (review). Technology and Culture, 24 (2):294–296, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891282/pdf.

Trimble:1989:AIP

[Tri89] William F. Trimble. Another Icarus: Percy Pilcher and the Quest for Flight by Philip Jarrett (review). Technology and Culture, 30(1):169–170, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889145/pdf.

Tropp:1984:CBP

[Tro84a] Henry S. Tropp. Charles Babbage: Pioneer of the Computer by Anthony Hyman (review). Technology and Culture, 25(3):651–653, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL http://www.jstor.org/stable/3104217;
https://muse.jhu.edu/pub/1/article/890005/pdf.

Trout:1984:JWH

[Tro84b] William E. Trout III. Juniper Waterway: A History of the Albemarle and Chesapeake Canal by Alexander Crosby Brown, and: Home on the Canal by Elizabeth Kytle (review). Technology and Culture, 25(3):658–659, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890009/pdf.

Tropp:1987:IST

[Tro87] Henry Tropp. The Inmos Saga: A Triumph of National Enterprise by Mick McLean, Tom Rowland, and: Looking Back to Tomorrow: Reflections on Twenty-five Years of Computers in New Zealand ed. by W. R. Williams (review). Technology and Culture, 28(4):891–893, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889336/pdf.

Tucker:1980:EBN

[Tuc80a] D. Gordon Tucker. Electricity before Nationalisation: A Study of the Development of the Electricity Supply Industry in Britain to 1948 by Leslie Hannah (review). Technology and Culture, 21 (2):254–257, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890688/pdf.

Tucker:1980:CPC

[Tuc80b] Richard Tucker. The Challenge Presented to Cultures by Science and Technology by Jean Ladrière, and: Science, Technology and the Social Order ed. by Ward Morehouse (review). Technology and Culture, 21(4):710–712, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890791/pdf.

Tucker:1982:TCL

[Tuc82a] Barbara M. Tucker. Twentieth Century Limited: Industrial Design in America, 1925–1939 by Jeffrey L. Meikle (review). Technology and Culture, 23(4):680–682, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890270/pdf.

Tucker:1982:DJN

[Tuc82b] D. Gordon Tucker. Diamond Jubilee of the Newcomen Society: The Science Museum, London, November 29, 1980. Technology and Culture, 23(1):73–77, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891105/pdf.

Tullos:1985:MMM

[Tul85] Allen Tullos. Miners, Millhands, and Mountaineers: Industrialization of the Appalachian South, 1880–1930 by Ronald D. Eller (review). Technology and Culture, 26(1):96–97, January 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889820/pdf.

Turner:1985:GCE

[Tur85] Robert D. Turner. “The Great CPR Exposition” at the Glenbow– Alberta Institute, Calgary, Alberta. Technology and Culture, 26 (2):257–261, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889765/pdf.

Turner:1989:BHG

[Tur89] Gerard L’E. Turner. A Brief History of Geomagnetism and a Catalog of the Collections of the National Museum of American History by Robert P. Multhauf, Gregory Good (review). Technology and Culture, 30(3):681–682, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889046/pdf.

Tuttle:1981:BIS

[Tut81] William M. Tuttle, Jr. The birth of an industry: The synthetic rubber “Mess” in World War II. Technology and Culture, 22 (1):35–67, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890797/pdf.

Tweedale:1986:MTC

[Twe86] Geoffrey Tweedale. Metallurgy and technological change: A case study of Sheffield specialty steel and America, 1830–1930. Technology and Culture, 27(2):189–222, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889582/pdf.

Tweedale:1988:DFT

[Twe88] Geoffrey Tweedale. “Days at the factories”: A tour of Victorian industry with the Penny magazine. Technology and Culture, 29(4):888–903, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889162/pdf.

Sun:1982:CIR

[tZS82] E tu Zen Sun. China’s Industrial Revolution: Politics, Planning, and Management, 1949 to the Present by Stephen Andors (review). Technology and Culture, 23(3):483–485, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891196/pdf.

Unger:1981:WCS

[Ung81] Richard W. Unger. Warships and cargo ships in Medieval Europe. Technology and Culture, 22(2):233–252, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890843/pdf.

Unger:1983:TWC

[Ung83] Richard W. Unger. Time, Work, and Culture in the Middle Ages by Jacques Le Goff (review). Technology and Culture, 24 (1):118–119, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891230/pdf.

Unger:1988:ZWW

[Ung88] Richard W. Unger. Zeewezen en wetenschap: De wetenschap en de ontwikkeling van de navigatietechniek in Nederland tussen 1585 en 1815 by C. A. Davids (review). Technology and Culture, 29 (4):949–950, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889186/pdf.

Upton:1983:BDS

[Upt83a] Dell Upton. Building the Dream: A Social History of Housing in America by Gwendolyn Wright (review). Technology and Culture, 24(2):283–285, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891276/pdf.

Upton:1983:HHP

[Upt83b] Dell Upton. Hearth and Home: Preserving a People’s Culture by George W. McDaniel (review). Technology and Culture, 24(2):281– 283, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891275/pdf.

Upton:1988:CRM

[Upt88] Dell Upton. The Colossus of Roads: Myth and Symbol along the American Highway by Karal Ann Marling (review). Technology and Culture, 29(3):702–703, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889259/pdf.

Usselman:1988:WIR

[Uss88] Steven W. Usselman. The World of the Industrial Revolution: Comparative and International Aspects of Industrialization ed. by Robert Weible (review). Technology and Culture, 29(4):966–968, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889197/pdf.

Ridout:1988:AML

[V88] Orlando Ridout V. The Architecture of Migration: Log Construction in the Ohio Country, 1750–1850 by Donald A. Hutslar (review). Technology and Culture, 29(2):302–304, April 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889278/pdf.

Vaczek:1981:DES

[Vac81] Louis Vaczek. Daedalus ed. by Stephen R. Graubard (review). Technology and Culture, 22(3):605–607, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890894/pdf.

Vaczek:1982:SSH

[Vac82] Louis Vaczek. The Search for Solutions by Horace Freeland Judson (review). Technology and Culture, 23(1):97–98, January 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891114/pdf.

Vaczek:1984:FTI

[Vac84] Louis Vaczek. Fox Talbot and the Invention of Photography by Gail Buckland (review). Technology and Culture, 25(1):115–116, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890087/pdf.

vanBerkel:1983:TLM

[vB83] K. van Berkel. Het technisch labyrint: Een maatschappijgeschiedenis van drie industriële revoluties [The technological labyrinth: a social history of three industrial revolutions] ed. by Maarten Pieterson (review). Technology and Culture, 24(4):672–673, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890127/pdf.

vanBerkel:1986:TBV

[vB86] K. van Berkel. Textielhistorische Bijdragen. Vol. 24 ed. by E. J. Fischer, et al. (review). Technology and Culture, 27(4):879–880, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889534/pdf.

vanderLinden:1983:EMS

[vdL83] Robert van der Linden. Eagles of Mitsubishi: The Story of the Zero Fighter by Jiro Horikoshi (review). Technology and Culture, 24(2):299–300, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891285/pdf.

Veysey:1989:AKG

[Vey89] Laurence Veysey. To Advance Knowledge: The Growth of American Research Universities, 1900–1940 by Roger L. Geiger (review). Technology and Culture, 30(2):499–501, April 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889108/pdf.

Victor:1984:IHT

[Vic84] Stephen Victor. The Industrial Heritage: The Third International Conference on the Conservation of Industrial Monuments ed. by Marie Nisser (review). Technology and Culture, 25(2):343–344, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890061/pdf.

Victor:1987:WWE

[Vic87] Stephen Victor. “Windows on the works” at the Eli Whitney Museum, Hamden, Connecticut. Technology and Culture, 28(3):675– 679, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889358/pdf.

Vietor:1983:NSO

[Vie83a] Richard Vietor. North Sea Oil and Environmental Planning: The United Kingdom Experience by Ian R. Manners (review). Technology and Culture, 24(3):528–530, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891413/pdf.

Vietor:1983:CRD

[Vie83b] Richard H. K. Vietor. Competition and Regulation: The Development of Oligopoly in the Meat Packing Industry by Mary Yeager (review). Technology and Culture, 24(2):274–275, April 1983. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891270/pdf.

Vigrass:1983:LIH

[Vig83] J. William Vigrass. Labyrinths of Iron: A History of the World’s Subways by Benson Bobrick (review). Technology and Culture, 24 (2):288–289, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891278/pdf.

Vincenti:1982:CVA

[Vin82] Walter G. Vincenti. Control-volume analysis: A difference in thinking between engineering and physics. Technology and Culture, 23(2):145–174, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891142/pdf.

Vincenti:1984:TKS

[Vin84] Walter G. Vincenti. Technological knowledge without science: The innovation of flush riveting in American airplanes, ca. 1930–ca. 1950. Technology and Culture, 25(3):540–576, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889992/pdf.

Vincenti:1986:DWP

[Vin86a] Walter G. Vincenti. The Davis wing and the problem of airfoil design: Uncertainty and growth in engineering knowledge. Technology and Culture, 27(4):717–758, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889497/pdf.

Vinikas:1986:DAW

[Vin86b] Vincent Vinikas. The Day after World War III by Edward Zuckerman (review). Technology and Culture, 27(3):616–618, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889558/
pdf.

Volti:1987:NMM

[Vol87] Rudi Volti. The National Motorcycle Museum, Birmingham, England. Technology and Culture, 28(1):96–98, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889448/pdf.

Volti:1988:CTT

[Vol88] Rudi Volti. China Takes Off: Technology Transfer and Modernization by E. E. Bauer (review). Technology and Culture, 29(1): 180–181, January 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888534/pdf.

vanSchelven:1980:DSB

[vS80] Arnout L. van Schelven. Dutch Shipbuilding before 1800: Ships and Guilds by Richard W. Unger (review). Technology and Culture, 21(3):488–491, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890716/pdf.

vonWeiher:1983:TGH

[vW83] Sigirid von Weiher. Technik-Geschichte: Historische Beiträge und neuere Ansätze ed. by Ulrich Troitzsch, Gabriele Wohlauf (review). Technology and Culture, 24(1):124–125, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891234.

Walsh:1981:CFF

[Wal81] Margaret Walsh. Cutting a Fashionable Fit: Dressmakers’ Drafting Systems in the United States by Claudia B. Kidwell (review). Technology and Culture, 22(4):802–803, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890956/pdf.

Wallace:1982:PCA

[Wal82] Michael Wallace. Preservation Comes of Age: From Williamsburg to the National Trust, 1926–1949 by Charles B. Hosmer, Jr. (review). Technology and Culture, 23(4):686–691, October 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890273/pdf.

Walker:1983:KKT

[Wal83] J. Samuel Walker. Kennedy, Khrushchev and the Test Ban by Glenn T. Seaborg(review). Technology and Culture, 24(2):302–303, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891287/pdf.

Walker:1987:BEL

[Wal87a] J. Samuel Walker. By the Bomb’s Early Light: American Thought and Culture at the Dawn of the Atomic Age by Paul Boyer (review). Technology and Culture, 28(2):393–394, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889442/pdf.

Wallace:1987:SCS

[Wal87b] Anthony F. C. Wallace. Silk City: Studies on the Paterson Silk Industry, 1860–1940 ed. by Philip B. Scranton (review). Technology and Culture, 28(2):365, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889420/pdf.

Walton:1988:ISI

[Wal88] Whitney Walton. Industrialisation and Social Inequality in 19th-Century Europe by Hartmut Kaelble (review). Technology and Culture, 29(3):690–691, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889245/pdf.

Walker:1989:NPE

[Wal89] J. Samuel Walker. Nuclear power and the environment: The Atomic Energy Commission and thermal pollution, 1965–1971. Technology and Culture, 30(4):964–992, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/888971/pdf.

Wang:1989:MAB

[Wan89] Zuoyue Wang. The Making of the Atomic Bomb by Richard Rhodes (review). Technology and Culture, 30(4):1078–1081, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889004/pdf.

Warner:1982:WSC

[War82] Deborah Jean Warner. Women and the Social Costs of Economic Development by Elizabeth Moen, et al. (review). Technology and Culture, 23(3):501, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891206/pdf.

Warner:1983:WHW

[War83] Deborah Jean Warner. Out to Work: A History of Wage-earning Women in the United States by Alice Kessler-Harris (review). Technology and Culture, 24(4):686–687, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890134/pdf.

Warner:1984:WSA

[War84] Deborah Jean Warner. Women Scientists in America: Struggles and Strategies to 1940 by Margaret W. Rossiter (review). Technology and Culture, 25(1):129–130, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890095/pdf.

Warner:1986:EIC

[War86a] Sam Bass Warner, Jr. English Industrial Cities of the Nineteenth Century: A Social Geography by Richard Dennis (review). Technology and Culture, 27(4):882–883, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889538/pdf.

Warren:1986:SBB

[War86b] Kenneth Warren. Steelmaking before Bessemer. Vol. 1: Blister Steel: The Birth of an Industry; vol. 2: Crucible Steel: The Growth of Technology by K. C. Barraclough (review). Technology and Culture, 27(3):624–626, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889569/pdf.

Warner:1987:NCS

[War87a] Deborah Jean Warner. Nineteenth-Century Scientific Instruments and Their Makers ed. by P. R. de Clercq (review). Technology and Culture, 28(4):855, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889312/pdf.

Warner:1987:WMP

[War87b] Deborah Jean Warner. Workshop on the Military and Postwar Academic Science — Johns Hopkins University, Baltimore, Maryland, April 17–18, 1986. Technology and Culture, 28(3):680–682, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889359/pdf.

Ward:1988:ARN

[War88] Barbara Mclean Ward. “After the revolution” and “New and different” — exhibitions at the National Museum of American History, Washington, D.C. Technology and Culture, 29(3):613–618, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889212/pdf.

Ward:1989:GNR

[War89a] James A. Ward. The Great Northern Railway: A History by Ralph W. Hidy, et al. (review). Technology and Culture, 30(4):1057–1058, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
888990/pdf.

Warner:1989:DCH

[War89b] Deborah Jean Warner. The Divided Circle: A History of Instruments for Astronomy, Navigation and Surveying by J. A. Bennett (review). Technology and Culture, 30(3):680–681, July 1989. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889045/pdf.

Watson:1980:EBQ

[Wat80] Sara Ruth Watson. The Eads Bridge by Quinta Scott, Howard S. Miller (review). Technology and Culture, 21(1):97–100, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890615/pdf.

Watson:1981:RBF

[Wat81] Sara Ruth Watson. The rate of bridge failure. Technology and Culture, 22(4):847–848, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890976/pdf.

Wayman:1989:MCC

[Way89] Michael L. Wayman. The Mexican Colonial Copper Industry by Elinore M. Barrett (review). Technology and Culture, 30(1): 156–157, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889137/pdf.

Weaver:1983:OWT

[Wea83] Bobby Weaver. Oil in West Texas and New Mexico: A Pictorial History of the Permian Basin by Walter Rundell, Jr. (review). Technology and Culture, 24(3):530–531, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891414/pdf.

Webb:1980:BWN

[Web80a] George E. Webb. Blowing on the Wind: The Nuclear Test Ban Debate, 1954–1960 by Robert A. Divine (review). Technology and Culture, 21(1):148–149, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890644/pdf.

Webb:1980:PEF

[Web80b] George E. Webb. Planetary Encounters: The Future of Unmanned Spaceflight by Robert M. Powers (review). Technology and Culture, 21(2):267–268, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890694/pdf.

Weber:1981:SGM

[Web81a] Wolfhard Weber. Die Schemnitzer Gedenkbibliothek von Miskolc in Ungarn by Lásló Zsámboki (review). Technology and Culture, 22(2):365–366, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890879/pdf.

Weber:1981:TGI

[Web81b] Wolfhard Weber. Technik und Gesellschaft im 19. und 20. Jahrhundert ed. by Reinhard Rürup (review). Technology and Culture, 22(2):363–365, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890878/pdf.

Weber:1981:TRS

[Web81c] Wolfhard Weber. Town in the Ruhr: A Social History of Bochum, 1860–1914 by David Crew (review). Technology and Culture, 22 (3):626–628, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/890905/pdf.

Weber:1984:KMT

[Web84] Wolfhard Weber. Karl Marx: Die technologisch-historischen Exzerpte ed. by Hans-Peter Müller, and: Karl Marx: Exzerpte über Arbeitsteilung, Maschinerie und Industrie ed. by Rainer Winkelmann (review). Technology and Culture, 25(3):648–651, July 1984.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890004/
pdf.

Webb:1985:FYS

[Web85a] David Webb. The First 25 Years in Space: A Symposium ed. by Allan A. Needed, and: Space Applications at the Crossroads: 21st Goddard Memorial Symposium ed. by John H. McElroy, E. Larry Heacock (review). Technology and Culture, 26(2):346–349, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889803/pdf.

Weber:1985:TRH

[Web85b] Wolfhard Weber. Technik Und Recht: Zur Historischen Entwicklung Rechtlicher Regelungen Im Bereich Der Technik — Verein Deutscher Ingenieure, Düsseldorf, February 21–22, 1985. Technology and Culture, 26(4):824–825, October 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889680/pdf.

Weber:1987:BTE

[Web87a] Wolfhard Weber. Blätter für Technikgeschichte ed. by Rolf Niederhuemer (review). Technology and Culture, 28(3):692, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889365/
pdf.

Weber:1987:BTF

[Web87b] Wolfhard Weber. Die Bewertung des Technischen Fortschritts: Eine systematische Übersicht der Theorien by Johan Hendrik Jacob van der Pot (review). Technology and Culture, 28(2):397–398, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889447/pdf.

Weil:1980:TTT

[Wei80a] Vivian M. Weil. A Thousand Thoughts on Technology and Human Values by Edward J. Gallagher (review). Technology and Culture, 21(4):700–701, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890787/pdf.

Weinstein:1980:STD

[Wei80b] Jay Weinstein. Science, Technology and Development. Essays in Honour of A. Rahman ed. by K. D. Sharma, M. A. Quershi (review). Technology and Culture, 21(1):153–156, January 1980. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/890647/pdf.

Weinstein:1981:ATD

[Wei81a] Jay Weinstein. Appropriate Technology for Development: A Discussion and Case Histories ed. by Donald D. Evans, Laurie Hogg Adler, and: Foundations of Gandhian Economics by Amritananda Das (review). Technology and Culture, 22(3):688–691, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890932/
pdf.

Weinstein:1981:DIP

[Wei81b] Jay Weinstein. In Defense of the Irrational Peasant: Indian Agriculture after the Green Revolution by Kusum Nair (review). Technology and Culture, 22(1):226–229, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890840/pdf.

Weil:1982:RFG

[Wei82] Vivian M. Weil. Responsibilities to Future Generations ed. by Ernest Partridge (review). Technology and Culture, 23(2):279–281, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891177/pdf.

Weible:1984:ICL

[Wei84] Robert Weible. The industrial city: Lowell Conference on Industrial History, University of Lowell, April 29–30, 1983. Technology and Culture, 25(1):103–108, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890083/pdf.

Weil:1986:ESE

[Wei86] Vivian Weil. Engineering in Society: Engineering Education and Practice in the United States by the Panel on Engineering Interactions with Society (review). Technology and Culture, 27(4): 840–842, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889506/pdf.

Weingart:1987:OTN

[Wei87a] Peter Weingart. Organized Technology: Networks and Innovation in Technical Systems by Wesley Shrum (review). Technology and Culture, 28(2):396–397, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889445/pdf.

Weinstein:1987:UAE

[Wei87b] Barbara Weinstein. Underdeveloping the Amazon: Extraction, Unequal Exchange, and the Failure of the Modern State by Stephen G. Bunker (review). Technology and Culture, 28(3):726–728, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889386/pdf.

Wells:1987:EHH

[Wel87a] Ellen B. Wells. An Early History of Horsemanship by A. Azzaroli (review). Technology and Culture, 28(2):348, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889409/pdf.

Welsh:1987:CEA

[Wel87b] Peter C. Welsh. The Craftsman in Early America ed. by Ian M. G. Quimby (review). Technology and Culture, 28(4):865–867, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889319/pdf.

Welsh:1989:SDY

[Wel89] Michael Welsh. The Southwestern Division: 50 Years of Service by D. Clayton Brown (review). Technology and Culture, 30(2):490– 491, April 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889102/pdf.

Wertime:1981:OAT

[Wer81] Theodore A. Wertime. The origin of agriculture and technology: Aarhus, Denmark, November 21–25, 1978. I. Introductory note: Whence early technology? East, West, or nowhere? Technology and Culture, 22(1):122–124, January 1981. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890801/pdf.

Wertime:1982:ISG

[Wer82] Theodore A. Wertime. Internationales Symposium zur Geschichte des Bergbaus und Hüttenwesens ed. by Eberhard Wächtler, Gisela-Ruth Engewald (review). Technology and Culture, 23(2):232–233, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891153/pdf.

Westbrook:1983:MIF

[Wes83] J. H. Westbrook. Men of Iron: Forbes and Adam by Kenneth T. Howell, Einar W. Carlson (review). Technology and Culture, 24(4):698–699, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890142/pdf.

Wesson:1984:FSC

[Wes84] Morgan Wesson. Foundations of the Stereoscopic Cinema: A Study in Depth by Lenny Lipton (review). Technology and Culture, 25(3):689–690, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890026/pdf.

Westbrook:1986:MTG

[Wes86a] J. H. Westbrook. Mills on the Tsatsawassa: A Guide for Local Historians by Philip L. Lord, Jr., Martha A. Costello (review). Technology and Culture, 27(3):632–634, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889577/pdf.

Westby:1986:AIE

[Wes86b] Robert Westby. Agricultural Innovation in the Early Islamic World: The Diffusion of Crops and Farming Techniques, 700– 1100 by Andrew M. Watson (review). Technology and Culture, 27 (2):293–294, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889593/pdf.

Wess:1987:IHT

[Wes87a] Robert C. Wess. Interface ’85: Humanities and Technology — Atlanta, Georgia, October 17–18, 1985. Technology and Culture, 28 (2):343–344, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889406/pdf.

Westbrook:1987:FGA

[Wes87b] J. H. Westbrook. A Field Guide to American Windmills by T. Lindsay Baker (review). Technology and Culture, 28(3):721–722, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889383/pdf.

Westbrook:1987:BRH

[Wes87c] J. H. Westbrook. The Bells of Russia: History and Technology by Edward V. Williams (review). Technology and Culture, 28(4):853–854, October 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889311/pdf.

Westfall:1989:FAB

[Wes89] Richard S. Westfall. Floods along the Bisenzio: Science and technology in the Age of Galileo. Technology and Culture, 30(4): 879–907, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888968/pdf.

Wheaton:1983:BAA

[Whe83] Bruce. R. Wheaton. Bringing Aerodynamics to America by Paul A. Hanle (review). Technology and Culture, 24(1):142–144, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891244/pdf.

Wheaton:1984:FSP

[Whe84] Bruce R. Wheaton. Fusion: Science, Politics, and the Invention of a New Energy Source by Joan Lisa Bromberg (review). Technology and Culture, 25(1):154–156, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/890110/pdf.

Wheaton:1987:MMS

[Whe87] Bruce R. Wheaton. The Man-Made Sun: The Quest for Fusion Power by T. A. Heppenheimer (review). Technology and Culture, 28(2):394–395, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889443/pdf.

White:1980:ILF

[Whi80] John H. White. Industrial locomotives: The forgotten servant. Technology and Culture, 21(2):209–216, April 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890674/pdf.

Whitaker:1981:ERA

[Whi81] James W. Whitaker. Electricity for Rural America: The Fight for the REA by D. Clayton Brown (review). Technology and Culture, 22(2):315–316, April 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890856/pdf.

White:1982:SBR

[Whi82a] John H. White. Safety with a bang: The railway torpedo. Technology and Culture, 23(2):195–201, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/891144/pdf.

White:1982:SME

[Whi82b] Lynn White, Jr. The Ship in the Medieval Economy, 600–1600 by Richard W. Unger (review). Technology and Culture, 23(2):223– 227, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891149/pdf.

White:1982:TFP

[Whi82c] William White, Jr. Technology and the Future: A Philosophical Challenge by Egbert Schuurman (review). Technology and Culture, 23(1):89–91, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891110/pdf.

White:1983:RJT

[Whi83a] J. H. White. The Railway Journey: Trains and Travel in the 19th Century by Wolfgang Schivelbusch (review). Technology and Culture, 24(3):515, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891404/pdf.

White:1983:WUH

[Whi83b] Roger B. White. Whither the urban history exhibit? The Peale Museum’s “Rowhouse”. Technology and Culture, 24(1):76–90, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891223/pdf.

White:1983:CSR

[Whi83c] John H. White, Jr. The California State Railroad Museum: A Louvre for locomotives. Technology and Culture, 24(4):644–654, October 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890123/pdf.

White:1984:TUT

[Whi84a] Roger B. White. The Taxicab: An Urban Transportation Survivor by Gorman Gilbert, Robert E. Samuels (review). Technology and Culture, 25(1):165–166, January 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890117/pdf.

White:1984:VSJ

[Whi84b] Roger B. White. The VW Story by Jerry Sloniger, and: The VW Beetle by Robin Fry (review). Technology and Culture, 25(2):365– 367, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890072/pdf.

Whiteley:1985:ADE

[Whi85] Nigel Whiteley. American Design Ethic: A History of Industrial Design to 1940 by Arthur J. Pulos (review). Technology and Culture, 26(2):313–315, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889785/pdf.

White:1986:SSG

[Whi86a] John H. White, Jr. Steam in the streets: The grice and long dummy. Technology and Culture, 27(1):106–109, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889637/
pdf.

Whiteley:1986:DSE

[Whi86b] Nigel Whiteley. Design since 1945 ed. by Kathryn B. Hiesinger, George H. Marcus (review). Technology and Culture, 27(2):334– 336, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889627/pdf.

Wik:1980:HFI

[Wik80] Reynold M. Wik. Henry Ford: “Ignorant Idealist” by David E. Nye (review). Technology and Culture, 21(4):686–688, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890779/
pdf.

Wik:1981:DWG

[Wik81] Reynold M. Wik. The Devil’s Wagon in God’s Country by Michael L. Berger (review). Technology and Culture, 22(1):177–178, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890814/pdf.

Wik:1982:MSR

[Wik82] Reynold M. Wik. The Modern Stentors: Radio Broadcasters and the Federal Government, 1920–1934 by Philip T. Rosen (review). Technology and Culture, 23(1):126–128, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891133/pdf.

Wilson:1980:WCA

[Wil80] Raymond L. Wilson. Wayne County: The Aesthetic Heritage of a Rural Area by Stephen W. Jacobs (review). Technology and Culture, 21(4):673–674, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/890773/pdf.

Wilson:1981:AVA

[Wil81] Raymond L. Wilson. Authentication in the Visual Arts: A Multidisciplinary Symposium ed. by H. L. C. Jaffe, et al. (review). Technology and Culture, 22(1):174–176, January 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/890812/pdf.

Williams:1983:SEB

[Wil83] Trevor I. Williams. The State and the Emergence of the British Oil Industry by Geoffrey Jones (review). Technology and Culture, 24(1):132–134, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891238/pdf.

Wilkins:1984:TWM

[Wil84a] Mira Wilkins. Third World Multinationals: The Rise of Foreign Investment from Developing Countries by Louis T. Wells, Jr. (review). Technology and Culture, 25(4):869–871, October 1984. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889964/pdf.

Williams:1984:SFA

[Wil84b] Hector Williams. Ships and Fleets of the Ancient Mediterranean by Jean Rougé (review). Technology and Culture, 25(2):323–325, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890051/pdf.

Williams:1984:AEH

[Wil84c] L. Pearce Williams. Albert Einstein, Historical and Cultural Perspectives: The Centennial Symposium in Jerusalem ed. by Gerald Holton and Yehuda Elkana (review). Technology and Culture, 25(3):699–700, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890031/pdf.

Williams:1984:BCS

[Wil84d] Rosalind Williams. The Birth of a Consumer Society: The Commercialization of Eighteenth-Century England by Neil McKendrick, et al. (review). Technology and Culture, 25(2):337–339, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890058/pdf.

Williams:1985:SOP

[Wil85a] L. Pearce Williams. Science Observed: Perspectives on the Social Study of Science ed. by Karin D. Knorr-Cetina, Michael Mulkay (review). Technology and Culture, 26(2):294–298, April 1985. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889776/pdf.

Williams:1985:RPD

[Wil85b] Michael Williams. Reckoners: The Prehistory of the Digital Computer, from Relays to the Stored Program Concept, 1935–1945 by Paul E. Ceruzzi (review). Technology and Culture, 26(2):339–341, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889799/pdf.

Wilde:1986:OSM

[Wil86a] Mark Wilde. The other side of the mountains: Mining at the Adirondack Museum. Technology and Culture, 27(2):285–290, April 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889589/pdf.

Wilson:1986:TCC

[Wil86b] David B. Wilson. Technology and the Character of Contemporary Life: A Philosophical Inquiry by Albert Borgmann (review). Technology and Culture, 27(4):907–908, October 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889571/pdf.

Wilde:1987:SPP

[Wil87a] Mark Wilde. Sweetness and Power: The Place of Sugar in Modern History by Sidney W. Mintz (review). Technology and Culture, 28(1):141–143, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/889464/pdf.

Williams:1987:AIV

[Wil87b] Michael Williams. Artificial Intelligence: The Very Idea by John Haugeland (review). Technology and Culture, 28(3):706–707, July 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889374/pdf.

Williams:1987:MSA

[Wil87c] Rosalind Williams. The Morality of Spending: Attitudes toward the Consumer Society in America, 1875–1940 by Daniel Horowitz (review). Technology and Culture, 28(2):375–376, April 1987. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/889427/pdf.

Wilde:1988:GYF

[Wil88] Mark Wilde. The Great Yellow Fleet: A History of American Railroad Refrigerator Cars by John H. White (review). Technology and Culture, 29(2):312–313, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889284/pdf.

Wilde:1989:PCL

[Wil89a] Mark Wilde. Prime Cut: Livestock Raising and Meatpacking in the United States, 1607–1983 by Jimmy M. Skaggs (review). Technology and Culture, 30(1):145–146, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889129/pdf.

Williams:1989:MUS

[Wil89b] James C. Williams. The Machine in the University: Sample Course Syllabi for the History of Technology and Technology Studies ed. by Terry S. Reynolds (review). Technology and Culture, 30(3):715–716, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889082/pdf.

Winters:1982:MAR

[Win82] Stanley B. Winters. The Modernization of Agriculture: Rural Transformation in Hungary, 1848–1975 ed. by Joseph Held (review). Technology and Culture, 23(2):245–248, April 1982. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/891160/pdf.

Winters:1987:ALM

[Win87] Stanley B. Winters. The Achievement of Lewis Mumford — New Jersey Institute of Technology, Newark, October 14–17, 1985. Technology and Culture, 28(1):106–112, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889452/pdf.

Winpenny:1989:TEI

[Win89] Thomas R. Winpenny. The Transfer of Early Industrial Technologies to America by Darwin H. Stapleton (review). Technology and Culture, 30(3):685–687, July 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889048/pdf.

Wise:1980:NRP

[Wis80] George Wise. A new role for professional scientists in industry: Industrial research at general electric, 1900–1916. Technology and Culture, 21(3):408–429, July 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890704/pdf.

Wise:1981:AEE

[Wis81] George Wise. Advances in Electronics and Electron Physics ed. by L. and C. Marton, and: Evolution of Naval Radio-Electronics and Contributions of the Naval Research Laboratory by Louis A. Gebhard (review). Technology and Culture, 22(4):812–816, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890962/pdf.

Wise:1984:YRT

[Wis84] George Wise. 70 Years of Radio Tubes and Valves: A Guide for Electronic Engineers, Historians, and Collectors by John W. Stokes (review). Technology and Culture, 25(4):884–885, October 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889974/pdf.

Withuhn:1983:DDW

[Wit83] William L. Withuhn. D-Day on the Western Pacific: A Railroad’s Decision to Dieselize by Virgil Staff (review). Technology and Culture, 24(3):524–526, July 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/891410/pdf.

Wolek:1983:ETI

[Wol83] Francis W. Wolek. Evaluating Technological Innovations by Bela Gold, et al. (review). Technology and Culture, 24(1):154–155, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891251/pdf.

Woods:1982:JMB

[Woo82] David L. Woods. John M. Brooke: Naval Scientist and Educator by George M. Brooke, Jr. (review). Technology and Culture, 23(4):673–676, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890266/pdf.

Woods:1986:AIM

[Woo86] David L. Woods. Arms and Independence: The Military Character of the American Revolution ed. by Ronald Hoffman, Peter J. Albert (review). Technology and Culture, 27(3):613–614, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889555/pdf.

Worthington:1981:AGE

[Wor81] Rick Worthington. Alternatives for Growth: The Engineering and Economics of Natural Resources Development ed. by Harvey Mc-Mains, Lyle Wilcox (review). Technology and Culture, 22(3):677– 680, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890928/pdf.

Workman:1986:AEW

[Wor86a] Michael E. Workman. The artifact as evidence: So what? — NMAH, Washington, D.C., November 29–30, 1984. Technology and Culture, 27(1):118–120, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889639/pdf.

Worthington:1986:GTW

[Wor86b] William Worthington. Growth through Welding: Perspectives on the History of a Norwegian Welding Firm — Norgas AS 1908– 1983 by Gunnar Nerheim (review). Technology and Culture, 27 (1):179–181, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889669/pdf.

Worster:1988:MEB

[Wor88] Donald Worster. The Magnificent Experiment: Building the Salt River Reclamation Project 1890–1917 by Karen L. Smith (review). Technology and Culture, 29(4):972–973, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/889200/pdf.

Wosk:1982:HMR

[Wos82] Julie Helen Wosk. Humanities and the machine: Responses to technology (State University of New York, Maritime College, Fort Schuyler, Bronx, April 3, 1981). Technology and Culture, 23 (3):435–438, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891184/pdf.

Wosk:1984:UES

[Wos84] Julie H. Wosk. A unique exhibition space: The Brooklyn Bridge Anchorage. Technology and Culture, 25(2):292–293, April 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890040/
pdf.

Wren:1989:FGM

[Wre89] Gary Wren. A Few Good Men from Univac by David E. Lundstrom (review). Technology and Culture, 30(1):173–174, January 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889147/
pdf.

Wright:1980:JMA

[Wri80] Monte D. Wright. The Jet Makers: The Aerospace Industry from 1945 to 1972 by Charles D. Bright (review). Technology and Culture, 21(1):135–136, January 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890637/pdf.

Wright:1983:MSS

[Wri83] Thomas Wright. Merehant Sailing Ships 1775–1815 by David R. MacGregor (review). Technology and Culture, 24(2):265–266, April 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891264/pdf.

Wright:1985:HAB

[Wri85a] Helena E. Wright. Historic America: Buildings, Structures, and Sites by Historic American Buildings Survey and Historic American Engineering Record (review). Technology and Culture, 26 (2):322–324, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889790/pdf.

Wright:1985:HFF

[Wri85b] Helena E. Wright. “Habs: The First Fifty Years (1933–1983)” — an Exhibition at the Library of Congress, Washington, D.C. Technology and Culture, 26(2):253–256, April 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/889764/pdf.

Wright:1986:PPN

[Wri86] Helena E. Wright. Pictures to Print: The Nineteenth-Century Engraving Trade by Anthony Dyson, and: Printmaking & Picture Printing: A Bibliographical Guide to Artistic & Industrial Techniques in Britain 1750–1900 by Gavin Bridson, Geoffrey Wakeman, and: A Century of American Printmaking, 1880–1980 by James Watrous (review). Technology and Culture, 27(3):629–632, July 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889574/pdf.

Yao:1988:TSN

[Yao88] Dennis Yao. Technology, Strategy and National Security ed. by Franklin D. Margiotta, Ralph Sanders (review). Technology and Culture, 29(2):330–331, April 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889296/pdf.

Yates:1981:AHM

[Yat81] W. Ross Yates. Alexander Holley and the Makers of Steel by Jeanne McHugh (review). Technology and Culture, 22(4):798–800, October 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
890954/pdf.

Yates:1982:HMC

[Yat82] W. Ross Yates. A History of Metals in Colonial America by James A. Mulholland (review). Technology and Culture, 23(3):495–497, July 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
891202/pdf.

York:1988:CAF

[Yor88] Neil L. York. The Craft Apprentice: From Franklin to the Machine Age in America by W. J. Rorabaugh (review). Technology and Culture, 29(4):915–916, October 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889166/pdf.

Young:1987:WMB

[You87a] Linda Young. Westralian mining brooches: Commemoration of a technology. Technology and Culture, 28(1):87–95, January 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/889446/
pdf.

Young:1987:TFM

[You87b] Otis E. Young, Jr. Technology on the Frontier: Mining in Old Ontario by Dianne Newell (review). Technology and Culture, 28 (2):361–362, April 1987. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/889417/pdf.

Zeek:1982:LHA

[Zee82] William C. Zeek. Lead in History and Art/Blei in Geschichte und Kunst by Vladimir W. Krysko (review). Technology and Culture, 23(4):665, October 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890261/pdf.

Zeek:1986:LLP

[Zee86] William Zeek. Lead and Lead Poisoning in Antiquity by Jerome O. Nriagu (review). Technology and Culture, 27(1):129–130, January 1986. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889643/pdf.

Ziegler:1989:RRE

[Zie89a] Charles A. Ziegler. RAND’s Role in the Evolution of Balloon and Satellite Observation Systems and Related U.S. Space Technology by Merton E. Davies, William R. Harris (review). Technology and Culture, 30(4):1081–1082, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889006/pdf.

Ziegler:1989:TPS

[Zie89b] Charles A. Ziegler. Technology and the process of scientific discovery: The case of cosmic rays. Technology and Culture, 30(4):

939–963, October 1989. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/888970/pdf.

Ziglar:1985:MNA

[Zig85] William L. Ziglar. Managing NASA in the Apollo Era by Arnold S. Levine (review). Technology and Culture, 26(3):671–672, July 1985. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
889752/pdf.

Ziglar:1988:CAM

[Zig88] William L. Ziglar. Chariots for Apollo: The Making of the Lunar Module by Charles R. Pellegrino, Joshua Stoff (review). Technology and Culture, 29(3):708–709, July 1988. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/889267/pdf.

Zimmerman:1984:PCH

[Zim84] Albright G. Zimmerman. Proceedings of the Canal History and Technology Symposium. Vol. 2: March 26, 1983 ed. by Lance E. Metz (review). Technology and Culture, 25(3):656–658, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/890008/
pdf.

Zuk:1980:TAT

[Zuk80] William Zuk. Tensile Architecture: From the Tent to the Bubble Dome by (Peter) Philip Drew (review). Technology and Culture, 21(4):691–693, October 1980. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/890782/pdf.

Zupko:1988:MMW

[Zup88] Ronald Edward Zupko. Measures and Men by Witold Kula (review). Technology and Culture, 29(1):141–142, January 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/888509/pdf.